A Complete Bibliography of Publications in Computer Physics Communications: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

10 June 2024 Version 1.37

Title word cross-reference

(1 + 1) [SP20, SLZ^{+}24]. (3 + 1) [DH20]. (zI − H)x = b [HKY^{+}21]. 1 [BGW^{+}22, JĎD^{+}21, LEE^{+}21a, MPH^{+}24, Sch21, TCSA21, WJB21, WKBW21, YYZ^{+}22, ZGW20]. 1/1 [ZJM^{+}21]. 2 [ALB22, BRAC23, BKRG22, CAC^{+}22, GUW^{+}22, LEE^{+}21a, LS22, NT24, RMCC22, RBFB20, SLDF^{+}21, SCL22, WWB22, WKBW21, YNV22]. 2/1 [ZJM^{+}21]. 3 [ATC^{+}23, BKP22, CZS^{+}21, CCW20, DFG^{+}23, DRZ^{+}21, DBH^{+}22, GTE21, GPM^{+}21, GSL^{+}23, GOS^{+}22, IGL^{+}24, JLW24, KWZ24, KNJ^{+}23, LG21b, LLZ^{+}23, LLH^{+}21b, MMP20, NUK^{+}22, NSY^{+}23, PMS^{+}20a, PGYF21, SBB^{+}24, WZC21, WT22, XFH^{+}22, XLG^{+}23, YNV22, YHY^{+}21, YKY^{+}22, ZL20]. 3 + 1 [Tor20]. 3j [XWYQ21]. ^{1}Σ [BGHC23]. ^{2} [YLH^{+}20]. ^{3} [UYS22]. _{3} [DSW^{+}23]. _{n} [FLW^{+}23]. B_{0} [BGW^{+}22]. C [FFTV23, NBCMH20]. δf [SNP^{+}20, YC20]. e [Dry21]. e^{+}e^{−} [ABWZ23, SY20]. f [BKRG22, CJD^{+}20, Ken24, MSU^{+}21, SBB^{+}24]. f = 1 [BKRG22]. F_{2} [ABF^{+}23a]. γ [BKS24b, Kut24]. HΦ [IKM^{+}24]. i [MEH21]. jj [GF21]. K [Bzo21, Ken24]. κ [SLIC24]. k · p [ZYL^{+}23]. Kω [HKY^{+}21]. L [FFTV23]. L^{2} [Yan24a]. M_{2} [Par21]. M_{n+1} [FLW^{+}23]. μ [Bie21]. N [ABB^{+}24, GSBN22]. ν [Bie21, GLW^{+}23]. O(a) [FFLR20]. O(N) [Kut20, LYX22]. O^{2} [KvH21]. p_{T} [AFJ^{+}23]. ∂ [MFS^{+}22]. q [MMP^{+}24]. R [BAB^{+}20, FMCB^{+}20, MBG^{+}20, PFG22, WBvdH20, BBV^{+}22, BL21]. S [BDdM20, CKT21]. S_{N} [LEE^{+}21a, WLL22]. SO(3) [ST23a]. SU(3) [DLD^{+}21]. T [BFCR24]. τ → 3π_{τ} [CV24]. ΘΦ [PTD20]. U(1) [SKM^{+}21]. X [ZWP^{+}22, ST23b]. Z′ [AES21].

-Adaptive [BBV^{+}22, BL21]. -component [ABB^{+}24]. -coupling [GF21]. -D [YYZ^{+}22, WZC21, XFH^{+}22, ZGW20]. -dimensional [DH20, SLZ^{+}24]. -explorer [AES21]. -from- [Bie21]. -function [BDdM20]. -improved [FFLR20]. -matrix [Kut24, WBvdH20, BAB^{+}20, FMCB^{+}20, MBG^{+}20]. -matrix-based [BFCR24]. -model [CKT21]. -point [LG21b]. -space [ZWP^{+}22]. -symbols [XWYQ21].

.Net [Gro22, GOS^{+}22].

/MOC [WLL22].

0D [DTC20]. 0D-2P [DTC20].

1 [Dau23, RDR^{+}20]. 1-D [Dau23]. 1-D/3-D [Dau23]. 1.0 [DCRF23, DMS^{+}22, OEI^{+}22, WBM^{+}24]. 1.2 [DP21]. 1.3 [Di 22]. 102 [IKM^{+}23]. 11-species [YKYK23]. 147 [MSHP20]. 147/3 [MSHP20]. 185 [AAT^{+}20]. 198 [Ano20c]. 1D [RDR^{+}20, YZHL22].

2 [KKM^{+}20, PSK^{+}24]. 2.0 [BMS^{+}23, BFMA^{+}24, CTPS22, CK23, FKS20, JZW^{+}22, KLMU21, Lat23, MSG^{+}21, NAZ^{+}22, RZY^{+}24, TRB^{+}20a, YKK^{+}23a]. 2.01 [KOT23]. 2.1 [ABB^{+}22]. 2.14 [BHH^{+}20]. 2.2 [SZY^{+}22]. 2.2.1 [SZY^{+}22]. 2/3D [PSK^{+}24]. 2020 [Ano20a, Ano20b, Ano20q, Ano20r, Ano20t, Ano20u, Ano20v, Ano20x, Ano20y]. 2021 [Ano21b, Ano21c, Ano21s, Ano21r, Ano21v, Ano21w, Ano21x]. 2022 [Ano22b, Ano22c, Ano22p, Ano22q, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x]. 2023 [Ano23a, Ano23b, Ano23c, Ano23p, Ano23q, Ano23s, Ano23r, Ano23t, Ano23u, Ano23w, Ano23v, Ano23x, CZB^{+}23, LZP^{+}24, MPZB^{+}24, SZT24, SVJ^{+}24, ZXT23]. 2024 [Ano24a, Ano24b, Ano24l, Ano24m, Ano24o, Ano24n, Ano24p, Ano24q, Ano24r, MBJB24]. 228 [SJY20]. 264 [MKPW22]. 269 [ZYX22]. 271 [YNV23]. 287 [Och23a]. 2D [YNV23, CAC^{+}22, JLL^{+}24, YZHL22]. 2D/1D [YZHL22]. 2D/3D [JLL^{+}24]. 2DMAT [MYM^{+}22]. 2HDECAY [KMS20b]. 2P [DTC20]. 2V [TCSA21].

3 [AKK^{+}24, DSSW22, DKM^{+}24, Hal21, KLM^{+}22, PMA^{+}21, SS21a, SLL^{+}24]. 3-D [Dau23]. 3.0 [HCP20, RSPJ21]. 3.01 [BKO20]. 3.1 [CGV^{+}22]. 30 [Pos22]. 3D [YNV23, JLL^{+}24, KAG24, PSK^{+}24, Yan23]. 3DO [GTE21].

3DO-SKMF [GTE21]. 3IDEA [YKYK23]. 3V [DFG^{+}23].

4 [EMM^{+}23]. 40 [DPR^{+}20]. 40-Feynman [DPR^{+}20]. 403.0 [DDM20].

5 [LOT^{+}20, LHG^{+}20]. 50th [SHB^{+}20]. 5D [Sit23].

6 [Gro22]. 6.0 [ABB^{+}24]. 6.5 [SZ24b]. 6D [CE22].

86 [LWC^{+}21].

9.3 [SMO20b]. 95 [LGBJ20]. 99-line [YJ24].

a-TDEP [BBB20]. AADIS [YZ20a]. AAfrag [KOT23]. Ab-Initio [vRCM21, BDP^{+}21, FN23, SLIC24, PP23]. ABINIT [LACL24, BCTS22, GAA^{+}20, BBB20]. absorbers [KJL^{+}23]. Accelerated [BRAC23, CDT22, MSY^{+}21, BMSP21, BMS^{+}23, CFLR24, Cos22, CESD^{+}23, DAC^{+}23, DRZ^{+}21, ECS23, FKK^{+}21, JMOC21, JLL^{+}24, KL23b, KMR22, KT23, KNJ^{+}23, LG23, MRH^{+}23, NT24, Nie20, Nie22a, RLW^{+}24, RZH22, RCW^{+}20, SXYD24, SCL22, VSM^{+}22, WWM^{+}22, WWJ^{+}22, WVK21, XBL^{+}20, XLG^{+}23, YLLW24, YLK21, YGSW21, ZPZH20, ZHI23]. Accelerating [CCM20, CCC20, CNS22, HSMR^{+}24, MKHT20, MPN^{+}21, SJ20]. Acceleration [GLB^{+}21, AUEO24, HLzY^{+}20, JS24, LAD^{+}21, LJQ^{+}22, RBWD^{+}24, SDL^{+}21, WZK^{+}24, zYMK^{+}21]. accelerator [ABK^{+}22, FA21, NBS^{+}20, TAY^{+}24, ZYG21]. accelerator-driven [ZYG21]. accelerators [HXS20]. acceptor [BMR^{+}24a]. accessible [AMK24, BBV23, MES^{+}24]. accuracy [AIZ23, GB22, Jab20, LMQ^{+}23, LWC^{+}21, Tan24, YM21]. Accurate [LVMGF^{+}23, ZRH21, BV21, CZWE23, DLD^{+}21, ENK24, Koz23, LN23, MT23, NRG24, PG23, TCY23, TLC^{+}21, VGGP^{+}21, XZLX20, Yan21b, ZCWY20]. Accurately [FN21, FO20, FO21, PPKK21]. ACFlow [Hua23]. acid [OSLC21]. acoustoelasticity [LLY^{+}22]. Acquisition [AAA^{+}20]. across [CCM20]. actin [HMM22]. acting [KT23]. actinides [TIG^{+}24]. actions [GNP20]. activation [DARJ23, PGS^{+}24]. active [Kür23, YYH21, ZLV23]. active-set [YYH21]. actuators [MOA24]. Adams [TSL21]. AdaPT [DHK^{+}21]. adaptability [WLL22]. Adaptable [DHK^{+}21]. adaptation [ZWG^{+}24]. adapted [GF21, GLSH21, MLD^{+}22]. Adaptive [MDZ24, NVCS23, SVSC20, SS22, TSL21, YWM23, ASC^{+}24, FAL20, FAL21, GP22, HMR22, KPR^{+}24, KSF^{+}22, MDR^{+}20, PD23, PG23, RCB^{+}20, SJP^{+}21b, Tan24, WJB21, WDMZ24, BBV^{+}22, BL21]. ADAQ [DIAA21]. ADBSat [SCMP^{+}22, SCR^{+}22b]. added [IKM^{+}24]. adding [KL22]. addition [GBH20, TCcN23]. additional [CCK23, GDB10, LG21b]. additive [SYFT23]. Additivity [WV22]. addressing [YCCW23]. ADG [ATRD21]. adhesive [QCZ23]. ADI [HPY21]. ADI-based [HPY21]. adiabatic [CGV^{+}22, SZT24, Xav22]. Adsorption [BRAC23, DvHSdS22, GAJK23, PRR23]. Advanced [KLM^{+}22, DFP^{+}20, FA20a, RMR^{+}22, WLL22, XLL^{+}22, LS21b]. advancements [OXOG23]. advection [PLF20]. aeroacoustics [FZD^{+}24, LIG^{+}24, MVAXP22]. aerodynamic [SCMP^{+}22, SCR^{+}22b]. aerosol [WNS^{+}21, WSK^{+}22, YFL22]. aerosols [MK22]. aerospace [BDR^{+}20]. aerothermodynamics [DFU20]. against [AES21]. agglomeration [FXQS21]. aggregate [Hal21]. Aggregates [FLK^{+}20]. aggregation [DRSZ23, YD20, ZJS^{+}20]. Agnostic [BHK^{+}21, Gon22]. AhKin [GRCT20]. AI [TAY^{+}24, ZHZ^{+}23]. AI-aided [ZHZ^{+}23]. AICON [FO20]. AICON2 [FO21]. aided [MG22, ZHZ^{+}23]. AIHFLTF [HY23]. AIMSim [BBV23]. air [SLL22, YKYK23]. Algebra [HW22, CB20, SCT21]. algebraic [ATC^{+}23, GF23, LWhK^{+}20, UZB22, UJ21, VÁFG^{+}22]. Algebras [FKS20]. Algorithm [HA23, RCS21, AAB^{+}21, AG21, AYWKL24, BKP22, Bar22a, Bel24, CL20, CZX^{+}21, DRSZ23, DC22, DWD^{+}24, GSBN22, IUJ21, Jab20, JLL^{+}24, KIK20, KRE22, Kut20, LBRW22, LMX^{+}21, LAD^{+}21, LLZ^{+}23, LJQ^{+}22, LM22, LFZ20, LNB23, MWJL23, MD20b, MDP22, MM22, NTO24, Pag24, PAZ^{+}22, PMS^{+}20a, PGS^{+}24, RSC^{+}22, Roh22, SN23, SDP24, SRE^{+}24, SJ20, SXW^{+}20, SKC21, SS21b, SWB^{+}23, TKS22, TKC^{+}21, TCY23, VPPQ21, WWJ^{+}22, WHSG24, YCC22, YYH21, Yan24a, YK21b, YNMR24, ZWP^{+}22, ZWG^{+}24, ZZC20, Zho23, dVAR^{+}24, MMC^{+}21]. algorithmic [CLEPF22, EML22, RBWD^{+}24, PCS^{+}23]. Algorithms [FFLR20, LWhK^{+}20, NR21, BBV^{+}22, AKL^{+}21, Bak23, BC21, BL21, CLY22, CFPS23, GHK20, KSDH23, LNP^{+}24, MGG^{+}20, SJP^{+}21b, TV24, TAY^{+}24, WZK^{+}24, XQ21]. ALICE [KvH21, NRG22, NRG24]. aligned [GBS^{+}20, SBZ23]. Aligning [Kür23]. alignment [KMN21, SJY18, SJY20]. all-atom [HQF^{+}20]. all-electron [GBR23, HLzY^{+}20, LKP24, SLZY21, SDL^{+}21]. all-in-one [GAJK23]. all-Mach [MA24]. Allen [CZY20, TCY23]. allowing [PTD20, RBV^{+}22]. alloys [CGZ^{+}20, CWG^{+}21, FDPT23, GFH23, MMCC^{+}22]. AllScale [OIA^{+}20]. almaBTE [RMCC22]. along [HMM22, MM23, Ryd20]. Alouette [Nie23]. ALPACA [HWAA22]. alpha [PCL24]. alteration [NLS24]. AlterBBN [AAHJ20]. alternative [AAHJ20, LF20]. aluminum [BCTS22]. Amazon [RMM21]. ambipolar [VGGP^{+}21]. AMD [OLNG21]. ameliorated [AGJ^{+}23]. AMFlow [LM23]. Amino [OSLC21]. AMMCR [MMC^{+}21]. amorphous [AMP^{+}21, KKB^{+}24]. AMP [YLH^{+}20]. amphiphiles [YD20]. Amplitude [fWL24b, BN20, BN23, KRL^{+}24]. Amplitudes [KMS20a, ADC^{+}21, GHL23, HJK^{+}24, PAM24]. amplituhedron [LM21b]. AMR [MOV21, PCS^{+}20, WA21]. AMR-based [PCS^{+}20]. ana cont [KH23]. analog [MKHT20]. analyse [HHMH^{+}22]. Analyser [ANU21]. analyses [GFD^{+}24, PP21]. Analysis [BHK^{+}21, CPL21, Dry21, GSM^{+}22, KAB^{+}21, KFC^{+}20, SBZ23, AMP^{+}21, AS22, ACD^{+}22, BDP^{+}21, BYL^{+}21, CdBMdAS^{+}21, COJ^{+}22, CGG21, CL21, DFP^{+}20, Eku24, EBBB22, FPSZ21, FA21, Gal22, GSL^{+}23, GvdBdGN24, GVV22, GHKW22, Gro22, HS22, HT24, HL21, Hor23b, HSO^{+}22, IUJ21, Jab22a, JKKN23, KJL^{+}23, KKB^{+}24, KDHL23, KLM^{+}22, LKP21, LLH21a, MMC^{+}22, MGG^{+}20, MZD^{+}20, MYM^{+}22, MRG22, NG21, NL23b, Nis22, PLT^{+}23, PPK22, RDH^{+}20, RZWW23, SKDH24, SMC^{+}22, SRML20, SKS24, SFC20, SLDF^{+}21, SCF20, SCMP^{+}22, SCR^{+}22b, Sva24, TPS^{+}24, WXL^{+}21, WLL22, WAYL23, WS23, Yev21, YLK^{+}20, ZXW^{+}22, ZDLS21, GOS^{+}22, GG24]. Analytic [WMM^{+}24, DML23, Hua23, KH23, KM22, MD22, PLT^{+}23]. Analytical [CLEP24, Flo24, Jab22a, Rod22, CLEPF23, Mau20, MZV23]. analyze [AMP^{+}21]. analyzer [YZ20a]. Analyzing [MMFdL21, DIKSN24, MKPW21, MKPW22, MRG22, YNV22, YNV23]. Anderson [NVCS23]. angle [LGK^{+}22, SMC^{+}22]. angled [AE23]. angles [MBCC23]. Angular [GF21, BFMA^{+}24, CCW20, CWJ21, JDS20, Pos22, ST23a]. angular-momentum [ST23a]. anharmonic [MPSK21]. anharmonicity [BBB20, CZB^{+}23]. Animating [GPD^{+}22]. Anisotropic [MD20a, MBH21, BCHE21, BVV22, BDR^{+}20, DRM20, FLK^{+}20, GBS^{+}20, LLH^{+}21b, MM23, PGYF21, YNV22, YNV23]. anisotropy [GHL^{+}22, RZWW23]. annealer [WWDM20]. annealing [RCS21, WWJ^{+}22]. Anniversary [SHB^{+}20]. Announcement [AAT^{+}20, ASPDL^{+}21, Pos22, SJP21a, Dio23, GST21, Jab22b, MBJB24, MPZB^{+}24, ZXT23]. annular [SDXY23]. ansatz [SZ23, JWK^{+}21]. ANT [SZT24]. anti [KTF22]. anti-neutrino [KTF22]. antineutrino [AG23]. antinuclei [KOT23]. antisymmetric [DEV20]. any [AKK^{+}24, KKSV24]. app [HCAH^{+}24]. Appell [ABF^{+}23a]. Application [BCTS22, DSW^{+}23, LBM^{+}23, WWB22, XSL^{+}22, XPF^{+}24, BB24a, BBA^{+}20, DS22, FKS20, GLB^{+}21, GG24, HQF^{+}20, LKP24, LLY^{+}22, LMMP23, MMC^{+}22, MYM^{+}22, MCB^{+}20, SPTPR21, SKS24, TCD20, YHH^{+}20, CMS22]. Applications [Jab24, AFB^{+}24a, BDR^{+}20, FRN^{+}23, Kan24, MMP20, OIA^{+}20, RDC^{+}20, TV24, Yan21a]. applied [AWV22, BDdM20, IUJ21, KFHR24, LJS21, TPCT22, YXX^{+}21]. applying [AMP^{+}21, Jia20]. approach [Ano20c, BOSM24, BB21, CPL21, CS22a, Che23, CB20, CLEPF22, Dau23, DR21, DVC^{+}22, DBR24, EST23, GBS^{+}20, GBD^{+}22, GQ22, GP22, HHT22, KK16, KFHR24, LOT^{+}20, LLH21a, MMM20, MSU^{+}21, MOA24, MSM24, NAZ^{+}21, PP23, PC21, RTRB21, TSAK21, WH23, WBvdH20, WYT23, Xav22, Xie23, YI22, YW21, YK21a, YD20, PBK21, SMO^{+}20a]. Approaches [TRN^{+}24, CGV^{+}22, GSLS20, LF20]. approximants [PPK22]. approximate [WWJ^{+}22]. approximation [AYI^{+}24, AKL^{+}21, BND22, CL22a, Dan23, JRS^{+}21, OYC24]. April [Ano20a, Ano21a, Ano22a, Ano23a, Ano24a]. Arbitrarily [CWJ21, GZW20, BAB^{+}20, GGCW24]. Arbitrary [OS24, ABD^{+}23, COW24, CT23, GF23, LH22, LH24, MSM24, SDP24, SHS22, Sij23, WGG20, Xie23]. arc [ZGW20, RSPJ21]. Architecture [YK21b, AGH21, FAL20, FAL21, Nor23, RCS21, RL21, ZWZ^{+}22].

Architecture-based [YK21b]. Architectures [LPSK20, DFG^{+}23, DRB22, EPM23, LJS21, OGL^{+}21, PNL^{+}24, PSK^{+}24, ZAW^{+}21]. arctangents [BTW20]. area [ZDG^{+}21]. ARGES [LS21b]. argon [BCF^{+}24]. arithmetic [SC20]. ARM [OLNG21]. arrangement [MRG22, RAJ^{+}24]. arrival [Hor23b]. arrival-time [Hor23b]. ARSENY [GSV23]. art [KMM21, KM20, KMS20b, SKM^{+}21]. ARTEMIS [JYL^{+}23, TDR^{+}20]. ARtificial [UMA21, LPSK20, SWZ23, YKYK23, KHKL24, PMSHG23, ZHI23]. artificial-neural-network [KHKL24]. ARUZ [KHR^{+}23]. aspects [JWK^{+}21]. assemblies [SDP24]. Assessing [DTGE21]. Assessment [HA21]. assignment [FN21]. assimilation [MMYU22]. assisted [HPRS23]. ASTI [MMYU22]. astroparticle [LFZ20]. astrophysical [RDC^{+}20]. asymmetric [SJY18, SJY20]. asymmetry [BMM21]. atmospheric [GSL^{+}23, HA21, RAJ^{+}24]. Atom [LCL^{+}23, CLZ^{+}21, FN23, HQF^{+}20, HLzY^{+}20, LHTP^{+}24, PBK21, Pos22, SCT21, YNMR24]. atom-atom [PBK21]. Atom-centered [LCL^{+}23, HLzY^{+}20]. atom-laser [FN23]. atomic [ČPF^{+}24, FMCB^{+}20, Gai20, GK21, GTE21, GLSH21, GOCSS^{+}23, GVV22, GSV23, Hor24, KRL^{+}24, LWS^{+}23, LU21, MMV^{+}24, RSPJ21, RPG^{+}20, TAB^{+}22, UZB22, YT23, YKY^{+}22, ZRH21]. Atomistic [Che23, ART^{+}20, CGZ^{+}20, CZTF23, IGL^{+}24, MKPW21, MKPW22, SKC21, WLF^{+}24, WBK^{+}24, YZ20a]. AtomREM [NA20]. atoms [BAB^{+}20, HY23, LWC^{+}21, RZY^{+}24, RZ23, SJP21a, SLLA22, TNL^{+}22, WLF^{+}24, ZLMH23]. A}Tools [YNV22, YNV23]. attenuation [GRCT20]. attice [FFTV23]. Attosecond [GM24, Koz23]. augmented [HM24, PLL24, WT22]. August [Ano20b, Ano21b, Ano22b, Ano23b, Ano24b]. auto [BS23, TAGC22a]. auto-correlated [BS23]. AutoEFT [HS24]. autoencoder [YLL^{+}22]. Automag [GO23]. automata [SYFT23]. Automated [AUO^{+}22, ATRD21, DSW^{+}23, GSBN22, HS24, HLM22, LJS21, VZ20, ZLV23, ABH^{+}23, GAJK23, HBM^{+}24, HGS20, LBM^{+}23, LEL^{+}22, MOA24, YLH^{+}20]. Automatic [DIAA21, DR21, MMM20, CGZ^{+}20, CZB^{+}23, GO23, XLG^{+}23]. automatically [KKSV24, WYZZ23]. automating [UMA21]. automation [GAS^{+}23]. automatized [WGGC22]. AutoPas [GSBN22]. autoregressive [BKS22]. autotuning [KMM21]. Auxiliary [MVK^{+}24, Bag22, GJA21, LM23, WYT23, YK21a]. auxiliary-field [GJA21]. averaged [MOMO24]. averaging [MPSK21]. avoiding [WGS^{+}22]. Aware [DS22]. AX [FLW^{+}23]. Axial [RSM21]. Axially [MSN^{+}22]. Axially-deformed [MSN^{+}22]. axisymmetric [BJS^{+}23, CEC^{+}24, HSB^{+}24]. azimuthal [LAD^{+}21].

B [CCK23, AGH21, WGG20]. B-spline [WGG20]. background [BKO20, MGC^{+}23]. backscattering [LGK^{+}22]. backward [MM22]. Baer [Dau23]. Baer-Nunziato [Dau23]. balanced [MA24]. balancing [MLT^{+}21, TPK^{+}21, ZHZG23]. Band [RMFB23, WXZH24, BFD22, BMI23, CL20, DSQ23, GLN23, IMB^{+}22, MFB23, SLZG20, SLL^{+}24, WZZS21]. bandstructure [LKP24]. Bang [AAHJ20]. Bardeen [PTD20]. Barna [RMCC22]. barycentric [WVK21]. baryon [BMM21, DH20]. Base [TL20]. based [AZH^{+}24, AAG^{+}20, ABGD^{+}20, ACD^{+}22, AMA^{+}20, ASC^{+}24, BCHE21, Bar22a, BFCR24, BCGT24, BTW20, CDD22, CLS22b, CZWE23, CTPS22, Dau23, DFU20, DP21, DSQ23, ENK24, FN21, FJ22, FYW23, FLW^{+}23, FMBD22, Gal22, GDK21, GAGO21, HPY21, HL21, HTL^{+}22, HT20, HWL^{+}23, HFP21, HWZ24, HWAA22, Hor24, HBM^{+}24, HLCD20, HHT^{+}24, IJVJ21, JLW24, JKSY22, KKLZ23a, KKLZ23b, KTF22, KRJ23, KPR^{+}24, KMN21, KSF^{+}22, KHR^{+}23, KR23, KL23b, KKB^{+}24, KM23, KRE22, KG21, LEE^{+}24, LGDF20, LAD^{+}21, LLC^{+}23, LWS^{+}23, LY24b, MHÅ21, MMCC^{+}22, MEDT^{+}23, MSU^{+}21, MMV^{+}24, MBCC23, MSM24, MRD23, NSY^{+}23, NJSY22, OIA^{+}20, OBGA24, OSK^{+}21, OLNG21, ONH^{+}20, PMS^{+}20a, PAL^{+}20, PM21, PCS^{+}20, QCZ23, RMCC22, RZY^{+}24, Di 22, SKEZ24a, SRC21, SDP24, SLB^{+}23, SBG23, SWTC23, TCY^{+}21, TPS^{+}24, TAB^{+}22, TLC^{+}21, UD24, VB22a, VV21, eSdSBST21, WAET22]. based [WPMK21, WVK21, WKJB23, XFH^{+}22, YFL22, YCCW23, YKK^{+}23a, YSX^{+}20, YZHL22, YKSH20, YMCF23, YK21b, YKK23b, YZZ^{+}23, ZWC^{+}20, Zha20, ZRZ^{+}21, ZWZ^{+}22, ZLY24, ZRH21, Zho23, ZHZG23, ZCS^{+}24, SKEZ24b]. bases [HY23, ZRH21]. Bashforth [TSL21]. Basic [Gro22, HW22]. Basics [BC22]. basis [DEV20, FASD20, HLzY^{+}20, JZW^{+}22, LM22, MSN^{+}22, MLD^{+}22, MSHP02, MSHP20, VCHH23, VTB^{+}21, WGG20, YJLW21]. Batched [HW22]. Bateman [CLEPF22]. bath [CLY22]. Bayesian [CV21, KKP22, HGS20, MTY^{+}22, NG21, TV24]. BaZrO [DSW^{+}23]. BCS [RSM21, Jia20]. BDDC [SFS22]. BDK [KMS20a]. BDSIM [NBS^{+}20]. beam [MGC^{+}23, PK24, TLC^{+}21, TCcN23]. beam-target [TLC^{+}21]. beams [HTL^{+}22, PMSHG23, SKEZ24a]. BEC2HPC [GTA21]. BECs [SCL22]. BEEC2.0 [YZW22]. behavior [JLL^{+}24]. behaviour [LP24, MES^{+}24]. Behler [vdHKB^{+}23]. Belle [BBB^{+}21]. belt [DRM20]. belts [NS20]. Benchmark [KFPV21, SNP^{+}20]. Benchmarking [ASA^{+}22, MGC^{+}23]. benchmarks [KPL^{+}21, WKR23, ZJM^{+}21, ZRH21]. BerkeleyGW [LACL24]. Berry [RCP^{+}24]. BESLE [GPM^{+}21]. Bessel [CCK23, GDB10]. beta [AG23]. Betaboltz [RCB21]. Bethe [JWK^{+}21, LG21a, SA23, ZAW^{+}21]. BetheSF [LG21a, LG21b]. between [BMREC21, BLN^{+}21, CXCZ23, Efr20, Efr21a, MRT^{+}22, Tan23, UZB22, YLL^{+}22, ZJM^{+}21]. beyond [AFJ^{+}23, COJ^{+}22, DDM20, KKSV24, MSG^{+}21, UMA21]. Bézier [ET24]. BGK [LSW^{+}20]. BGSDC [TBAR21]. Bhatnagar [FJ22]. bi [BA24, LLH^{+}21b]. bi-anisotropic [LLH^{+}21b]. bi-directional [BA24]. bi-Lebedev [LLH^{+}21b]. biasing [PYT^{+}24]. BiconeDrag [SPTPR21]. bifold [Kar23a, Kar23a]. Big [AAHJ20]. Big-Bang [AAHJ20]. billion [CLZ^{+}21]. BIMBAMBUM [SSPF24]. bimetric [Tor20]. bimEX [Tor20]. binary [LLT^{+}23, LLQ^{+}23, LB24, NDFL24, VHBK21, ZLLM23, ZCWY20]. Binding [DSQ23, JPJ^{+}23, KVSC21, LZK^{+}23, NKP20, WPMK21, ZYLY22, NKP20].

BinPo [MFB23]. bio [COJ^{+}22]. bio-shield [COJ^{+}22]. biofilm [OCE^{+}23]. Biofilm.jl [OCE^{+}23]. biomedical [MVF20]. biomolecular [DCZ23]. biophysical [PNL^{+}24]. Biot [SSH^{+}23]. bipartite [CL22a]. Bipolar [WSK^{+}22]. Bit [Fis24]. Bit-twiddling [Fis24]. Black [BLN^{+}21]. Black-box [BLN^{+}21]. blackout [GSL^{+}23]. blade [KKY24]. BLAS [HW22]. blende [HLCD20]. blends [BMR^{+}24a]. Bloch [GLB^{+}21, MONW21, RCP^{+}24, RJ21]. block [Cos22, RAJ^{+}24, UJ21, ZJ23]. block-pressure-velocity [UJ21]. blocking [AWV22]. blood [ZMC23, ZHZ^{+}23]. blowup [HWZ24]. Bluues2 [SOH^{+}23]. BNP [CCG21]. board [AAG^{+}24, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano21o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k]. bob [SPTPR21]. Body [GBJ^{+}21, ATRD21, DC22, JDS20, KWK^{+}21, MAJ20, NTO24, OS24, PPKK21, RMM21, Xav22, XBL^{+}20]. body-of-revolution [NTO24]. Bogner [CCK23]. Bogoliubov [ATRD21, CLS^{+}22a, MSN^{+}22, SKDH24]. Boltzmann [ASC^{+}21, AHP21, AS24b, BBA^{+}20, CK20b, CKT21, DACA^{+}22, DCZ23, JMOC21, KFPV21, KBSH20, KAS24, LFL^{+}24, LSS24, MWJL23, MD21, NL23a, PPR^{+}21, PLSB22, QCZ23, STA20, WJB21, WWB22, WHB21, WA21, WKR23, YCCW23, YSI^{+}24, YM21, YD20, ZPZH20, ZPL^{+}24]. bond [HLCD20]. bond-orbital [HLCD20]. bonded [VXT^{+}23]. boom [EOR21]. boosting [CFL^{+}22]. Bootstrapped [ZJ23]. BORAY [sXBkB^{+}22]. Boris [DML23, ZK20]. Born [SOH^{+}23, XZLX20]. Bose [BKRG22, KRG21, MBA21, RVM^{+}21, SKDH24]. boson [BHH^{+}20, DDM20, KKM^{+}20, SZ23, GKIB21]. bosons [AKK^{+}24, ABB^{+}22]. BoSS [GKIB21]. both [RVRT22]. bounce [GNP20]. bound [GTMB21, MZ22]. Boundaries [LM21b, FBMD20, GPN20, NR21, TS23, XSM22]. Boundary [GPM^{+}21, ATC^{+}23, ELSV24, FYW23, FMHH24, GVV22, HPAW21, KD23, MD20a, MVF20, NL23a, PLL24, PCL24, RZH22, STA20, SP20, SXYD24, TCD20, TTM22, WXY20, WXZH24, Yan21a, YSI^{+}24, YSX^{+}20, YXX^{+}21, YJ24, ZPZH20, Zha20, ZPL^{+}24]. boundary-cascaded [ZPZH20]. bounded [DAC^{+}23, HPAW21]. Boussinesq [TL20, TSL21]. Boussinesq-type [TL20]. BOUT [DKM^{+}24, SDXY23, WXW^{+}21, ZSqXY21]. box [BLN^{+}21]. brackets [CLVV22, Efr21b, Efr23, SMGK21]. Braginskii [BJL^{+}24]. Bratu [Ano20c, KK16]. Brazilian [CCG21]. breakdown [CHA21]. breast [BSG^{+}21]. Breit [HA23, WBvdH20]. BREMS [Pos22]. bremsstrahlung [EMM^{+}23, Pos22]. bridge [WHSG24]. bridge-inspired [WHSG24]. bridging [Yan23]. Brillouin [Dür23]. BROADCAST [PCS^{+}23]. broadening [Bul21]. Brownian [DG20]. Brute [JRG21, XSM22]. brute-force [XSM22]. Brute-forcing [JRG21]. BSHF [WGG20]. BSM [BBH^{+}23, PAM24]. BSMArt [GJ24]. BSMPT [BMM21]. BTE [RMCC22]. BTE-Barna [RMCC22]. bubble [SSPF24]. bubbly [LMHL20]. build [IGL^{+}24]. building [MPH^{+}24]. bulk [DAA^{+}24, SLDF^{+}21, WGS^{+}22, MMC^{+}22]. buried [CML^{+}24]. Burn [ZYG21]. Burn-up [ZYG21]. Burnup2.0 [ZYG21].

C [ADC^{+}21, Dan23, DH22, DLD^{+}21, FXZT21, FTG23, GCK21, Kar23b, KMU^{+}23, Koz23, MVF20, SW23, UMA21]. C# [Gro22]. C-code [GCK21]. C1 [ZJM^{+}21]. Cache [AWV22]. caching [PWC24]. CAD [OBGA24]. CAD-based [OBGA24]. Cadabra [CFPS23]. CAGE [RHW^{+}21]. Cahn [CZY20, KSF^{+}22, LM21a, LHC20, TCY23, Yan21b]. Cahn-type [TCY23]. CALANIE [MD20a]. Calc [BMR^{+}23, BMR^{+}24b]. calculate [DZZ21, Efr20, GTE21, LL23, PB23]. Calculating [BGH22, GBH20, AAHJ20, BGHC23, CL20, CSW24, FO20, FO21, GO23, GBR23, HTL^{+}22, HHT^{+}24, Jab20, KSG22, LKP24, LLS^{+}21, LLZ^{+}22, MKPW21, MKPW22, RPG^{+}20, SLZG20, SP23, YKSH20]. Calculation [BRHT21, Efr21a, Efr21b, SLLA22, SMGK21, UZB22, WKBW21, ZYG21, AUEO24, BRAC23, BMR^{+}24b, BND22, CL21, DCRF23, DS20, ENK24, FASD20, GTMB21, JKST22, KvdW20, KKLZ23a, KKLZ23b, Kar23a, KT23, KM20, KMS20b, LYC20, LEL^{+}22, LHG^{+}20, MMC^{+}21, MMP^{+}24, Mar22, MPZB^{+}24, NAZ^{+}21, NAZ^{+}22, Och23a, Och23b, OYC24, Ols23, Pos22, SJP21a, TMC22, TB20, UO20, WGGC22, XOTI22, BMR^{+}23]. calculations [ADW^{+}23, AAMY23, AJW^{+}21, AM21a, BHH^{+}20, BR20a, BLN^{+}21, CLY22, CFBRE24, CMS22, Cas24, ČPF^{+}24, Dan24, DvHSdS22, Fon21, GM20, HY23, Hor24, IJVJ21, Jab22a, JPJ^{+}23, KCS22, LVB22, LWS^{+}23, LLS^{+}21, LG23, MBJB24, MBC^{+}24, MAJ20, MSHP02, MSHP20, MDR^{+}20, NVCS23, DARJ23, Pot24, RDV^{+}20, RCP^{+}24, RG21, RSPJ21, RTRB21, SY20, SMB20, SLL^{+}24, SXW^{+}20, SDL^{+}21, SC22, SOH^{+}23, UMA21, VCF22, VvBTH20, WZZ^{+}23, WXZH24, YT23, YHY^{+}21, YLH^{+}20, ZRH21, ZHM21]. calculator [ABH^{+}23, FCTFR20, Kut24, MBTB21, SKEZ24a, UAS22]. calculus [BSK^{+}22, CLEPF22, KMR22]. calorimeters [ABF^{+}23b]. Camassa [ZYXC24]. can [SWZ23]. capabilities [KGT22]. capability [FA20a, MCB^{+}20]. capacitively [JĎD^{+}21]. capacity [LHWX24]. capsules [TKC^{+}21]. capture [TAE^{+}21]. capturing [LF24]. Caravel [ADC^{+}21]. carbon [PK24, PWC24]. cardiovascular [AFB^{+}24a]. Carlo [ABGD^{+}20, ASPDL^{+}21, ACD^{+}22, AJW^{+}21, BRAC23, BB21, BCF^{+}24, BKO20, BKS24b, BIK^{+}21, BKBL22, BKG^{+}23, CCM20, DRSZ23, DDM20, DdCAG23, ELSV24, FXQS21, GJA21, GAGO21, GHK20, HT24, HLMB24, HQF^{+}20, Hua23, HSMR^{+}24, Ilt21, JWW^{+}23, JKKN23, JĎD^{+}21, KOT23, Kal20, Koł22, Koł23, LEE^{+}24, LMWW24b, LAC21, LTMK21, LLL24, LC24, LHWX24, MMM23, MT23, NL23b, Nie22a, OBL^{+}21, PAZ^{+}22, PLT^{+}23, PYT^{+}24, PPR^{+}21, PC21, RSC^{+}22, RCB21, Roh22, SNP^{+}20, SJP^{+}21b, SHS22, SNG20, SLIC24, TAY^{+}24, WAN^{+}22, WSRO24, WZPW20, YLK^{+}20, ZHS^{+}20, ZDLS21]. Carlo-effective [PC21]. carlomat 4.0 [Koł22]. carrier [TQGE23].

Cartesian [HG22, LEE^{+}21a, RVRT22]. cascade [CLZ^{+}21, SZY^{+}22]. Cascaded [AHP21, ZPZH20]. cascades [BKO20, EES24]. case [ABFP24, Jia20, TPS^{+}24]. catalysts [RSC^{+}22]. Catalytic [KBSH20]. Causal [Bel24, CDT22]. cavitation [SSPF24]. cavities [BBA^{+}20, DKRSR22, WYHW24]. CBM [GBJ^{+}21, SCF20]. CCA [TCY^{+}21]. CCS [KWK^{+}21]. CDFTPY [VCF22]. CDPDS [KR23]. CDT [BLG24]. cdugksFOAM [WLZ^{+}24]. Celebrating [SHB^{+}20]. Celeris [TL20]. Cell [FAL21, YYC^{+}23, ASC^{+}24, BCHE21, BC21, BTK24, CTZW23, DBH^{+}22, FN21, FYM^{+}22, LMX^{+}21, LAD^{+}21, LM20, MFS^{+}22, Mar22, MRH^{+}23, MLT^{+}21, NTO24, OCR^{+}22, PSMRS^{+}23, PGYF21, RC23, SXYD24, SWB^{+}23, TWR21, Ume22, XSM22, XHY^{+}24, YC20, YWX^{+}23, ZK20, ZDSS23, ZLS^{+}22, DFG^{+}23, DWD^{+}24, FAL20, FA21, KG21, OGL^{+}21, SAC^{+}21]. cell-centered [PGYF21]. Cell/Monte [JĎD^{+}21]. CellListMap.jl [Mar22]. cells [BMP^{+}24, LLR23, ZHM^{+}24]. cellular [SYFT23]. Center [SMO^{+}20a, BKP22, BJS^{+}23, GDS^{+}21, LM22, MHP23, SSB^{+}23, XQ21]. center-evolving [LM22]. centered [HLzY^{+}20, LCL^{+}23, PGYF21]. centering [ZLS^{+}22]. central [AAG^{+}24, For22, GST21, WGG20]. centrality [YZL^{+}23]. CepGen [For22]. cepstral [EBBB22]. cerebral [dMMLOS20]. Cerman [MHÅ21]. CFD [ADF^{+}22, PCS^{+}23, SJWL22, AFB^{+}24a]. CGMF [TSJ^{+}21]. Chain [HSMR^{+}24, HQF^{+}20, LWhK^{+}20, LTMK21, MMM23, VFS23, WGS^{+}22, EGKH24]. chains [EGKH24, LF20]. chalcogenides [SKS24]. chambers [TAGC22a]. Chandrasekhar [Jab20]. change [RSD20, SRT^{+}20]. channel [CGV^{+}22, GOS^{+}22]. channeling [Nie20]. channels [ABWZ23]. Chaos [GBJ^{+}21]. chaotic [SRML20]. character [YZ20a]. characterisation [ANU21]. characteristic [XDF20]. characteristics [CL21, CTPS22]. characteristics/diamond [CL21]. characteristics/diamond-difference [CL21]. characterization [HMM22, SLBR22, vTDGCR21]. charge [BTK24, DBM^{+}24, FN21, KKLZ23a, KKLZ23b, OCR^{+}22, RC23, SWB^{+}23, ZPL^{+}21]. charge-conserving [FN21]. Charged [KKSV24, TGGC23, AAD^{+}24, BFL^{+}22, SLP^{+}22, SLLA22, SA23]. charges [FH22]. charging [WSK^{+}22, YFL22]. Chebyshev [BND22, LS22]. chemi [Nis23]. chemi-topology [Nis23]. chemical [AHM^{+}23, FCTFR20, FBC^{+}21, MPZB^{+}24, RCX24, RHW^{+}21, SVJ^{+}24, ŻTR^{+}22, CV21]. chemically [CZWE23, Gal22]. chemicals [BBV23]. cheminformatics [BBV23]. Chemistry [GM24, BLN^{+}21, CNS22, Koz23, OCE^{+}23, PM21, RAJ^{+}24, XFGS24]. Chern [GBR23]. chi [GST21]. chi-square [GST21]. chip [DHK^{+}21]. CHIPR [RV20, RV21, Xav22]. chiral [LU21, YLL^{+}22]. Chirally [HHT22]. choice [SRK^{+}24]. Chromodynamics [KK20]. Chrono [ZTV^{+}24]. cij [LDW^{+}21]. CIMBA [Ilt21]. circuit [OV23, YWX^{+}23]. circuits [WHSG24, dBBVA20]. circular [NG21]. CITDSE [FN23]. CKBIT [CV21]. CLAS12 [TAGC22a, TAGC22b, TGGC23, TGIM23]. class [FYW23, GGCW24, YZW21]. Classical [TNL^{+}22, GSV23, KF23, PSK^{+}24, SLLA22, VCF22, XQ21]. Classification [SLL^{+}21, BBA^{+}20, DR21, Nis23]. classifiers [KHKL24, VEHCM21]. Clifford [KF23]. Clinamen2 [WBK^{+}24]. clinical [BSG^{+}21]. CLIP [UO20]. cloak [YWM23]. closure [BJL^{+}24, LMMP23]. clots [ZHZ^{+}23]. cloud [RMM21]. CLT [ZJM^{+}21]. clumps [OAP^{+}24]. Cluster [Nis23, ZZC20, CHY^{+}24, DRSZ23, HAM^{+}20a, PPKK21, YX24, ZWG^{+}24]. cluster-cluster [DRSZ23]. Clustering [DR21]. cm [FYM^{+}22]. CMInject [WAK22]. CMS [TPS^{+}24]. CNOK [SW23]. CNUCTRAN [BOSM24]. coagulation [WNS^{+}21]. coarse [MBCC23, NA20, OCK^{+}24, ZKZ^{+}24, Zho23]. coarse-grained [MBCC23, OCK^{+}24, ZKZ^{+}24]. Coarsest [EVFRHR23]. Coarsest-level [EVFRHR23]. coaxial [YMCF23]. Code [OGL^{+}21, RSM21, VXT^{+}23, YX24, ZYG21, ASS^{+}24, AAG^{+}20, ABGD^{+}20, ARLDG24, AMA^{+}20, AAHJ20, ASC^{+}24, BJL^{+}24, BFD22, BFCR24, BCTS22, BDA^{+}20, BCSS24, CR20, CFBRE24, CZA^{+}23, CAC^{+}22, CAWK22, CZS^{+}21, COJ^{+}22, CTZW23, CZB^{+}23, CKC^{+}21, CESD^{+}23, Dan23, DMS^{+}22, DIK^{+}23, DFU20, DSQ23, DBH^{+}22, DHE^{+}24, Dry21, DBdFdSR21, EBBB22, EES24, FMCB^{+}20, FFTV23, FAL20, FAL21, GRCT20, GCK21, GBD^{+}22, GLN23, Hua24, JKST22, Jia20, Kal20, Kar23a, KAB^{+}21, KWZ24, KJL^{+}23, KBH^{+}24, KMU^{+}23, KDK23, KT23, KSIL22, Koz23, KAG24, Kul20, LEE^{+}24, LOT^{+}20, LG23, LVMGF^{+}23, LDW^{+}21, LJS21, MG22, MGX24, MFB23, MPS^{+}24, MBC^{+}24, MRH^{+}23, MSU^{+}21, MHP23, MAMK21, MXH^{+}24, MDR^{+}20, NUK^{+}22, NSY^{+}23, NBS^{+}20, NL23b, NDFL24, Nis22, Nor23, Och23a, Och23b, OAP^{+}24, OLNG21, PP23, PLSB22, PWC24, RMR^{+}22, RCP^{+}24, RV20]. code [RV21, RMFB23, SLK23, SMA24, SP23, SHW^{+}21, SPF21, SZNW23, SBB^{+}24, SW23, TSJ^{+}21, TMC22, TCSD24, UÁEPGBP24, WKBW21, WXL^{+}21, WXX24, XSM22, sXBkB^{+}22, XHY^{+}24, YHH^{+}20, YGSW21, YLK^{+}20, YJ24, ZMC23, ZJM^{+}21, ZDSS23, ZXW^{+}22, NUK^{+}22]. code-generation [LJS21]. codes [BMREC21, BPMMP24, DRM20, LSF23, MLT^{+}21, MT23, PSMRS^{+}23, SRT^{+}20, SAC^{+}21, TWR21]. coefficient [LN23]. coefficients [DLD^{+}21, Efr20, Efr21a, EBBB22, GF21, MBTB21, NRKA22, ÖAÖ24, PDD24, TAE^{+}21]. COGENT [KDK23]. coherent [Jia20, KMBP24, MR22, MS24, Roh22]. coherent-pair [Jia20]. cold [BGW^{+}22]. ColESo [Bak23]. collaborative [DBBP23, WLL22]. Collection [Bak23, Pot24]. collective [ASA^{+}22, GLW^{+}23]. collider [GPD^{+}22, YZW22]. colliders [JWW^{+}23]. collinear [GOCSS^{+}23]. collision [BDK^{+}23, BKS24b, CKT21, CLZ^{+}21, KSJ^{+}22, KDK23, SNP^{+}20, WJB21, WWB22]. collisionality [BJL^{+}24]. collisionality-extended [BJL^{+}24]. collisionless [NBB^{+}21]. collisions [Ada22, BCT20, BKBL22, CJD^{+}20, For22, GBJ^{+}21, GSV23, JĎD^{+}21, MBH21, MEH21, NRG24, SLLA22, WLF^{+}23, YZL^{+}23]. collocated [KFC^{+}20]. collocation [FYW23, YNMR24]. colloidal [BMT^{+}20, BLM^{+}22, DG20]. color [GG24]. combination [ADH^{+}20, BRAC23, HT24]. combinatorial [JWK^{+}21]. combined [ACD^{+}22, HK24, LF24, LZP^{+}24]. combining [LACL24, SHW^{+}21, ZYML24]. combustion [Di 22]. COMET [YLK21]. Comm [AAT^{+}20, Ano20c, MSHP20, SJY20]. command [Kut24]. command-line [Kut24]. commensurate [NNMJ22]. Comment [CCK23]. Commics [PRS^{+}20]. common [DAA^{+}24]. Commun [FAL21, MKPW22, Och23a, YNV23, ZYX22]. communication [GSL^{+}23, KMM21]. communications [BGR23]. community [YCC22]. compact [PLL24]. Comparative [SKM^{+}21]. Comparing [DAA^{+}24]. Comparison [HWZ24, NCF^{+}23, SLZ^{+}24, TV24, TM24, EVMP20, GSLS20]. Compatibility [GOCSS^{+}23]. Compatible [Sij23, BCD^{+}21, OBC^{+}24]. compiler [VB22a]. Complete [CS22b, YK21b]. complete-search [YK21b]. complex [ASPDL^{+}21, ABD^{+}23, CZS^{+}21, CL22a, DIKSN24, GFH23, HMR22, KHKL24, KDIN^{+}23, LLT^{+}23, LLQ^{+}23, LLH^{+}21b, MD20b, NR21, SN23, SXYD24, SFC20, SBP20, WZZS21, WHSK24, WZK^{+}24, YMCF23]. complex-valued [CL22a]. complexity [GJA21]. component [ABB^{+}24, BSC^{+}21, DKM^{+}24, HWL^{+}23, YWTK23, Yev21, ZYXC24]. component-based [HWL^{+}23]. composed [KAG24]. composites [BDR^{+}20, FLK^{+}20, SRK^{+}24]. compound [NRK^{+}21]. compounds [AMP^{+}21]. comprehensive [GWA^{+}23, MPS^{+}24]. compressible [Bak23, BMSP21, BBA23, BSC^{+}21, Dau23, DAC^{+}23, GSL24, GB22, HMR22, HWL^{+}23, HJGL22, LF24, LJH^{+}23, LIG^{+}24, LJS21, MA24, PCS^{+}23, RLW^{+}24, RCW^{+}20, SPLD20, Tan24]. compression [LMHUR23, TW21b, VPS23, WXZH24]. compressional [YYC^{+}23]. compressive [LLH21a]. Comput [AAT^{+}20, Ano20c, FAL21, MKPW22, MSHP20, Och23a, SJY20, YNV23, ZYX22]. Computation [CT23, LY22, ADC^{+}21, BV21, Bag22, DLD^{+}21, GTA21, GST21, KF23, KGT22, KYH24, KD23, LG21a, LM23, LY24b, LLR23, MG22, MBTB21, PBK21, RZ23, UÁEPGBP24, WZZS21, YYZ^{+}22]. Computational [AMA^{+}20, OSLC21, PRS^{+}20, VHBK21, AFB^{+}24a, BMREC21, BB24a, BMR^{+}24a, BBA23, DBBP23, EL24, FZD^{+}24, FLW^{+}23, HPAW21, HPP23, LWS^{+}23, LSF23, LP24, MOA24, MSM24, NAZ^{+}21, RGS^{+}21, SHB^{+}20, Tan23, TW21b, TRB20b, WMM^{+}24, WNS^{+}21, WSK^{+}22, XFGS24, YNMR24, ZHM^{+}24, Gal22, OLNG21]. computationally [AGJ^{+}23, BKS24a, RDZ^{+}20]. computations [CLVV22, CDD22, CGR21, DNG^{+}20, MMM20, RDR^{+}20, RHLTG^{+}22, Tor20]. compute [Ham20b, KKSV24, KMM21, LBM^{+}23, MSH22, MFB23, MONW21, PAM24, RVRT22, SFBG20]. computed [WKJB23]. Computer [AGK^{+}23, Jia20, PBK21, SCT21, CZB^{+}23, Dry21, EOR21, MPZB^{+}24, NAZ^{+}22, OYC24]. computers [EVFRHR23]. Computing [DEV20, TGBM22, TAE^{+}21, BOSM24, CR20, CZS^{+}21, CCC20, CGV^{+}22, EBNS22, GAS^{+}23, GSV23, HYL^{+}22, HTL^{+}22, HXS20, HHVB21, HA23, Hid21, Hua24, KIK20, KPL^{+}21, LGBJ20, LJS21, SMO^{+}20a, SLB^{+}23, SCF20, TBAR21, Ume22, WXY20, WXL^{+}21, WICA22, Xie23, YS22, YJLW21, dMMLOS20]. concentrated [PLSB22]. concentrations [BOSM24, TQGE23]. concurrent [AS24a, ZWC^{+}20]. condensate [BKRG22, Jia20, KRG21, RVM^{+}21]. condensates [MBA21, RPG^{+}20, SKDH24]. condensed [Jab22b, Jab24, Sch21, Tan23]. condition [CdBMdAS^{+}21, MMP20, NL23a, RZH22]. conditions [CMS22, FMHH24, GVV22, HMM22, MD20a, MVF20, PLL24, PCL24, STA20, SP20, WXY20, XZLX20, Yan21a]. conduction [Bel24, MM22]. conductivity [AS24a, FO20, HYL^{+}22, KMG^{+}20, LDGN24, MMC^{+}21, MPN^{+}21, WLCF21]. confidence [CTPS22]. configuration [AAMY23, GF21, LWS^{+}23, PRR23]. configurational [BRAC23]. configurationally [GFH23]. configurationally-complex [GFH23]. configurations [CW22, IUJ21, LDGN24, MMM23, sXBkB^{+}22]. confined [WGS^{+}22]. confinement [CW22, LLH21a, TKC^{+}21, XSL^{+}22, ZLL^{+}24]. conformal [Bzo21, HHT22, OKBM23]. conical [SPTPR21]. conjugate [AUEO24]. connecting [SZ24a]. Conservation [MM23, Yev21, HWAA22, LY24a, OCR^{+}22, RC23]. Conservative [FMHH24, CCW20, DTC20, Gon22, TCSA21, TCY23, WKR23, YZ20b]. conserved [CZY20, Nog21b, PYT^{+}24, ZWP^{+}22]. conserving [BC21, BTK24, FN21, PMK^{+}23, Sij23, SWB^{+}23, TIG^{+}24]. considerations [CHA21]. Consistent [vRCM21, BN20, BN23, FH22, KM22, Kut20]. consistently [YWTK23]. constant [FCTFR20, Pag24]. constants [CT23, LLS^{+}21, LLZ^{+}22, LEL^{+}22, NL23b, PB23]. constitutive [LLY^{+}22, MEC^{+}24]. Constrained [GHK20, KWK^{+}21, LVMGF^{+}23, DR21]. constraining [TRN^{+}24]. Constraint [SE24]. constraints [AAHJ20, KM22]. constructing [RBWD^{+}24, Xav22, ZYL^{+}23]. Construction [HTS^{+}21, Kor23, NNMJ22, SKS24, HS24, Kan24, Ols23]. Contact [MBCC23, ZTV^{+}24]. Contact-based [MBCC23]. Contacts [Gul20, MBCC23]. content [May21]. context [EVMP20]. continental [Org22]. continuation [FBMD20, FMBD22, Hua23, KH23, KM22]. continuity [AIZ23, FN21]. continuous [HK24, LS22, MKHT20, SNG20]. continuous-time [MKHT20, SNG20]. Continuum [SSB^{+}23, GDS^{+}21, STRF^{+}20, TAB^{+}22, TCSD24, VV21, ZWZ^{+}22]. Contraction [Dju20, YK21b]. Control [AAA^{+}20, BB21, Cas24, COW24, JM24, MOV21, RHW^{+}21, RBWD^{+}24, RRM^{+}23]. controlled [OS24]. Controlling [BKB^{+}21]. CONUNDrum [GM20]. convection [HA21, RSD20, SRT^{+}20, XDF20]. convective [KKPC23]. convergence [ABFP24, DTGE21]. convergent [WLF^{+}23]. conversion [ACKB23, PPK22, ŻTR^{+}22]. Converter [ACKB23]. convex [LHG^{+}20, Yan24a]. convolutional [TAGC22a]. CoolFOAM [FCSP20]. cooling [EBNS22]. CoolProp [FCSP20]. Cooper [PTD20]. coordinate [BR20a, CLS^{+}22a, MSU^{+}21, WH23, WXX24, YJ24]. coordinate-scaled [BR20a]. coordinate-space [CLS^{+}22a]. coordinates [ALB22, BJS^{+}23, CZB^{+}23, SSB^{+}23]. copolymer [CZY20, LM21a]. copolymer/homopolymer [LM21a]. core [CXCZ23, CL21, KMM21, OGL^{+}21, SDL^{+}21]. corepresentations [LZYY23]. cores [WWM^{+}22]. corona [DBR24]. coronaChargingFoam [YFL22]. corrected [AKK^{+}24, KKM^{+}20, SA23]. correction [CS22a, MD20a]. corrections [BKG^{+}23, KM20, KMS20b, WLCF21]. correlated [AMK24, BS23, HAM^{+}20a, SP23]. correlation [Bzo21, PSK^{+}24]. CorrelationFunctions.jl [PSK^{+}24]. corresponding [CGV^{+}22]. Corrigendum [AAT^{+}20, Ano20c, MKPW22, MSHP20, Och23a, ZYX22]. cortex [dMMLOS20]. COSE [GLW^{+}23]. cosmic [Mau20, MNS^{+}24, SBG23, VPPQ21]. cosmic-ray [Mau20]. cosmological [Bar22a, VPPQ21]. cosmologies [AAHJ20, Kar23b]. cosmology [CFPS23]. Cost [RL21]. Cost-efficient [RL21]. Coulomb [AB21, CXCZ23, RZ23, SSB^{+}23, TRN^{+}24, TNL^{+}22]. counting [GvdBdGN24, ZWG^{+}24]. COUP [AKK^{+}24, KKM^{+}20]. Coupled [KR23, ASU^{+}21, BKRG22, BGW^{+}22, CZY20, CGV^{+}22, GTMB21, JLL^{+}24, JĎD^{+}21, KRG21, KMBP24, KSF^{+}22, Koz23, KFPV21, KBSH20, LM21a, MEC^{+}24, MPH^{+}24, MVAXP22, MBA21, OCC23, PTH24, PBC^{+}24, PPKK21, RVM^{+}21, STRF^{+}20, TWR21, VGGP^{+}21, Xav22, XPF^{+}24, ZCWY20, Zho23]. coupled-channel [CGV^{+}22]. coupled-cluster [PPKK21]. Coupling [MEDT^{+}23, RSC^{+}22, YLK^{+}20, AG21, BRHT21, CCC20, Dau23, Gai20, GF21, GOCSS^{+}23, HTL^{+}22, JMOC21, LLC^{+}23, LACL24, PDD24, SHRK22, TLC^{+}21, WXZH24, XLL^{+}22, YD20, Gai20]. couplings [HLM22]. covariation [HL21]. CP [LEE^{+}21a]. CP3d [GWA^{+}23]. CPC [SHB^{+}20]. CPL [STRF^{+}20]. CPU [DMS^{+}22, Dür23, GB22, HG22, PSK^{+}24, RCX24, SBG23]. CPUs [CZTF23]. Crank [LNB23]. creation [WAN^{+}22]. criterion [MHÅ21]. critical [GO23, ZYG21]. criticality [PAZ^{+}22]. Cross [SJY18, SJY20, AS22, AYB24, BFMA^{+}24, DS20, Hor24, Jab24, OSK^{+}21, PYT^{+}24, SY20, SLLA22, ST23b, WLL22]. Cross-platform [SJY18, SJY20, OSK^{+}21]. cross-section [WLL22]. cross-sections [AYB24]. crossings [GSV23]. crosslinking [ZBS^{+}23]. Crystal [GK21, LS23, Dan24, EGLK20, FPSZ21, Kor23, PRR23, RZ23, TCY23, YML^{+}24, ZHS^{+}20, dSOZ22, CZA^{+}23]. crystalline [GBR23, MPSK21, WLF^{+}24, YLH^{+}20, vTDGCR21]. CRYSTALpytools [CZA^{+}23]. crystals [MSH22, SLL^{+}24, SPF21, TRB^{+}20a, WZZS21, WXZH24, ZYL^{+}23]. CSEM [Yan23]. CSL [XSM22]. CSPlib [DIKSN24]. cubic [Ilt21]. CUDA [BKP22, BLM^{+}22, DS22, JRG21, JS24, NS20, STA20, SKEZ24a, SKEZ24b, WKJB23, XHY^{+}24, YW21]. CUDA-Aware [DS22]. CUDA-based [SKEZ24b, SKEZ24a]. cuFFT [HT24]. cumulative [GST21]. CUR [LMHUR23]. current [EBBB22, FN21, SWB^{+}23, VTB^{+}21]. currents [VPS23]. Curvature [KWK^{+}21]. curve [ET24, ZLV23]. curved [PMS20b, YXX^{+}21]. curves [DZZ21]. curvilinear [ALB22, MSM24, PD23]. custom [Kan23, Kan24]. customizable [Mar22, ZTV^{+}24]. cutoff [Mar22]. CVODE [PTH24]. cycle [VTB^{+}21]. cyclic [LF20]. Cyclically [CGR21]. cyclotron [DFP^{+}20, FYM^{+}22, WXX24]. cylindrical [Flo24, LEE^{+}21a, WXX24].

D [YYZ^{+}22, ALB22, ATC^{+}23, BKP22, BRAC23, BGW^{+}22, CAC^{+}22, CZS^{+}21, CCW20, Dau23, DFG^{+}23, DRZ^{+}21, DBH^{+}22, GTE21, GPM^{+}21, GUW^{+}22, GSL^{+}23, GOS^{+}22, IGL^{+}24, JLW24, JĎD^{+}21, KWZ24, KNJ^{+}23, LEE^{+}21a, LLZ^{+}23, LS22, LLH^{+}21b, MMP20, MPH^{+}24, NUK^{+}22, NSY^{+}23, NT24, PMS^{+}20a, PGYF21, RMCC22, RBFB20, Sch21, SP20, SLDF^{+}21, SCL22, SBB^{+}24, TCSA21, WJB21, WWB22, WKBW21, WZC21, WT22, WWDM20, XFH^{+}22, XLG^{+}23, YNV22, YHY^{+}21, YKY^{+}22, ZGW20, ZL20]. D-2V [TCSA21]. D-3V [DFG^{+}23]. D-VIRTUS [MMP20]. D-Wave [WWDM20]. D/3 [Dau23]. damping [CFW^{+}23]. DAMQT [KLM^{+}22]. Darcy [Yan21b]. Dark [BHK^{+}21, ABB^{+}24, EMM^{+}23, LSZ23, May21, PAM24, BCD^{+}21, GUAD22, OBC^{+}24]. DarkNews [AZH^{+}24]. DarkPACK [PAM24]. Data [AAA^{+}20, CFW^{+}23, HCAH^{+}24, HSO^{+}22, KvH21, MMYU22, MYM^{+}22, SMC^{+}22, ACD^{+}22, AES21, ANU21, BYL^{+}21, BS23, CGG21, DR21, DFP^{+}20, Dry21, GvdBdGN24, Hua23, JKKN23, KvdW20, KIK20, KFHR24, LM22, LSF23, OSE^{+}20, PNL^{+}24, PP21, RDH^{+}20, RdSH^{+}24, SPTPR21, SJ20, SRML20, SLIC24, SCF20, SHRK22, TRN^{+}24, TC24, WBF^{+}24, WKJB23, WYZZ23, XLG^{+}23]. Data-analysis [MYM^{+}22, CGG21]. Data-driven [CFW^{+}23, HCAH^{+}24, HSO^{+}22, SMC^{+}22, LSF23, WYZZ23]. data-structures [SJ20]. datasets [BBV23]. day [BR20a]. DCA [HAM^{+}20a]. De-noising [TAGC22a]. Debugging [SAN24]. Debye [DRZ^{+}21, XFH^{+}22]. Debye-type [XFH^{+}22]. decay [AG23, ABWZ23, ABH^{+}23, Hor23b, LWhK^{+}20, LF20, Sat21, TSJ^{+}21, VFS23]. decays [AKK^{+}24, CV24, DDM20, KKM^{+}20, KKSV24, KM20, KMS20b]. December [Ano20d, Ano21c, Ano22c, Ano23c]. decomposition [GHL23, KFHR24, LGDF20, LAD^{+}21, LY24b, MRD23, OCR^{+}22, TGS^{+}20, VPPQ21, YYY22, ZLY24]. decompositions [CKT21, RYS^{+}24]. decoupled [LM21a, Yan21b]. decoupling [ZWZ^{+}22]. Deep [GUAD22, MM22, TSAK21, TGIM23, YYZ^{+}22, ZGW20, BKG^{+}23, KKY24, MLZ^{+}23, ZWC^{+}20, LWC^{+}21]. DeepBHCP [MM22]. DeepFlame} [MLZ^{+}23]. DeePKS [CZWE23]. DeePKS-kit [CZWE23]. defect [Ham20b, Pot24, TQGE23]. defects [AM21a, DIAA21, TR22, vTDGCR21]. defined [KKSV24]. definition [YZL^{+}23]. deflation [Kub24]. deformation [AE23, SN23]. deformed [MSN^{+}22, VvBTH20]. degeneracy [TRN^{+}24]. degree [LMWW24a, WJB21]. DeHNSSo [WHSK24]. Delft [WHSK24]. delta [Bul21, CJD^{+}20, MM23]. delta- [CJD^{+}20]. DEM-Engine [ZTV^{+}24]. Deneb [YKK23b]. denoising [HGS20]. dense [BCTS22, BRHT21, KHR^{+}23, KKL21, MHK24, zYMK^{+}21]. densities [WKBW21]. Density [MVK^{+}24, QJ21, BGW^{+}22, BW23, CPL21, CZWE23, DMS^{+}22, DH20, FASD20, FWL^{+}24a, GM20, GBR23, GBD^{+}22, HFP21, IJVJ21, IUJ21, JRS^{+}21, Kar23a, KLM^{+}22, LYC20, LR22, MHK24, MDP22, MDR^{+}20, Pag24, SLE^{+}22, SXW^{+}20, SLZY21, SHW^{+}21, TMC22, VCF22, YT23, ZLLM23]. density-functional [GBD^{+}22, SXW^{+}20, SLZY21, TMC22]. density-in [Kar23a]. density-in/dependent [Kar23a]. DensityTool [LR22]. dependence [BAB^{+}20, YZL^{+}23]. dependencies [AS24a]. Dependent [BBB20, MVK^{+}24, BAB^{+}20, Dio23, GSLS20, GBD^{+}22, HSB^{+}24, Hor23a, JRS^{+}21, KPST21, Kar23a, KKPC23, KM23, LY22, LY24b, NLS24, SGS23, Scr22a, WJB21, WWB22, WHB21, WBvdH20, XSL^{+}22, YSMBA23]. depletion [BKB^{+}21]. Deploying [LSF23]. deposition [HTL^{+}22, KPR^{+}24, SWB^{+}23, TLC^{+}21]. derivation [NYN^{+}21, PSW23, PSW23]. derivatives [GOST23, Rod22]. derived [CMJC21, RG21]. DescMAP [ZLV23]. described [MBCC23]. description [GSV23]. descriptor [ZLV23]. descriptors [HJM^{+}20]. design [AGMFGE23, AS24a, CZX^{+}21, HSB^{+}24, KL23b, MMCC^{+}22, MOA24, RdSH^{+}24, VB22a, XLG^{+}23, YWM23]. designs [HK24]. detailed [RCX24, SPF21]. detect [KIK20]. detection [EGKH24, JKST22, KRJ23, LMQ^{+}23, PNL^{+}24]. detector [CGG21, Kan24, NRG22, NRG24, TAGC22b]. detectors [ABGD^{+}20, KTF22]. determinants [AJDS^{+}21]. determination [HSO^{+}22, LN23, NBC^{+}24, PWC24, ZLV23]. determine [DEdM24, HHMH^{+}22, MDP22]. deterministic [CHA21]. detonation [PD23, SWTC23]. detonationFoam [SWTC23]. detrended [GHKW22]. develop [WAET22]. developed [JMOC21, LHWX24, MG22]. Developing [WRM^{+}24, CZWE23]. Development [BKP22, COJ^{+}22, CKC^{+}21, CLEPF22, DBV^{+}24, JKSY22, KMN21, LBM^{+}23, LIG^{+}24, MGX24, OBC^{+}24, DARJ23, VLS22, ZMC23, ZDSS23, ZYG21, DBBP23, DSW^{+}23, HWL^{+}23, OIA^{+}20]. developments [GAA^{+}20]. devices [BMI23, BM22, HSB^{+}24, KK20, KNJ^{+}23, RMCC22, RdSH^{+}24, ZHM^{+}24]. dfemtoolz [MVF20]. DFMSPH22 [GCK21]. DFT [DMS^{+}22, GOCSS^{+}23, LG23, MSG^{+}21, MDR^{+}20, RCP^{+}24, VMRFC23]. DFT-FE [DMS^{+}22, MDR^{+}20]. DG [PMK^{+}23, YKK23b]. diagnostic [PSMRS^{+}23]. diagnostics [JSS^{+}24]. diagonalization [ACS^{+}22, KF23, LQ21, LQGL21]. diagonally [KBSL22]. diagram [BMI23]. diagrammatic [MSY^{+}21]. diagrams [ATRD21, Gro23, MSH22, MRD23, PCL24, RUV20, WSRO24]. diamond-difference [CL21]. DIAPHANE [RDC^{+}20]. Diatomic [RV21, YZ20b, BGHC23]. Diatomic-py [BGHC23]. diblock [CZY20]. dichotomy [TCY^{+}21, YX24]. dichroism [NG21]. Dicke [VCHH23]. dielectric [MHÅ21]. difference [CL21, Cos22, CESD^{+}23, GMZ^{+}20, KBSH20, SP20, WDMZ24, Zho23, dSOZ22]. difference/spectral [GMZ^{+}20]. difference/spectral-Galerkin [GMZ^{+}20]. differences [BGH22, GBH20]. different [HWZ24, KFC^{+}20, PLL24, SLZ^{+}24]. differentiable [AS24b, BBA23, FSMM24, MFS^{+}22, XLG^{+}23]. Differential [ASW20, AS22, ASU^{+}21, DEdM24, Jab24, LWhK^{+}20, Lee21b, ZLY24]. differentiation [MMM20, RCP^{+}24, PCS^{+}23]. DiffExp [Hid21]. diffraction [HMYH22, HSO^{+}22, KYH24]. diffractive [ST23b]. diffuse [LY24a, YWTK23]. diffuse-interface [YWTK23]. diffusion [BB24b, DRM20, GBS^{+}20, GHL^{+}22, GMZ^{+}20, HSMR^{+}24, KBB21, KHR^{+}23, LN23, PTH24, PGYF21, SRML20, SS21b, TWW22, WSK^{+}22, XDF20, YFL22, ZYX21, ZYX22]. diffusive [MP21]. dihedral [MBCC23]. dilepton [BKBL22]. dimension [GSLS20, GG24, ZJ23, dMMLOS20]. Dimensional [RBV^{+}22, AHP21, BV21, BM22, BKO20, BR20b, BSK^{+}22, CLS^{+}22a, CL21, Dan24, DS20, DH20, DKRSR22, GMZ^{+}20, HT24, Hid21, KLD^{+}21, KM23, KMD^{+}21, LW24, LMQ^{+}23, LHTP^{+}24, MD20b, MSH22, MFB23, MOMO24, MYM^{+}22, MCP23, NA20, OCE^{+}23, PMK^{+}23, PD23, SRT^{+}20, SLZ^{+}24, SMC^{+}22, SXYD24, TTM22, UÁEPGBP24, WXZH24, WZPW20, ZDG^{+}21, Zho23]. dimensionless [RDZ^{+}20]. dimensions [EPM23, JRS^{+}21, SRE^{+}24]. dimuonium [UAS22]. dimuonium-matter [UAS22]. dipolar [YSMBA23]. dipole [Arn20, KR23]. Dirac [FKK^{+}21, GOST23, Kut24, LY22, MN21, SJP21a]. Direct [IJVJ21, RV21, ZYXC24, BOSM24, BMSP21, GUAD22, GWA^{+}23, JKST22, NUK^{+}22, WWZ^{+}23, YFL22, YCCW23, ZPZH20]. direct-addressing [YCCW23]. direct-drive [NUK^{+}22]. Direct-Fit [RV21]. Direct/split [ZYXC24]. direction [ALB22]. directional [BA24]. DIRQFAM [BN20, BN23]. discharges [DBR24, MMP20]. disconnected [KDHL23, LMWW24a, WSRO24]. discontinuity [CAC^{+}22]. Discontinuous [CJ21, DGM20, SVSC20, SDBS24, FRN^{+}23, FMHH24, HWL^{+}23, HWZ24, IPSW24, JKSY22, KSJ^{+}22, KMD^{+}21, MGX24, MPQ^{+}22, QWZ^{+}21, SMA24, SS21b, YXX^{+}21, ZJS^{+}20]. Discovering [TC24]. Discrete [LSW^{+}20, MYMK^{+}21, ZTV^{+}24, BSK^{+}22, CFBRE24, GLB^{+}21, HK24, HTL^{+}22, JGJ20, KCS22, KMR22, LHC20, LHG24, ONH^{+}20, QCZ23, SRC21, YW21, Zha20, ZWP^{+}22, ZZ21a, KWZ24]. discrete-event [SRC21]. discrete-time [JGJ20]. discretisation [SBZ23, WJB21]. discretization [DMS^{+}22, EOR21, GBS^{+}20, LC24, PMK^{+}23, SVSC20, SFS22, TKS22]. disjoining [NLS24]. disk [RDR^{+}20]. Disks [DG20]. dislocation [YZ20a]. disordered [MBCC23, TRB20b, YLLW24]. dispatch [SAC^{+}21]. dispersion [KR23, LLL24, MSH22, Pöt20, PPK22, RVRT22]. dispersive [DNG^{+}20, DRZ^{+}21, XFH^{+}22]. displacement [GLSH21, XSM22]. disruption [HEF21]. dissipation [KFC^{+}20, Kul20]. Dissipative [DVC^{+}22, DHE^{+}24, JCM20, CGSO20, DC22, DH20, UD24, XBL^{+}20, BKB^{+}21]. distance [Zha20]. distances [TGBM22]. distinguishability [OV23]. distorted [OYC24, PGYF21]. distorted-wave [OYC24]. Distributed [NRG22, GMZ^{+}20, KvH21, KFHR24, RSC^{+}22, TPS^{+}24, zYMK^{+}21]. distributed-order [GMZ^{+}20]. Distribution [BJS^{+}23, BCSS24, CCMR21, GST21, KKB^{+}24, KAG24, NBC^{+}24, PK24, SXW^{+}20, YZ20a]. distributions [BKS24a, BFL^{+}22, BFMA^{+}24, KPST21, Pos22, SJ21, SMC^{+}22, Xie23]. Disturbance [LJH^{+}23, HJGL22]. Divergence [LGDF20, MOV21, YZW21]. Divergence-free [LGDF20, YZW21]. divertor [YHH^{+}20]. divided [BGH22, GBH20]. DIZET [AGK^{+}23]. DLBFoam [TPK^{+}21]. DMFT [MSY^{+}21]. DMFTwDFT [SHW^{+}21]. dmscatter [GJJN23]. DNMOGA [WYHW24]. DNS [LBS^{+}23]. DoFun [HCP20]. Domain [SDBS24, AS24a, DGM20, DRZ^{+}21, ET24, HPAW21, KD23, LY24b, MRD23, PNL^{+}24, PPK22, RRC^{+}24, TTM22, YZW21, YYY22, Yan23, ZLY24, ZWG^{+}24]. Domain-specific [SDBS24]. domains [CZS^{+}21, NR21]. DoNOF [PM21]. donor [BMR^{+}24a]. donor/acceptor [BMR^{+}24a]. Doppler [GRCT20]. dot [GLB^{+}21]. dots [PC21]. Double [VB22b, ABF^{+}23a, GCK21, Kar23a, TNL^{+}22]. double-folded [Kar23a]. Doublet [KM20, KMS20b]. down [Gon22]. down-sampling [Gon22]. DP [ZWC^{+}20]. DP-GEN [ZWC^{+}20]. DPD [LBM^{+}23]. Dragon [MHK24]. DRalgo [EST23]. Drawing [Gro23]. DREAM [HEF21]. Drell [AFJ^{+}23]. Drift [ZSqXY21, TAGC22a]. drive [NUK^{+}22]. driven [AE23, BM20, BGW^{+}22, CFW^{+}23, HCAH^{+}24, HSO^{+}22, LKK23, LLH21a, LSF23, PMSHG23, PLT^{+}23, SMC^{+}22, SBP20, WYZZ23, ZYG21]. Driver [BMREC21]. driving [FDPT23]. DRM [YKK23b]. DRM-DG [YKK23b]. droplet [AE23]. Droplets [PLSB22]. Drude [LW24]. DSc [Dry21]. DScribe [HJM^{+}20]. DSMC [ZWZ^{+}22]. DSQSS [MYMK^{+}21]. dual [CJZ21, NTO24, WXZH24, WVK21, ZTV^{+}24]. dual-GPU [ZTV^{+}24]. duct [OSE^{+}20]. Duffing [BDdM20]. DUGKS [KWZ24]. DUGKS-GPU [KWZ24]. Duguet [CCK23]. dump [Ham20b]. dune [BDR^{+}20]. during [Dan24, MRG22, ZDG^{+}21]. dust [LLT^{+}23]. dye [NI22]. Dynamic [MLT^{+}21, MRD23, BBB20, DRM20, KGT22, MRH^{+}23, PCS^{+}20, SXW^{+}20, TPK^{+}21, YD20, ZHZG23]. dynamical [CE22, DWD^{+}24, GLSH21, HTL^{+}22, JGJ20, MBH21, MEH21, TC24, SHW^{+}21]. dynamically [RCB^{+}20]. Dynamics [AMA^{+}20, DVC^{+}22, DHE^{+}24, JCM20, MKPW21, MKPW22, YD20, vRCM21, AFB^{+}24a, BB24a, BMR^{+}24a, BBA23, BCTS22, CLY22, CGSO20, CLS22b, CZTF23, CHY^{+}24, CDT22, CL22b, DTGE21, DC22, Dio23, DH20, FFTV23, GVV22, Ham20b, HyLF23, HWZ24, KF23, KCJX24, KMM21, LLT^{+}23, LLQ^{+}23, LN23, LF24, LSF23, LVMGF^{+}23, LJS21, MHP23, MHK24, MBCC23, MRG22, NRKA22, NS20, NI22, OEI^{+}22, OAP^{+}24, PCS^{+}20, Sch21, SKC21, SC22, SSPF24, SSD^{+}22, SJY18, SJY20, TM24, UD24, VXT^{+}23, VRI24, VCHH23, WNS^{+}21, WSK^{+}22, Wor20, WS23, XBL^{+}20, XQ21, XLKX21, ZHI23, ZJS^{+}20, ZPL^{+}21, BKB^{+}21, LWC^{+}21, OLNG21]. dynamos [TPCT22]. DynDen [DTGE21]. dyons [MCP23].

EAST [BYWW23]. easy [MD20b, WKBW21]. easy-to-implement [MD20b]. easy-to-use [WKBW21]. EasyScan HEP [SZ24a]. ECCPA [SLLA22]. ECOGEN [SPLD20]. ecology [OCE^{+}23]. ECPSSR [Hor24]. ECRad [DFP^{+}20]. Eddy [CJ21, VPS23, VTB^{+}21, YCC22]. eddy-current [VTB^{+}21]. edge [GUW^{+}22, RSHS24, SDXY23, SBZ23, WPMK21, WRM^{+}24, YWM23]. EDIpack [ACS^{+}22]. Editorial [Sco21, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano21o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k]. education [DBBP23]. effect [MD20a, YYC^{+}23]. Effective [BBB20, DPR^{+}20, DRR^{+}24, HLCD20, DZZ21, EST23, GTE21, HS24, NYN^{+}21, PC21, RTRB21, TM24]. Effects [LLT^{+}23, LLQ^{+}23, ASA^{+}22, BC21, CW22, DBM^{+}24, GVV22, JGJ20, VGGP^{+}21]. Efficacious [GLSH21]. Efficacy [MLD^{+}22]. Efficiency [KBSL22, TPCT22, HTL^{+}22, KRJ23, PP21, SKEZ24a, Tan24, XHY^{+}24]. Efficient [AJDS^{+}21, BP21, BW23, CE22, CZY20, DFG^{+}23, EPM23, EL24, GHKW22, HG22, JĎD^{+}21, KPST21, KCS22, KBB21, LG21a, LM21a, MK22, Mar22, PPK22, SRC21, SLZY21, SNG20, SFC20, TAY^{+}24, UÁEPGBP24, WGS^{+}22, YWTK23, Zeb22, ZCWY20, ZX23, AG21, AGJ^{+}23, BND22, CDD22, DEdM24, DBdFdSR21, EGLK20, GRCT20, GHL^{+}22, GP22, JZW^{+}22, JPJ^{+}23, KWZ24, KYH24, LBRW22, LLH21a, LAD^{+}21, LXY^{+}21, LLZ^{+}23, LLL24, LLZ^{+}22, LVMGF^{+}23, LSS24, MD21, MVF20, MRN20, PSMRS^{+}23, PPKK21, PWD22, RZH22, RZ23, RL21, STA20, SFKC22, SS21b, SWB^{+}23, WYT23, WS23, YSX^{+}20, ZPL^{+}24, ZHM21, dVAR^{+}24]. efficiently [YZW21]. EFIT [BYWW23]. EFT [CFL^{+}22]. eigenfunctions [BV21]. Eigenmode [DNG^{+}20]. eigenmodes [FBA^{+}20]. eigenproblems [zYMK^{+}21]. eigensolver [FKK^{+}21, zYMK^{+}21]. eigensolvers [AAMY23, FWL^{+}24a]. eigenspectrum [LG21a]. eigenvalue [DNG^{+}20, LLH^{+}21b, MG22, WHSG24, ZJ23]. eigenvalues [BV21, IMB^{+}22]. Eilmer [GDJG23]. Einstein [BKRG22, KRG21, LB24, MBA21, RVM^{+}21, SKDH24]. ElasT [LVB22]. ELAStic [FLK^{+}20, KC21, BFD22, BDA^{+}20, CT23, EL24, Jab22b, Jab24, LLS^{+}21, LLZ^{+}22, LEL^{+}22, LLY^{+}22, MD20a, NAZ^{+}22, ONH^{+}20, PB23, RZWW23, SJP21a, SLLA22, SLDF^{+}21, YNV22, YNV23, ZZZ^{+}20, NAZ^{+}21]. Elastic3rd [LLS^{+}21]. elasticity [GPM^{+}21, WGGC22, XZLX20]. ElasTool [EL24, LEL^{+}22]. elastoplastic [BR20b]. ElecTra [GLN23]. electric [AE23, WLCF21, YI22]. Electrical [HMM22]. electrified [Di 22]. electro [YC20]. electro-magnetic [YC20]. electrochemical [ZHM^{+}24]. electrode [PK24]. electrodynamics [FH22, GPN20]. electrolyzers [ZHM^{+}24]. electromagnetic [BVV22, BKO20, DRZ^{+}21, EES24, LOT^{+}20, LKP21, NTO24, OGL^{+}21, SWB^{+}23, WAYL23, YR21, ZLS^{+}22]. electromagnetics [WMM^{+}24]. electroMicroTransport [GDK21]. electromigrative [GDK21]. Electron [ABGD^{+}20, LACL24, Ada22, ASA^{+}22, AJDS^{+}21, BCF^{+}24, COW24, CW22, DRM20, DFP^{+}20, DVC^{+}22, FMCB^{+}20, FYM^{+}22, GBR23, GSV23, HAM^{+}20a, HMYH22, HM24, HA23, HEF21, Hor24, HBM^{+}24, HLzY^{+}20, Jab24, KKB^{+}24, LKP24, LSS24, MMP^{+}24, MFB23, MGC^{+}23, MAMK21, MPH^{+}24, MZD^{+}20, Nie22a, Pos22, SFBG20, SLE^{+}22, SLL22, SLZY21, SDL^{+}21, SP23, ST23b, TCcN23, TGIM23, WLF^{+}23, YZW22, ZPL^{+}21]. electron-atom [Pos22]. electron-impact [FMCB^{+}20]. electron-ion [ST23b]. electron-phonon [MMP^{+}24, ZPL^{+}21]. electron-positron [HA23]. electron/positron [HBM^{+}24]. Electronic [BLN^{+}21, BSS^{+}23, BGR23, DSQ23, DKRSR22, FASD20, GWPW21, GLN23, HY23, HTH^{+}20, HLzY^{+}20, JPJ^{+}23, MBG^{+}20, NVCS23, PAL^{+}20, PPKK21, PTD20, RL21, TCSD24, YHY^{+}21, zYCD^{+}20, ZHM21]. electronically [CMJC21]. electronically-derived [CMJC21]. electrons [CXCZ23, DTC20, Dan24, Jab22b, Jab24, MBG^{+}20, SJP21a, SLK23, ZYX21, ZYX22]. electrophoretic [KPR^{+}24]. electrophysiology [VSM^{+}22]. electrostatic [ATC^{+}23, CGR21, KG21, LSW^{+}20, YWX^{+}23, ZYX21, ZYX22]. electrostatics [FAL20, FAL21, LYX22, SOH^{+}23]. Electroweak [KM20, AJW^{+}21, BMM21, KMS20b]. Element [Sha21, ZTV^{+}24, ALB22, BM20, BCHE21, ČPF^{+}24, DMS^{+}22, EPM23, Flo24, GPM^{+}21, IGL^{+}24, JMOC21, KSDH23, KBB21, KM23, KD23, LLC^{+}23, LHC20, MPS^{+}24, MVF20, MDR^{+}20, MCB^{+}20, OCR^{+}22, ONH^{+}20, OKBM23, QCZ23, RSD20, SRT^{+}20, SKDH24, SLC^{+}22, SBB^{+}24, TCD20, TTM22, WZZS21, WWZ^{+}23, WXZH24, XDF20, XLG^{+}23, YT23, YWM23, YJ24, ZZ21a, VB22a]. element-based [LLC^{+}23]. element/multipole [YT23]. Elementary [PDD24]. Elements [GSL24, AIZ23, BGH22, DBV^{+}24, LZ21, RZ23, TIG^{+}24, TWW22]. ellipses [CTPS22]. elliptic [BSK^{+}22, BDR^{+}20, WW21]. ELMAG [BKO20]. ELPA2 [zYMK^{+}21]. elsepa [SJP21a]. ELSI [zYCD^{+}20]. Embedded [AAG^{+}24, TW21a]. embedding [AMK24, QJ21]. EMcLAW [MOV21]. emerging [DRB22]. emission [CML^{+}24]. Empathes [BFD22]. empirical [CWG^{+}21, KVSC21, MMCC^{+}22, NA20, Pot24]. employing [VMRFC23]. empowered [MLZ^{+}23]. emulsions [GUAD22, PLSB22]. enable [XQ21]. enabled [GSL24]. Enabling [OBGA24, TDR^{+}20]. encapsulated [Nie23]. encoded [ZPL^{+}24]. encoders [TAGC22a]. Encoding [GPN20]. end [MFS^{+}22]. end-to-end [MFS^{+}22]. ended [YMCF23]. endpoint [LP24]. energetic [JSS^{+}24, LJQ^{+}22]. energetic-particle [JSS^{+}24]. energetics [TQGE23]. Energies [AJW^{+}21, GTMB21, GBJ^{+}21, Pos22]. Energies-Updates [AJW^{+}21]. energization [WXX24]. Energy [BMR^{+}23, BTK24, PMK^{+}23, AGJ^{+}23, ABB^{+}22, BRAC23, BC21, BCCM^{+}24, CZWE23, CGV^{+}22, Dan24, EES24, FXZT21, FTG23, Gar21, GZW20, GGCW24, GFD^{+}24, HMYH22, HTL^{+}22, HSO^{+}22, IJVJ21, IUJ21, JS22, KSF^{+}22, Kut20, LM21a, LW24, MD20a, Man24, MBTB21, MSHP02, MSHP20, NRK^{+}21, NA20, NYN^{+}21, NBB^{+}21, PRR23, Pöt20, RG21, RC23, RV20, RV21, SRC21, SE24, SLL22, SMC^{+}22, SVJ^{+}24, Sij23, TV24, TLC^{+}21, UO20, VZ20, VvBTH20, WZZ^{+}23, Xav22, XZLX20, YWTK23, Yan24a, ZWC^{+}20, ZCWY20, ZDLS21]. energy-angle [SMC^{+}22]. Energy-conserving [PMK^{+}23, BC21]. energy-momentum [Pöt20]. energy-preserving [GGCW24, LW24]. energy-stable [KSF^{+}22, YWTK23, Yan24a]. energy-strain [XZLX20]. Engine [GBJ^{+}21, GLW^{+}23, RCB^{+}20, ZTV^{+}24]. Engineering [ANU21, FYW23]. enhance [MPS^{+}24]. enhanced [BBC^{+}24, CWG^{+}21, MLT^{+}21, RdPS24, SBB^{+}24, XLKX21]. enhancement [TTM22]. Enhancing [KHKL24, MCB^{+}20, SLL22]. ensemble [SHRK22, TNL^{+}22]. ensembles [GLB^{+}21, VCHH23, WGS^{+}22]. Ensign [CE22]. entangled [CL22a, ECS23]. Entanglement [HPP23]. entanglements [KDHL23]. entropy [MMCC^{+}22]. entry [GSL^{+}23]. environment [GAA^{+}20, OIA^{+}20, OSE^{+}20, Scr22a]. environments [ART^{+}20]. ePDFpy [KKB^{+}24]. EPI [MMP^{+}24]. epidemic [CF21]. EPW [LACL24]. eQE [MSG^{+}21]. EQMO [ZHM21]. equal [Efr23]. Equation [GMPG^{+}21, LS21b, LY22, OSE^{+}20, RS20, ZHM21, AIZ23, ATC^{+}23, BB24b, BAB^{+}20, Bul21, BBA^{+}20, CS22a, CKT21, CCW20, CWJ21, DSSW22, ENK24, FN21, GSLS20, GLPG^{+}23, GMZ^{+}20, HXS20, HWZ24, KMR22, KPL^{+}21, KFPV21, LZ21, LHC20, LFL^{+}24, LY24b, LSS24, OEI^{+}22, OKBM23, PPR^{+}21, SVSC20, SGS23, SP20, SLZ^{+}24, SCL22, TTM22, TS23, WHB21, WXY20, WZC21, WKR23, XSL^{+}22, YSMBA23, YNMR24, YJ24, ZAW^{+}21, dVAR^{+}24, ZXT22, ZXT23]. equations [ASU^{+}21, AGH21, BV21, BKRG22, BCGT24, CR20, CLEPF23, CLEP24, DGM20, DH22, DEdM24, GTA21, GHL^{+}22, GGCW24, HPY21, HWL^{+}23, HCP20, JM24, JZW^{+}22, JRS^{+}21, KRG21, KBB21, KSF^{+}22, KMD^{+}21, Koz23, KRE22, LWhK^{+}20, LGDF20, LW24, MSN^{+}22, MOV21, PMK^{+}23, PLL24, PSW23, PG23, Pöt20, QWZ^{+}21, RJ21, TSL21, VFS23, WGG20, WYZZ23, XDF20, XPF^{+}24, YZW21, YXX^{+}21, YKY^{+}22, ZCP^{+}22, ZLY24, ZX23, ZZZ^{+}20]. equilibria [BJS^{+}23]. Equilibrium [MMP20, BYWW23, Dau23, DVC^{+}22, HPAW21, LBM^{+}23, LZ21, MES^{+}24, MEC^{+}24, SYFT23, WJB21, WWB22, WHB21, YKYK23, ZCS^{+}24, SGM^{+}20]. ERCS08 [Hor24]. ERCS24 [Hor24]. ERNIE [AG23]. erosion [YGSW21]. Erratum [FAL21, SJY20, YNV23]. error [JKKN23, MZ22, MNS^{+}24, SSH^{+}23, THH21]. errors [RVRT22]. ERSN [LEE^{+}24]. ERSN-OpenMC-Py [LEE^{+}24]. ERYA [MMC^{+}22]. ERYA-Bulk [MMC^{+}22]. ERYA-Profiling [MMC^{+}22]. escape [NA20]. ESPRESSO [HHT^{+}24, MONW21]. Essentially [MGX24]. estimate [EBBB22]. estimating [GO23, GBJ^{+}21, GG24, SKC21, WV22, YMCF23]. Estimation [VEHCM21, AM21b, LMWW24b, LWV20, MPN^{+}21, NG21]. Euler [GWA^{+}23, HWL^{+}23, YXX^{+}21]. Eulerian [LMHL20, MSM24, ODR^{+}22, Sij23, TKC^{+}21, THH21, WWM^{+}22]. EUTERPE [KBH^{+}24]. evaluate [FLW^{+}23]. Evaluating [MYKC23, NVC20, BA24, BTW20, Bzo21]. Evaluation [ABD^{+}23, GOST23, PSK^{+}24, ABF^{+}23a, Arn20, ATRD21, BC20, HAA^{+}20, NSU20, PBC^{+}24, SRML20, SSV22, TR22, WW21, XZLX20]. evaporation [AYB24, HyLF23, OSE^{+}20]. event [AAA^{+}21, AZH^{+}24, ACD^{+}22, AG23, BIK^{+}21, BKG^{+}23, For22, GPD^{+}22, HQF^{+}20, JWW^{+}23, LTMK21, PNL^{+}24, SRC21, SJP^{+}21b, Sjö20, SBP20, VEHCM21, YLK21, ZDLS21]. event-chain [HQF^{+}20, LTMK21]. event-driven [SBP20]. event-time-frame [GPD^{+}22]. events [DRSZ23, Hor23b, SKC21]. EvGen [BKG^{+}23]. Evolution [OSLC21, ZDG^{+}21, BCGT24, GF23, HFP21, MZ22, VGGP^{+}21, ZZ21b]. Evolutionary [DR21, KMG^{+}20, HTS^{+}21, WBK^{+}24]. evolving [LM22, OS24]. EW [SY20]. ewN2HDECAY [KM20]. Exact [GF23, ACS^{+}22, Bak23, BR20a, BBA^{+}20, GQ22, KW23, LG21a, LQ21, LQGL21, LHTP^{+}24, Pag24, RC23, Tor20]. Exactly [Man24]. exactness [FFLR20]. ExaHyPE [RCB^{+}20]. example [YYZ^{+}22, ZGW20]. examples [CLVV22, SOH^{+}23]. exascale [DRB22, DFU20, ZDSS23]. exascale-oriented [DFU20]. exchange [BR20a, CSW24, XLKX21]. excitations [LL23, UÁEPGBP24]. excited [NRK^{+}21, QJ21, RHW^{+}21, RBWD^{+}24]. excitonic [DSQ23]. excitons [PC21]. exclusive [For22, ST23b]. expanded [RSPJ21]. Expanding [LTT^{+}24, Ber24, FFTV23]. Expansion [HJJ^{+}22, LHZ20, SNG20, YT23, Zho23]. expansions [ABD^{+}23, FMBD22, Hid21]. experiment [SCF20, YLK21]. experimental [BC20, MYM^{+}22]. experimentally [MES^{+}24]. Experiments [KMN21, AGJ^{+}23, BDP^{+}21, BCD^{+}21, DG20, OBC^{+}24, ZDLS21]. Explicit [QWZ^{+}21, Bel24, CJ21, DS22, PMK^{+}23, YZW21]. Exploiting [RZ23]. Exploration [HPRS23, MMCC^{+}22]. explorer [AES21, LSZ23]. Exploring [AS24a, DG20]. explosion [TRN^{+}24]. exponential [CCK23, EOR21, GDB10, GBH20, IJVJ21, QWZ^{+}21, YNMR24]. express [ZGZW23]. expression [ENK24]. extendable [KVSC21]. Extended [BCTS22, BJL^{+}24, BMM21, IJVJ21, KKSV24, TSL21]. Extending [KL23a]. Extensible [ZGZW23, GAGO21]. extension [DSSW22, HYL^{+}22, KMS20a, LHG^{+}20, LLR23, MCMS24, RMCC22, SVSC20, SDXY23]. extensions [CK23, MSY^{+}21]. extensive [CNB^{+}23]. exterior [BSK^{+}22, KMR22]. external [LG21b, YWX^{+}23]. Extraction [PK24, LS21a]. extragalactic [BKO20]. extrapolation [KW23, YLL^{+}22]. extreme [CGSO20, GH21, HXS20, LYZL24]. extreme-scale [HXS20]. EZ [SWB^{+}23].

FabSim3 [GAS^{+}23]. facilitate [LSF23]. Facilitating [FDPT23, WXL^{+}21]. FACt [RPG^{+}20]. factor [CCK23, GDB10, SZNW23]. factorised [HHT22]. factorization [CCC20]. factors [MKPW21, MKPW22]. false [Sat21]. far [DVC^{+}22, WJB21, WHB21]. far-from-equilibrium [WHB21]. far-off-equilibrium [DVC^{+}22]. far-out-of-equilibrium [WJB21]. FaSE [TW21a]. FaSE-GLoBES [TW21a]. Fast [CLY22, DCZ23, HAA^{+}20, Ilt21, IUJ21, KF23, Pag24, WZC21, WOP^{+}20, BV21, BDGS21, CGR21, CB20, ENK24, FBA^{+}20, GJ24, GJJN23, GB20, KKLZ23a, KIK20, KT23, KYH24, MSH22, OCE^{+}23, STA20, SSD^{+}22, TCD20, TTM22, TPK^{+}21, TLC^{+}21, TBAR21, WVK21, WXX24, Xie23, XOTI22, YLLW24]. Fast-QSGS [YLLW24]. fast-update [XOTI22]. faster [TB20]. FaVAD [vTDGCR21]. FCC [TCY23]. FDTD [KGN^{+}21, LKP21, LW24, XPF^{+}24]. FE [DMS^{+}22, MDR^{+}20, UO20]. FE-CLIP [UO20]. FeAmGen.jl [fWL24b]. feature [ABK^{+}22]. feature-rich [ABK^{+}22]. features [LTT^{+}24, SLK23, SMO20b, WHSK24]. February [Ano20q, Ano21p, Ano22p, Ano23p, Ano24l]. FEL [ASA^{+}22]. FELINE [SLC^{+}22]. FEM [EVMP20, XLG^{+}23]. FEMPAR [BM20]. FEMS [Sha21]. femto [Koz23]. femto- [Koz23]. Fermi [GOST23, MN21, YNMR24]. fermion [DDM20]. fermionic [XOTI22]. fermions [Dür23]. ferroelectric [KNJ^{+}23, ZWC22]. ferromagnetic [LL23]. ferromagnetic/non [LL23]. ferromagnetic/non-magnetic [LL23]. FerroX [KNJ^{+}23]. few [WLL22]. few-group [WLL22]. FeynCalc [SMO20b]. FeynGame [HKL20]. FeynGKZ [ABBD23]. FeynGrav [Lat23]. FeynHiggs [BHH^{+}20]. Feynman [ABBD23, ABD^{+}23, BMT23, CS22b, DPR^{+}20, DRR^{+}24, FR20, Gro23, Hid21, LM23, Nog21a, Nog21b, NVC20, SC20, SSV22, SZ24b, WW21, fWL24b, WBM^{+}24]. FeynMaster [FR20]. FeynMG [MCMS24]. FeynRules [MCMS24]. FFT [Cos22]. FFT-accelerated [Cos22]. FFTW [DSSW22]. fibre [AAB^{+}21]. fibrogenesis [ZKZ^{+}24]. fictious [Yan23]. fidelity [BMSP21, CKC^{+}21, KSC^{+}23, MCB^{+}20, RdPS24, XLL^{+}22]. Field [DPR^{+}20, DRR^{+}24, SHW^{+}21, AIZ23, AE23, BN20, BN23, Che23, CZY20, CCL^{+}22, DH22, DKRSR22, EGLK20, EST23, FDPT23, FFTV23, Flo24, FFLR20, GK21, GJA21, GBS^{+}20, GNP20, HS24, KKLZ23a, KKLZ23b, KLD^{+}21, KLMU21, KNJ^{+}23, KW23, LGBJ20, LMX^{+}21, LCL^{+}23, LHTP^{+}24, MM23, MT23, NRG22, NRG24, TCY23, WMA^{+}22, WYT23, XPF^{+}24, Yan21a, YK21a, YYY22, YML^{+}24, ZCWY20, Zha20, ZDG^{+}21]. field-aligned [GBS^{+}20]. field-split [YYY22]. fields [BKO20, CMS22, GH21, MYKC23, MPQ^{+}22, RHW^{+}21, RBWD^{+}24, YI22, YYZ^{+}22, ZKZ^{+}24]. FIESTA5 [SSV22]. filaments [HMM22]. file [LG21a]. film [BMP^{+}24, RMR^{+}22]. filter [BDGS21, CJ21, NDFL24, WHSG24]. filtering [LYC20, PP21, XPF^{+}24]. final [ABWZ23, BOSM24, DDM20, SAS20]. find [MRT^{+}22, NKP20]. FindBounce [GNP20]. finder [EGKH24]. finding [Bar22a, BBB^{+}21, ZWG^{+}24]. Fine [WWZ^{+}23]. Fine-grained [WWZ^{+}23]. Fingerprinting [SLL^{+}21]. Finite [BC21, GSL24, KSDH23, Sha21, SLC^{+}22, AIZ23, ALB22, ADF^{+}22, BM20, BCHE21, BN20, BN23, ČPF^{+}24, Che23, Cos22, CESD^{+}23, DMS^{+}22, Dau23, DS22, DVC^{+}22, EPM23, FZD^{+}24, GJA21, GCWZ20, GMZ^{+}20, IGL^{+}24, IJVJ21, JMOC21, KBB21, KLMU21, KM23, KFC^{+}20, KBSH20, LLC^{+}23, MVF20, MSM24, MDR^{+}20, OCR^{+}22, OKBM23, PGYF21, RSD20, SRT^{+}20, SKDH24, SP20, SKC21, SBB^{+}24, VGGP^{+}21, WZZS21, WWZ^{+}23, WDMZ24, WH23, XDF20, XZLX20, XLKX21, XLG^{+}23, YT23, YI22, YWM23, Zho23, ZLS^{+}22, dSOZ22, PLSB22, VB22a]. finite-difference [Cos22, SP20]. finite-element [DMS^{+}22, IGL^{+}24, MDR^{+}20, RSD20, SRT^{+}20]. finite-order [ZLS^{+}22]. FINite-size [PLSB22]. finite-sized [Che23]. finite-temperature [GJA21]. finite-time [SKC21, XLKX21]. finite-volume [ADF^{+}22, Dau23, FZD^{+}24, VGGP^{+}21]. FiniteFieldSolve [Man24]. FIPI [GB20]. FIRE [PGS^{+}24, SZ24b]. FIRE6 [SC20]. FireFly [KL20, KKL21]. First [Dan24, Och23a, Och23b, YI22, AM21a, BCTS22, DEdM24, KvdW20, KD23, LLS^{+}21, LLZ^{+}22, MBTB21, MPN^{+}21, PB23, SDL^{+}21, TAE^{+}21, YJLW21]. first- [KD23]. First-principles [Dan24, Och23a, Och23b, YI22, AM21a, BCTS22, KvdW20, LLS^{+}21, LLZ^{+}22, MPN^{+}21, PB23, SDL^{+}21, YJLW21]. Fission [TSJ^{+}21, NRKA22]. Fit [RV21]. FITEVT [Hor23b]. fits [BS23]. fitting [ADdM20, BDGS21, CAC^{+}22, KHKL24, RV20, RV21]. five [SMGK21]. five-particle [SMGK21]. fix [DACA^{+}22]. Fixed [MBCC23, Ano20c, BCD^{+}21, KK16, OBC^{+}24, VB22b]. fixed-point [Ano20c, KK16]. FLAME [ART^{+}20]. flames [CEC^{+}24]. flat [SJ21]. flavor [KKSV24]. Flavour [TW21a]. FLEKS [CTZW23]. flexible [ABK^{+}22, CTZW23, MSM24, MDP22, SXYD24, TAB^{+}22, WKBW21, WOP^{+}20, WS23, YTC^{+}20, YMCF23]. flexible-structure [MSM24]. FlexibleDecay [ABH^{+}23]. FlexibleSUSY [KKSV24]. flight [WBF^{+}24]. floating [SSH^{+}23]. flow [AGH21, BSC^{+}21, CJZ21, CL22a, CNS22, FRN^{+}23, GDJG23, GZW20, KWZ24, KBSL22, KT23, KRE22, KBSH20, LIG^{+}24, LM23, MA24, MLZ^{+}23, NJSY22, OSK^{+}21, QCZ23, RLW^{+}24, SSPF24, SFS22, Tan24, TPK^{+}21, TGS^{+}20, WA21, XBL^{+}20, YCC22, Yan21b, YYY22, YCCW23, YKK23b, YZ20b, ZMC23, ZHZG23]. flows [AHP21, BBV^{+}22, Bak23, BL21, BMSP21, BBA23, BKS24a, CAWK22, CS22a, Cos22, CESD^{+}23, Dau23, DAC^{+}23, DS22, DBV^{+}24, FA20b, FJ22, FTZ^{+}23, FMBD22, GWA^{+}23, GB22, GP22, HRG^{+}22, HJGL22, KFHR24, Kul20, LMHL20, LJH^{+}23, LMQ^{+}23, LHG24, MEC^{+}24, MD21, OCC23, RCX24, RCW^{+}20, SPLD20, SXYD24, SFC20, TO21, VPPQ21, WT22, WICA22, YWTK23, ZYML24]. fluctuation [CT23, GHKW22]. fluctuations [RPG^{+}20]. Fluid [AMA^{+}20, OLNG21, AHP21, AFB^{+}24a, AGH21, BB24a, BBA23, CJZ21, CFW^{+}23, DIK^{+}23, DACA^{+}22, DH20, EPM23, FA20a, FA20b, GB20, HEF21, JYL^{+}23, JMOC21, KKPC23, KP23, KM23, LBS^{+}23, LY24a, LSF23, LHG^{+}20, LH22, LH24, LJS21, MBH21, MEH21, MPH^{+}24, MSM24, NJSY22, DARJ23, PG23, SFC20, eSdSBST21, VSM^{+}22, WICA22, WXZH24, WNS^{+}21, WSK^{+}22, WYT23, YSX^{+}20, ZSqXY21]. fluid-kinetic [HEF21]. fluid-structure [LBS^{+}23, YSX^{+}20]. fluid-structure-electrophysiology [VSM^{+}22]. fluid-structure-interaction [JMOC21]. fluid-surfactant [WYT23]. fluids [FBMD20, KSDH23, LBM^{+}23, MES^{+}24, BBA23]. FLUNED [DARJ23]. fluorescence [NI22]. FluTAS [CESD^{+}23]. Flux [DBV^{+}24, AWV22, CS22a, MBTB21, MSU^{+}21, RCW^{+}20, SKS24, TWW22, VÁFG^{+}22, VLS22, WH23, BBV^{+}22]. flux-coordinate [MSU^{+}21]. fly [BMREC21]. fmas [MD22]. foaming [ASC^{+}21]. foci [CMS22]. Fock [CLS^{+}22a, MSN^{+}22, RSM21, AYI^{+}24, WGG20, Zeb22]. focused [CMS22, SKEZ24a]. Fokker [TCSA21, TKC^{+}21, ZLL^{+}24, DTC20, KSJ^{+}22]. folded [Kar23a]. folding [GCK21, ZJS^{+}20]. following [WA21]. force [BGW^{+}22, BB21, EGKH24, HGS20, LCL^{+}23, LM20, MYKC23, XSM22, YKSH20, ZKZ^{+}24, EGKH24]. Force-chain [EGKH24]. force-chains [EGKH24]. forced [BDdM20]. forcefields [KMG^{+}20]. Forcer [RUV20]. forces [CMJC21, DACA^{+}22, FDPT23, KT23, NVC20, WT22, YI22, ZTV^{+}24]. forcing [JRG21, ZPZH20, Zha20]. Forecasting [CLS22b]. form [HKY^{+}21, KSJ^{+}22, LHC20, RUV20]. formalism [BJL^{+}24, CT23, Jab22a, SP23, ZCS^{+}24]. Formalization [CHA21]. format [GPD^{+}22]. formation [AMP^{+}21, JLW24, KLD^{+}21, RDZ^{+}20]. formed [MRG22]. formula [SA23, WLCF21]. formulation [ALB22, AGH21, TWR21]. formulations [PBC^{+}24, VTB^{+}21]. Förster [BMR^{+}23]. Fortnet [vdHKB^{+}23]. FORTRAN [BKRG22, Hor23b, KRG21, RPG^{+}20, NSU20, XHY^{+}24, YSMBA23]. FORTRESS [BKRG22, KRG21]. forward [BCT20, MD22, SL22]. four [DDM20, HYL^{+}22, RUV20]. four-fermion [DDM20]. four-loop [RUV20]. four-phonon [HYL^{+}22]. Fourier [ALB22, CCW20, FBMD20, FMBD22, KPST21, KL23b, LAD^{+}21, SBB^{+}24, YW21, ZYXC24]. Fourier-enhanced [SBB^{+}24]. FourPhonon [HYL^{+}22]. Fourth [PG23, KBB21, PB23]. Fourth-order [PG23, KBB21, PB23]. FPGA [KMN21, KHR^{+}23, KK20]. FPGA-based [KHR^{+}23]. FPGAs [WZK^{+}24]. Fractal [GSM^{+}22, Gro22, GG24, dMMLOS20, GOS^{+}22, GG24]. fraction [CCL^{+}22, GHL23]. fractional [CZS^{+}21, CLEPF22, CLEP24, GMZ^{+}20, HPY21, JLL^{+}24, MD20b]. fractional-step [HPY21]. fractionation [LWhK^{+}20]. fractioning [BBH^{+}24]. fractions [HvM22, TGBM22]. fracture [ZL20]. fragment [PWC24, TSJ^{+}21]. fragmentation [KL22, Roh22]. frame [GPD^{+}22, PP21]. framework [ADC^{+}21, ASU^{+}21, ACD^{+}22, ABF^{+}23b, BMREC21, BC20, BN20, BN23, CE22, CGG21, CK23, Dau23, FZD^{+}24, FSMM24, GAGO21, GOCSS^{+}23, GKT^{+}24, HAM^{+}20a, HWL^{+}23, HEF21, JYL^{+}23, JSS^{+}24, JKKN23, KWK^{+}21, KTF22, KSF^{+}22, KVSC21, KSC^{+}23, KGT22, KNJ^{+}23, LBS^{+}23, LYZL24, MEC^{+}24, MPS^{+}24, MW21, MPH^{+}24, MVF20, MYM^{+}22, NL23a, NT24, RSC^{+}22, RGS^{+}21, STRF^{+}20, UMA21, WXW^{+}21, WICA22, WAK22, WAET22, XFH^{+}22, YHY^{+}21, ZSqXY21, ZHZG23, OBK^{+}20]. frameworks [DAA^{+}24, VÁFG^{+}22]. FRAPCON [YLK^{+}20]. Free [VvBTH20, AMP^{+}21, BRAC23, BMR^{+}23, BW23, ELSV24, GM20, HPAW21, LGDF20, LMQ^{+}23, LMMP23, MS24, MM23, MHK24, NBC^{+}24, OKBM23, RG21, SE24, UO20, WZZ^{+}23, WA21, YZW21, ZYML24, Zho23, ZHZG23]. free-boundary [HPAW21]. free-surface [LMQ^{+}23, WA21, ZYML24]. frequency [DNG^{+}20, KM23, KD23, MMP20]. frequency-dispersive [DNG^{+}20]. FRET [BMR^{+}23]. FRET-Calc [BMR^{+}23]. freud [RDH^{+}20]. friendly [AMP^{+}21, BMR^{+}24b, HGS20, MMCC^{+}22, WXL^{+}21]. fringe [MRG22]. frontal [KMG^{+}20]. frontier [LNP^{+}24]. FSEI [VSM^{+}22]. FSEI-GPU [VSM^{+}22]. fuel [KJL^{+}23, NUK^{+}22, YLK^{+}20, ZHM^{+}24]. fuels [LLL24]. Fugaku [IKM^{+}23]. Full [GLN23, PMS^{+}20a, DRZ^{+}21, EGLK20, FBA^{+}20, FYM^{+}22, Ken24, LKP24, MG22, MSU^{+}21, SDXY23, SBB^{+}24, Tan23, WMA^{+}22, YHH^{+}20, YCC22, YML^{+}24]. full- [Ken24, MSU^{+}21, SBB^{+}24]. Full-band [GLN23]. full-field [EGLK20, WMA^{+}22, YML^{+}24]. full-orbit [FBA^{+}20, YHH^{+}20]. full-potential [Tan23]. full-scale [YCC22]. Fully [BCD^{+}21, MAMK21, MSM24, BBA23, DTC20, KSF^{+}22, LHC20, OBC^{+}24, Yan21b, YLH^{+}20]. fully-coupled [KSF^{+}22]. fully-decoupled [Yan21b]. fully-differentiable [BBA23]. FUMILIM [SAS20, SAN24]. function [AUEO24, AJDS^{+}21, ABF^{+}23a, Ano20c, BND22, BDdM20, Bul21, CZ21, CDT22, CLEP24, GCWZ20, GBH20, HY23, HFP21, Jab20, KKB^{+}24, LWS^{+}23, LM22, PPKK21, SWZ23, ZCS^{+}24]. Functional [HCP20, MVK^{+}24, WBK^{+}24, BW23, CZWE23, DMS^{+}22, FASD20, FWL^{+}24a, GM20, GBR23, GBD^{+}22, IPSW24, IJVJ21, LYC20, MHK24, MDR^{+}20, PSW23, PM21, SXW^{+}20, SLZY21, SHW^{+}21, TMC22, VCF22, YT23]. Functional-style [WBK^{+}24]. functionalities [WMA^{+}22]. functionality [BBC^{+}24, BC22]. functionalization [AUO^{+}22]. functionals [MSG^{+}21]. functions [ABFP24, Bag22, BA24, BGH22, Ber24, Bzo21, CCMR21, CCK23, CGV^{+}22, GDB10, GST21, HLzY^{+}20, HG22, IKM^{+}24, KK16, KHKL24, KL20, KKL21, Kor23, LP24, MSHP02, MSHP20, PSK^{+}24, Pot24, RBV^{+}22, SLE^{+}22, SLZG20, UZB22, Xav22]. fundamentally [SWZ23]. fusion [AYB24, DFP^{+}20, GBS^{+}20, HSB^{+}24, LLH21a, NUK^{+}22, NSY^{+}23, RdSH^{+}24, SMA24, TKC^{+}21, TBAR21, TWW22, Xie23, YGSW21, ZDSS23, ZLL^{+}24]. fusion-evaporation [AYB24]. future [Sjö20]. fuzzy [GG24].

G0W0 [RDV^{+}20]. gains [TPCT22]. galactic [Mau20]. Galerkin [CJ21, DGM20, FRN^{+}23, FMHH24, GMZ^{+}20, HWL^{+}23, HWZ24, IPSW24, JKSY22, KSJ^{+}22, KMD^{+}21, LS22, MGX24, QWZ^{+}21, SVSC20, SMA24, SS21b, SDBS24, YXX^{+}21]. gamma [Fis24, GST21]. GammaCHI [GST21]. Gamow [MAJ20]. Garfield [DBM^{+}24]. Gas [KWZ24, FJ22, FTZ^{+}23, LF24, LSW^{+}20, LHG24, MEC^{+}24, MGC^{+}23, RCB21, WZZ^{+}23, YZ20b, ZLLM23, ZWP^{+}22]. gaseous [ABGD^{+}20, SWTC23]. gases [ZWZ^{+}22]. gauge [FFTV23, GHK20]. gauge-Higgs [GHK20]. Gaunt [ÖAÖ24]. Gauss’ [OCR^{+}22]. Gaussian [RBV^{+}22, BTG22, HG22, PPK22, SKEZ24a]. GDML [ACKB23]. Geant [EMM^{+}23]. GEANT4 [LGK^{+}22, MK22, BCD^{+}21, KTF22, Kan23, Kan24, OBC^{+}24]. Gebremariam [CCK23]. Geliosphere [SBG23]. GEM [AFB^{+}24b]. GEN [ZWC^{+}20]. Genarris [TRB^{+}20a]. GenASiS [BC22, CB23]. gene [ENK24, CJD^{+}20, MSU^{+}21]. GENE-X [MSU^{+}21]. GenEOS [ENK24]. General [VFS23, XZLX20, ASPDL^{+}21, ABF^{+}23b, BFD22, BAB^{+}20, CFPS23, DCRF23, DC22, GK21, HT20, Koł22, LQ21, PWD22, RBV^{+}22, RV20, RV21, RVRT22, RMFB23, Wor20, XDF20, ZWC22, CLEPF22, DBdFdSR21]. general-purpose [ASPDL^{+}21, ABF^{+}23b, DCRF23, HT20, ZWC22]. Generalisation [BCGT24]. generalised [LP24, NSU20, TGS^{+}20]. Generalization [LQGL21, BDdM20]. Generalized [GCWZ20, HvM22, KAG24, MCP23, RZH22, TCY23, GOST23, YWX^{+}23, SOH^{+}23]. generate [KWK^{+}21, MMM23]. Generating [May21, TIG^{+}24, CB20, WBM^{+}24].

Generation [AYB24, fWL24b, ATRD21, CKGW22, FPSZ21, GKT^{+}24, Ilt21, KKY24, LJS21, NL23b, Nog21a, PMSHG23, SKEZ24a, WBF^{+}24, WLL22, WGS^{+}22, YLLW24, ZWC^{+}20, ZLV23]. generator [AAA^{+}21, AZH^{+}24, Ada22, AG23, BCT20, BKG^{+}23, DPR^{+}20, DRR^{+}24, For22, GUW^{+}22, JWW^{+}23, Koł23, KMS20a, LHZ20, Sjö20, TRB^{+}20a]. generators [BIK^{+}21, LAC21, LWS^{+}23]. generic [For22, HyLF23, ZDLS21]. genetic [CZX^{+}21, LFZ20, YNMR24, ZZ21b]. GenEvaPa [HyLF23]. GENIE [AAG^{+}20]. GENIE-based [AAG^{+}20]. Gennes [SKDH24]. geofluids [ENK24]. Geometric [TS23, KAG24, LH22, LH24, PBC^{+}24, WHSK24, XQ21]. geometrical [PMS^{+}20a]. geometries [COW24, CW22, HMR22, LHG^{+}20, OBGA24, SN23, SFKC22, SBP20, WAN^{+}22, WGS^{+}22]. geometry [ASPDL^{+}21, CZS^{+}21, JKSY22, KRC^{+}20, KBH^{+}24, LEE^{+}21a, LOT^{+}20, MG22, MSH22, SDBS24, TRN^{+}24, WZZS21]. geophysical [SGY24]. GeoTaichi [SGY24]. get [SLZY21]. GETaLM [Ada22]. GetDP [DNG^{+}20]. GFCCLib [PPKK21]. GGA [MSG^{+}21]. ghost [JGJ20, SXYD24]. Ginzburg [Hor23a]. GITR [YGSW21]. given [GO23]. gKdV [HWZ24]. GKZ [ABBD23]. glass [JRG21]. glasses [BPMMP24, VvBTH20]. Glauber [SW23]. GLE [Gro23]. Global [FA21, LW24, PLT^{+}23, ADdM20, KBH^{+}24, LOT^{+}20, LSS24, OGL^{+}21, RV20, RV21, YGSW21]. Globally [KP23]. GLoBES [TW21a, TW21a]. Godunov [MOV21]. goodness [ADdM20]. Gordon [GGCW24, GMZ^{+}20]. Gordon-type [GMZ^{+}20]. governed [JM24]. governing [WYZZ23]. GPEs [JLL^{+}24]. gpGPU [PNL^{+}24]. GPGPUs [CGSO20]. GPI-Space [BBH^{+}24]. GPIC [XHY^{+}24]. GPR [RBV^{+}22]. GPSD [KAG24]. GPSD-3D [KAG24]. GPU [AGH21, AUEO24, CCMR21, CCC20, CFLR24, CESD^{+}23, DMS^{+}22, DAC^{+}23, DS22, DFG^{+}23, DRZ^{+}21, DFU20, DHE^{+}24, EGLK20, EPM23, ECS23, GSL24, GB22, HPY21, HL21, HTL^{+}22, HLzY^{+}20, HG22, JS24, JMOC21, JĎD^{+}21, KKLZ23a, KWZ24, KL23b, KMR22, KT23, KNJ^{+}23, LG23, LJQ^{+}22, LHG24, LMHUR23, MWJL23, MVK^{+}24, MBC^{+}24, MHK24, NT24, Nie20, Nie22a, NRG24, OCK^{+}24, PSK^{+}24, RLW^{+}24, RCX24, RZH22, RBWD^{+}24, RCW^{+}20, RAJ^{+}24, SYFT23, SXYD24, SCL22, SBG23, TLC^{+}21, TMH^{+}23, VSM^{+}22, WWM^{+}22, WWJ^{+}22, WVK21, XBL^{+}20, XHY^{+}24, XLG^{+}23, YCCW23, Yan23, YKK^{+}23a, YLLW24, YLK21, zYMK^{+}21, ZMC23, ZPZH20, ZCP^{+}22, ZHI23, ZTV^{+}24, ZPL^{+}24]. GPU-accelerated [CFLR24, CESD^{+}23, ECS23, JMOC21, KL23b, KMR22, KT23, KNJ^{+}23, NT24, RLW^{+}24, RZH22, RCW^{+}20, SCL22, WWM^{+}22, WWJ^{+}22, WVK21, XBL^{+}20, XLG^{+}23, YLK21, ZHI23]. GPU-acceleration [JS24, zYMK^{+}21]. GPU-enabled [GSL24]. GPU-offload [SYFT23]. GPU-parallelized [KKLZ23a]. GPUs [CZ21, Ein20, HL21, MKHT20, NS20, ODR^{+}22, RBFB20, WA21, WS20, WZPW20, ZW20, ZDSS23]. Grad [AIZ23, ELSV24, HSB^{+}24, LZ21, SVSC20]. Grad-Shafranov [HSB^{+}24, LZ21]. Gradient [RdPS24, AUEO24, CL22a, GZW20, KL23b, WHSG24]. gradient-based [KL23b]. Gradient-enhanced [RdPS24]. grading [TIG^{+}24]. grain [XSM22]. grained [AHM^{+}23, MBCC23, OCK^{+}24, WWZ^{+}23, ZKZ^{+}24]. graining [NA20]. granular [EGKH24, RMR^{+}22, TLC^{+}21, YLLW24, ZZ21a]. graph [GF23, MMV^{+}24, Nog21a, RSC^{+}22, ZYML24]. graph-based [MMV^{+}24]. graph-theoretical [RSC^{+}22]. graphene [CSBF23, SRK^{+}24]. GraphIAST [DvHSdS22]. graphic [DC22, AGH21]. graphical [DvHSdS22, YR21]. GRASP2018 [SLE^{+}22]. Gravitational [CFPS23]. gravity [BLG24, MCMS24]. gray [FXQS21, JLW24]. Green [Ano20c, AUEO24, CDT22, GCWZ20, KK16, PPKK21, ZCS^{+}24]. Greenwood [Bul21]. grid [BC21, BW23, DFG^{+}23, Flo24, GUW^{+}22, MSM24, MSHP02, MSHP20, NDFL24, TCSA21]. grid-based [MSM24]. Gridap [VB22a]. grids [CAC^{+}22, DGM20, LH22, LH24, LJS21, PLF20, SRML20, SS21b, TS23, VB22b, YXX^{+}21]. GROMACS [RG21]. Gross [FJ22, BKRG22, CR20, CCW20, CWJ21, GTA21, KRG21, KMR22, KPL^{+}21, YSMBA23]. ground [CR20, GO23, KW23]. ground-state [KW23]. Group [LS21b, NL23b, WV22, CFBRE24, CPL21, DH22, Fon21, IPSW24, LCZ^{+}21, Sij23, WLL22, YS22]. GroupMath [Fon21]. groups [LZYY23]. growth [Dan24, WNS^{+}21]. gSeaGen [AAG^{+}20]. GSGPEs [CR20]. GSGPEs-v1.1 [CR20]. GSvit [KGN^{+}21]. GUI [KKB^{+}24, Soz21]. GUI4dft [Soz21]. guide [JDS20, KSIL22, WAYL23]. guided [ZHS^{+}20]. guiding [BKP22, BJS^{+}23, MHP23, XQ21]. Gutzwiller [PWD22]. gVOF [LH22]. GW [HM24, Kut20, SLZG20, LACL24]. Gyrofluid [Ken24]. Gyrokinetic [OGL^{+}21, YYC^{+}23, CJD^{+}20, JKSY22, KSJ^{+}22, KBH^{+}24, LOT^{+}20, MSU^{+}21, SBB^{+}24, YC20, ZDSS23, ZYX21, ZYX22].

H [AKK^{+}24, AYI^{+}24, KKM^{+}20]. H-COUP [AKK^{+}24, KKM^{+}20]. H-wave [AYI^{+}24]. hacks [Fis24]. HADOKEN [CW22]. Hadron [BDK^{+}23, For22, SZY^{+}22]. hadronic [KOT23]. half [LH24, TMH^{+}23]. half-spaces [LH24]. Hall [TWR21, VGGP^{+}21]. Hall-MHD [TWR21]. Halo [BHK^{+}21, FBA^{+}20]. Hamiltonian [GQ22, KF23, LHZ20, NKP20]. Hamiltonians [LQ21]. handle [KL23a]. Handling [WYHW24]. handyG [NSU20]. hard [SWZ23]. Hardware [WZK^{+}24, CCM20, EOR21, MRH^{+}23, TAY^{+}24, WWM^{+}22]. hardware-accelerated [MRH^{+}23]. Harmonic [WHSK24, LSS24, MSN^{+}22, PMSHG23, RZY^{+}24, SKEZ24a, SMGK21, dMMLOS20]. harmonic-oscillator [SMGK21]. harmonically [MPSK21]. harmonics [DEV20, Efr20, Efr21a]. Hartree [MSN^{+}22, RSM21, AYI^{+}24, CLS^{+}22a, WGG20]. HDMR [RBV^{+}22]. HEAPS [MMCC^{+}22]. heart [VSM^{+}22]. heat [AS24a, Bel24, CJZ21, KKPC23, LFL^{+}24, MBTB21, MMYU22, MM22, SKS24, XLL^{+}22]. heat-transport [MBTB21]. heaters [SS22]. heavy [AZH^{+}24, BKBL22, MBH21, MEH21, NUK^{+}22, OEI^{+}22, TIG^{+}24, YZL^{+}23, YZW22]. heavy-ion [MBH21, MEH21, YZL^{+}23]. heavy-quarkonium [OEI^{+}22]. HEJ [ABB^{+}22]. Hele [Yan21b]. heliosphere [MNS^{+}24, SBG23]. helium [FN23, LHTP^{+}24, VB22b]. helium-like [FN23, LHTP^{+}24]. Hellmann [NVC20]. Helmholtz [LGDF20, OCR^{+}22, OKBM23, VPPQ21]. help [SWZ23]. HemeLB [ZMC23]. HEP [KMN21, PP21, WZK^{+}24]. HEP-Frame [PP21]. HepLib [FXZT21, FTG23]. HEPLike [BC20]. HepMC3 [BIK^{+}21]. Hermes [DKM^{+}24]. Hermes-3 [DKM^{+}24]. Hermite [KMD^{+}21, PMK^{+}23]. Hermite-discontinuous [KMD^{+}21]. Hermitian [FKK^{+}21, zYMK^{+}21, YT22]. Heterogeneous [GB22, MDDI21, RCX24, Ryd20, BB24b, Hal21, LLZ^{+}23, LDGN24, LJS21, RSC^{+}22, SCF20, WWZ^{+}23]. heterostructures [Che23]. hfbtho [MSN^{+}22]. Hfszeeman [LGBJ20]. Hibridon [ADW^{+}23]. hidden [GSV23, SCT21, WYZZ23]. Hierarchical [BKS22, LDGN24]. Higgs [KM20, KMS20b, AKK^{+}24, BHH^{+}20, BMM21, DDM20, GHK20, KKM^{+}20, KKSV24, KMU^{+}23, KM20, KMS20b]. Higgs-boson [BHH^{+}20, DDM20]. HiggsBounds [BBH^{+}23]. HiggsSignals [BBH^{+}23]. HiggsTools [BBH^{+}23]. High [AALK20, ABB^{+}22, BPMMP24, BDR^{+}20, ČPF^{+}24, HSA22, KPL^{+}21, LMWW24a, LMQ^{+}23, NT24, PLL24, RBV^{+}22, RBFB20, XLL^{+}22, AGJ^{+}23, AFJ^{+}23, BBV^{+}22, ADF^{+}22, BCCM^{+}24, BL21, BTK24, BMSP21, BBA23, BCTS22, BLM^{+}22, BSC^{+}21, CZTF23, CKC^{+}21, CKGW22, CWJ21, Dan24, DBV^{+}24, DFU20, DS20, EGLK20, EVMP20, EPM23, EES24, FJ22, FXZT21, FTG23, FRN^{+}23, GBS^{+}20, GZW20, GB22, GHL^{+}22, GKT^{+}24, GAS^{+}23, GGCW24, GFD^{+}24, HMYH22, HA23, HWL^{+}23, HSO^{+}22, IUJ21, Jab20, KAB^{+}21, KMN21, KSC^{+}23, KGT22, LF24, LLZ^{+}22, Man24, MMCC^{+}22, MT23, MCB^{+}20, NJSY22, Nie22a, NAZ^{+}21, OCC23, OSK^{+}21, PMSHG23, PCS^{+}23, PFG22, QWZ^{+}21, RDH^{+}20, RZY^{+}24, RdPS24, RCW^{+}20, SRC21, SMC^{+}22, SGY24, SSV22, SDBS24, TCY23, Tan24, TV24, VLS22, WXY20, WXL^{+}21, WGGC22, XPF^{+}24, XHY^{+}24, YJLW21, YK21b, YKK23b, ZGZW23, ZDLS21, Zho23]. high- [AFJ^{+}23]. High-accuracy [LMQ^{+}23, GB22]. High-degree [LMWW24a]. high-dimensional [DS20, SMC^{+}22]. high-efficiency [Tan24, XHY^{+}24]. high-efficient [LLZ^{+}22]. High-energy [ABB^{+}22, BCCM^{+}24, Dan24, HMYH22, HSO^{+}22, Man24]. high-entropy [MMCC^{+}22]. High-fidelity [XLL^{+}22, BMSP21, CKC^{+}21, KSC^{+}23, MCB^{+}20, RdPS24]. high-intensity [HA23, Nie22a]. high-level [ZGZW23]. High-order [ČPF^{+}24, BBV^{+}22, ADF^{+}22, BBA23, BSC^{+}21, CWJ21, DBV^{+}24, DFU20, EVMP20, EPM23, FJ22, FRN^{+}23, GBS^{+}20, GZW20, GHL^{+}22, GKT^{+}24, GGCW24, HWL^{+}23, LF24, PMSHG23, PCS^{+}23, QWZ^{+}21, RZY^{+}24, RCW^{+}20, SDBS24, TCY23, VLS22, WXY20, YK21b, YKK23b, Zho23]. High-performance [AALK20, BPMMP24, BDR^{+}20, HSA22, KPL^{+}21, BLM^{+}22, CZTF23, MT23, OSK^{+}21, SGY24, SSV22, YKK23b]. high-precision [DS20, KAB^{+}21, KGT22]. high-pressure [BTK24]. high-resolution [EGLK20]. high-speed [BL21, CKGW22, DBV^{+}24, OCC23]. high-throughput [NAZ^{+}21, PFG22, WXL^{+}21, WGGC22]. Higher [LLY^{+}22, BA24, HTL^{+}22, LLZ^{+}22, SLL22, SP20, YS22]. higher-energy [SLL22]. Higher-order [LLY^{+}22, LLZ^{+}22, SP20]. highly [DRM20, LAD^{+}21, LDGN24, OLNG21, PCS^{+}20, TWR21, WYT23, YSX^{+}20, ZPL^{+}24]. highly-coupled [TWR21]. highly-efficient [ZPL^{+}24]. highly-scalable [OLNG21]. HighPT [AFJ^{+}23]. Hilbert [SKM^{+}21]. Hilliard [Yan21b, KSF^{+}22, LM21a, LHC20]. HiPACE [DBH^{+}22]. histogram [Sit23]. HL [TPS^{+}24]. HL-LHC [TPS^{+}24]. HLS [WZK^{+}24]. HODG [HWL^{+}23]. Hodge [VPPQ21]. Holm [ZYXC24]. HoloGen [CKGW22]. hologram [CKGW22, MZD^{+}20]. homogeneous [LHTP^{+}24, YZHL22]. homopolymer [LM21a]. HONPAS [SXW^{+}20]. HORSE3D [FRN^{+}23]. Householder [dBBVA20]. hp [MCB^{+}20, TMC22]. HPC [ARLDG24, GTA21, OLNG21]. hPIC [SMC^{+}22]. hPIC2 [MRH^{+}23]. HSMA [LYX22]. HTR [DFU20, DP21, Di 22]. HTR-1.2 [DP21]. HTR-1.3 [Di 22]. Hubbard [TMC22]. Hubble [TSAK21]. human [ZMC23]. human-scale [ZMC23]. hundred [CLZ^{+}21]. hundred-billion-metal-atom [CLZ^{+}21]. hunting [JDS20]. Hybrid [AAB^{+}21, AAMY23, CJ21, LJQ^{+}22, OBL^{+}21, YT23, ASC^{+}24, DMS^{+}22, GB22, GHK20, GP22, HPAW21, LL23, MRH^{+}23, RCS21, SKYQ21, SXW^{+}20, TWR21, WXX24, YW21, YD20, ZYML24, ZLS^{+}22]. Hybridizable [SVSC20]. hydraulic [ZL20]. hydrocarbons [MRG22]. hydrodynamic [CSBF23, DACA^{+}22, KD23, SLC^{+}22, WDMZ24]. hydrodynamically [CZY20]. hydrodynamically-coupled [CZY20]. Hydrodynamics [AMA^{+}20, AYWKL24, FRN21, KP23, LMQ^{+}23, LYZL24, MXH^{+}24, ODR^{+}22, RZH22, Sij23, ZCWY20, ZRZ^{+}21, ZHZG23]. hydrogen [SCT21]. hydrogenic [FN23]. hydrogeology [Org22]. hydrophilicity [SD24]. Hyper [GSM^{+}22, BA24, Gro22, GOS^{+}22, GG24]. Hyper-Fractal [GSM^{+}22, Gro22, GOS^{+}22, GG24]. hyper-radial [BA24]. Hyperbolic [TWW22, JKSY22, MD20b, PMS^{+}20a]. hyperfine [BGHC23, LGBJ20]. hyperfine-induced [LGBJ20]. hypergeometric [ABF^{+}23a, ABBD23, ABFP24, Ber24]. hyperparameter [TV24]. hyperplanes [LH24]. HyperRAF [BA24]. hypersonic [DFU20, GDJG23, Di 22, DP21]. hyperspectral [GG24]. hyperspherical [DEV20, Efr20, Efr21a]. hyperthermal [JS22]. HYPIC [WXX24].

IAST [DvHSdS22]. ICSFoam [OCC23]. Ideal [DvHSdS22]. identical [XSM22]. Identification [KSDH23, WLF^{+}24, WPMK21, CZB^{+}23, Ham20b, LNP^{+}24, MRG22, TAGC22b]. identifying [PRR23, WYZZ23]. II [ARLDG24, ATRD21, BBB^{+}21, BN23, CFPS23, NCF^{+}23]. III [HLCD20, Nog21b]. III-nitride [HLCD20]. IIM [WT22]. ill [SWZ23]. ill-posedness [SWZ23]. Illustration [NDFL24]. Image [OBK^{+}20, CEC^{+}24, KLD^{+}21, MCP23, OSK^{+}21]. image-based [OSK^{+}21]. images [GOS^{+}22, GG24, PSK^{+}24]. imaginary [KCS22]. imaging [ANU21, Yan24b]. IMEX [YWTK23]. immersed [BCHE21, TGBM22, YSX^{+}20, ZPZH20, Zha20]. immersed-boundary [YSX^{+}20]. immersed-finite-element [BCHE21]. immersive [TL20].

Impact [GAA^{+}20, FMCB^{+}20]. IMPC [ZYG21]. IMPC-Burnup2.0 [ZYG21]. IMPGYRO [YHH^{+}20]. imple [VMRFC23]. implement [MD20b]. Implementation [BN20, BN23, CJD^{+}20, EML22, KDK23, KGT22, KM22, KAS24, LEE^{+}21a, NG21, NBB^{+}21, OV23, YCCW23, YD20, ZRH21, CZTF23, DS22, Dür23, EGLK20, FWL^{+}24a, GBR23, JĎD^{+}21, KRE22, LYC20, LHG24, LNB23, MWJL23, MR22, Mar22, MVK^{+}24, MZL^{+}21, MHP23, MDP22, PWD22, PM21, PGS^{+}24, RG21, RMFB23, RAJ^{+}24, RCS21, SYFT23, SRE^{+}24, SNG20, SDBS24, VÁFG^{+}22, WZK^{+}24, XOTI22, Yan23, YML^{+}24, ZZZ^{+}20]. Implementations [YLIO22, HG22, MMFdL21, RBFB20]. implemented [DACA^{+}22, LYX22, SN23]. Implementing [DRB22]. Implicit [TWR21, BCHE21, BTK24, CL22b, CCW20, DTC20, DBV^{+}24, DBR24, EVMP20, KBB21, KBSL22, MAMK21, OCC23, PMK^{+}23, RDR^{+}20, RC23, SN23, SXYD24, TKS22, VLS22, YZ20b]. implosion [NUK^{+}22, TKC^{+}21]. imposition [MVF20]. improper [MBCC23]. Improved [BBC^{+}24, DACA^{+}22, HK24, Jab20, CML^{+}24, FFLR20, LY24a, MZ22, SXYD24]. Improvements [AAD^{+}24, CK23, EVFRHR23, KKL21, RBWD^{+}24, SMO20b, WMA^{+}22]. Improving [BGR23, KRJ23, PP21]. impulse [HMM22, OYC24]. impurities [YHH^{+}20]. impurity [ACS^{+}22, MSY^{+}21, NSY^{+}23, YHH^{+}20]. in/dependent [Kar23a]. including [BBB20, KOT23, KM20, KMS20b, MOV21]. inclusive [BKG^{+}23, ZDLS21]. incompressible [CS22a, CJZ21, Cos22, FBMD20, FRN21, HPY21, KBSL22, KRE22, LJH^{+}23, ODR^{+}22, TGS^{+}20, WT22, YWTK23]. incremental [KFHR24]. independent [ADW^{+}23, BV21, FMCB^{+}20, MSU^{+}21, WH23]. index [BMR^{+}24b, MWJL23]. induced [LGBJ20, OYC24, Roh22, SJY18, SJY20]. Industrialization [BCCM^{+}24]. inelastic [BKG^{+}23, GSV23, MBJB24, SMB20]. inequality [YYH21]. inertia [OS24]. inertial [LLH21a, NUK^{+}22, TKC^{+}21, ZLL^{+}24]. inexact [SLC^{+}22]. Inference [CV21, LM22, BB24a, HL21, KKP22]. infinite [RRC^{+}24]. influence [AS24a]. Informed [LY22, RRC^{+}24, SE24, WBF^{+}24]. infrastructure [CZA^{+}23, zYCD^{+}20]. INGRID [GUW^{+}22]. initial [AS24a, FYW23, SRML20]. initialization [NR21]. initialize [CCL^{+}22]. Initio [vRCM21, BDP^{+}21, JPJ^{+}23, MMC^{+}21, FN23, KSG22, MLD^{+}22, SLBR22, SLIC24, ZGZW23, IMB^{+}22, KMM21, LWC^{+}21, NYN^{+}21, PP23, RTRB21, YLH^{+}20, ZPL^{+}21, TDR^{+}20, YHY^{+}21]. injection [WAK22]. Inline [TW21b]. inorganic [ZHS^{+}20]. input [GHL23]. inputs [GBH20]. insight [RBV^{+}22]. insights [BMI23]. inspired [NDFL24, WHSG24]. instabilities [MD20b, YYC^{+}23]. instability [LY24a]. installed [XOTI22]. instruction [PNL^{+}24]. insulator [MCP23]. insulators [YI22]. Integral [KLMU21, SC20, CKT21, FN21, GOST23, MN21, SSV22, SZ24b, VPS23, VTB^{+}21, XFH^{+}22, YSI^{+}24, YZHL22]. Integrals [HY23, ABBD23, ABD^{+}23, Arn20, BA24, Bzo21, DS20, DEdM24, Flo24, Hid21, HG22, Lee21b, LM23, PBC^{+}24, UZB22, WW21, WBM^{+}24]. integrated [Yan24b]. Integrating [KMM21]. integration [Bar22b, BMT23, CL20, CCK23, CNS22, DBR24, EML22, GDB10, HPY21, HW22, KBSL22, KMR22, Koł23, LC24, MPQ^{+}22, PMK^{+}23, STA20, SLZ^{+}24, TSL21, TPCT22, WZPW20, WBM^{+}24, dVAR^{+}24]. integration-by-parts [WBM^{+}24]. Integrations [ZW20, CZ21]. Integrative [FYM^{+}22]. integrator [CL22b, EOR21, QWZ^{+}21, TBAR21]. integrator/WENO [EOR21]. integrators [TM24, ZK20]. Integro [ASW20]. Integro-Differential [ASW20]. Intel [OLNG21]. intelligence [PMSHG23]. intense [BAB^{+}20, Tan23]. intensity [Dan23, HA23, Nie22a]. intensive [BKS24a]. interacting [LJQ^{+}22, MBG^{+}20]. interaction [AAMY23, FBA^{+}20, GCK21, HHVB21, HLCD20, JMOC21, LBS^{+}23, MMP^{+}24, MSM24, Nie22a, RZY^{+}24, SNG20, UAS22, VSM^{+}22]. interaction-expansion [SNG20]. interactions [ASA^{+}22, AB21, BFI^{+}21, CGR21, CSW24, DG20, FN23, Gar21, HA23, KOT23, Kar23a, LMX^{+}21, MLD^{+}22, MRH^{+}23, NBS^{+}20, NBB^{+}21, PWD22, PFG22, Tan23, YKSH20, YGSW21, ZPL^{+}21]. interactive [GUW^{+}22, KKB^{+}24, TL20]. Interatomic [SRK^{+}24, DRB22, DBdFdSR21, HTS^{+}21, KMG^{+}20, MPN^{+}21, SKS24, WAET22]. Interface [AACE^{+}21, BMREC21, KMY^{+}23, AKW21, BLN^{+}21, DvHSdS22, HWAA22, LY24a, PFG22, TDR^{+}20, VMRFC23, WXL^{+}21, YWTK23, YHY^{+}21, ZDG^{+}21]. interfaces [DTGE21, GB20, MCP23, WT22, XSM22, YR21]. interfacial [SPTPR21, UO20]. intermediate [LGBJ20, MLD^{+}22]. intermediate-mass [MLD^{+}22]. Intermediates [RG21]. intermetallic [AMP^{+}21]. internal [BBA^{+}20, CZB^{+}23]. interoperability [BMREC21]. InterPhon [YHY^{+}21]. Interpolation [KKL21, YLL^{+}22, DS20, FMHH24, Ilt21, KOT23, KKLZ23a, KKLZ23b, KFC^{+}20, WVK21]. interpolation-based [KKLZ23a, KKLZ23b]. interpretable [TC24]. interpretation [BSS^{+}23, RMFB23, TRB20b]. interpreting [Hua24]. intersection [LH24]. interstitial [WLF^{+}24]. intrinsic [JGJ20]. Introducing [PSK^{+}24]. introduction [BM20, WOP^{+}20]. intuitive [YR21]. invariance [CLS22b]. invariant [ZYXC24]. invariant-preserving [ZYXC24]. invariants [GBR23, LS23]. Inverse [Sva24, AG23, HL21, TCD20, XLG^{+}23]. inversion [DRZ^{+}21, GST21, MN21]. inverted [CJZ21]. Investigating [ATC^{+}23]. Investigation [SCF20, MVK^{+}24, MZD^{+}20, WBvdH20]. InvFD [MN21]. involving [BA24]. ion [Ada22, BFT20, FYM^{+}22, GSV23, JS22, KLD^{+}21, LS21a, MBH21, MEH21, NUK^{+}22, PK24, PFG22, SNP^{+}20, SMC^{+}22, ST23b, TWR21, TBAR21, WXX24, Xie23, YZL^{+}23, KOF21]. ion-ion [GSV23]. ion-material [PFG22]. ionization [SZNW23, TNL^{+}22]. ions [BKBL22, SJP21a]. IR [KSG22]. IRAND [KRJ23]. irradiated [CHY^{+}24]. irradiations [MVK^{+}24]. irreducible [GWPW21, IMB^{+}22, LCZ^{+}21, MONW21]. irregular [TS23]. IrRep [IMB^{+}22]. Irvsp [GWPW21]. Ising [BPMMP24, HL21, OBL^{+}21, RBFB20, WZC^{+}24, YM21]. isoconversional [RHLTG^{+}22]. isomeric [WGS^{+}22]. isotherm [BRAC23]. isothermal [FA20b, LHG24, RHLTG^{+}22]. isotopes [KAB^{+}21]. ISPH [FRN21, MRN20, ZYML24]. Issue [GM24, SZY^{+}22]. Issues [LS21a]. ITER [COJ^{+}22]. iterated [WW21]. iteration [HPAW21]. iterations [Ano20c, KK16]. iterative [DCZ23, FWL^{+}24a, GQ22, KKY24, MRN20, SGS23, WLCF21]. ITVOLT [SGS23]. IV [MSN^{+}22].

Jacobian [LMMP23, Zho23]. Jacobian-free [LMMP23, Zho23]. January [Ano20r, Ano21q, Ano22q, Ano23q, Ano24m]. Jas4pp [CGG21]. JaSTA [Hal21]. JaSTA-3 [Hal21]. JAX [BBA23, XLG^{+}23]. JAX-FEM [XLG^{+}23]. JAX-Fluids [BBA23]. JefiGPU [ZCP^{+}22]. Jefimenko [ZCP^{+}22]. JeLLyFysh [HQF^{+}20]. jet [BP21, LNP^{+}24, LIG^{+}24, Roh22]. jet-fragmentation [Roh22]. jet-tagging [LNP^{+}24]. JFNK [HA21]. Jiezi [ZCS^{+}24]. JIT [VB22a]. joint [HFP21]. Jones [SHRK22]. Josephson [Gul20]. JRAF [Bag22]. Julia [Bag22, BA24, FASD20, VB22a, fWL24b]. July [Ano20s, Ano21r, Ano22r, Ano23r, Ano24n]. junction [BMI23]. June [Ano20t, Ano21s, Ano22s, Ano23s, Ano24o]. just [May21]. Jx [YKSH20].

Kalman [AGJ^{+}23, BDGS21]. KANTBP [CGV^{+}22]. Kawasaki [CZY20]. Kernel [STA20, BND22, LGDF20, TWW22]. kernels [EPM23]. KGMf [KFPV21]. KinBot [VZ20]. kinematical [Dan23]. Kinetic [KWZ24, BB21, CLEP24, DIKSN24, FBC^{+}21, HEF21, KDIN^{+}23, LSW^{+}20, LHG24, LWV20, MAMK21, RSC^{+}22, SKC21, ZWP^{+}22, ZYX21, ZYX22]. Kinetics [CV21, CLEPF23, FCTFR20, MPH^{+}24, RHLTG^{+}22, RCX24, DdCAG23]. kink [ZJM^{+}21]. Kira [KLMU21]. kit [WGGC22, MAMK21, CZWE23]. KKMC [AJW^{+}21]. KKMCee [JWW^{+}23]. Klein [GGCW24]. KLIFF [WAET22]. KM3NeT [AAA^{+}20, AAG^{+}20, AAG^{+}24]. kMap.py [BYL^{+}21]. knockout [OYC24, SW23]. Kohn [JZW^{+}22, LG23, YT23]. Kokkos [EVMP20]. Krook [FJ22]. Krylov [HKY^{+}21, WPMK21, Zho23]. KSSOLV [JZW^{+}22]. Kubo [Bul21, SP23, WLCF21]. Kutta [KBSL22, KFC^{+}20].

lab [DHK^{+}21, DG20]. LaBCof [NL23a]. Lac [BB24b]. laden [GWA^{+}23, GB20]. Lagrange [GWA^{+}23, SFBG20, WVK21]. Lagrange-mesh [SFBG20]. Lagrangian [CS22b, Ein20, MSM24, SRE^{+}24, Sij23, THH21, TC24, WWM^{+}22]. lagrangians [May21, May21]. Laguerre [HY23, PCL24]. lamellar [FLK^{+}20]. Laminar [VLS22]. Laminar-Turbulent [VLS22]. LAMMPS [CT23, CL22b, DCRF23, DACA^{+}22, Ham20b, JCM20, KAS24, LYX22, TAB^{+}22, YLIO22]. Lanczos [GBD^{+}22, ZJ23]. Landau [AM21b, CFW^{+}23, Hor23a, MDP22, SJ21, ZSqXY21]. landscape [LSZ23]. landscapes [NA20]. Langevin [JGJ20, LLT^{+}23, TM24]. Langtry [VLS22]. Laplace [ATC^{+}23, CLEPF23, WZC21]. Laplacian [CZS^{+}21]. lapse [DG20]. Large [CJ21, WA21, YCC22, YJLW21, ARLDG24, BC22, BDR^{+}20, CB23, CMJC21, EVFRHR23, FWL^{+}24a, GOST23, HSA22, KSG22, KKPC23, KFHR24, KLM^{+}22, LBRW22, LKP21, LZK^{+}23, LG23, LGK^{+}22, Man24, MDR^{+}20, OCK^{+}24, PLT^{+}23, PBC^{+}24, RL21, SKM^{+}21, VTB^{+}21, ZJ23, ZW20]. large-dimension [ZJ23]. Large-scale [WA21, YJLW21, BC22, BDR^{+}20, CB23, EVFRHR23, FWL^{+}24a, HSA22, KKPC23, KFHR24, LBRW22, LKP21, LZK^{+}23, LG23, MDR^{+}20, PLT^{+}23, RL21]. laser [BAB^{+}20, EBNS22, FN23, HA23, LLH21a, LMX^{+}21, MXH^{+}24, Nie22a, PMSHG23, RZY^{+}24, SJY18, SJY20, Tan23]. laser-electron [Nie22a]. laser-induced [SJY18, SJY20]. laser-plasma [HA23]. laser-produced [MXH^{+}24]. lasers [ZLMH23]. latex [RDZ^{+}20]. Lattice [BBB20, DBdFdSR21, KBSH20, MD21, NL23a, PLSB22, YD20, vRCM21, AHP21, AS24b, CL22b, DACA^{+}22, Dür23, EVFRHR23, FFTV23, FFLR20, FKK^{+}21, GLSH21, IKM^{+}23, JMOC21, Kub24, KAS24, LMWW24b, LLL24, MWJL23, MBC^{+}24, MOY^{+}22, QCZ23, RZ23, STA20, SKM^{+}21, WA21, WSRO24, YCCW23, ZPZH20, ZPL^{+}24, ASC^{+}21, KK20]. lattice-Boltzmann [DACA^{+}22]. lattices [BRAC23]. LAVA [DCRF23]. laws [HWAA22, OCR^{+}22, Yev21]. layer [CML^{+}24, MPH^{+}24, OKBM23, MAMK21]. layered [WZC21]. layers [NTO24, Pöt20]. lb [DACA^{+}22]. lb/fluid [DACA^{+}22]. LBcuda [BLM^{+}22]. LBfoam [ASC^{+}21]. LBsoft [BMT^{+}20, BLM^{+}22]. LCLS [NCF^{+}23]. LCLS-II [NCF^{+}23]. LCPAO [YI22]. leading [ABB^{+}22, PBC^{+}24]. leapfrog [TKS22]. learned [MYKC23, RBV^{+}22]. Learning [GUAD22, TGGC23, VEHCM21, ZKZ^{+}24, AYB24, AUO^{+}22, CFL^{+}22, CWG^{+}21, CZWE23, HPRS23, HJM^{+}20, KRJ23, KKY24, LAC21, LCL^{+}23, LDGN24, MLZ^{+}23, MGG^{+}20, MPN^{+}21, SE24, SKS24, ST23b, TV24, TAGC22b, TSAK21, TGIM23, UD24, WAET22, YYZ^{+}22, ZWC^{+}20, ZGW20]. learning-based [CZWE23, UD24]. Lebedev [LLH^{+}21b]. Leffler [CLEP24]. left [VSM^{+}22]. Legolas [CK23]. Lehmann [KCS22]. Lennard [SHRK22]. Lennard-Jones [SHRK22]. LEP [AJW^{+}21]. LEP/SLC [AJW^{+}21]. LeptogeneSiS [GMPG^{+}21, GLPG^{+}23]. Lepton [AJW^{+}21, AZH^{+}24, AAD^{+}24, JWW^{+}23, KKSV24, LNP^{+}24]. LeptonInjector [AAA^{+}21]. leptons [SLP^{+}22]. LeptonWeighter [AAA^{+}21]. leptoquarks [CS22b]. LES/DNS [LBS^{+}23]. less [AMA^{+}20, BR20a, Pos22]. Level [YZZ^{+}23, AHM^{+}23, EVFRHR23, GSBN22, HWAA22, Hua24, KSG22, NT24, TGIM23, ZGZW23]. level-3 [TGIM23]. Level-set [YZZ^{+}23, HWAA22]. levels [CGV^{+}22]. LEVIS [MPQ^{+}22]. LHC [AES21, TPS^{+}24]. LHCb [AAB^{+}21, BDGS21]. LIBAMI [EML22]. libdlr [KCS22]. libEMM [Yan23]. libFastMesh [FZD^{+}24]. Libra [Lee21b]. Library [HJM^{+}20, ART^{+}20, ABK^{+}22, AS24b, BM20, BB24a, BFI^{+}21, BIK^{+}21, CK20a, CCL^{+}22, DH22, DLD^{+}21, FXZT21, FTG23, FPSZ21, GSBN22, HTH^{+}20, HKY^{+}21, KRC^{+}20, KKPC21, MEDT^{+}23, MRT^{+}22, MT23, NJSY22, NBCMH20, Nie22b, OCC23, Par21, PPKK21, RDC^{+}20, SVJ^{+}24, SLDF^{+}21, SZ24b, STRF^{+}20, SJWL22, WAN^{+}22, Yan23, YKK^{+}23a, YMCF23, ZRZ^{+}21].

Lie [FKS20]. LieART [FKS20]. lifetime [LU21]. lifetimes [GTMB21]. lifex [AFB^{+}24a]. lifex-cfd [AFB^{+}24a]. ligand [Ryd20]. Light [Hal21, LSNRD20, BKO20, KOT23, MR22, MS24, MLD^{+}22, YYZ^{+}22]. light- [MLD^{+}22]. lightweight [RGS^{+}21]. like [FN23, LHTP^{+}24, Scr22a]. likelihood [BC20]. likelihoods [AKW21]. LIMAO [SJY20, SJY18]. limiters [VÁFG^{+}22]. Lindblad [OEI^{+}22]. line [Kut24, YJ24]. Linear [HW22, KCJX24, ZJM^{+}21, AGMFGE23, AIZ23, BB21, DNG^{+}20, FLW^{+}23, GPM^{+}21, HA23, HMSV23, LBS^{+}23, MG22, Man24, NBC^{+}24, PAL^{+}20, PBC^{+}24, RVRT22, RAJ^{+}24, SLB^{+}23, SLZY21, SJY18, SJY20, Yan21b, Yan24a, Zeb22]. linear- [SJY18, SJY20]. Linear-scale [KCJX24]. linearization [FYW23]. linearization-based [FYW23]. linearized [GMZ^{+}20, HM24, KDK23, PPR^{+}21]. linearly [CCW20]. linearly-implicit [CCW20]. lines [MM23]. LIon [BFT20]. Liouville [GBD^{+}22]. Liouville-Lanczos [GBD^{+}22]. Liouvillian [DEdM24]. liquid [BCF^{+}24, BMI23, RSD20, SRT^{+}20, UO20, XZLX20, dSOZ22]. liquids [MHÅ21, VCF22]. LisbOn [DdCAG23]. LISE [JRS^{+}21]. list [Mar22]. lithium [WXW^{+}21]. lived [ADH^{+}20]. load [LWS^{+}23, MLT^{+}21, TPK^{+}21, ZWZ^{+}22, ZHZG23]. load-decoupling [ZWZ^{+}22]. loads [KD23]. Local [IPSW24, LY24a, DCZ23, JLL^{+}24, JRS^{+}21, LTMK21, LW24, NVC20, PWD22, SKM^{+}21]. localised [MSHP20]. locality [VHBK21]. localized [CF21, HLzY^{+}20, Kor23, MSHP02]. locally [OKBM23, SBZ23, ZPL^{+}24]. locally-aligned [SBZ23]. locally-conformal [OKBM23]. logarithms [ABB^{+}22]. logderivatives [BTW20]. logic [AAG^{+}24, KMN21]. Long [ADH^{+}20, HG22, MP21]. Long-lived [ADH^{+}20]. long-range [HG22]. long-term [MP21]. loop [ADC^{+}21, AKK^{+}24, FQRR22, HQRR20, KKM^{+}20, KM20, KMS20b, RUV20]. loops [LMWW24a]. Lorentz [FH22]. Lorenzetti [ABF^{+}23b]. Low [Kub24, VPS23, CE22, CL22a, CJ21, Gar21, JS22, JĎD^{+}21, Kul20, LHG24, MD21, NYN^{+}21, VV21, ZLLM23]. low-density [ZLLM23]. low-dissipation [Kul20]. low-energy [Gar21, NYN^{+}21]. Low-mode [Kub24]. low-order [CJ21]. low-pressure [JĎD^{+}21]. Low-rank [VPS23, CE22]. low-speed [LHG24]. low-temperature [VV21]. lower [RBV^{+}22]. lower-dimensional [RBV^{+}22]. lowest [BKG^{+}23]. lowest-order [BKG^{+}23]. lp [FLW^{+}23]. lubrication [NLS24, SLC^{+}22]. luminescence [TR22]. luminosity [Ada22]. LWR [NL23b].

M [Xav22, LM22]. M-CHIPR [Xav22]. m-NLP [LM22]. M1 [LMMP23]. M3D [ZJM^{+}21]. M3D-C1 [ZJM^{+}21]. MAAT [AMP^{+}21]. Mach [MA24]. Machine [CWG^{+}21, TGGC23, VEHCM21, AYB24, AUO^{+}22, CK20b, CZWE23, HPRS23, HJM^{+}20, KRJ23, KHR^{+}23, LAC21, LCL^{+}23, MGG^{+}20, MYKC23, MPN^{+}21, RBV^{+}22, SE24, SKS24, ST23b, TV24, TAGC22b, UD24, WAET22, YM21]. machine-learned [MYKC23, RBV^{+}22]. machine-learning [LCL^{+}23, MGG^{+}20, MPN^{+}21, SKS24]. machines [WWDM20]. macromolecular [KDHL23]. macroscopic [PMSHG23]. made [AMK24, GFH23]. MADHAT [BHK^{+}21]. MAELAS [NAZ^{+}22, NAZ^{+}21]. MAFIA [MRG22]. MAFIA-MD [MRG22]. MagGene [ZZ21b]. magnet [HK24]. magnetic [AIZ23, BTK24, BM22, BKO20, CNB^{+}23, DKRSR22, GO23, GBS^{+}20, HHVB21, LL23, LGBJ20, LZYY23, LU21, LHTP^{+}24, MZD^{+}20, NRG22, NRG24, PTD20, RMR^{+}22, Yan21a, YYC^{+}23, YC20, YKSH20, YLL^{+}22, ZYLY22, ZYL^{+}23, ZZ21b]. magnetic-field- [LGBJ20]. MagneticKP [ZYL^{+}23]. MagneticTB [ZYLY22]. magnetism [GOCSS^{+}23]. magnetized [KOF21, MMP20, sXBkB^{+}22, ZSqXY21]. Magneto [VGGP^{+}21, DIAA21, NAZ^{+}22, WDMZ24, NAZ^{+}21]. magneto-elastic [NAZ^{+}22, NAZ^{+}21]. magneto-hydrodynamic [WDMZ24]. magneto-optical [DIAA21]. Magneto-thermal [VGGP^{+}21]. magnetohydrodynamic [MGX24]. magnetohydrodynamics [LJQ^{+}22, LMMP23, SMA24]. magnetostatics [ALB22]. maintainability [LTT^{+}24]. MAISE [HTS^{+}21]. major [WMA^{+}22]. Majorana [WPMK21]. Mammography [BSG^{+}21]. MAN [WAYL23]. management [CCC20, LYZL24]. manifold [LJQ^{+}22]. manifolds [CB23, XQ21]. manipulation [Ols23]. manufacturing [SYFT23]. Many [GBJ^{+}21, SDL^{+}21, AJDS^{+}21, ATRD21, DC22, MAJ20, NA20, OGL^{+}21, PPKK21, RMM21, WS20, Xav22, XBL^{+}20]. many- [OGL^{+}21]. Many-Body [GBJ^{+}21, ATRD21, DC22, MAJ20, PPKK21, RMM21, XBL^{+}20]. Many-core [SDL^{+}21]. many-dimensional [NA20]. many-electron [AJDS^{+}21]. manybody [CT23]. manycore [RL21]. map [SRML20]. Maple [JWK^{+}21, PMS20b]. mapped [MPSK21]. mapping [SFC20, Zeb22]. March [Ano20u, Ano21t, Ano22t, Ano23t, Ano24p]. Markov [HSMR^{+}24, MMM23]. MARLEY [Gar21]. Marlics [dSOZ22]. MARTY [UMA21]. MASAP [KRL^{+}24]. mass [KAB^{+}21, Kub24, LM23, MP21, MLD^{+}22, MDDI21, PC21, RTRB21]. masses [ABD^{+}23, Efr23, HLM22]. Massive [SRML20, KKPC21, LYZL24, YKK^{+}23a]. Massively [FWL^{+}24a, KHR^{+}23, AS24b, BBH^{+}24, CHY^{+}24, Cos22, DMS^{+}22, KMM21, KSC^{+}23, MDR^{+}20, RYS^{+}24, ZAW^{+}21]. Massively-parallel [KHR^{+}23]. massless [RUV20]. Master [FFLR20, ZXT22, ZXT23, BB24b]. Master-field [FFLR20]. matched [NTO24, OKBM23, Pöt20]. matching [HLM22]. material [ASPDL^{+}21, DCRF23, MRH^{+}23, MVF20, MYKC23, MYM^{+}22, NYN^{+}21, PP23, PFG22, Sij23, TRB20b, YGSW21]. Materials [AMP^{+}21, SLL^{+}21, AMK24, CL20, CHY^{+}24, EGKH24, EL24, Eku24, GLN23, HTL^{+}22, HJM^{+}20, KKB^{+}24, LDGN24, ONH^{+}20, PB23, RMCC22, SA23, SLDF^{+}21, TAB^{+}22, TLC^{+}21, TQGE23, UÁEPGBP24, VMRFC23, WWB22, WLF^{+}24, YNV22, YNV23, YTC^{+}20, YLH^{+}20, ZYLY22, ZGZW23, ZWC22]. Mathematica [ABBD23, ABFP24, Ber24, Bzo21, FKS20, Fon21, Hid21, HCP20, LM23, PSW23, Tor20, Xav22]. mathematical [TRB20b]. Mathematics [CB23]. MATLAB [CR20, JZW^{+}22, YJ24, dMMLOS20, KMBP24]. matrices [CGV^{+}22, Fis24, HT24, MMM23, MRT^{+}22]. Matrix [WBvdH20, BGH22, BFCR24, BAB^{+}20, CPL21, CCC20, FMCB^{+}20, KIK20, Kut24, MBG^{+}20, QJ21, RZ23, RMFB23, SLZY21, SKM^{+}21, XFGS24]. matrix-product-state [SKM^{+}21]. Matsubara [EML22]. matter [ABB^{+}24, BRHT21, Jab22b, Jab24, LSZ23, May21, MHK24, NBS^{+}20, Pag24, PAM24, Sch21, UAS22, BCD^{+}21, GUAD22, OBC^{+}24]. MAX [FLW^{+}23]. MAXIM [YR21]. maximally [Kor23]. maximally-localized [Kor23]. maximization [AS24a]. maximize [MRT^{+}22]. Maxwell [DGM20, GLB^{+}21, KMD^{+}21, LW24, LLH^{+}21b, MOV21, PMK^{+}23, QWZ^{+}21, RJ21, YZW21, YKY^{+}22, ZLL^{+}24]. May [Ano20v, Ano21u, Ano22u, Ano23u, Ano24q]. maze [Ryd20]. mbsolve [RJ21]. MCBTE [PPR^{+}21]. MCC [WXX24]. MCGPU [HLMB24]. MCGPU-PET [HLMB24]. MCMC [HL21]. MCMC-based [HL21]. MCNNTUNES [LAC21]. MCNOX [Hua24]. mcrtFOAM [FXQS21]. MCS [NL23b, YLK^{+}20]. MD [CLZ^{+}21, MRG22]. MD2D [LN23]. mdapy [WS23]. Mead} [PP23]. Mean [SHW^{+}21, BN20, BN23]. mean-field [BN20, BN23]. means [Kul20, SJ20, UZB22]. measure [DRSZ23]. measurement [HTN21, MPSK21, WBF^{+}24]. measurements [KAB^{+}21, MYM^{+}22]. measuring [HPP23]. mechanical [EL24, LZP^{+}24, Ols23, SLDF^{+}21]. mechanically [KvdW20]. Mechanics [Sha21, LYZL24]. Mechanics-oriented [Sha21]. mechanism [Kar22, LLY^{+}22]. mechanisms [DIKSN24, FBC^{+}21, KDIN^{+}23, RAJ^{+}24]. mechanistic [XLG^{+}23]. MechElastic [SLDF^{+}21]. media [BVV22, Hoh24, LLH^{+}21b, MS24, OSK^{+}21, SFS22, WZC21, XFH^{+}22, YCCW23, YLLW24, ŻTR^{+}22]. Medical [GSM^{+}22, GOS^{+}22]. medium [DRZ^{+}21, Roh22, Tan23]. Meek [MHÅ21]. MELASA [FLK^{+}20]. melt [CZY20]. membrane [ZBS^{+}23]. Memory [MD21, EVMP20, LYZL24, MLT^{+}21, ZWZ^{+}22]. Memory-efficient [MD21]. Menter [VLS22]. MercuryDPM [WOP^{+}20, OAP^{+}24]. merging [BP21, DWD^{+}24, LXY^{+}21]. Merlin [ABK^{+}22]. mesh [AMA^{+}20, ASC^{+}24, BBA^{+}20, FXQS21, FAL20, FAL21, GKT^{+}24, KKY24, LW24, MM23, NSY^{+}23, PG23, SN23, SFBG20, SJP^{+}21b, WDMZ24, ZDSS23, Zho23]. mesh-adaptive [SJP^{+}21b]. mesh-agglomeration [FXQS21]. mesh-free [MM23]. mesh-less [AMA^{+}20]. mesh-refined [LW24]. meshes [CS22a, KSF^{+}22, PGYF21, PD23, RVRT22, TGBM22]. meshing [RSHS24]. meshless [JLL^{+}24]. meso [TAB^{+}22]. meson [KL23a]. mesoscale [AALK20]. metal [CXCZ23, CLZ^{+}21, SRK^{+}24]. metals [CWG^{+}21, HM24, MOV21]. metamaterials [LW24]. metastructures [HSA22]. metasurface [KL23b, YWM23]. Metasurfaces [YR21]. Metasurfaces-oriented [YR21]. Method [MD21, PLT^{+}23, RLW^{+}24, STA20, ZTV^{+}24, AUEO24, AMA^{+}20, ATC^{+}23, BA24, BCHE21, BN20, BND22, BN23, BDdM20, ČPF^{+}24, CHA21, CJZ21, CHY^{+}24, CL21, CB20, CF21, CLEPF23, CCW20, DZZ21, DCZ23, FJ22, FLW^{+}23, FBMD20, FYM^{+}22, GTE21, GLSH21, GRCT20, GB20, GGCW24, HPY21, HPAW21, HTL^{+}22, HFP21, HKY^{+}21, HA21, HJGL22, HGS20, JKSY22, JĎD^{+}21, KKLZ23a, KBB21, KSJ^{+}22, KR23, KL23b, KFC^{+}20, KMD^{+}21, KYH24, Kut20, LEE^{+}21a, LBM^{+}23, LKP21, LKK23, LXY^{+}21, LKP24, LF24, LLL24, LLZ^{+}22, LJH^{+}23, LS22, LDGN24, LMHUR23, LVMGF^{+}23, LY24b, LNB23, MWJL23, MSH22, MBG^{+}20, MZL^{+}21, MSM24, MVAXP22, MOV21, MCP23, MRD23, MNS^{+}24, NG21, Och23a, Och23b, ONH^{+}20, OKBM23, PB23, PWD22, QCZ23, RBV^{+}22, RC23, RV20, RV21, RCW^{+}20, RVRT22, RCS21, SFBG20, SJP^{+}21b, SLZY21, SXYD24, SNG20, SLC^{+}22, SCMP^{+}22, SCR^{+}22b, SAC^{+}21]. method [SS21b, TCD20, TTM22, TS23, TLC^{+}21, TPCT22, eSdSBST21, WZZS21, WZC21, WWM^{+}22, WA21, WLCF21, WLL22, WXZH24, WPMK21, WS20, WVK21, WNS^{+}21, WYZZ23, XDF20, XSL^{+}22, XFH^{+}22, XPF^{+}24, XLKX21, YT23, YS22, YWM23, YSX^{+}20, YC20, YXX^{+}21, YZHL22, YKK23b, YJ24, ZPZH20, Zha20, ZPL^{+}24, ZYML24, ZZ21a, ZYX21, ZYX22, Zho23, ZHM21, ZLS^{+}22, ASC^{+}21, GSL24, LMHUR23]. method-based [KR23]. Methodology [SCMP^{+}22, WBF^{+}24, WZK^{+}24, CML^{+}24, SFKC22]. Methods [LSNRD20, AYB24, BBV^{+}22, BCGT24, CL20, CJ21, DGM20, ELSV24, FYW23, GCWZ20, HAM^{+}20a, HWL^{+}23, HWZ24, IKM^{+}24, JMOC21, KAB^{+}21, KSF^{+}22, KLMU21, KFC^{+}20, KvH21, LHG^{+}20, LH22, LH24, MGX24, MCB^{+}20, NCF^{+}23, PMK^{+}23, PLL24, PAL^{+}20, PM21, SL22, SLZ^{+}24, SKM^{+}21, SDBS24, UZB22, VPS23, YLIO22, YZW21, YWTK23, YK21b, YZZ^{+}23, ZYXC24]. Metric [BTW20]. metrology [CDD22, HTN21]. Metropolis [SJ20]. MeV [Pos22]. MFC [BSC^{+}21]. MFDFA [GHKW22]. MGOS [KRC^{+}20]. MHD [CK23, FMBD22, MG22, SDXY23, TKS22, TWR21]. MHFEM [SFS22]. Mi3 [HL21]. Mi3-GPU [HL21]. Micro [LLY^{+}22]. microfluidics [AALK20]. micromagnetic [RMR^{+}22, SE24]. micromagnetics [PRS^{+}20]. micrOMEGAs [ABB^{+}24]. micron [LU21]. microparticles [DHK^{+}21, KT23]. microrheology [DG20]. microscale [KM23, MEC^{+}24]. microscope [KLD^{+}21]. Microscopic [Gul20, YYC^{+}23]. microscopy [AUO^{+}22, DG20, HGS20]. microstructures [EGLK20]. Migration [Gro22]. million [LWC^{+}21, RZ23]. MiMeS [Kar22]. Minimal [AAT^{+}20, AKK^{+}24, AAT^{+}14, KKM^{+}20, LXY^{+}21, May21, SSH^{+}23, STRF^{+}20, KM20]. minimal-lagrangians [May21]. minimisation [PGS^{+}24]. minimization [IJVJ21, IUJ21, SE24, SAS20, SAN24]. minimum [NA20]. Minkowski [BMT23]. Minnesota [ZXT22, ZXT23]. minor [EGKH24]. Mirheo [AALK20]. Misalignment [Kar22]. missions [GSL^{+}23]. MISTER [COW24]. MISTER-T [COW24]. mitigate [MD20a]. MiTMoJCo [Gul20]. MITNS [KOF21]. Mittag [CLEP24]. Mittag-Leffler [CLEP24]. mixed [LHC20]. mixing [Nog21a, NVCS23]. mixtures [LM21a, LB24, SLK23, ZLLM23]. ML [LNP^{+}24]. MLAnalysis [GFD^{+}24]. MLQD [UD24]. mobile [LLC^{+}23]. mobility [MMC^{+}21]. MOC [LLZ^{+}23, WLL22]. MOCOKI [BB21]. Modal [WAYL23, KL23b]. mode [BDK^{+}23, Kub24, ZJM^{+}21]. Model [BHK^{+}21, DPR^{+}20, DRR^{+}24, Gul20, RBV^{+}22, ASU^{+}21, AFJ^{+}23, AHM^{+}23, BCHE21, BGW^{+}22, BB21, BTK24, BCTS22, CFLR24, CL20, CZY20, CKT21, Dau23, DRB22, ET24, FJ22, FBA^{+}20, HLCD20, JLW24, JPJ^{+}23, KKSV24, KLD^{+}21, KDK23, KM23, KW23, KAS24, LL23, LLH21a, LY24a, LGK^{+}22, MMC^{+}21, May21, MBCC23, NYN^{+}21, NLS24, NRG24, OBL^{+}21, RZY^{+}24, RBFB20, SYFT23, SKYQ21, SZY^{+}22, SZ23, SOH^{+}23, SJWL22, SW23, TKS22, TCY^{+}21, TCY23, TPK^{+}21, TPCT22, TNL^{+}22, VLS22, VEHCM21, WZZ^{+}23, WYT23, WZC^{+}24, YZL^{+}23, YYH21, YK21a, YX24, YML^{+}24, YWX^{+}23, ZYLY22, ZGW20, ZSqXY21, Bie21, DDM20, HQRR20, KM20, KMS20b, MAJ20, UMA21]. Model-Agnostic [BHK^{+}21]. ModelFLOWs [HCAH^{+}24]. ModelFLOWs-app [HCAH^{+}24]. Modeling [BB24b, RSHS24, Sha21, YD20, AHP21, AG21, ART^{+}20, BFL^{+}22, CFW^{+}23, CML^{+}24, EGLK20, FA20a, FLK^{+}20, GUW^{+}22, HSB^{+}24, HyLF23, HRG^{+}22, JYL^{+}23, KRE22, KNJ^{+}23, LMX^{+}21, LZP^{+}24, LSF23, LFL^{+}24, MK22, MHK24, MMV^{+}24, MYKC23, MT23, MOMO24, NRKA22, NI22, NLS24, PMA^{+}21, PG23, RCX24, RRC^{+}24, SFKC22, SMC^{+}22, TAB^{+}22, VB22b, WLZ^{+}24, Yev21, YMCF23, ZGZW23, ZKZ^{+}24, ZHZ^{+}23]. modelling [BVV22, DRM20, HCAH^{+}24, LS21a, MBG^{+}20, MVAXP22, Org22, OXOG23, TDR^{+}20, TRB20b, WNS^{+}21, WSK^{+}22, Yan23, Yan24b]. Models [GFH23, RMR^{+}22, AKK^{+}24, AES21, BRAC23, CZWE23, CSBF23, EVMP20, FSMM24, FA21, GZW20, GHK20, HTS^{+}21, HSMR^{+}24, Jab24, KPR^{+}24, KC21, KVSC21, LLC^{+}23, LSS24, Mau20, MD22, MPH^{+}24, MDDI21, NRK^{+}21, OSE^{+}20, PCL24, PBC^{+}24, RDZ^{+}20, SZ24a, SBG23, SKM^{+}21, YWTK23, YM21, YKYK23, ZWC^{+}20, ZYL^{+}23, Zho23]. moderate [GOST23]. moderate/large [GOST23]. modern [EOR21, FFTV23, HAM^{+}20a, UMA21]. modernization [VXT^{+}23]. modes [HRU22, JGJ20, WPMK21]. modification [BGW^{+}22]. Modified [FA20a, BCGT24, HLCD20, YK21a]. modular [GRCT20, MOMO24, PAM24, SC20, WKJB23]. modulation [MNS^{+}24, SBG23]. module [BGHC23, GK21, GBJ^{+}21, GSM^{+}22, GOS^{+}22, HYL^{+}22, LN23, MKPW21, MKPW22, MEH21, WHSG24]. modules [BDR^{+}20]. Moiré [NNMJ22]. Molecular [LWC^{+}21, MKPW21, MKPW22, MRG22, SOH^{+}23, AHM^{+}23, Bag22, BMREC21, BCTS22, CMJC21, CGR21, CDT22, DTGE21, FGCN^{+}21, GDS^{+}21, Ham20b, HyLF23, KHR^{+}23, KRC^{+}20, KMM21, KLM^{+}22, LN23, LZP^{+}24, LVMGF^{+}23, MES^{+}24, MHK24, MBCC23, NBC^{+}24, Ols23, PCS^{+}20, RZY^{+}24, SMO^{+}20a, SKC21, SC22, SSD^{+}22, SZNW23, TCSD24, TRB^{+}20a, VCF22, VXT^{+}23, VvBTH20, Wor20, WS23, ZHI23, ZJS^{+}20, YD20]. molecule [PRR23, PNL^{+}24, SHRK22, Xav22]. molecules [BGHC23, BFMA^{+}24, BAB^{+}20, Hua24, KSG22, MBJB24, MBG^{+}20, NI22, RZY^{+}24, RV20, RV21, SJP21a, SMB20, TRN^{+}24, ZLMH23]. MolSSI [BMREC21]. MolTwister [Ols23]. Moment [SJP^{+}21b, CFW^{+}23, HFP21, PLF20]. Moment-preserving [SJP^{+}21b]. moments [Ume22, LMHUR23]. momentum [CCW20, CWJ21, JDS20, KPST21, Pöt20, ST23a, WJB21]. monitor [Ada22]. monodisperse [KAG24]. monokinetic [LMHL20]. Monte [Kal20, ABGD^{+}20, ASPDL^{+}21, ACD^{+}22, AJW^{+}21, BRAC23, BB21, BCF^{+}24, BKO20, BKS24b, BIK^{+}21, BKBL22, BKG^{+}23, CCM20, DRSZ23, DDM20, DdCAG23, ELSV24, FXQS21, GJA21, GAGO21, GHK20, HT24, HLMB24, HQF^{+}20, Hua23, HSMR^{+}24, Ilt21, JWW^{+}23, JKKN23, KOT23, Koł22, Koł23, LEE^{+}24, LMWW24b, LAC21, LTMK21, LLL24, LC24, LHWX24, MMM23, MT23, NL23b, Nie22a, OBL^{+}21, PAZ^{+}22, PLT^{+}23, PYT^{+}24, PPR^{+}21, PC21, RSC^{+}22, RCB21, Roh22, SNP^{+}20, SJP^{+}21b, SHS22, SNG20, SLIC24, TAY^{+}24, WAN^{+}22, WSRO24, WZPW20, YLK^{+}20, ZHS^{+}20, ZDLS21]. Monte-Carlo [BKS24b, BKG^{+}23, GAGO21, JĎD^{+}21, RCB21, SJP^{+}21b, ZDLS21]. morphology [ANU21, RDZ^{+}20]. mosaic [SPF21]. Moshinsky [Efr21b]. most [PRR23]. motion [GSV23, ZHM21, dVAR^{+}24]. movements [GTE21]. moving [FH22, HMR22, MDZ24, TCSA21, WDMZ24]. moving-grid [TCSA21]. MP [FN23]. MP-CITDSE [FN23]. MPI [LYZL24, YW21, ZHZG23]. MPI-based [ZHZG23]. MPI-CUDA [YW21]. MPS [eSdSBST21, XFGS24]. MPS-VQE [XFGS24]. MQCT [MBJB24, SMB20]. mqdtfit [Pot24]. MRADSIM [ACKB23]. MRADSIM-Converter [ACKB23]. ms2 [FGCN^{+}21]. MSGCorep [LZYY23]. MSSM [BHH^{+}20, KMU^{+}23]. MSTor [CZB^{+}23]. MTASpec [KSG22]. MTP [MPN^{+}21]. MTP/ShengBTE [MPN^{+}21]. muffin [PAL^{+}20]. muffin-tin [PAL^{+}20]. MuFinder [HHMH^{+}22]. Multem [SLL^{+}24]. Multi [CZX^{+}21, CJD^{+}20, DS22, DKM^{+}24, JWW^{+}23, LLC^{+}23, MDDI21, OCK^{+}24, ADC^{+}21, AG21, BP21, BSC^{+}21, BW23, CZ21, CTZW23, COW24, CFW^{+}23, Cos22, DRZ^{+}21, DFU20, EGLK20, FJ22, FRN^{+}23, GDS^{+}21, GDJG23, GNP20, HPY21, HSA22, KMM21, KDK23, KMD^{+}21, KMG^{+}20, LSS24, MVK^{+}24, MBC^{+}24, MHK24, MPH^{+}24, Nor23, OGL^{+}21, PMK^{+}23, PWD22, PG23, RMR^{+}22, RZ23, SNP^{+}20, SPMS23, Sij23, TPCT22, WZPW20, Xav22, XLL^{+}22, XPF^{+}24, Yan21a, YFL22, YKK^{+}23a, YKK23b, ZPZH20, ZW20, ZRZ^{+}21, FLK^{+}20]. multi-architecture [Nor23]. multi-block [Cos22]. multi-center [GDS^{+}21]. Multi-component [DKM^{+}24, BSC^{+}21]. multi-core [KMM21, OGL^{+}21]. multi-dimensional [KMD^{+}21, PMK^{+}23, WZPW20]. multi-direct-forcing [ZPZH20]. multi-electron [COW24]. multi-field [GNP20]. multi-fluid [MPH^{+}24, PG23]. multi-function [CZ21]. Multi-GPU [DS22, OCK^{+}24, DRZ^{+}21, DFU20, EGLK20, HPY21, MVK^{+}24, MBC^{+}24, MHK24]. multi-GPU-based [YKK^{+}23a]. multi-GPUs [WZPW20, ZW20]. multi-grid [BW23]. multi-group [Sij23]. multi-harmonic [LSS24]. multi-ion-species [SNP^{+}20]. multi-jet [BP21]. multi-loop [ADC^{+}21]. multi-material [Sij23]. multi-million [RZ23]. multi-moment [CFW^{+}23]. Multi-objective [CZX^{+}21, KMG^{+}20, SPMS23]. multi-orbital [PWD22]. multi-phase [BSC^{+}21, FLK^{+}20]. Multi-photon [JWW^{+}23]. multi-physical [YFL22, YKK23b]. Multi-physics [LLC^{+}23, AG21, FRN^{+}23, GDJG23, XLL^{+}22, XPF^{+}24, ZRZ^{+}21].

Multi-Rate [MDDI21]. multi-region [Yan21a]. multi-resolution [ZRZ^{+}21]. multi-scale [BSC^{+}21, CTZW23, FJ22, TPCT22]. Multi-species [CJD^{+}20, KDK23]. multi-stable [HSA22]. multi-state [Xav22]. multi-term [LSS24]. multi-threaded [AG21]. multi-timescale [RMR^{+}22]. multicenter [HG22]. multichannel [Koł23, Pot24]. Multicolor [Ume22]. multiconfiguration [LWS^{+}23]. multiconfigurational [Hua24]. multicoordinate [LKK23]. multidimensional [Bar22b, KMBP24, RBV^{+}22, Yan24b]. Multidisciplinary [WKR23]. multielectron [BAB^{+}20]. multifractal [GHKW22]. multigrid [EVFRHR23, FKK^{+}21, TS23, UJ21]. multigroup [MXH^{+}24]. MultiHypExp [Ber24]. multilayered [BVV22]. multilayers [LL23]. Multilevel [ELSV24, YYY22, WSRO24]. multiloop [Lee21b]. multiobjective [NCF^{+}23]. multiphase [CESD^{+}23, FA20b, RLW^{+}24, SPLD20, TO21, YCCW23]. Multiphysics [ARLDG24, MEDT^{+}23, SPLD20, Zho23]. Multiple [Jab22b, ZK20, Ber24, CCM20, CGSO20, DC22, JM24, KDHL23, ODR^{+}22, PNL^{+}24, SZ23, WA21, KOF21]. Multiple-Ion [KOF21]. multiple-thread [PNL^{+}24]. multiply [KMM21]. multipole [CB20, KKLZ23a, TCD20, WZC21, WVK21, YT23]. Multipolynomial [LMWW24b]. multiresolution [GP22, HWAA22]. Multiscale [LWV20, SD24, AS24a, FTZ^{+}23, LFL^{+}24, SGY24, Tan23, ZHZ^{+}23]. multistable [MOA24]. multistructural [CZB^{+}23]. Multithreaded [LTMK21, LLR23]. multiUQ [TO21]. multivariable [ABFP24]. multivariate [AKL^{+}21, BBH^{+}24, Ber24, EBBB22]. MultivariateApart [HvM22]. multizone [LLZ^{+}23]. MuMuPy [UAS22]. MUNHECA [EES24]. munuSSM [Bie21]. muon [BFI^{+}21, EES24, HHMH^{+}22]. muon-nuclear [BFI^{+}21]. muon-rich [EES24]. muphy [ARLDG24]. MUSES [SMA24]. mVMC [XOTI22]. MXAN [BDP^{+}21]. MXE [MP21]. MyElas [WGGC22].

N [Haz23, NUK^{+}22, VMRFC23, LNB23, SGM^{+}20]. N-on-Equilibrium [SGM^{+}20]. N-shaped [LNB23]. Nano [SLIC24, WMA^{+}22]. Nano- [SLIC24]. nano-optics [WMA^{+}22]. nanobem [Hoh24, HRU22]. nanodrop [SD24]. nanoheteroepitaxial [DZZ21]. NanoNET [KVSC21]. nanoparticle [WAK22]. nanoparticles [HT24]. Nanophotonic [HRU22, Hoh24]. nanophotonics [BFCR24]. nanoporous [XBL^{+}20]. nanoscale [KGN^{+}21, MMM20, MP21, TRB20b]. nanostructures [PWC24, RZ23, SHS22]. nanosystems [MVK^{+}24]. nanowire [CW22]. nanowires [CSW24]. natural [PM21, RSD20, SRT^{+}20, SLE^{+}22]. natural-orbital-functional-based [PM21]. Nauticle [HT20]. Navier [WHSK24, BMS^{+}23, DAC^{+}23, HPY21, HWL^{+}23, JM24, KSF^{+}22, MOMO24, ZX23]. nccrFOAM [MEC^{+}24]. NCrystal [CK20a, KC21]. near [UYS22, ZWZ^{+}22]. near-continuum [ZWZ^{+}22]. nearest [WZC^{+}24]. nearshore [TL20]. NeatIBP [WBM^{+}24]. neBEM [DBM^{+}24]. needed [Jab22a]. neighbor [WZC^{+}24]. Nek5000 [MPS^{+}24]. NekMesh [GKT^{+}24].

Nektar [LBS^{+}23, LIG^{+}24, MCB^{+}20]. Nelder [PP23]. NEMOH [KD23]. neoclassical [SNP^{+}20]. Nernst [ZX23]. NESSi [SGM^{+}20]. net [DH20, ZLY24]. Network [LPSK20, BRAC23, BB24a, CDD22, CGZ^{+}20, DRB22, ET24, HTS^{+}21, KHKL24, LY24b, MM22, MOY^{+}22, WZC^{+}24, YKYK23, ZHI23, ZYML24]. network-based [LY24b]. Networks [LY22, SWZ23, AM21b, BKS22, KKP22, LSF23, RRC^{+}24, Rod22, ZLY24, vdHKB^{+}23]. Neural [LY22, LPSK20, SWZ23, BRAC23, BB24a, BKS22, CGZ^{+}20, DRB22, ET24, HTS^{+}21, KHKL24, KKP22, LSF23, LY24b, MM22, RRC^{+}24, Rod22, WZC^{+}24, YKYK23, ZHI23, ZYML24, ZLY24, vdHKB^{+}23]. neural-network [WZC^{+}24]. neutral [AZH^{+}24, SLP^{+}22, TKS22]. neutrino [AAA^{+}21, AZH^{+}24, AAG^{+}20, ASW22, Gar21, KTF22, Pag24]. neutrino-nucleus [AZH^{+}24]. Neutrinos [EES24, GLW^{+}23]. neutron [CK20a, CKC^{+}21, CL21, CLEPF23, CLEP24, KJL^{+}23, KC21, MKPW21, MKPW22, VGGP^{+}21, WBF^{+}24, WLL22, XSL^{+}22]. neutrons [BCT20]. Newly [IKM^{+}24, LHWX24]. Newton [AG21, LVMGF^{+}23, SLC^{+}22, Zho23]. Newtonian [AHP21]. Next [AAT^{+}14, AAT^{+}20, ABB^{+}22, WZC^{+}24, KM20]. next-nearest [WZC^{+}24]. next-to-leading [ABB^{+}22]. Next-to-Minimal [AAT^{+}20, AAT^{+}14, KM20]. NIC [RHW^{+}21]. NIC-CAGE [RHW^{+}21]. Nicolson [LNB23]. nitride [HLCD20]. NLO [SY20]. NLOX [FQRR22, HQRR20]. NLP [LM22]. NNPred [LSF23]. noble [TNL^{+}22]. Nodal [WNS^{+}21, SS21b, XSL^{+}22, YZHL22, Zho23, ZLS^{+}22]. nodal-staggered [ZLS^{+}22]. node [GSBN22]. node-level [GSBN22]. nodes [PP23, YZHL22, PP23]. noise [LMWW24a, TCcN23]. noising [TAGC22a]. noisy [KIK20]. nominal [SFBG20]. Non [DNG^{+}20, JLL^{+}24, Kar23b, KKY24, MGX24, NA20, AHP21, AKK^{+}24, ADW^{+}23, Dau23, DFP^{+}20, DH20, Efr23, FMBD22, GBS^{+}20, GOCSS^{+}23, JDS20, KKM^{+}20, LBS^{+}23, LHG^{+}20, MES^{+}24, MEC^{+}24, NVC20, PCS^{+}20, QCZ23, RHLTG^{+}22, RG21, RVRT22, SYFT23, SJ21, SLB^{+}23, SBZ23, VXT^{+}23, VB22b, WLCF21, XSM22, ZYLY22, ZYL^{+}23, ZZ21a, ZCS^{+}24]. non-aligned [SBZ23]. non-bonded [VXT^{+}23]. non-collinear [GOCSS^{+}23]. non-convex [LHG^{+}20]. Non-empirical [NA20]. non-equilibrium [MES^{+}24, MEC^{+}24, SYFT23, ZCS^{+}24]. non-flat [SJ21]. non-identical [XSM22]. non-isothermal [RHLTG^{+}22]. Non-iterative [KKY24, WLCF21]. Non-linear [DNG^{+}20, LBS^{+}23, RVRT22, SLB^{+}23]. Non-local [JLL^{+}24, NVC20]. non-magnetic [LL23, ZYLY22, ZYL^{+}23]. non-minimal [AKK^{+}24, KKM^{+}20]. non-Newtonian [AHP21]. Non-Oscillatory [MGX24]. non-pairwise [RG21]. non-periodic [FMBD22]. non-reactive [ADW^{+}23]. non-spherical [QCZ23, ZZ21a]. Non-standard [Kar23b]. non-thermal [DFP^{+}20]. non-uniform [PCS^{+}20, VB22b]. non-vanishing [JDS20]. non-zero [DH20]. nonadiabatic [SZT24]. noncentral [GST21]. nonequilibrium [FTZ^{+}23]. nonequispaced [YW21]. Nonintrusive [TGS^{+}20]. Nonlinear [KSJ^{+}22, MEC^{+}24, BJL^{+}24, BCGT24, CLS22b, DTC20, FBA^{+}20, GTA21, GGCW24, GMZ^{+}20, Hua24, JLL^{+}24, KBB21, LDGN24, NS20, SMA24, SCL22, WXY20, WHSK24, WKR23, ZJM^{+}21]. nonlocal [AB21, BDA^{+}20, CLVV22, CSW24, MZV23]. Nonrad [TAE^{+}21]. nonradiative [TAE^{+}21]. nonsequential [TNL^{+}22]. nontrivial [Pöt20]. nOOn [BCT20]. NORA [Sit23]. norm [TIG^{+}24]. norm-conserving [TIG^{+}24]. normalizing [BKS24a]. nouveau [PGS^{+}24]. novel [Ano20c, JLL^{+}24, KK16, LC24, LNB23, PMS^{+}20a, SFKC22, SCMP^{+}22, SCR^{+}22b, TCY^{+}21, WLL22, YZW21, Yan21b, ZHZG23]. novelties [ADH^{+}20]. November [Ano20w, Ano21v, Ano22v, Ano23v]. NSC [Kar23b]. NTMpy [ASU^{+}21]. nuclear [AAMY23, BFI^{+}21, COJ^{+}22, CCG21, DR21, GUAD22, GSV23, Hor23b, KJL^{+}23, LLC^{+}23, LLL24, MAJ20, TBAR21, Zho23]. nuclear-decay [Hor23b]. nucleation [Bar22a, WNS^{+}21]. nuclei [GCK21, MLD^{+}22]. Nucleon [BDA^{+}20, Kar23a, MLD^{+}22, OYC24, SW23]. nucleosynthesis [AAHJ20]. nucleus [AZH^{+}24, BDA^{+}20, GJJN23, MLD^{+}22, NRK^{+}21]. nuclide [BOSM24, LWhK^{+}20]. nudged [BFD22]. number [ATRD21, CdBMdAS^{+}21, HFP21, LLT^{+}23, MD21]. numbers [CXCZ23, Nog21b]. numeric [HLzY^{+}20]. Numerical [AS22, Arn20, Bar22b, CKT21, GMZ^{+}20, HJK^{+}24, KOF21, KRE22, MPQ^{+}22, NS20, SZ23, SSV22, WJB21, WWB22, WHB21, WW21, ADC^{+}21, Bak23, BMSP21, BTG22, CFBRE24, CZY20, FA21, GSLS20, GOCSS^{+}23, GWA^{+}23, GHL^{+}22, GB20, KSF^{+}22, KPL^{+}21, KFC^{+}20, Kul20, LM21a, LS21a, MPS^{+}24, MN21, NSU20, PG23, SGY24, THH21, VB22b, WAK22, Yan21b, YWM23, ZCWY20, ZRH21]. Numerically [LHTP^{+}24]. Nunziato [Dau23]. NuSD [KTF22]. nuSQuIDS [ASW22].

O [NUK^{+}22]. O-SUKI-N [NUK^{+}22]. Object [BC22, CB23, JKST22]. Object-oriented [BC22, CB23, JKST22]. objective [CZX^{+}21, KMG^{+}20, SPMS23]. objects [GTE21]. observable [YZL^{+}23]. observables [CNB^{+}23, KKSV24, PAM24]. observatories [AAA^{+}21]. obtain [GWPW21, VMRFC23]. obtaining [AUEO24, DBdFdSR21]. ocean [GG24]. OCTAVE [MN21]. October [Ano20x, Ano21w, Ano22w, Ano23w]. octree [KSF^{+}22, TS23]. OECP [YMCF23]. off [AB21, DVC^{+}22, MPH^{+}24, MAMK21]. offload [SYFT23]. Ohmic [SZ23, VGGP^{+}21]. Ohta [CZY20]. Oldroyd [AGH21]. Oldroyd-B [AGH21]. Olsson.wl [ABFP24]. on-site [CXCZ23]. on-the-fly [BMREC21]. one [AKK^{+}24, BV21, FQRR22, GAJK23, GSLS20, GSV23, Hid21, HQRR20, KKM^{+}20, KM20, KMS20b, LC24, MWJL23, OCE^{+}23, SFBG20, TMH^{+}23]. one-dimensional [BV21, Hid21, OCE^{+}23]. one-electron [GSV23, SFBG20]. one-loop [AKK^{+}24, FQRR22, HQRR20, KKM^{+}20, KM20, KMS20b]. one-point [LC24]. one-step [MWJL23]. OPAL [ASA^{+}22]. OPAL-FEL [ASA^{+}22]. Open [GDK21, HMYH22, HKY^{+}21, KD23, WNS^{+}21, WSK^{+}22, AFB^{+}24a, ASU^{+}21, ADF^{+}22, ASC^{+}21, BC20, BM22, BLN^{+}21, BMT^{+}20, BFI^{+}21, BSC^{+}21, CAWK22, CZS^{+}21, CPL21, COW24, CSW24, CW22, CKGW22, DIK^{+}23, DNG^{+}20, DFU20, DAA^{+}24, DBBP23, FTZ^{+}23, FH22, GDJG23, GKT^{+}24, GFD^{+}24, HSB^{+}24, HXS20, HWL^{+}23, HLMB24, HRG^{+}22, Hua23, HHT^{+}24, KGN^{+}21, KGT22, LEE^{+}24, LMHL20, LH22, MLZ^{+}23, MBTB21, MMP^{+}24, MFB23, MONW21, MAJ20, MVF20, MVAXP22, MRN20, NBC^{+}24, Nor23, DARJ23, PM21, PWC24, RZWW23, RHW^{+}21, RJ21, SPLD20, SHB^{+}20, SHW^{+}21, SWTC23, TPK^{+}21, XOTI22, YKSH20, YMCF23, YKK23b, zYCD^{+}20, ZRZ^{+}21, ZJS^{+}20, ZCS^{+}24]. open-ended [YMCF23]. Open-source [GDK21, HMYH22, HKY^{+}21, KD23, WNS^{+}21, WSK^{+}22, AFB^{+}24a, ADF^{+}22, ASC^{+}21, BM22, BLN^{+}21, BMT^{+}20, BFI^{+}21, BSC^{+}21, CZS^{+}21, COW24, CSW24, CW22, CKGW22, DIK^{+}23, DFU20, DAA^{+}24, FTZ^{+}23, FH22, GDJG23, GKT^{+}24, GFD^{+}24, HSB^{+}24, HXS20, HWL^{+}23, HLMB24, HRG^{+}22, HHT^{+}24, KGT22, LEE^{+}24, LMHL20, LH22, MLZ^{+}23, MBTB21, MMP^{+}24, MFB23, MONW21, MVF20, MVAXP22, NBC^{+}24, Nor23, DARJ23, PM21, PWC24, RZWW23, RHW^{+}21, RJ21, SPLD20, SHB^{+}20, SHW^{+}21, SWTC23, TPK^{+}21, XOTI22, YKSH20, YMCF23, YKK23b, ZRZ^{+}21, ZCS^{+}24]. OpenACC [HT24, EVMP20]. OpenDust [KT23]. OpenFOAM [CS22a, FCSP20, FA20a, FA20b, KRE22, KG21, MEC^{+}24, NJSY22, OCC23, SN23, SWTC23, TPK^{+}21, TGS^{+}20, UJ21, VV21, Yan21a, YFL22, Zha20, ŻTR^{+}22]. OpenFOAM-based [NJSY22]. OpenFOAM(R) [KFC^{+}20, MDDI21, OXOG23]. OpenFSI [YSX^{+}20]. openFuelCell2 [ZHM^{+}24]. OpenMC [LEE^{+}24, LEE^{+}24]. OpenMP [HJGL22, MBA21, SYFT23, YSMBA23, EVMP20]. OpenMPI [DS22, PTH24]. OpenNTP [LEE^{+}21a]. OpenSANS [MZV23]. OpenSBLI [LJS21]. Operating [CTPS22, KRC^{+}20]. operation [HBM^{+}24]. operations [BBV23, CB23]. operator [FKK^{+}21, HS24, KSJ^{+}22, KDK23, SNP^{+}20, WJB21, WWB22]. operators [CB20, GBS^{+}20, HG22, RVRT22, YT22]. Operon [BB24b]. OPSimTool [Kan23]. optical [AGMFGE23, AB21, BMP^{+}24, BRHT21, CMS22, DIAA21, HMSV23, Kan23, MD22, OV23, SKEZ24b, SLB^{+}23, UÁEPGBP24]. optics [KGN^{+}21, WMA^{+}22]. Optimal [SRK^{+}24, BB21, Cas24, COW24, Gai20, JM24, KKY24, LY24b, RHW^{+}21, RBWD^{+}24, RRM^{+}23, RDZ^{+}20, Tan24]. Optimisation [TWW22, Bar22a, TV24]. optimised [TIG^{+}24]. optimistic [NBCMH20]. optimization [AS24a, CdBMdAS^{+}21, CZX^{+}21, FLW^{+}23, FBC^{+}21, HTS^{+}21, KL23b, KMG^{+}20, KM22, LDGN24, MTY^{+}22, NCF^{+}23, RdPS24, SPMS23, VHBK21, WYHW24, WBK^{+}24, WWJ^{+}22, YK21b, YNMR24, ZZC20]. optimizations [HK24]. Optimized [RAJ^{+}24, VTB^{+}21, XOTI22, YSI^{+}24, CL20, CCL^{+}22, FZD^{+}24, LYZL24, RdSH^{+}24, TMH^{+}23]. Optimizing [HTL^{+}22, SHRK22, WSRO24, CB20, Gon22]. OptiSMOKE [FBC^{+}21]. optoelectronic [DSQ23, VMRFC23]. Orb [LOT^{+}20]. ORB5 [NBB^{+}21, OGL^{+}21]. orbit [BKRG22, BJS^{+}23, BRHT21, FBA^{+}20, GOCSS^{+}23, KRG21, MBA21, RVM^{+}21, YHH^{+}20]. orbital [BW23, GM20, HLCD20, MHK24, PAL^{+}20, PWD22, PM21, RZY^{+}24, ZRH21]. orbital-free [BW23, GM20, MHK24]. orbitals [GOCSS^{+}23, SLE^{+}22, ZXW^{+}22]. orbits [JDS20, JSS^{+}24, MPQ^{+}22]. order [AIZ23, BBV^{+}22, ADF^{+}22, BBA23, BSC^{+}21, BKG^{+}23, ČPF^{+}24, CJ21, CWJ21, DBV^{+}24, DFU20, DEdM24, EVMP20, EPM23, FJ22, FRN^{+}23, GBS^{+}20, GZW20, GHL^{+}22, GKT^{+}24, GGCW24, GMZ^{+}20, HWL^{+}23, HCAH^{+}24, KBB21, KSF^{+}22, KD23, LM21a, LF24, LLS^{+}21, LLZ^{+}22, LLY^{+}22, MDP22, PMSHG23, PLL24, PB23, PLF20, PG23, PCS^{+}23, QWZ^{+}21, RZY^{+}24, RCW^{+}20, SP20, SDBS24, TCY23, Tan24, TSL21, VLS22, WXY20, XPF^{+}24, YS22, Yan21b, YK21b, YKK23b, ZCWY20, Zho23, ZZZ^{+}20, ZLS^{+}22]. ordering [TCY23]. ordinary [DEdM24, MCP23]. organic [BMR^{+}24a]. orientation [AE23, OS24, SJY18, SJY20]. oriented [ACD^{+}22, BC22, CB23, DFU20, JKST22, Sha21, Soz21, YR21, YK21b]. orthogonal [RYS^{+}24]. orthogonality [ÖAÖ24]. orthonormal [DEV20]. oscillating [SPTPR21]. oscillation [GLW^{+}23, Pag24]. oscillations [Dan23]. Oscillator [Efr23, BDdM20, Efr21b, MSN^{+}22, SKEZ24b, SMGK21]. oscillators [FH22]. Oscillatory [MGX24]. Oslo [MZL^{+}21]. osmo [FFTV23]. OSSCAR [DBBP23]. otebook [VMRFC23]. other [ADH^{+}20, KKL21, MS24, ZHM^{+}24]. OTSLM [LSNRD20]. out-of-equilibrium [WWB22]. overhead [BGR23, VHBK21]. overlaps [AJDS^{+}21]. overrelaxation [BJL^{+}24]. overset [CS22a].

P [VMRFC23, SHS22, VMRFC23]. P-TRANS [SHS22]. PACIAE [SZY^{+}22, YZL^{+}23]. Package [GNP20, PMS20b, AGMFGE23, ASU^{+}21, ASPDL^{+}21, ACS^{+}22, ABBD23, AYI^{+}24, AGK^{+}23, ASC^{+}21, Bag22, BA24, BYWW23, BBH^{+}24, BFT20, Ber24, BWW20, Bie21, BM22, BCD^{+}21, Bzo21, CMS22, CDD22, CZWE23, COW24, CSW24, CW22, DKRSR22, EBNS22, EST23, FASD20, FH22, Fon21, GAJK23, GK21, Gal22, GST21, GvdBdGN24, GBR23, HyLF23, HHVB21, HPP23, Hid21, HGS20, JWK^{+}21, JCM20, JRS^{+}21, JPJ^{+}23, KH23, KRL^{+}24, KDHL23, Kür23, Lee21b, LKK23, LCL^{+}23, LZK^{+}23, LYX22, LCZ^{+}21, LM23, LZYY23, LH22, MD22, MP21, MHK24, MOMO24, MTY^{+}22, NG21, NKP20, OBC^{+}24, PAL^{+}20, PSW23, PTD20, RHLTG^{+}22, RHW^{+}21, RZY^{+}24, RYS^{+}24, Sat21, SGM^{+}20, SXW^{+}20, SAS20, Sit23, SAN24, TR22, Tor20, TMH^{+}23, TO21, UD24, VCF22, VB22a, WZZ^{+}23, Wor20, WZPW20, WBM^{+}24, XBL^{+}20, YSX^{+}20, ZYLY22, ZYL^{+}23, ZPL^{+}21, dSOZ22, vdHKB^{+}23]. packages [ABFP24, SHW^{+}21]. packet [Dio23]. Padé [PPK22]. Pair [AJW^{+}21, BKBL22, JWW^{+}23, Jia20, KKB^{+}24]. PairDiag [LQ21]. PairDiagSph [LQGL21]. pairing [GQ22, LQ21, LQGL21]. pairs [ABWZ23]. pairwise [HAA^{+}20, Mar22, RG21]. PAMCARS [ZHS^{+}20]. panel [SCMP^{+}22, SCR^{+}22b]. PANNA [LPSK20]. paper [GDK21]. paper-based [GDK21]. Parabolic [OSE^{+}20, ASU^{+}21, BCGT24]. paradigm [MMV^{+}24]. Parallel [AHM^{+}23, ASC^{+}24, BR20b, DS20, HMR22, Ham20b, HW22, JM24, SRT^{+}20, ACS^{+}22, AS24b, BM20, BKP22, BGW^{+}22, BBH^{+}24, BB24b, BMT^{+}20, CZS^{+}21, CCC20, CJZ21, CHY^{+}24, Cos22, DMS^{+}22, DIKSN24, EVMP20, FWL^{+}24a, GTMB21, GAGO21, HJGL22, KWZ24, KF23, KHR^{+}23, KKPC21, KDIN^{+}23, KMM21, KSC^{+}23, KFHR24, LBRW22, LLZ^{+}23, LYZL24, LMHUR23, LNB23, MD20b, MW21, MAMK21, MDR^{+}20, Nor23, OIA^{+}20, RCB^{+}20, RS20, RYS^{+}24, SXW^{+}20, SLZY21, SBG23, WWZ^{+}23, WLZ^{+}24, XLKX21, YS22, YCC22, Yan23, YML^{+}24, YKY^{+}22, ZAW^{+}21, ZWZ^{+}22, ZLMH23, ZHZG23, AACE^{+}21, BBH^{+}24]. Parallel-in-time [JM24]. paralleling [DRZ^{+}21]. Parallelisation [UJ21]. parallelism [MLT^{+}21]. Parallelization [DBM^{+}24, TCD20, DFG^{+}23, GB22, RCX24, YSI^{+}24, ZWP^{+}22]. Parallelized [BSK^{+}22, CGR21, KKLZ23a, LKP21]. parameter [GJ24, HPRS23, LWV20, SMC^{+}22, SZ24a, TSAK21, ZW20]. parameterizations [Koł23]. parameters [CL20, CFL^{+}22, DBdFdSR21, FA21, GOST23, HHVB21, HLCD20, Jab22a, MMCC^{+}22, NKP20, SKC21, SHRK22, TMC22, TNL^{+}22]. parametric [RUV20, SKEZ24b]. parametrised [TGS^{+}20]. Parametrization [OKBM23]. Parametrization-free [OKBM23]. parametrized [AFB^{+}24b, BDGS21]. ParamGULP [DBdFdSR21]. PARCE [OSLC21]. PaReDiSo [PTH24]. Pareto [KMG^{+}20]. Pareto-frontal [KMG^{+}20]. PARIS [AACE^{+}21]. Parrinello [vdHKB^{+}23]. PARSEC [LYC20]. PARSIFAL [AFB^{+}24b]. parsimonious [SJ20]. Part [PLT^{+}23]. PArthENoPE [GdSPC22]. Partial [Pos22, BBH^{+}24, GHL23, HvM22, OV23, SJP21a, ZLY24]. Partial-wave [Pos22, SJP21a]. partially [MR22, MS24, MM23]. Particle [AMA^{+}20, ANU21, DHK^{+}21, FAL21, JCM20, KPR^{+}24, YYC^{+}23, ABK^{+}22, AYWKL24, ASC^{+}24, BCHE21, BKB^{+}21, BC21, BTK24, BFL^{+}22, CGSO20, CTZW23, DC22, DBH^{+}22, DWD^{+}24, Efr20, Efr21a, Efr23, FN21, FJ22, FTZ^{+}23, FRN21, FA21, FYM^{+}22, GWA^{+}23, GH21, GSBN22, GB20, HT20, HFP21, JSS^{+}24, KP23, KMU^{+}23, KSC^{+}23, LBRW22, LG21a, LMX^{+}21, LAD^{+}21, LLT^{+}23, LM20, LMQ^{+}23, LYZL24, MK22, Mar22, May21, MRH^{+}23, MHP23, MLT^{+}21, MMYU22, MPQ^{+}22, MBE^{+}21, NTO24, NSY^{+}23, NR21, NBS^{+}20, NBB^{+}21, NRG24, ODR^{+}22, OCR^{+}22, OBGA24, OAP^{+}24, PAM24, PYT^{+}24, PSMRS^{+}23, PK24, PCS^{+}20, RDH^{+}20, RZH22, RC23, SRC21, SFKC22, SDP24, SLL22, STRF^{+}20, SWB^{+}23, SMGK21, SBP20, TV24, THH21, TWR21, TAGC22b, TGGC23, TAB^{+}22, TAY^{+}24, Ume22, VEHCM21, VRI24, WWM^{+}22, WOP^{+}20, WS20, XBL^{+}20, XQ21, XHY^{+}24, YC20, YGSW21]. particle [YWX^{+}23, YZZ^{+}23, ZK20, ZRZ^{+}21, ZDSS23, ZZC20, ZZ24, ZHZG23, ZLS^{+}22, ATRD21, DFG^{+}23, DWD^{+}24, FAL20, FA21, JĎD^{+}21, KG21, OGL^{+}21, SAC^{+}21]. Particle-based [KPR^{+}24, HT20, TAB^{+}22]. Particle-In-Cell [FAL21, YYC^{+}23, ASC^{+}24, BCHE21, BC21, BTK24, CTZW23, DBH^{+}22, FN21, FYM^{+}22, LMX^{+}21, LAD^{+}21, MRH^{+}23, MLT^{+}21, NTO24, OCR^{+}22, PSMRS^{+}23, RC23, SWB^{+}23, TWR21, Ume22, XHY^{+}24, YC20, YWX^{+}23, ZK20, ZDSS23, ZLS^{+}22, DFG^{+}23, DWD^{+}24, FAL20, FA21, OGL^{+}21, SAC^{+}21, JĎD^{+}21, KG21]. Particle-in-Cell/Monte-Carlo [JĎD^{+}21]. particle-laden [GWA^{+}23, GB20]. particle-laser [LMX^{+}21]. particle-matter [NBS^{+}20]. Particle-number [ATRD21]. particle-swarm [ZZC20]. particles [AGJ^{+}23, ADH^{+}20, FBA^{+}20, Kür23, LJQ^{+}22, LHWX24, Nie20, PCL24, RDZ^{+}20, SLLA22, SA23, SAC^{+}21, TCcN23, YFL22, ZZ21a]. Particlization [MEH21]. particulate [QCZ23]. partition [LNB23]. partitioning [Xav22]. pARTn [PGS^{+}24]. Parton [CCMR21, SZY^{+}22]. parts [WBM^{+}24]. PaScaL TCS [KKPC23]. PaScaL TDMA [KKPC21, YKK^{+}23a]. passage [KKY24]. Past [Sjö20]. patch [ASC^{+}24]. patch-based [ASC^{+}24]. Path [PBC^{+}24, KDHL23, Sva24]. pathological [HMM22]. paths [NA20]. pattern [HT24, Kan24]. patterns [JLW24]. Pauli [WBvdH20, XQ21]. Pauling [ZHS^{+}20]. Paulo [CCG21]. PBCAVE [BBA^{+}20]. PDE [OIA^{+}20]. PDFFlow [CCMR21]. PDMC [PLT^{+}23]. Peak [ZWG^{+}24]. peakon [ZYXC24]. PENELOPE [GAGO21]. penetration [BKB^{+}21]. Peng [ZLLM23]. PenRed [GAGO21, OBGA24]. pentaquark [GBJ^{+}21]. Perfectly [Pöt20, NTO24, OKBM23]. Performance [LKP21, TBAR21, WLZ^{+}24, AALK20, BPMMP24, BBC^{+}24, BLM^{+}22, BDR^{+}20, CZTF23, DRB22, DIKSN24, FAL20, FAL21, GAS^{+}23, HSA22, KDIN^{+}23, KMM21, KPL^{+}21, MT23, OSK^{+}21, OLNG21, RBFB20, SRE^{+}24, SGY24, SSV22, WRM^{+}24, YJLW21, YML^{+}24, YKK23b, YLK^{+}20, ZMC23]. performant [BB24a, RLW^{+}24]. performing [SC22, VCF22]. Peridynamic [LFL^{+}24]. peridynamics [LC24]. Periodic [GVV22, HPP23, FMBD22, JDS20, KAG24, LXY^{+}21, MD20a]. peripheral [BCT20, BKBL22]. Perl [HBM^{+}24]. Perl-based [HBM^{+}24]. permaFoam [OXOG23]. Permafrost [OXOG23]. permanent [HK24]. permittivity [YMCF23]. permutation [YLIO22, Zeb22]. perspective [LC24]. Perturbation [RTRB21, ATRD21, GBD^{+}22, LACL24, SLZY21, SDL^{+}21, TMC22, Gal22]. perturbations [YYC^{+}23]. Perturbo [ZPL^{+}21]. PET [HLMB24]. PETOOL [OSE^{+}20]. Pfaffians [XOTI22]. pfd [BBH^{+}24]. pfd-parallel [BBH^{+}24]. PFLOPS [IKM^{+}23, LWC^{+}21]. pgm [WZZ^{+}23]. pGrAdd [WV22]. PHARE [ASC^{+}24]. phase [AYI^{+}24, Bar22a, BMM21, BBA23, BND22, BSC^{+}21, CAWK22, CFLR24, CZY20, Dau23, EST23, ENK24, FDPT23, KMN21, KBSL22, Koł23, KNJ^{+}23, MMFdL21, MA24, NBB^{+}21, PBK21, RSD20, SRT^{+}20, SZ23, SFS22, TCSA21, TCY23, TQGE23, VB22b, WT22, WYT23, Yan21b, YK21a, ZCWY20, ZDG^{+}21, YLK21, FLK^{+}20]. phase-change [RSD20, SRT^{+}20]. phase-field [CZY20, FDPT23, KNJ^{+}23, TCY23, WYT23, YK21a, ZCWY20, ZDG^{+}21]. Phase-I [YLK21]. phase-space [NBB^{+}21, TCSA21]. phases [AMP^{+}21, FLW^{+}23, KvdW20]. PhasicFlow [Nor23]. phenomena [MP21]. phenomenology [BBH^{+}23]. phonon [HYL^{+}22, LACL24, LFL^{+}24, MMP^{+}24, SHS22, SLIC24, WZZ^{+}23, YHY^{+}21, ZPL^{+}21]. phononic [WXZH24]. phonons [PPR^{+}21]. photo [RHW^{+}21, RBWD^{+}24]. photo-excited [RHW^{+}21, RBWD^{+}24]. photoelectric [WSK^{+}22]. photoelectrochemical [BMI23]. photoelectron [BFMA^{+}24, Jab22a, TB20]. photoemission [BYL^{+}21, RMFB23]. photoinjector [NCF^{+}23]. Photoionization [BFMA^{+}24, HBM^{+}24, TCSD24]. photoluminescence [BMR^{+}24a]. Photon [Kal20, AAD^{+}24, JWW^{+}23, Kan23, KMS20a, LHWX24, OV23]. photonic [DNG^{+}20, KR23, MSH22, SLL^{+}24, WZZS21]. photons [MBG^{+}20]. Phys [AAT^{+}20, Ano20c, FAL21, MKPW22, MSHP20, Och23a, SJY20, YNV23, ZYX22]. Physalis [WS20]. PHYSBO [MTY^{+}22]. physical [AM21b, JYL^{+}23, KKSV24, WAN^{+}22, YFL22, YKK23b]. phYsicist [UMA21]. Physics [ACD^{+}22, LY22, RRC^{+}24, AG21, AGJ^{+}23, ABK^{+}22, BLN^{+}21, BC22, CB23, CGG21, DR21, EBNS22, FXZT21, FTG23, FRN^{+}23, GDJG23, GFD^{+}24, Kal20, KPR^{+}24, Koz23, LLC^{+}23, LP24, LFZ20, PAM24, PBC^{+}24, RGS^{+}21, RSPJ21, SRC21, SDP24, SE24, SHB^{+}20, SZ24a, SSD^{+}22, SCT21, TV24, TW21b, VEHCM21, WBF^{+}24, WLL22, WAET22, XLL^{+}22, XPF^{+}24, XHY^{+}24, ZRZ^{+}21, ZDLS21]. physics-based [SDP24, WAET22]. Physics-Informed [LY22, RRC^{+}24, SE24, WBF^{+}24]. physiologically [ZHZ^{+}23]. physiologically-significant [ZHZ^{+}23]. PIC [Gon22, LOT^{+}20, MT23, MBE^{+}21, NDFL24, WXX24, YC20]. PIC-MCC [WXX24]. picFoam [KG21]. PICLS [SBB^{+}24]. pICNIK [RHLTG^{+}22]. PIGE [MMC^{+}22]. pikoe [OYC24]. Pilgrim [FCTFR20]. pilot [OIA^{+}20]. pin [Zho23]. pin-by-pin [Zho23]. Pineline [BCCM^{+}24]. pipe [XLL^{+}22]. pipelined [PP21]. pipelines [WAK22, WKJB23]. Pitaevskii [BKRG22, CR20, CCW20, CWJ21, GTA21, KRG21, KMR22, KPL^{+}21, YSMBA23]. PittPack [HXS20]. PIV [NT24]. pKa [SOH^{+}23]. pkgPENGEOM [ASPDL^{+}21]. planar [ZL20]. Planck [ZLL^{+}24, ZX23, DTC20, KSJ^{+}22, TCSA21, TKC^{+}21]. plane [BR20a, FASD20, FWL^{+}24a, HM24, JZW^{+}22, MSHP02, MSHP20, YJLW21]. plane-wave [FWL^{+}24a, JZW^{+}22, MSHP02, MSHP20]. plasma [ARLDG24, BYWW23, BGW^{+}22, Bel24, BL21, CTZW23, DIK^{+}23, DML23, DKM^{+}24, GUW^{+}22, HA23, KT23, LAD^{+}21, LLT^{+}23, LLQ^{+}23, LSW^{+}20, MMP20, MRH^{+}23, PG23, RSHS24, RdSH^{+}24, SMC^{+}22, SKYQ21, SSD^{+}22, TKS22, Ume22, WRM^{+}24, sXBkB^{+}22, YGSW21, ZDSS23, ZGW20, ZSqXY21]. plasma-based [LAD^{+}21]. plasma-material [MRH^{+}23, YGSW21]. plasma-neutral [TKS22]. plasmas [BJL^{+}24, BTK24, DFP^{+}20, JĎD^{+}21, KOF21, MXH^{+}24, MMYU22, SMA24, VV21, WLF^{+}23, YC20, YKY^{+}22, ZLL^{+}24]. plasmonic [MZV23]. plasticity [EGLK20, YML^{+}24]. Plateau [LY24a]. platform [BBV23, BLN^{+}21, DBBP23, MLZ^{+}23, OSK^{+}21, PMA^{+}21, SJY18, SJY20, ZWC^{+}20]. platforms [BB24b, CCM20, SCF20, YJLW21]. plethora [FR20]. PLQ [BMR^{+}24a]. PLQ-sim [BMR^{+}24a]. plugin [PGS^{+}24]. PLUTO [MHP23]. PML [XFH^{+}22, ZZZ^{+}20]. pMuTT [LWV20]. Poincaré [SRML20]. point [Ano20c, AM21a, CFBRE24, CLEPF23, CLEP24, DIAA21, FH22, Ham20b, KK16, LG21b, LC24, SSH^{+}23, VZ20]. point-group [CFBRE24]. PointGroupNRG [CFBRE24]. Pointing [PLT^{+}23]. points [BBA^{+}20]. Poisson [BBA^{+}20, CLVV22, DCZ23, HXS20, LSW^{+}20, SLZ^{+}24, SBB^{+}24, TS23, YSI^{+}24, YKY^{+}22, YJ24, ZX23]. Pol [BDdM20]. POLALMM [SFBG20]. polarimeter [CV24]. polarizabilities [SFBG20]. polarizability [Kut20]. polarization [MOV21]. polarizations [NTO24]. Polarized [SY20, BKS24b, BAB^{+}20]. polaron [SZ23]. Pole [KHKL24, AS22]. Pole-fitting [KHKL24]. polyamide [ZBS^{+}23]. polyatomic [LHZ20]. polyatomics [MPZB^{+}24]. polygonal [SVSC20]. polyhedron [Nis22]. PolyHoop [VRI24]. polylogarithms [Ber24, NSU20]. polymer [HT24, SJ20, VvBTH20, ZBS^{+}23]. Polymers [HPP23, ECS23, LVMGF^{+}23, NBC^{+}24]. polynomial [BND22, CCK23, GDB10, LMWW24a, LYC20, LP24]. polynomials [PWC24]. Polyrate [MPZB^{+}24]. polytope [LH24]. PolyWeight [NBC^{+}24]. ponderomotive [BGW^{+}22]. pools [KJL^{+}23]. POP [ZWC22]. population [HFP21, RCS21]. pore [KAG24, OSK^{+}21]. poromechanical [CCC20]. porous [OSK^{+}21, SFS22, YCCW23, ŻTR^{+}22]. porousMultiphaseFoam [HRG^{+}22]. port [BLM^{+}22]. portability [DRB22]. Portable [Dür23, BB24a, DIKSN24, DBH^{+}22, KDIN^{+}23, MDP22, RDC^{+}20, SRE^{+}24, WRM^{+}24, YSX^{+}20, YML^{+}24]. Portobello [AMK24]. posed [ZZ24]. posedness [SWZ23]. position [PMS^{+}20a]. positive [LS21a, SJP21a]. positivity [PGYF21, XDF20]. positivity-preserving [PGYF21]. positron [Dry21, HMYH22, HA23, HSO^{+}22, HBM^{+}24, YZW22]. positrons [MBG^{+}20, SJP21a]. possible [BBA^{+}20]. Post [AM21a, CEC^{+}24, Ham20b, HCAH^{+}24, LR22, MPSK21, WGGC22]. Post-processing [AM21a, CEC^{+}24, Ham20b, HTH^{+}20, HCAH^{+}24, LR22, WGGC22]. post-processor [MPSK21]. Potential [BBB20, LWC^{+}21, NRK^{+}21, BDA^{+}20, CCG21, DSW^{+}23, DBdFdSR21, FMBD22, GCK21, LKP24, NA20, PBK21, RV20, RV21, SKS24, SVJ^{+}24, SSPF24, Tan23, TNL^{+}22, VZ20, Xav22, ZWC^{+}20, ZHI23]. potential-based [FMBD22]. potentials [AB21, BCGT24, CWG^{+}21, CT23, Dio23, HAA^{+}20, KWK^{+}21, Kar23a, LG21b, MPN^{+}21, SRK^{+}24, WGG20, WAET22]. PotLib [SVJ^{+}24]. Potts [KW23]. power [SKS24, ZZ21a]. powered [SGY24]. powers [SA23]. PPR [RdSH^{+}24]. Practical [AKL^{+}21, LACL24, KSIL22]. pre [HTH^{+}20, YZZ^{+}23]. pre-processing [YZZ^{+}23]. pre/post [HTH^{+}20]. pre/post-processing [HTH^{+}20]. precise [MPSK21]. Precision [BHH^{+}20, DS20, KAB^{+}21, KGT22]. preconditioners [YYY22]. preconditioning [LJH^{+}23]. predict [KLD^{+}21]. Predicting [Di 22, ST23b, CW22, LDGN24, RHW^{+}21, ZHS^{+}20]. prediction [ZZ21b, ZZC20]. predictions [BCCM^{+}24, ECS23]. predictive [ZZ21a]. preliminary [WZK^{+}24]. preparation [GAJK23]. presence [AB21, THH21]. present [Sjö20]. preserving [CWJ21, GGCW24, HA21, LW24, PGYF21, SJP^{+}21b, XDF20, XQ21, ZYXC24]. pressure [BTK24, CS22a, JĎD^{+}21, NJSY22, NLS24, UJ21, XZLX20]. pretrained [ZLY24]. primitive [Sva24]. principal [EGKH24, LS23, Yev21]. principle [Jia20]. Principles [NI22, AM21a, BCTS22, Dan24, KvdW20, LLS^{+}21, LLZ^{+}22, MBTB21, MPN^{+}21, Och23a, Och23b, PB23, SDL^{+}21, TAE^{+}21, YI22, YJLW21]. printing [JLW24]. probabilities [DRSZ23]. Probability [PYT^{+}24, BKS24a, PLT^{+}23]. Probability-conserved [PYT^{+}24]. probe [AUO^{+}22, AES21, YMCF23]. problem [Ano20c, CdBMdAS^{+}21, ELSV24, KK16, LLH^{+}21b, SSB^{+}23]. problems [AG21, ACS^{+}22, ATC^{+}23, BKP22, BSK^{+}22, BDR^{+}20, CJZ21, DNG^{+}20, FYW23, HAM^{+}20a, JRG21, JM24, KKPC23, LYZL24, LF20, MM22, PGYF21, PPKK21, RCB^{+}20, Scr22a, SGY24, SLC^{+}22, TGS^{+}20, VPS23, VHBK21, VTB^{+}21, WWZ^{+}23, WHSK24, YYY22, ZJ23]. procedure [CEC^{+}24, KM23]. Process [RBV^{+}22, GAJK23, HA23]. processes [CF21, DAA^{+}24, For22, GPD^{+}22, HMSV23, HQRR20, KKSV24, KMS20a, MBG^{+}20, MT23, RCB21, SC22, WBvdH20, ZHM^{+}24]. Processing [AGH21, AM21a, CEC^{+}24, DC22, Ham20b, HTH^{+}20, HCAH^{+}24, KvH21, LR22, RdSH^{+}24, SPTPR21, WGGC22, WICA22, WKJB23, YZZ^{+}23]. processor [MPSK21, OLNG21]. processor-based [OLNG21]. Procrustes [MRT^{+}22]. produced [MXH^{+}24]. product [CCK23, GDB10, SKM^{+}21, TWW22, XFGS24, YT22]. Production [AJW^{+}21, AZH^{+}24, BKBL22, GBJ^{+}21, JWW^{+}23, KOT23, YZW22]. Professor [Haz23]. profiling [Dry21, MMC^{+}22]. Program [AJW^{+}21, BBA^{+}20, Dio23, Efr20, AKK^{+}24, ADW^{+}23, BOSM24, BDP^{+}21, BN20, BN23, BYL^{+}21, BKBL22, Cas24, CGZ^{+}20, CGV^{+}22, DDM20, FO20, FO21, Gai20, GTMB21, GM20, GJJN23, GFD^{+}24, GSV23, Hor23b, Hor24, HHMH^{+}22, HHT^{+}24, KKM^{+}20, Koł22, KM20, KMS20b, LGBJ20, LZP^{+}24, LQ21, LQGL21, MBJB24, MSN^{+}22, MONW21, MPZB^{+}24, MZ22, NBC^{+}24, NAZ^{+}22, OYC24, PBK21, Roh22, RUV20, SLP^{+}22, SMO^{+}20a, SFBG20, SLZG20, SMB20, SLL^{+}24, SZT24, WGG20, WLZ^{+}24, WAYL23, fWL24b, Xav22, XWYQ21, YLLW24, YLH^{+}20, ZLLM23, ZXT22, ZBS^{+}23, ZXT23, ZJS^{+}20, ZZ21b, DBdFdSR21]. programmed [SKC21, SC22]. programming [EVMP20, ENK24, FSMM24, FLW^{+}23, Par21]. programs [BKRG22, FN23, KSDH23, KRG21, SZ24a, YSMBA23]. project [GAA^{+}20, BMREC21]. projected [ATRD21]. projection [ST23a]. prolate [SSB^{+}23]. Prompt [PYT^{+}24]. propagate [SLP^{+}22]. propagation [AAD^{+}24, ASW22, Koz23, MHÅ21, MR22, MS24, Mau20, MD22, MZL^{+}21, NRG24]. propagator [LG21a, LG21b, Nog21a, PBC^{+}24, RUV20]. proper [RYS^{+}24, TGS^{+}20]. Properties [LY22, LPSK20, BSS^{+}23, BBB20, BTW20, BRHT21, DCRF23, DIAA21, DSQ23, EL24, ENK24, FO21, FGCN^{+}21, FLK^{+}20, GLN23, KKPC23, KLM^{+}22, LBM^{+}23, MMP^{+}24, Mar22, MPSK21, NAZ^{+}21, NAZ^{+}22, SMO^{+}20a, SLK23, SP23, SLDF^{+}21, TR22, VMRFC23, WGGC22, WV22, YNV22, YNV23, YKYK23, ZLLM23]. property [AUEO24, MM23, MMM20, MYKC23, SJWL22]. Prophecy4f [DDM20]. PROPOSAL [AAD^{+}24]. protein [BBA^{+}20, HL21, OCK^{+}24, Ryd20, ZJS^{+}20]. Proteins [HPP23, MBCC23]. Protocol [OSLC21, KMN21]. proton [Ada22, GBJ^{+}21, OYC24]. proton-proton [GBJ^{+}21]. Prototyping [TPS^{+}24]. provider [FQRR22, HQRR20]. providing [AIZ23]. pseudo [CHA21, CCW20, FRN21, LBRW22, MBCC23, ZYXC24, ZLS^{+}22]. pseudo-deterministic [CHA21]. pseudo-improper-dihedral [MBCC23]. pseudo-spectral [CCW20, FRN21, LBRW22, ZYXC24, ZLS^{+}22]. PseudoDojo [TIG^{+}24]. pseudopotentials [BTW20, NVC20, SPMS23, TIG^{+}24]. pseudospectral [RC23]. PSGen [Koł23]. PSTD [LKP21]. PSTD-FDTD [LKP21]. PSTGF [FMCB^{+}20]. public [AAHJ20, SLP^{+}22]. pulse [CML^{+}24, MD22, WMM^{+}24]. pulsed [Nie22a]. PULSEE [CNB^{+}23]. pulses [BAB^{+}20, Tan23]. PUMAS [Nie22b]. pure [SSD^{+}22]. pure-Python [SSD^{+}22]. purpose [ASPDL^{+}21, ABF^{+}23b, DCRF23, DC22, HT20, Koł22, LVMGF^{+}23, Wor20, ZWC22]. pusher [DML23, GH21, ZZ24]. PV [MFS^{+}22]. PWDFT.jl [FASD20]. PXLink [ZBS^{+}23]. py [BFT20, Kür23, PP23, BYWW23, LEE^{+}24, BGHC23, MD22]. Py-EFIT [BYWW23]. py-fmas [MD22]. PY-Nodes [PP23]. pyAPEP [GAJK23]. PYATB [JPJ^{+}23]. PyBEST [BLN^{+}21, BBC^{+}24]. PyCharge [FH22]. PyCSP [Gal22]. pyDSM [ECS23]. pyerrors [JKKN23]. PyFitit [MGG^{+}20]. PyFocus [CMS22]. PyFrac [ZL20]. Pyg4ometry [WAN^{+}22]. pyGDM [WMA^{+}22]. pyhf [AKW21]. pyHMA [MPSK21]. PyHoLo [MZD^{+}20]. PyLCP [EBNS22]. PyLlama [BVV22]. pyMCD [LKK23]. PyMikor [Bar22b]. PyOECP [YMCF23]. PyOpenCL [MR22]. PyPhotonics [TR22]. PyProcar [HTH^{+}20, LTT^{+}24]. PyR [SS21a]. Pyrough [IGL^{+}24]. pySecDec [HJJ^{+}22, HJK^{+}24]. PySPOD [RYS^{+}24]. PyStructureFactor [SZNW23]. pyTDGL [Hor23a]. PYTHIA [KL22, Sjö20]. Python [AZH^{+}24, AGMFGE23, ABGD^{+}20, AYI^{+}24, AM21a, AS24b, BYWW23, BVV22, Bie21, BGHC23, BLN^{+}21, BYL^{+}21, CZA^{+}23, CMS22, CFPS23, DCRF23, DBdFdSR21, EBNS22, ECS23, FH22, FPSZ21, Gal22, GHKW22, HS22, HHVB21, HTH^{+}20, HQF^{+}20, Hor23a, JKST22, JPJ^{+}23, JKKN23, Kar23a, KH23, KKB^{+}24, KVSC21, LEE^{+}24, LKK23, LN23, LDW^{+}21, LWV20, MD22, MRT^{+}22, NT24, PP23, Pot24, RHLTG^{+}22, RGS^{+}21, RSPJ21, SLP^{+}22, Sch21, SLB^{+}23, SP23, SSD^{+}22, SLDF^{+}21, SZNW23, TR22, TQGE23, TMH^{+}23, VCF22, WAN^{+}22, WZZ^{+}23, WBK^{+}24, WAK22, WMA^{+}22, WV22, ZCS^{+}24]. Python-based [AZH^{+}24, ABGD^{+}20, KKB^{+}24, LEE^{+}24, SLB^{+}23]. Pythonic [BLN^{+}21]. PyTorch [JS24]. pywidgets [VMRFC23]. PyWolf [MR22, MS24]. PyXtal [FPSZ21].

Q [ZWC22]. Q-POP-Thermo [ZWC22]. QBMM [LMHL20]. QCD [Dür23, EVFRHR23, FFLR20, FKK^{+}21, HLM22, IKM^{+}23, KM20, KMS20b, Kub24, LMWW24b, MBC^{+}24, WSRO24]. QCT [Dju20]. QDD [DVC^{+}22]. QDT [KMBP24]. QED [BKG^{+}23, MT23]. QEHeat [MBTB21]. qeirreps [MONW21]. QERaman [HHT^{+}24]. QISG [BPMMP24]. QMeS [PSW23]. QMeS-Derivation [PSW23]. QMMM [LZP^{+}24]. qocttools [Cas24]. QOptCraft [AGMFGE23]. QPAD [LAD^{+}21]. QPC [ZLMH23]. QPC-TDSE [ZLMH23]. Qprop [TB20]. QS [UYS22]. QSGS [YLLW24]. QSW MPI [MW21]. QTRAJ [OEI^{+}22]. quadratic [Par21]. Quadrature [LP24, GDS^{+}21, GCWZ20, HFP21, LC24, MM23]. quadrature-based [HFP21]. quadratures [ST23a]. Quantics [Wor20]. quantification [KMG^{+}20, TO21]. quantitative [BDP^{+}21, Jab22a, MMC^{+}22, MGG^{+}20]. quantities [KKSV24]. Quantum [AMK24, ASW20, AYWKL24, DHE^{+}24, HTN21, KPL^{+}21, KK20, MONW21, MYMK^{+}21, UYS22, dBBVA20, vRCM21, AGMFGE23, ADW^{+}23, ACS^{+}22, BPMMP24, BTG22, BLG24, CLY22, CNB^{+}23, CPL21, Cas24, CDD22, COW24, DH22, Dio23, GJA21, GF23, GLB^{+}21, HAM^{+}20a, Hua23, KSDH23, KMBP24, LZP^{+}24, MW21, MAJ20, MOY^{+}22, Nog21b, OV23, PM21, PC21, Pot24, RMM21, RBWD^{+}24, RRM^{+}23, SLLA22, SDL^{+}21, SZ23, SNG20, TMH^{+}23, UD24, VCHH23, WWDM20, WWJ^{+}22, Wor20, WZC^{+}24, WHSG24, XFGS24, YJLW21, ZCS^{+}24, DVC^{+}22, HTN21, HHT^{+}24]. quantumfdtd [DSSW22]. Quark [AJW^{+}21, IKM^{+}23, JWW^{+}23, Dju20]. quarkonium [OEI^{+}22, YZW22]. Quasi [KW23, LY24b, OCR^{+}22, BR20b, DBH^{+}22, LAD^{+}21, MCP23, YZHL22]. Quasi-exact [KW23]. Quasi-Helmholtz [OCR^{+}22]. Quasi-optimal [LY24b]. quasi-static [BR20b, DBH^{+}22, LAD^{+}21]. quasi-transport [YZHL22]. quasi-two [MCP23]. quasicrystals [Yan24a]. quasiharmonic [LDW^{+}21]. quasiparticle [BN20, BND22, BN23, SLZG20]. quasistationary [CF21]. QuDPy [SLB^{+}23]. quench [KCJX24]. quenching [BMR^{+}24a]. Questaal [PAL^{+}20]. QuGIT [BTG22]. quick [SCR^{+}22b]. quickly [FO20, FO21, ZYL^{+}23]. QuOCS [RRM^{+}23]. qvasp [YTC^{+}20].

Rabi [MBA21, RVM^{+}21]. Rabi-coupled [MBA21, RVM^{+}21]. Racah [DLD^{+}21]. radial [BA24, LGDF20, LM22, NTO24, SLE^{+}22]. radiation [ASPDL^{+}21, Arn20, COJ^{+}22, CZX^{+}21, DRM20, DFP^{+}20, GAGO21, LLH21a, LMMP23, MXH^{+}24, NS20, Nie20, OBGA24, PLT^{+}23, PSMRS^{+}23, RDC^{+}20, Sij23, WAN^{+}22]. radiations [Roh22]. radiative [BKG^{+}23, FXQS21, KSIL22, SJWL22]. radio [KM23, MMP20, PSMRS^{+}23]. radio-frequency [KM23, MMP20]. radioactive [PMS^{+}20a]. RadLib [SJWL22]. Raman [HHT^{+}24, KSG22]. random [AYI^{+}24, BND22, KW23, TRB^{+}20a, RBV^{+}22]. random-field [KW23]. randomized [LMHUR23]. randomly [LDGN24, YYZ^{+}22]. range [GSBN22, HG22, VXT^{+}23]. rank [CE22, CL22a, VPS23]. RANS [VLS22]. RANSBox [MOMO24]. Raphson [AG21]. Rapid [NSU20, BMI23, MOA24]. rapidly [TPCT22]. rare [DRSZ23, SJP^{+}21b, SKC21]. rare-event [SJP^{+}21b]. rarefied [CAWK22, MEC^{+}24]. rarefiedMultiphaseFoam [CAWK22]. Rate [MDDI21, FCTFR20, KFC^{+}20, SC22]. rates [HYL^{+}22, LGBJ20, MES^{+}24, MPZB^{+}24, SZNW23]. rational [AKL^{+}21, DEdM24, KL20, KKL21]. rationalization [BWW20]. RationalizeRoots [BWW20]. ray [CML^{+}24, FXQS21, GSL^{+}23, Hua24, Jab22a, MKPW21, MKPW22, Mau20, SHS22, SPF21, SBG23, WWM^{+}22, WKJB23, sXBkB^{+}22]. ray-tracing [FXQS21, SHS22]. Rayleigh [LY24a]. rays [MNS^{+}24]. RC [BKG^{+}23].

RDM [PFG22]. Re [YC20, OS24]. re-orientation [OS24]. Re-splitting [YC20]. reacting [Gal22, MLZ^{+}23, NJSY22, RCX24, TPK^{+}21]. reaction [AYB24, BB24b, FLW^{+}23, GMZ^{+}20, MPZB^{+}24, PTH24, XDF20]. reaction-diffusion [BB24b]. reaction-diffusion-wave [GMZ^{+}20]. reactions [AHM^{+}23, OYC24, SW23]. reactive [AS22, ADW^{+}23, CNS22, GP22, MPH^{+}24, MRG22]. reactive-flow [CNS22]. reactivities [Xie23]. reactor [AG23, LLC^{+}23, SBZ23, WLL22, XLL^{+}22, Zho23]. reactor-scale [SBZ23]. reactors [TBAR21]. ready [SMB20]. Real [MVK^{+}24, NJSY22, WKJB23, BW23, DVC^{+}22, DHE^{+}24, HLMB24, LKP24, MSHP02, MSHP20, OSE^{+}20, SLZY21, SCF20]. Real-fluid [NJSY22]. real-space [BW23, DVC^{+}22, LKP24, SLZY21]. Real-Time [MVK^{+}24, WKJB23, DVC^{+}22, DHE^{+}24, HLMB24, SCF20]. realistic [CMS22, KGN^{+}21, YCC22]. realizable [PLF20]. Receiver [CTPS22]. reciprocity [WXZH24]. Reconnoitering [LSZ23]. Reconstructing [KL20, YYZ^{+}22]. Reconstruction [DBV^{+}24, OBK^{+}20, AAB^{+}21, AWV22, ADdM20, ABF^{+}23b, BYWW23, BCSS24, GOS^{+}22, LNP^{+}24, PMS^{+}20a, RCW^{+}20, SWZ23, SS21b, TGGC23, TRN^{+}24, TWW22, VLS22, YLK21, BBV^{+}22]. record [BIK^{+}21]. recording [RMR^{+}22]. recreation [CEC^{+}24]. rectangular [MSH22]. recurrences [SRML20]. recursive [EGKH24, XFH^{+}22]. redistribution [RCS21, YGSW21]. reduced [BGR23, HCAH^{+}24, PPR^{+}21, YYH21, ZSqXY21]. reduced-space [YYH21]. Reducing [GJA21, LWS^{+}23]. reduction [CHA21, KLMU21, PLT^{+}23, RUV20, SZ24b, VHBK21, SC20]. reductions [GHL23]. redundant [CZB^{+}23]. refined [LW24]. Refinement [OSLC21, ASC^{+}24, FAL20, FAL21, PG23]. reflection [Arn20, CGV^{+}22, Dan24, HMYH22, HSO^{+}22]. reflective [KYH24]. reformulated [LSW^{+}20]. reformulations [HWZ24]. refractive [BMR^{+}24b]. Regge [AS22]. regime [BMT23]. regimes [LSW^{+}20, MMFdL21, YZ20b]. region [HJGL22, LJH^{+}23, Yan21a]. regions [ADH^{+}20, ABFP24, HJJ^{+}22]. Regression [RBV^{+}22, LM22]. regressors [KHKL24]. regularization [SKS24]. regularized [KRE22]. reinforced [SRK^{+}24]. reinforcement [KKY24]. related [KLM^{+}22, MMP^{+}24, WW21, WGGC22]. relation [LLY^{+}22, LB24, MEC^{+}24]. relations [WBM^{+}24]. relationships [ÖAÖ24]. Relativistic [KKLZ23b, SLE^{+}22, Bag22, BN20, BN23, BAB^{+}20, DTC20, DSSW22, DH20, GK21, KKLZ23a, Kul20, LMX^{+}21, LMMP23, MGC^{+}23, MHP23, Nie20, UZB22, YZL^{+}23, YNMR24, YKY^{+}22, ZZ24]. relativity [CFPS23, Tor20]. Relaxation [PGS^{+}24, NNMJ22]. relaxed [ZPZH20]. release [FGCN^{+}21]. relevant [YGSW21]. reliability [AM21b, LTT^{+}24]. reliable [LBM^{+}23, ZWC^{+}20]. ReMKiT1D [MPH^{+}24]. removal [GBH20, Hor24]. ReneSANCe [BDK^{+}23, SY20]. ReneSANCe-v1.0.0 [SY20]. ReneSANCe-v1.3.0 [BDK^{+}23]. Renormalisation [LS21b, DH22, IPSW24]. renormalization [CFBRE24, CPL21, JS24, YS22]. reordering [Ume22]. replica [RCS21, XLKX21, YLIO22]. replica-permutation [YLIO22]. Representation [BFL^{+}22, FKS20, RBV^{+}22, CLS^{+}22a, Jab24, KCS22, SXYD24]. representations [ABFP24, GWPW21, Gon22, IMB^{+}22, LCZ^{+}21, MONW21]. representing [MM23, RBV^{+}22]. Reproducibility [RDV^{+}20]. repulsion [CXCZ23]. resampling [MBE^{+}21]. research [AMA^{+}20, Di 22, DP21]. resistive [ZJM^{+}21]. resolution [EGLK20, LU21, SLL22, ZRZ^{+}21]. Resolved [WS20, CHY^{+}24, LR22]. resolving [GSL24, GB22]. Resonance [BMR^{+}23, AS22, CNB^{+}23, FYM^{+}22, HRU22, HHT^{+}24, WXX24]. resonances [GTMB21]. resonators [Hoh24, JYL^{+}23, WAYL23]. RESPACK [KMY^{+}23, NYN^{+}21]. response [ACD^{+}22, BM22, BND22, SLZY21]. responses [SLB^{+}23]. REST [ACD^{+}22]. REST-for-Physics [ACD^{+}22]. restricted [YM21]. restructuring [TDR^{+}20]. results [KOT23]. resummation [ABB^{+}22]. Rethinking [SWZ23]. reticulated [SFKC22]. retQSS [SFKC22]. reuse [CCC20]. reverse [BRAC23, HT24]. reversible [KP23]. revertible [Nie23]. revised [SLL^{+}24, ZZC20]. Revisiting [YZL^{+}23]. revolution [NTO24]. revolutions [GdSPC22]. REvolver [HLM22]. reweighting [Kub24, SJP^{+}21b]. Reynolds [MD21, MOMO24]. Reynolds-averaged [MOMO24]. RF [BGW^{+}22, SS22]. RF-transpond [BGW^{+}22]. RGB [GOS^{+}22]. RGE [DH22]. RHDLPP [MXH^{+}24]. RHEED [DZZ21, Dan23, HMYH22]. rheology [ECS23]. rheometer [SPTPR21]. RHMC [Kub24]. RI [BMR^{+}24b]. RI-Calc [BMR^{+}24b]. rich [ABK^{+}22, EES24]. RichardsFoam [Org22]. RichardsFoam3 [Org22]. Riemann [CK20b, LMMP23]. Riemann-Theta [CK20b]. Rigid [OAP^{+}24, SXYD24]. Rigidity [HS22]. rigidPy [HS22]. rigorous [MZL^{+}21]. RMC [LHWX24]. RMT [BAB^{+}20]. Robust [AACE^{+}21, SBP20, IUJ21]. ROC2.wl [ABFP24]. rocking [DZZ21]. rocks [XBL^{+}20]. Rode [MMC^{+}21]. rogue [WXY20]. ROOT [ACD^{+}22, TPS^{+}24]. ROOT-based [ACD^{+}22, TPS^{+}24]. roots [BWW20]. Rosenbluth [KSJ^{+}22]. RoseNNa [BB24a]. rotating [BJL^{+}24, GTA21, MBA21, TPCT22]. rotation [CCW20, CWJ21, LZ21, ZYXC24]. rotation-two-component [ZYXC24]. rotational [BGHC23, GTMB21, WGS^{+}22, dVAR^{+}24]. rotational-vibrational [GTMB21]. rough [IGL^{+}24, YYZ^{+}22]. routine [MN21]. routines [HW22, KOT23, SSH^{+}23]. RPCs [DBM^{+}24]. RS [RBV^{+}22]. RS-HDMR-GPR [RBV^{+}22]. RTGW2020 [PWD22]. rules [CS22b, DPR^{+}20, DRR^{+}24, Nog21b, ZHS^{+}20]. runaway [DTC20, HEF21]. Runge [KBSL22, KFC^{+}20]. running [HLM22, WLZ^{+}24]. Rutherford [LGK^{+}22]. Rydberg [MMV^{+}24]. RydIQule [MMV^{+}24].

S [CCK23, VMRFC23, LW24, MZV23, YK21a]. S-FDTD [LW24]. S-SAV [YK21a]. S3D [MEH21]. SAFARI [JS22]. safety [KJL^{+}23]. Salpeter [ZAW^{+}21]. samples [IGL^{+}24, ZDLS21]. Sampling [CK20b, KvH21, RBV^{+}22, AM21b, Gon22, HSMR^{+}24, SJP^{+}21b, TCcN23, XLKX21]. SANCphot [BKS24b]. SANTA [KJL^{+}23]. São [CCG21]. Sar [ST23b]. Sarkas [SSD^{+}22]. SAT [MKHT20]. satellite [GG24]. satellites [SCMP^{+}22, SCR^{+}22b]. SATLAS2 [GvdBdGN24]. saturated [HRG^{+}22]. saturated/unsaturated [HRG^{+}22]. saturation [UYS22]. SAV [GZW20, WYT23, YK21a]. Savart [SSH^{+}23]. SBETHE [SA23]. scalability [BGR23]. Scalable [LYC20, PPKK21, DTC20, DC22, HAA^{+}20, KKPC21, MAJ20, OCK^{+}24, OLNG21, RLW^{+}24, YCC22]. scalar [ABH^{+}23, BBH^{+}23, CS22b, FFTV23, MCMS24, WYT23, YK21a]. scalar-tensor [MCMS24]. scale [BSC^{+}21, BC22, BDR^{+}20, CB23, CGSO20, CTZW23, EVFRHR23, FJ22, FWL^{+}24a, GSL24, GB22, HXS20, HRG^{+}22, HSA22, KCJX24, KKPC23, KFHR24, LBRW22, LKP21, LZK^{+}23, LG23, LU21, MDR^{+}20, OSK^{+}21, PLT^{+}23, RL21, SBZ23, TPCT22, VTB^{+}21, WA21, YCC22, YJLW21, YGSW21, ZMC23]. scale-resolving [GSL24, GB22]. scale-separated [TPCT22]. scaled [BR20a, YJ24]. scales [TAB^{+}22]. scaling [BR20a, CXCZ23, RDZ^{+}20, SLZY21, Zeb22]. scan [SZ24a]. scanning [AUO^{+}22, ZW20]. scans [GJ24]. scatterer [WZZS21]. Scattering [AB21, AZH^{+}24, ADW^{+}23, BFCR24, BDA^{+}20, BTW20, BKG^{+}23, GTMB21, GJJN23, Hal21, HYL^{+}22, HJK^{+}24, HBM^{+}24, Jab22b, Jab24, JS22, KC21, KRL^{+}24, MBJB24, PBK21, RCB21, Roh22, SJP21a, SMB20, YYZ^{+}22]. SCELib4.0 [SMO^{+}20a]. SCELT [MG22]. scheme [ADdM20, CZY20, FN21, Gai20, GDS^{+}21, GMZ^{+}20, Kul20, KBSH20, LF24, LSW^{+}20, LHC20, LHG24, LLH^{+}21b, MM23, MA24, MRN20, PLF20, PGYF21, Sij23, Tan23, Tan24, THH21, WJB21, WDMZ24, XZLX20, Yan21b, YZ20b, ZCWY20, ZWP^{+}22, ZPL^{+}24, KWZ24]. schemes [CWJ21, GZW20, KBSL22, LM21a, LW24, SP20, SBZ23, ZX23]. Schrieffer [PTD20]. Schrödinger [BV21, BCGT24, BAB^{+}20, DSSW22, GTA21, GSLS20, GGCW24, Koz23, LY24b, Pöt20, RS20, SGS23, SP20, SLZ^{+}24, Scr22a, SCL22, WXY20]. Schrödinger-like [Scr22a]. Schrödinger-type [Pöt20]. science [DBBP23, HJM^{+}20, XLG^{+}23]. sciences [BMREC21]. scientific [PMA^{+}21, PP21, WICA22]. scintillating [AAB^{+}21]. Scott [Haz23, JLW24]. scrape [MPH^{+}24, MAMK21]. scrape-off [MPH^{+}24, MAMK21]. scripts [HBM^{+}24]. SCTBEM [YJ24]. SDE [MNS^{+}24]. sdfibm [Zha20]. sDMD [ZJS^{+}20]. Seamless [ANU21]. search [BFD22, GUAD22, LS23, VZ20, YK21b, ZHS^{+}20]. searches [WBK^{+}24]. searching [Gai20, LKK23, PP23, XSM22]. Second [KSF^{+}22, AIZ23, DEdM24, KBB21, KD23, LM21a, PLF20, SKEZ24a, WJB21, Yan21b, ZCWY20, ZZZ^{+}20]. second-degree [WJB21]. second-harmonic [SKEZ24a]. Second-order [KSF^{+}22, KBB21, KD23, LM21a, PLF20, Yan21b, ZZZ^{+}20]. secondary [KOT23, NG21]. section [Jab24, PYT^{+}24, SY20, ST23b, WLL22]. sections [AS22, AYB24, BFMA^{+}24, Hor24, SLLA22]. sector [BHH^{+}20, BP21]. sectors [KKM^{+}20, BMM21]. seeding [AAB^{+}21]. seeker [NA20]. segmented [KTF22]. seismic [RRC^{+}24, Yan24b]. SeismicNet [RRC^{+}24]. selection [GSBN22, WSRO24, ZLV23]. selective [UJ21]. Self [Kut20, LDGN24, LM20, vRCM21, BN20, BN23, FH22, WLF^{+}24, WGS^{+}22]. self-avoiding [WGS^{+}22]. Self-Consistent [vRCM21, Kut20, BN20, BN23, FH22]. Self-force [LM20]. self-interstitial [WLF^{+}24]. Self-optimization [LDGN24]. Semi [Ein20, Mau20, BTK24, BAB^{+}20, BKG^{+}23, HPAW21, KBB21, KVSC21, MMCC^{+}22, PLT^{+}23, RRC^{+}24, SN23, SRE^{+}24, TKS22, MZV23]. semi-analytic [PLT^{+}23]. Semi-analytical [Mau20, MZV23]. semi-bounded [HPAW21]. semi-empirical [KVSC21, MMCC^{+}22]. semi-implicit [BTK24, KBB21, SN23, TKS22]. semi-inclusive [BKG^{+}23]. semi-infinite [RRC^{+}24]. Semi-Lagrangian [Ein20, SRE^{+}24]. semi-relativistic [BAB^{+}20]. semiclassical [GLB^{+}21]. semiconductor [BMI23]. semiconductor-liquid [BMI23]. semiconductors [AM21a, DIAA21, HM24]. semiflexible [WGS^{+}22]. semimetals [SLBR22]. semitransparent [FXQS21]. sensitivity [FA21, HSO^{+}22, PCS^{+}23, SMC^{+}22]. sensors [MMV^{+}24]. separated [TPCT22]. separations [GDK21]. September [Ano20y, Ano21x, Ano22x, Ano23x, Ano24r]. Seq [WYZZ23]. Seq-SVF [WYZZ23]. sequential [Par21, Yan23]. serial [KFHR24]. Series [Jab24, ABFP24, ABD^{+}23, EBBB22, Hid21]. Serpent [Kal20]. server [BMR^{+}23, BMR^{+}24b]. SESCA [NG21]. set [CNB^{+}23, CS22b, FA20b, FN23, HWAA22, JZW^{+}22, XHY^{+}24, YYH21, YJLW21, YZZ^{+}23]. SFQEDtoolkit [MT23]. Shafranov [AIZ23, ELSV24, HSB^{+}24, LZ21, SVSC20]. shallow [DS22, STA20]. Sham [JZW^{+}22, LG23, YT23]. SHAPE [ANU21, PCL24, WYHW24, ZTV^{+}24, ANU21]. shaped [LNB23]. shared [EVMP20, MLT^{+}21, ZWZ^{+}22]. shared-memory [EVMP20, MLT^{+}21]. sharp [HWAA22]. sharp-interface [HWAA22]. Shaw [Yan21b]. shear [MES^{+}24, SPTPR21]. Sheared [FMHH24]. sheaths [KM23]. Shell [MAJ20]. ShengBTE [HYL^{+}22, MPN^{+}21]. shield [COJ^{+}22]. shielding [CZX^{+}21]. shift [GRCT20, PBK21, WSRO24]. shifted [HKY^{+}21]. shifts [SOH^{+}23]. shock [LF24, MDZ24]. shock-capturing [LF24]. shooting [JM24]. short [CML^{+}24, GSBN22, VXT^{+}23]. short-range [GSBN22, VXT^{+}23]. shortcut [ZHS^{+}20]. Shortest [KDHL23]. SHORYUKEN [CSW24]. shot [TCcN23]. shot-noise [TCcN23]. Shower [LAC21, BP21, Roh22]. Showers [ABF^{+}23b]. SIDES [BDA^{+}20]. SIDIS [BKG^{+}23]. SIDIS-RC [BKG^{+}23]. SIESTA [Soz21]. signal [ADH^{+}20, ABF^{+}23b, MD22]. signals [ET24, JKST22]. signed [TGBM22, Zha20]. significance [CTPS22]. significant [ZHZ^{+}23]. silver [SKS24]. Sim [BMP^{+}24, BMR^{+}24a, HMYH22]. sim-trhepd-rheed [HMYH22]. Simflowny [PMA^{+}21]. similarity [BBV23, MRT^{+}22]. Simple [CF21, GJ24, LXY^{+}21, WXY20, AES21, GFH23, LBM^{+}23, MWJL23, MBC^{+}24, PC21, Sat21, WT22, Xie23, ZZZ^{+}20, ASW20]. SimpleBounce [Sat21]. simplex [DBV^{+}24, PP23]. simplification [ANU21, Gal22, SZ24b]. simplified [CLEPF23]. Simplifier [LS21b]. simplifying [XWYQ21]. simulate [BFI^{+}21, DBM^{+}24, GSBN22, LGK^{+}22, MD20b, SHS22, ZWZ^{+}22]. simulated [MMFdL21, TRN^{+}24]. SIMULATeQCD [MBC^{+}24]. Simulating [BLG24, DG20, Gar21, PMSHG23, BMR^{+}24a, BFT20, BM22, MHÅ21, MR22, MP21, PD23, Sch21, SCL22, SJY18, SJY20, TLC^{+}21, TMH^{+}23, XFH^{+}22, ZYML24, ZCS^{+}24]. Simulation [CMJC21, CJ21, EMM^{+}23, KHR^{+}23, RMM21, RdSH^{+}24, WXW^{+}21, ABGD^{+}20, ARLDG24, ASPDL^{+}21, AFB^{+}24b, ASC^{+}21, Bak23, BOSM24, BCF^{+}24, BMP^{+}24, BMSP21, BB24b, BKO20, BMT^{+}20, BLM^{+}22, BKS24b, BCD^{+}21, BYL^{+}21, BKBL22, CNB^{+}23, CCM20, CE22, CLZ^{+}21, CSBF23, DML23, DBR24, Ein20, EOR21, FTZ^{+}23, FGCN^{+}21, FN23, FYM^{+}22, GLW^{+}23, GWA^{+}23, GBD^{+}22, GB20, GP22, HT20, HMSV23, HSA22, JS22, JĎD^{+}21, KTF22, Kan23, KF23, KMBP24, KGT22, Kul20, KNJ^{+}23, LBS^{+}23, LKP21, LLT^{+}23, LLC^{+}23, LZK^{+}23, LY24a, LF24, LLL24, LJQ^{+}22, LWC^{+}21, MES^{+}24, MW21, MHK24, MAMK21, MRG22, MDZ24, NS20, NSY^{+}23, Nie20, Nie22a, NRG22, OBC^{+}24, OV23, PMA^{+}21, PK24, PFG22, QCZ23, RLW^{+}24, RSD20, RDH^{+}20, RMCC22, RCB21, RS20, SRT^{+}20, SD24, SRC21, SFKC22, SNP^{+}20, SJ20, SKYQ21, SS22, SLIC24, SPF21]. simulation [STRF^{+}20, SWTC23, Tan24, TL20, TAB^{+}22, TCcN23, eSdSBST21, VCHH23, WICA22, WLF^{+}24, WAK22, WLF^{+}23, XBL^{+}20, XHY^{+}24, YCC22, YYC^{+}23, YSX^{+}20, YC20, YR21, ZMC23, ZK20, ZHI23, ZBS^{+}23, ZZ21a, ZYX21, ZYX22, ZJS^{+}20, ZGW20, dSOZ22, SGM^{+}20]. simulations [AHP21, AALK20, BL21, BTK24, BCTS22, BR20b, BC22, BGR23, CLY22, CB23, CGSO20, CTZW23, CZTF23, CHY^{+}24, CF21, Cos22, CJD^{+}20, CNS22, Dan23, DTGE21, DWD^{+}24, DAA^{+}24, DKM^{+}24, FN21, FDPT23, FRN^{+}23, FFTV23, FH22, FFLR20, GAJK23, GSL24, Gon22, GB22, GAS^{+}23, GMZ^{+}20, Hal21, Ham20b, HTL^{+}22, HyLF23, HEF21, IGL^{+}24, JGJ20, KPR^{+}24, KWZ24, KCJX24, KBSL22, KSJ^{+}22, KP23, KGN^{+}21, KPL^{+}21, KMG^{+}20, KYH24, LBRW22, LBM^{+}23, LAD^{+}21, LZP^{+}24, LM20, MMM23, MKPW21, MKPW22, MLZ^{+}23, MPS^{+}24, MBH21, MEH21, MVF20, MBE^{+}21, NJSY22, Nor23, OCK^{+}24, OGL^{+}21, OCC23, OBGA24, OSK^{+}21, PCS^{+}20, RSC^{+}22, RCB^{+}20, RSHS24, RdPS24, RL21, SLL22, SJ21, SDXY23, Sij23, SBZ23, SWB^{+}23, TKC^{+}21, TSJ^{+}21, TPK^{+}21, TNL^{+}22, Ume22, VPPQ21, VXT^{+}23, VV21, VGGP^{+}21, VSM^{+}22, VvBTH20, WWM^{+}22, WDMZ24, WA21]. simulations [WOP^{+}20, WMA^{+}22, WS20, WWJ^{+}22, Wor20, WRM^{+}24, WS23, YFL22, YJLW21, YTC^{+}20, YWX^{+}23, ZDSS23, ZLL^{+}24, ZSqXY21]. Simulator [AACE^{+}21, BSG^{+}21, FCTFR20, HMYH22, HLMB24, MFS^{+}22, SKEZ24b, SGY24, XFGS24, ZTV^{+}24, ZL20]. simultaneous [KF23]. simultaneously [TQGE23]. Single [DFG^{+}23, PAZ^{+}22, BW23, CFLR24, CL20, LG21a, PNL^{+}24, SSPF24, SHRK22, SW23, TWR21, SMO^{+}20a]. single-file [LG21a]. single-grid [BW23]. single-instruction [PNL^{+}24]. single-ion [TWR21]. single-molecule [PNL^{+}24]. single-nucleon [SW23]. single-phase [CFLR24]. Single-step [PAZ^{+}22]. Singular [Gal22, KFHR24, LP24, WT22, BBH^{+}24]. Singular/GPI [BBH^{+}24].

Singular/GPI-Space [BBH^{+}24]. SIRF [OBK^{+}20]. site [CXCZ23, ZLV23]. sites [HHMH^{+}22]. six [SRE^{+}24]. size [HFP21, KAG24, PLSB22, WBM^{+}24]. size-velocity [HFP21]. sized [Che23]. skew [LF24]. skew-symmetric-splitting [LF24]. SKMF [GTE21]. Sky3D [ASS^{+}24]. SkyAx [RSM21]. Skyrme [CLS^{+}22a, MSN^{+}22]. skyrmion [IUJ21]. slabs [MCP23]. SLC [AJW^{+}21]. slender [LBS^{+}23]. SLEPc [DNG^{+}20]. SLIMP [RZY^{+}24]. slip [SN23]. SLJCompact [BMI23]. slope [NLS24]. slope-dependent [NLS24]. Slow [XQ21, LJQ^{+}22]. small [LLC^{+}23, ST23b, WBM^{+}24, ZLMH23]. small-size [WBM^{+}24]. SMATool [Eku24]. SmeftFR [DPR^{+}20, DRR^{+}24]. Smirnov [Efr21b]. SMIwiz [Yan24b]. SModelS [AKW21, ADH^{+}20]. Smoothed [AMA^{+}20, JCM20, AYWKL24, FRN21, KP23, LMQ^{+}23, LYZL24, ODR^{+}22, RZH22, ZRZ^{+}21, ZHZG23]. smoother [AGJ^{+}23]. smoothing [SL22]. SOD2D [GSL24]. Soft [SPMS23, TNL^{+}22, VRI24, MOA24]. SOFTSUSY [AAT^{+}20, AAT^{+}14]. Software [BWW20, CTPS22, GBR23, KvdW20, MHÅ21, Nis22, Sha21, AMP^{+}21, AAG^{+}24, AAD^{+}24, ACKB23, ASC^{+}21, BMR^{+}23, BMR^{+}24b, BMT^{+}20, CNB^{+}23, COW24, CSW24, CW22, DvHSdS22, DIKSN24, DKRSR22, EGKH24, FLK^{+}20, GAJK23, GO23, GPM^{+}21, HAM^{+}20a, HGS20, KWK^{+}21, KSG22, KDIN^{+}23, LEE^{+}24, MMP^{+}24, MGG^{+}20, MZL^{+}21, MZD^{+}20, MOMO24, MYM^{+}22, NKP20, PAM24, PRR23, RDH^{+}20, RHW^{+}21, SHB^{+}20, SHS22, SJY18, SJY20, TL20, TO21, VB22a, WV22, WS23, YKSH20, YR21, YMCF23, ZPL^{+}21, ZCS^{+}24, ZLV23, vdHKB^{+}23, vTDGCR21]. SOL [MAMK21, YHH^{+}20]. SOL/divertor [YHH^{+}20]. solar [BMP^{+}24, MFS^{+}22]. solar-cell [MFS^{+}22]. Solid [PTD20, AMP^{+}21, ONH^{+}20, PB23, RSD20, SRT^{+}20, UO20, WXZH24, XZLX20]. solid-fluid [WXZH24]. solid-liquid [RSD20, SRT^{+}20]. solidification [SYFT23, ZDG^{+}21]. solids [BR20b, BFI^{+}21, DSQ23, FXQS21, GBR23, KRL^{+}24, Och23a, Och23b, RDV^{+}20, WGGC22]. Soliton [HWZ24]. solitons [GMZ^{+}20, ZYXC24]. solute [HRG^{+}22]. Solution [AIZ23, DvHSdS22, AE23, BBA^{+}20, CKT21, CLEPF23, CLEP24, Flo24, GSLS20, GQ22, LF20, LSS24, MSN^{+}22, MPN^{+}21, MNS^{+}24, OKBM23, PPR^{+}21, QWZ^{+}21, YNMR24, CLEPF22]. solutions [AMP^{+}21, Bak23, GF23, LWhK^{+}20, LHTP^{+}24, NI22, SSB^{+}23, SCL22, VFS23, YZ20b]. solvation [DCZ23]. solve [BKRG22, CLEPF23, DH22, HAM^{+}20a, WGG20]. Solver [ASW20, GKIB21, MYMK^{+}21, TCSA21, AFB^{+}24a, ADF^{+}22, BMSP21, BMS^{+}23, BMI23, BBA23, BSC^{+}21, CLS^{+}22a, Cos22, DTC20, DAC^{+}23, DS22, DFP^{+}20, DBV^{+}24, DP21, DdCAG23, EGLK20, FXQS21, FTZ^{+}23, FRN^{+}23, FMBD22, FA21, GTA21, GSL^{+}23, GDJG23, GWA^{+}23, GLPG^{+}23, GHL^{+}22, HPAW21, HXS20, HWAA22, IKM^{+}23, JMOC21, JKSY22, Kar22, KR23, KKPC23, KGN^{+}21, KFPV21, KG21, KD23, LMHL20, LMX^{+}21, LIG^{+}24, MMP20, MEC^{+}24, MEDT^{+}23, MSY^{+}21, MKHT20, MOY^{+}22, MBA21, MZV23, OEI^{+}22, OXOG23, OCE^{+}23, PTH24, PD23, RVM^{+}21, Di 22, RJ21, RCW^{+}20, RAJ^{+}24, STA20, SGS23, SSPF24, SLC^{+}22, SBB^{+}24, SWTC23, UYS22, VLS22, VV21, WWB22, WHB21, WWZ^{+}23, WKR23, WHSG24, XOTI22, XLG^{+}23, Yan21a, YFL22, YCCW23, YKK23b, ZLMH23, ZWC22, ŻTR^{+}22, GMPG^{+}21, KOF21, WHSK24, DFU20]. Solvers [JRS^{+}21, BW23, CB23, EVMP20, EPM23, FA20a, FA20b, KKPC21, LMMP23, OIA^{+}20, YSI^{+}24, zYCD^{+}20, YKY^{+}22]. Solving [BAB^{+}20, LZ21, ZAW^{+}21, ZGW20, ASU^{+}21, ABBD23, ATC^{+}23, BJL^{+}24, BKP22, BDR^{+}20, HWL^{+}23, JZW^{+}22, KRG21, KSF^{+}22, LQ21, Man24, MM22, Scr22a, TS23, XDF20, XSL^{+}22, YZW21, YKK^{+}23a, YSMBA23, YJ24, ZLY24]. SOM [KM22]. SOMAFOAM [VV21]. Sommerfeld [Arn20]. Sommerfeld-type [Arn20]. sonic [EOR21]. sonic-boom [EOR21]. SootImage [CEC^{+}24]. sooting [CEC^{+}24]. source [AFB^{+}24a, ASU^{+}21, ADF^{+}22, ASC^{+}21, BC20, BM22, BLN^{+}21, BMT^{+}20, BFI^{+}21, BSC^{+}21, CAWK22, CZS^{+}21, COW24, CSW24, CW22, CKGW22, DIK^{+}23, DFU20, DAA^{+}24, FTZ^{+}23, FH22, FYM^{+}22, GDK21, GDJG23, GKT^{+}24, GFD^{+}24, HMYH22, HSB^{+}24, HXS20, HWL^{+}23, HLMB24, HRG^{+}22, HKY^{+}21, Hua23, HHT^{+}24, KGN^{+}21, KGT22, Koz23, KD23, LEE^{+}24, LMHL20, LH22, MLZ^{+}23, MBTB21, MMP^{+}24, MFB23, MONW21, MVF20, MVAXP22, MRN20, NBC^{+}24, Nor23, PMS^{+}20a, DARJ23, PM21, PWC24, RZWW23, RHW^{+}21, RJ21, SPLD20, SHB^{+}20, SHW^{+}21, SWTC23, THH21, TPK^{+}21, WNS^{+}21, WSK^{+}22, XBL^{+}20, XOTI22, YKSH20, YMCF23, YKK23b, ZRZ^{+}21, ZJS^{+}20, ZCS^{+}24]. sources [COJ^{+}22]. Space [MYMK^{+}21, BW23, CLS^{+}22a, DBM^{+}24, DVC^{+}22, GJ24, GMZ^{+}20, HHT22, JLL^{+}24, KKLZ23a, KKLZ23b, Koł23, LKP24, LCZ^{+}21, LZYY23, LR22, MS24, MMP20, MSHP02, MSHP20, NBB^{+}21, SLZY21, SZ24a, TCSA21, XHY^{+}24, YYH21, ZW20, ZWP^{+}22, YKY^{+}22, BBH^{+}24]. space- [LR22]. space-charge [DBM^{+}24, KKLZ23a, KKLZ23b]. space-fractional [JLL^{+}24]. SpaceGroupIrep [LCZ^{+}21]. spaces [HPRS23, LH24, SKM^{+}21, Zeb22]. sparse [DGM20, DFG^{+}23, KKL21, NDFL24]. SPARTACUS [FTZ^{+}23]. Spatial [HJGL22, BC21, MM23, XPF^{+}24]. spatial-grid [BC21]. spatially [CHY^{+}24, KFHR24]. spatiotemporal [GPN20, PSMRS^{+}23]. SPEC [MPQ^{+}22]. Special [GH21, GM24, LVMGF^{+}23]. species [CJD^{+}20, KDK23, SNP^{+}20, WV22, YKYK23]. specific [SDBS24, ZDG^{+}21]. spectra [CML^{+}24, GBD^{+}22, Hua24, HHT^{+}24, KSG22, LGK^{+}22, MGG^{+}20, Pos22, RZY^{+}24, RMFB23, TB20]. Spectral [CdBMdAS^{+}21, CZS^{+}21, CCW20, FRN21, GTA21, LBRW22, LZ21, LS22, MPS^{+}24, MCB^{+}20, NI22, RYS^{+}24, SL22, SLZG20, SLL22, SWZ23, WLL22, WKR23, ZYXC24, ZLS^{+}22, GSL24]. spectral-element [MPS^{+}24]. spectral-Galerkin [GMZ^{+}20]. spectral/hp [MCB^{+}20]. spectrometer [SPF21]. spectroscopic [CK23]. spectroscopy [Jab22a, KMBP24]. spectrum [KMU^{+}23, NS20, SKS24]. SpectrumSDT [GTMB21]. specular [Dan24]. speed [BL21, CKGW22, DBV^{+}24, KMN21, LHG24, OCC23]. Speedup [BR20a]. spent [KJL^{+}23]. SPH [CFLR24, HMR22, MEDT^{+}23, MDZ24, OSK^{+}21]. SPH-based [MEDT^{+}23]. SPHERA [AMA^{+}20]. sphere [YYH21]. spheres [KAG24]. spherical [CCK23, DHK^{+}21, GDB10, GCK21, LEE^{+}21a, LQGL21, PCL24, QCZ23, SDBS24, TKC^{+}21, ZZ21a, dMMLOS20]. spheroidal [SSB^{+}23]. spheropolyhedral [QCZ23]. spheropolyhedral-based [QCZ23]. SPHinXsys [ZRZ^{+}21]. Spin [RVM^{+}21, BKRG22, BPMMP24, BRHT21, BCSS24, CZTF23, CL22b, FSMM24, GF23, GOCSS^{+}23, GBD^{+}22, IUJ21, JRG21, KRG21, KL22, KL23a, LL23, LR22, MBA21, QJ21, SZ23, TMH^{+}23, UYS22, VCHH23, Yev21, VMRFC23]. spin- [BKRG22]. Spin-1 [RVM^{+}21, KRG21, MBA21]. spin-boson [SZ23]. spin-glass [JRG21]. spin-half [TMH^{+}23]. spin-one [TMH^{+}23]. spin-orbit [BKRG22, BRHT21, GOCSS^{+}23, KRG21]. spin-orbit- [MBA21, RVM^{+}21]. spin-resolved [LR22]. spin-wave [GBD^{+}22]. Spinney [AM21a]. spinning [OS24]. spinor [SCL22]. Spinsim [TMH^{+}23]. SPIRAL [dVAR^{+}24]. SPIRED [BCSS24]. spline [WGG20]. Splines [KWK^{+}21]. split [SAC^{+}21, YYY22, ZYXC24]. splitting [BCGT24, DWD^{+}24, LF24, Yan24a, YC20]. SporTran [EBBB22]. SPP2 [CCG21]. square [BRAC23, BWW20, GST21]. squared [PAM24]. SQuIDS [ASW20]. SRC [COJ^{+}22]. SRC-UNED [COJ^{+}22]. stability [FFLR20, PCS^{+}23, SKDH24, SFC20, TQGE23, WHSK24]. Stabilized [AGH21, YK21a]. stabilized-scalar [YK21a]. stable [BVV22, GZW20, HSA22, KSF^{+}22, LWhK^{+}20, LM21a, PRR23, TM24, YWTK23, Yan24a, ZCWY20, ZZ24]. staggered [ZLS^{+}22]. Stan [Haz23]. standalone [AAB^{+}21, RL21]. Standard [Bie21, AFJ^{+}23, KHKL24, Kar23b, KKSV24, DPR^{+}20, DRR^{+}24, DDM20, HQRR20, UMA21]. standardized [BMREC21]. stars [VGGP^{+}21]. state [AS22, ABWZ23, CR20, ENK24, GO23, KMM21, KM20, KMS20b, KW23, LWS^{+}23, PTD20, SKM^{+}21, Xav22, YZ20b]. state-of-the-art [KM20, KMS20b, SKM^{+}21]. state-to-state [AS22]. states [BFD22, BTG22, CF21, DDM20, GTA21, GF21, GWPW21, GTMB21, LKK23, LU21, LR22, MDP22, PTD20, QJ21, SLZG20, TCSD24, WZC^{+}24, XFGS24, YLL^{+}22]. static [BR20b, DBH^{+}22, JRS^{+}21, LAD^{+}21]. Stationary [GTA21, MDZ24, VZ20]. statistical [BKS22, Gon22, MZL^{+}21, MNS^{+}24, Ols23]. steady [HJGL22, LJH^{+}23, YYY22, YZ20b]. stellarator [HK24, KBH^{+}24]. step [AYB24, HPY21, MWJL23, PAZ^{+}22, SL22, ACKB23]. stepping [CJZ21, NBCMH20, WJB21, ZX23]. steric [DCZ23]. SternheimerGW [SLZG20]. stiffness [XSL^{+}22]. stochastic [CHY^{+}24, FJ22, FTZ^{+}23, KM22, MW21, RdPS24, WLF^{+}23]. stochasticity [GTE21]. Stokes [WHSK24, ZX23, BMS^{+}23, DAC^{+}23, HPY21, HWL^{+}23, JM24, KSF^{+}22, LGDF20, LB24, MOMO24, PLL24, WT22]. Stopping [SA23, HHMH^{+}22]. strain [Che23, LLZ^{+}22, XZLX20]. strain-stress [LLZ^{+}22]. Strategies [MBE^{+}21, AWV22, CZX^{+}21, KMM21, VXT^{+}23]. strategy [TCSA21, WLZ^{+}24, ZWZ^{+}22]. stratified [Hoh24]. STREAM [CKC^{+}21, CL21]. streamer [MHÅ21]. STREAmS [BMSP21, BMS^{+}23]. STREAmS-2.0 [BMS^{+}23]. Strength [Eku24]. stress [CT23, EGKH24, LLZ^{+}22, ZRH21]. stress-fluctuation [CT23]. stretched [RVRT22]. string [KL22]. StringSpinner [KL22, KL23a]. Strong [HMSV23, AG21, BGR23, GHL^{+}22, MT23, RZY^{+}24, ZLMH23]. strong-field [MT23]. strongly [BDR^{+}20, GBS^{+}20]. structural [BDP^{+}21, NNMJ22]. Structure [BLN^{+}21, BGHC23, BGR23, ČPF^{+}24, CWJ21, DSQ23, FASD20, FPSZ21, GK21, GO23, HTS^{+}21, HY23, HTH^{+}20, HSO^{+}22, HLzY^{+}20, JMOC21, JPJ^{+}23, LBS^{+}23, LS23, LLT^{+}23, MKPW21, MKPW22, MFB23, MSM24, NG21, NVCS23, PAL^{+}20, PPKK21, SZNW23, TCY23, TRB^{+}20a, VSM^{+}22, WZZS21, WBK^{+}24, WXZH24, XQ21, YLLW24, YSX^{+}20, YHY^{+}21, zYCD^{+}20, ZZ21b, ZZC20, ZHM21]. structure-preserving [CWJ21, XQ21]. Structured [LSNRD20, LJS21, MBCC23, PMSHG23]. structures [AS22, ASPDL^{+}21, DNG^{+}20, HLCD20, IMB^{+}22, LBS^{+}23, LXY^{+}21, LLQ^{+}23, MYM^{+}22, MRG22, RL21, SLZG20, SJ20, TDR^{+}20, ZHS^{+}20, vTDGCR21]. studies [CGG21, GLSH21, LSF23, LFZ20, SZ23]. Studio [NKP20]. Study [RDZ^{+}20, THH21, AGMFGE23, NS20, NBB^{+}21, SKM^{+}21, XLL^{+}22, ZDG^{+}21, ZBS^{+}23]. studying [BPMMP24, May21, SC22]. style [Ham20b, WBK^{+}24]. SU3lib [DLD^{+}21]. sub [SZ23, YLIO22, ZYG21]. sub-critical [ZYG21]. sub-Ohmic [SZ23]. sub-permutation [YLIO22]. subatomic [AGJ^{+}23]. subdomains [SVSC20]. Subprograms [HW22]. Subsidiary [GKIB21]. subspace [HKY^{+}21]. Subsystem [MSG^{+}21]. subtraction [LMWW24a, LM20]. successful [YHH^{+}20]. successive [BJL^{+}24]. SudoDEM [ZZ21a]. suite [ADW^{+}23, BPMMP24, HBM^{+}24, KLM^{+}22, MEC^{+}24, MBG^{+}20, RDH^{+}20, RRM^{+}23, SSD^{+}22]. SUKI [NUK^{+}22]. summation [XWYQ21]. Sunway [CLZ^{+}21]. super [TIG^{+}24]. super-heavy [TIG^{+}24]. supercomputer [RCS21]. supercomputers [GB22, KMM21, Kul20]. superconducting [BM22, JYL^{+}23, WYHW24]. superfluid [JRS^{+}21, VB22b]. superlattices [NNMJ22]. supermultiplet [PDD24]. SuperScreen [BM22]. Supersonic [BMS^{+}23, BBV^{+}22, RCX24]. superstructure [PTD20]. supersymmetric [KMU^{+}23, Bie21]. Supervised [WZC^{+}24]. Support [CZ21, WWDM20, ZW20, SPF21]. supporting [ABF^{+}23b]. Surf [PMS20b]. surface [CFLR24, HSO^{+}22, LGDF20, LMQ^{+}23, PRR23, PCL24, SXYD24, SVJ^{+}24, WA21, WNS^{+}21, ZYML24, ZHZG23]. surfaces [CGR21, IGL^{+}24, Org22, PMS20b, RV20, RV21, SD24, SDP24, TGBM22, VZ20, WGS^{+}22, XDF20, YYZ^{+}22]. surfactant [WYT23, YK21a, ZCWY20]. Surfing [PMS20b]. surrogate [SMC^{+}22]. survey [WRM^{+}24]. SuSpect3 [KMU^{+}23]. Susskind [Dür23]. SVF [WYZZ23]. SWANLOP [AB21]. swarm [TV24, ZZC20]. swift [SA23]. swifter [ZGZW23]. switchable [MBCC23]. Symbolic [GDB10, CLS22b, CB20, MG22, PSW23, CCK23]. symbols [XWYQ21]. symmetric [DEV20, LF24, SJY18, SJY20, zYMK^{+}21, Zeb22]. symmetric- [SJY18, SJY20]. symmetries [CFBRE24]. Symmetrization [YT22]. Symmetry [IMB^{+}22, TW21a, ALB22, FPSZ21, GF21, GLSH21, HA21, Kor23, MLD^{+}22, PDD24, SKM^{+}21, SCT21, ZXW^{+}22]. symmetry-adapted [GF21, GLSH21, MLD^{+}22]. symmetry-preserving [HA21]. Synergistic [OBK^{+}20]. synthesis [dBBVA20]. synthesizing [FLW^{+}23]. synthetic [JSS^{+}24, WBF^{+}24]. System [AAA^{+}20, CLY22, HBM^{+}24, KRC^{+}20, KvH21, LM21a, LLT^{+}23, LSW^{+}20, MMYU22, PYT^{+}24, PK24, UJ21, eSdSBST21, WXX24, Yan21b, Yev21, ZCWY20, ZYXC24, ZYG21]. system-bath [CLY22]. Systems [MYMK^{+}21, SGM^{+}20, AGMFGE23, ARLDG24, ABBD23, BKS22, BMT^{+}20, BLM^{+}22, CR20, CPL21, CMJC21, COW24, DZZ21, DHK^{+}21, DVC^{+}22, FA21, Gal22, GTMB21, GF23, GLB^{+}21, GSBN22, HPP23, IJVJ21, IUJ21, KMBP24, KHR^{+}23, KKPC21, KSC^{+}23, KDHL23, KAG24, KLM^{+}22, LHZ20, LG21a, Lee21b, LLZ^{+}23, LQGL21, MD20b, Man24, MFB23, MP21, MAJ20, MOY^{+}22, OCK^{+}24, Ols23, OLNG21, QJ21, RSD20, RMR^{+}22, RHW^{+}21, RMM21, RdSH^{+}24, RBWD^{+}24, SRT^{+}20, SFKC22, SFBG20, SVJ^{+}24, SP23, TC24, TMH^{+}23, TRB20b, WPMK21, XZLX20, YKK^{+}23a, YM21].

T [CCK23, COW24]. tackling [PPKK21]. tagged [LG21a]. tagged-particle [LG21a]. tagger [Ada22]. tagging [LNP^{+}24]. Taichi [SGY24]. Taichi-powered [SGY24]. Taihulight [CLZ^{+}21]. tails [AFJ^{+}23]. takes [PGS^{+}24]. Talmi [Efr21b]. tapir [GHL23]. target [BKS24a, BCD^{+}21, NUK^{+}22, OBC^{+}24, SHRK22, TLC^{+}21, YK21b]. target-oriented [YK21b]. Task [DP21, CCC20, DFU20, Di 22]. Task-based [DP21, DFU20, Di 22]. Tau [LNP^{+}24]. TAUOLA [ABWZ23, Nie23]. TauRunner [SLP^{+}22]. TB2J [HHVB21]. TBPLaS [LZK^{+}23]. TC [Och23a, Och23b]. TChem [KDIN^{+}23]. TDDFT [DHE^{+}24]. TDEP [BBB20]. TDHF [ASS^{+}24]. TDSE [ZLMH23]. TE [SS21a]. tearing [ZJM^{+}21]. technical [KFHR24, NKP20]. technique [MG22, MDZ24, NDFL24, NBB^{+}21, PAL^{+}20, WMM^{+}24, XFH^{+}22, YMCF23, PGS^{+}24]. techniques [DG20, GLB^{+}21, KHKL24, RdSH^{+}24, SFC20, VPS23, YZZ^{+}23]. telescopes [AAG^{+}20]. Temperature [BBB20, ASU^{+}21, Bar22a, BCTS22, GO23, GJA21, GCWZ20, IUJ21, KKPC23, LLQ^{+}23, SKC21, SC22, VV21, XZLX20]. temperature-dependent [KKPC23]. tempering [MMFdL21]. Temporal [SL22, ZLY24]. temporally [KFHR24]. TeNeS [MOY^{+}22]. TENO [LF24, Tan24]. tension [CFLR24]. Tensor [CDD22, MOY^{+}22, CKT21, JS24, LG23, MCMS24, OS24, TWW22, WICA22, YS22, YK21b]. Tensor-network [CDD22]. tensor-product [TWW22]. TensorAlloy [CGZ^{+}20]. TensorBNN [KKP22]. TensorFlow [KKP22, WICA22]. TEPPP [HPP23]. term [LSS24, MP21, XZLX20]. terms [Ber24, Hid21, IUJ21, MD22, RBV^{+}22]. ternary [KAS24]. tessellation [Nis22, SDP24]. TETHYS [CSBF23]. tetragonal [LHZ20]. tetrahedral [YXX^{+}21]. tetratomic [RV21]. TFmix [SLK23]. their [ABFP24]. Theoretical [UMA21, CLEPF22, Jab24, RSC^{+}22]. theories [DCZ23, HS24, MCMS24]. Theory [DvHSdS22, FKS20, MVK^{+}24, ATRD21, BCCM^{+}24, BW23, DMS^{+}22, DH22, EST23, FASD20, FWL^{+}24a, Fon21, Gai20, GF23, GM20, GBD^{+}22, Hor24, HLzY^{+}20, KKSV24, LACL24, LYC20, Man24, MHK24, MDR^{+}20, QJ21, SLZY21, SHW^{+}21, TMC22, VCF22, YT23, YKSH20, DPR^{+}20, DRR^{+}24, SHW^{+}21]. thermal [CK20a, CT23, DFP^{+}20, EST23, FO20, FCTFR20, HYL^{+}22, HTL^{+}22, KMG^{+}20, LDGN24, MMM20, MPN^{+}21, RMCC22, VGGP^{+}21, ZGW20, ŻTR^{+}22, PLSB22]. Thermo [FLW^{+}23, ZWC22]. Thermo-lp [FLW^{+}23]. Thermo4PFM [FDPT23]. thermochemical [LWV20, WV22, YKYK23]. Thermochemistry [LWV20]. Thermodynamic [SLK23, AMP^{+}21, FDPT23, FGCN^{+}21, KvdW20]. thermodynamics [FLW^{+}23, ZWC22]. thermoelastic [LVB22]. thermoelasticity [LDW^{+}21]. thermoelectric [CL20]. thermophysicalModels [NJSY22]. thermosolutal [ZDG^{+}21]. Theta [CK20b]. thin [BMP^{+}24, RMR^{+}22]. thin-film [BMP^{+}24]. third [LLS^{+}21, TSL21]. third-order [LLS^{+}21]. Thomas [YNMR24]. those [MBCC23]. thread [PNL^{+}24, RAJ^{+}24]. thread-block [RAJ^{+}24]. threaded [AG21]. Three [CLS^{+}22a, CL21, AHP21, ASU^{+}21, BKO20, BR20b, BSK^{+}22, Efr20, Efr21a, GOS^{+}22, GMZ^{+}20, JDS20, JRS^{+}21, KLD^{+}21, LMQ^{+}23, MD20b, MSH22, PK24, PD23, SRT^{+}20, SXYD24, TTM22, WXZH24, YWTK23, ZDG^{+}21, Zho23]. three-body [JDS20]. three-channel [GOS^{+}22]. three-component [YWTK23]. Three-dimensional [CLS^{+}22a, CL21, AHP21, BKO20, BR20b, BSK^{+}22, GMZ^{+}20, KLD^{+}21, LMQ^{+}23, MD20b, MSH22, PD23, SRT^{+}20, SXYD24, TTM22, WXZH24, ZDG^{+}21, Zho23]. three-electrode [PK24]. three-particle [Efr20, Efr21a]. three-temperature [ASU^{+}21]. throughput [NAZ^{+}21, PFG22, RDH^{+}20, WXL^{+}21, WGGC22]. TIFF [Ken24]. Tight [DSQ23, JPJ^{+}23, KVSC21, LZK^{+}23, NKP20, WPMK21, ZYLY22, NKP20]. tight-binding [JPJ^{+}23, KVSC21, LZK^{+}23, NKP20, WPMK21, ZYLY22, NKP20]. Time [DRZ^{+}21, FMCB^{+}20, Hor23a, LY22, MVK^{+}24, PNL^{+}24, RS20, TKS22, ZYX21, ZYX22, ADW^{+}23, BV21, BAB^{+}20, CJZ21, DGM20, DS22, DVC^{+}22, DHE^{+}24, Dio23, DG20, DBR24, EBBB22, ET24, GPD^{+}22, GSLS20, GF23, GBD^{+}22, GMZ^{+}20, HSB^{+}24, HW22, HLMB24, Hor23b, JM24, JGJ20, JRS^{+}21, KCS22, KBSL22, KP23, KMR22, KM23, Koz23, LS22, LY24b, MZ22, MKHT20, PMK^{+}23, PPK22, QWZ^{+}21, RSC^{+}22, SL22, SGS23, Scr22a, SNG20, SKC21, SCF20, TTM22, TSL21, WJB21, WWB22, WHB21, WBF^{+}24, WKJB23, WBvdH20, XSL^{+}22, XLKX21, YZW21, YSMBA23, ZX23]. time-dependence [BAB^{+}20]. Time-Dependent [MVK^{+}24, Hor23a, LY22, BAB^{+}20, Dio23, GSLS20, GBD^{+}22, HSB^{+}24, JRS^{+}21, KM23, LY24b, SGS23, Scr22a, WJB21, WHB21, WBvdH20, XSL^{+}22, YSMBA23]. Time-discretization [TKS22]. Time-domain [DRZ^{+}21, PNL^{+}24, ET24, TTM22, YZW21]. time-evolution [GF23]. time-explicit [DS22]. Time-independent [FMCB^{+}20, ADW^{+}23, BV21]. time-lapse [DG20]. time-of-flight [WBF^{+}24]. Time-parallel [RS20]. time-reversible [KP23]. time-space [GMZ^{+}20]. time-spectral [LS22, SL22]. time-stepping [ZX23]. time-warp [RSC^{+}22]. TimeEvolver [MZ22]. times [LTMK21]. timescale [RMR^{+}22]. tin [PAL^{+}20]. tinie [DKRSR22]. tip [AUO^{+}22]. Tiresia [TCSD24]. tissue [VRI24]. TLBfind [PLSB22]. TMDICE [Roh22]. TMM [BMP^{+}24]. TMM-Sim [BMP^{+}24]. TNQMetro [CDD22]. Tokamak [YYC^{+}23, BYWW23, BJS^{+}23, HEF21, JKSY22, LZ21, MGX24, MAMK21, MPH^{+}24, RSHS24, SKYQ21, SDXY23, SBZ23, WDMZ24, YC20]. TokaMaker [HSB^{+}24]. tokamaks [BFL^{+}22, JSS^{+}24, MG22, NSY^{+}23]. tomography [BYL^{+}21, HTN21, RZY^{+}24, WKJB23]. tomosynthesis [BSG^{+}21]. Tool [BHK^{+}21, BLN^{+}21, AMP^{+}21, AFJ^{+}23, AES21, BMM21, BMR^{+}24a, BMP^{+}24, BKS24b, CSBF23, EGKH24, FLW^{+}23, FGCN^{+}21, FLK^{+}20, GHL23, HSB^{+}24, HT20, HRG^{+}22, IGL^{+}24, Kan23, KKB^{+}24, KMY^{+}23, LLS^{+}21, LR22, MMCC^{+}22, MZD^{+}20, MRG22, NYN^{+}21, Ols23, DARJ23, RCB21, RJ21, SLBR22, SPLD20, SLB^{+}23, SZ24a, SRML20, SCMP^{+}22, TDR^{+}20, TAB^{+}22, UO20, WGGC22, YNV22, YNV23, ZHM^{+}24, ZDLS21, dMMLOS20, Dju20]. tool-kit [WGGC22]. Toolbox [CV21, LWV20, OSE^{+}20, ASW22, BTG22, CKGW22, FBC^{+}21, GDK21, HTN21, HRU22, Hoh24, JZW^{+}22, KMBP24, LSNRD20, PCS^{+}23, RSD20, RZWW23, RSPJ21, RPG^{+}20, SKDH24, Sch21, VCHH23, Yan24b]. toolchain [OIA^{+}20]. toolkit [AFB^{+}24b, BVV22, DCRF23, DIKSN24, EL24, Eku24, GAS^{+}23, Hua23, KDIN^{+}23, LVB22, LEL^{+}22, MES^{+}24, OSK^{+}21, SC22, TQGE23, WMA^{+}22, YTC^{+}20]. Tools [CLVV22, DBBP23, FR20, HCAH^{+}24, KvdW20, LHG^{+}20]. TopoAna [ZDLS21]. topography [NBCMH20]. topological [ATC^{+}23, GBR23, LU21, MCP23, VRI24, HPP23]. topologies [GHL23]. topology [AS24a, Nis23]. tops [SJY18, SJY20]. TORCWA [KL23b]. toroidal [CMS22, LOT^{+}20, LZ21, MG22, MMYU22]. toroidally [BJL^{+}24]. toroidally-rotating [BJL^{+}24]. torsional [CZB^{+}23]. Total [MSHP02, MSHP20, HMYH22, HSO^{+}22, MD20a]. Total-energy [MSHP02, MSHP20]. total-reflection [HMYH22, HSO^{+}22]. Townsend [MHÅ21]. TPMD [SC22]. tqix [HTN21]. tqix.pis [VCHH23]. trace [LMWW24b]. tracing [FXQS21, GSL^{+}23, SHS22, WWM^{+}22, sXBkB^{+}22]. Track [BBB^{+}21, AAB^{+}21, BDGS21, TAGC22b, SP23]. tracker [AAB^{+}21]. Tracking [DHK^{+}21, AGJ^{+}23, ABK^{+}22, KSC^{+}23, LBRW22, NSY^{+}23, NBS^{+}20, PCL24, SRC21, SBP20, WWM^{+}22, YGSW21]. tractable [RDZ^{+}20]. traction [HGS20]. traditional [TCY^{+}21]. Training [BKS24a, WZC^{+}24, vdHKB^{+}23]. trajectories [BFT20, SKC21, SC22, SZT24, TBAR21]. TRANS [SHS22, ZLY24]. Trans-Net [ZLY24]. transcendental [BA24]. transcorrelated [Och23a, Och23b]. Transfer [BMR^{+}23, CJZ21, FXQS21, KKPC23, KSIL22, MDDI21, NBB^{+}21]. transferability [MYKC23]. transferable [SPMS23, ZLY24]. transfers [HRG^{+}22]. transform [CLEPF23, KIK20]. transformation [IJVJ21, KF23, Lee21b, PP21, RMFB23, SMGK21, YW21, YJ24]. transformations [ABFP24, Efr20, Efr21a, MRT^{+}22, YKY^{+}22, dBBVA20]. transformed [MSN^{+}22]. transforms [BJS^{+}23, KPST21]. transient [AE23, ET24, Ryd20]. Transition [VLS22, BMM21, BFD22, CXCZ23, LKK23, LGBJ20, MMM23, MMFdL21, VB22b]. transitions [Bar22a, EST23, GSV23, SZ23, VRI24]. TRANSLATE [BCF^{+}24]. transmission [Arn20, CGV^{+}22, SLL^{+}24]. transmutation [LF20]. Transparent [SP20]. transpond [BGW^{+}22]. Transport [KOF21, NRKA22, AUEO24, ABGD^{+}20, ASPDL^{+}21, BGW^{+}22, BCF^{+}24, BSS^{+}23, CK20a, CPL21, CHA21, CL20, CKC^{+}21, CL21, DFP^{+}20, DKRSR22, EBBB22, FO21, GAGO21, GLN23, Jab24, Kal20, LBM^{+}23, LLL24, LFL^{+}24, LHWX24, MK22, MBTB21, MMM20, MGC^{+}23, MP21, MAMK21, MMYU22, NSY^{+}23, OBGA24, PLT^{+}23, PYT^{+}24, PPR^{+}21, RDC^{+}20, SHS22, SP23, SLIC24, TAY^{+}24, WAN^{+}22, WXW^{+}21, XDF20, XSL^{+}22, YHH^{+}20, YYH21, YJLW21, YZHL22, ZLLM23, ZPL^{+}21, ZCS^{+}24]. transport-property [AUEO24]. TransROTA [BJL^{+}24]. transverse [KPST21]. trapped [BFT20]. traps [BTK24]. traversal [WVK21]. TRAVOLTA [RBWD^{+}24]. t}re [ST23b]. treams [BFCR24]. treatment [RZH22, YXX^{+}21]. tRecX [Scr22a]. Tree [CFL^{+}22, SJ20, WVK21]. treecode [CGR21, KKLZ23b]. trends [KIK20]. trhepd [HMYH22, HMYH22, HSO^{+}22]. trial [MMM23]. trials [BSG^{+}21]. triangular [AIZ23, BRAC23, SS21b, YXX^{+}21]. triangulated [TGBM22]. Triatomic [RV20, GTMB21, Xav22]. trichalcogenides [CXCZ23]. tridiagonal [KKPC21, YKK^{+}23a]. trigger [TGIM23]. triggering [ABF^{+}23b]. trigonal [LHZ20]. TRIM [PFG22]. trions [PC21]. triple [AFB^{+}24b, Bzo21]. triple- [Bzo21]. triple-GEM [AFB^{+}24b]. TripleK [Bzo21]. triply [LXY^{+}21]. TRIQS [KM22]. TRIQS/SOM [KM22]. triSurfaceImmersion [TGBM22]. Tropical [BMT23]. truncated [HHT22]. Tsinghua [ZXT22, ZXT23]. TTCF4LAMMPS [MES^{+}24]. TTDFT [LG23]. Tucker [LG23]. TUMME [ZXT22, ZXT23]. tuneable [GTE21]. tungsten [YHH^{+}20]. Tuning [BSS^{+}23, LAC21]. Tunneling [Gul20, SZNW23]. tunnels [Ryd20]. turboMagnon [GBD^{+}22]. TurboPy [RGS^{+}21]. turbulence [FA20a, Ken24, KPL^{+}21, LBRW22, LF24, MSU^{+}21, SKYQ21, SDXY23, SBZ23, ZYX21, ZYX22, ZSqXY21]. turbulent [BMSP21, BMS^{+}23, FA20b, GB22, KWZ24, KKPC23, NS20, VLS22]. TURTLE [NBCMH20]. tutorial [BM20, SOH^{+}23]. tutorial-driven [BM20]. twiddling [Fis24]. twisted [Kub24]. twisted-mass [Kub24]. Twister [NNMJ22]. Two [HT24, JYL^{+}23, KM20, LF20, LHTP^{+}24, AYB24, ABFP24, BBA23, BM22, CAWK22, CCK23, Dan24, Dau23, DKRSR22, EPM23, ENK24, GDB10, GCK21, KWK^{+}21, KBSL22, KM23, KMS20a, LW24, MA24, MBJB24, MFB23, MYM^{+}22, MCP23, SRT^{+}20, SLZ^{+}24, SMB20, SSB^{+}23, SFS22, UÁEPGBP24, WT22, Yan21b, ZYXC24, KMS20b]. two-body [KWK^{+}21]. two-center [SSB^{+}23]. Two-dimensional [HT24, LHTP^{+}24, BM22, Dan24, DKRSR22, KM23, MFB23, MYM^{+}22, SRT^{+}20, UÁEPGBP24]. Two-fluid [JYL^{+}23]. Two-Higgs-Doublet [KM20, KMS20b]. two-phase [BBA23, CAWK22, Dau23, ENK24, KBSL22, MA24, SFS22, WT22, Yan21b]. two-photon [KMS20a]. two-step [AYB24]. twoWayGPBEFoam [LMHL20]. type [Arn20, BCGT24, CZY20, GMZ^{+}20, Pöt20, TCY23, TL20, XFH^{+}22, ZDLS21].

U [GOCSS^{+}23]. UCNS3D [ADF^{+}22]. udkm1Dsim [Sch21]. UKRmol [HBM^{+}24, MBG^{+}20]. UKRmol-scripts [HBM^{+}24]. ultra [BCT20, BKBL22, EES24]. ultra-peripheral [BCT20, BKBL22]. ultrafast [Hua24, Sch21, SLB^{+}23, Tan23, WKJB23, WBvdH20, ZPL^{+}21]. ultrashort [BAB^{+}20, MD22]. ULYSSES [GMPG^{+}21, GLPG^{+}23]. unbinding [Ryd20]. uncertainty [KMG^{+}20, MZL^{+}21, SMC^{+}22, TRN^{+}24, TO21, VEHCM21]. Unconditionally [Yan24a, GZW20, LM21a, ZCWY20, ZZ24]. unconditionally-stable [ZZ24]. underlying [RZ23]. UNDI [BFI^{+}21]. UnDiFi [CAC^{+}22]. UnDiFi-2D [CAC^{+}22]. UNED [COJ^{+}22]. unfolding [RMFB23]. Unified [KWZ24, ZWP^{+}22, FJ22, FTZ^{+}23, LSW^{+}20, LHG24, STA20, XFH^{+}22]. Uniform [JLW24, PCL24, NR21, PCS^{+}20, VB22b]. unipolar [WSK^{+}22, YFL22]. uniqueness [MNS^{+}24]. Unit [AAA^{+}20, AGH21]. unitarity [ADC^{+}21]. unitary [GH21]. units [DC22, WICA22]. Universal [CXCZ23, GLPG^{+}23, GMPG^{+}21]. universe [BMM21, FFTV23]. University [ZXT22, ZXT23]. Unleashing [ZZ21a]. Unlocking [RYS^{+}24]. unoccupied [SLZG20]. Unravelling [VPPQ21]. UNRES [OCK^{+}24]. unsaturated [HRG^{+}22]. unscented [AGJ^{+}23]. unsplit [LH22, LH24, MOV21]. unstable [KvdW20, KAB^{+}21, LWhK^{+}20]. unsteady [CJZ21, SFS22, YYY22]. unstructured [ADF^{+}22, CAC^{+}22, EPM23, KFC^{+}20, NTO24, NSY^{+}23, PLF20, TGBM22, ZDSS23]. unsupervised [WYZZ23]. Up-sampling [TCcN23]. Upcgen [BKBL22]. Update [IKM^{+}24, AJDS^{+}21, ABWZ23, GvdBdGN24, HJGL22, LJH^{+}23, XOTI22]. Updated [CTPS22, Hor24, SPTPR21, SLL^{+}24, SZY^{+}22]. Updates [AJW^{+}21, FQRR22]. Upgraded [NUK^{+}22, PMA^{+}21, YZW22]. URANOS [DAC^{+}23]. urban [YCC22]. usage [CHA21]. use [LSS24, MG22, TPS^{+}24, WKBW21]. used [KIK20]. User [SMB20, WAYL23, BMR^{+}24b, DvHSdS22, HGS20, KKSV24, MMCC^{+}22, WXL^{+}21, YR21]. user-defined [KKSV24]. user-friendly [HGS20, MMCC^{+}22, WXL^{+}21]. User-ready [SMB20]. users [YTC^{+}20]. usine [Mau20]. Using [MMM23, TAGC22b, ABBD23, Ano20c, ASC^{+}21, BOSM24, BRAC23, BGH22, BKB^{+}21, BB24b, CE22, CGSO20, CFLR24, CL20, CZB^{+}23, CKT21, CT23, CJ21, CLEPF22, CLEPF23, CLEP24, CNS22, DMS^{+}22, Dau23, DS22, DC22, Dry21, FASD20, FMBD22, GF23, GBD^{+}22, GAS^{+}23, HT24, HW22, HSO^{+}22, HLzY^{+}20, HG22, Ilt21, JSS^{+}24, JS24, JPJ^{+}23, KWK^{+}21, KK16, KWZ24, KF23, KCS22, KSJ^{+}22, KKY24, KPL^{+}21, Kor23, KKP22, LOT^{+}20, LMX^{+}21, LZ21, LJQ^{+}22, LM22, LVMGF^{+}23, MMC^{+}21, MSN^{+}22, MGG^{+}20, MBG^{+}20, MAJ20, MKHT20, MPSK21, MPQ^{+}22, Nis22, OIA^{+}20, Och23a, Och23b, PMSHG23, PBK21, PP23, Par21, PNL^{+}24, PCS^{+}23, PFG22, Di 22, RCW^{+}20, SFBG20, SKC21, SC22, SFC20, SP23, SLC^{+}22, SS21b, TAGC22a, TGGC23, TMC22, WGG20, WXW^{+}21, WXL^{+}21, WWM^{+}22, WA21, WXZH24, WS20, WKJB23, Xav22, XHY^{+}24, YJLW21, YLL^{+}22]. using [ZLV23, dBBVA20]. utilitarian [BC22]. Utility [DBdFdSR21]. utilizing [WHSG24].

v [AGK^{+}23]. v.9.0.0 [AMA^{+}20]. v1.0.0 [BM20, SY20]. v1.1 [CR20, GKIB21]. v1.2 [ADH^{+}20]. v1.3.0 [BDK^{+}23]. V2 [LG21b, AAHJ20, BMM21]. v2.0 [OSE^{+}20]. v2.0.0 [BN23]. v2107 [GDK21, HRG^{+}22]. v3 [DRR^{+}24]. v3.0 [EL24]. v4.0 [MSN^{+}22]. v5 [ZW20]. v5.1 [CZ21]. vacancies [WLF^{+}24]. vacuum [Sat21]. valence [CXCZ23]. validating [MMFdL21]. validation [CEC^{+}24, DIK^{+}23, DARJ23, SCR^{+}22b, ZYG21]. value [ATC^{+}23, FYW23, KFHR24]. valued [CL22a]. values [GOST23, Jab20]. Vanderbilt [TIG^{+}24]. vanishing [JDS20]. variable [WYT23, YK21a, PBK21]. variables [ABFP24, Par21]. variance [CHA21, PLT^{+}23, PPR^{+}21]. variance-reduced [PPR^{+}21]. variant [WYT23, YK21a]. variants [Hor24]. Variational [WHSG24, YYH21, AGH21, Jia20, PC21, SZ23, XSL^{+}22, XOTI22, XFGS24, YZHL22]. Variationally [RG21]. variations [MM23]. various [CW22, HWZ24, SHW^{+}21, sXBkB^{+}22]. varRhoTurbVOF [FA20a, FA20b]. varying [SD24]. VASP [DCRF23, GWPW21, LR22, MPSK21, WXL^{+}21, YTC^{+}20]. VASPKIT [WXL^{+}21]. Vector [FMBD22, ONH^{+}20, ABB^{+}22, CV24, KL23a, TW21b, WWDM20]. Vector-based [ONH^{+}20]. vector-meson [KL23a]. vectorial [CMS22]. vectorised [EPM23]. vectorized [CNS22]. VegasFlow [CCM20]. VELAS [RZWW23]. velocity [HFP21, KFC^{+}20, UJ21, VPPQ21, Xie23]. VENUS [MPQ^{+}22]. VENUS-LEVIS [MPQ^{+}22]. Verification [DIK^{+}23, MGC^{+}23, SCR^{+}22b, SBB^{+}24, Bak23, CJD^{+}20, KDK23]. verified [GAS^{+}23, LHWX24]. versatile [BVV22, BMP^{+}24, KKPC23, WLL22]. Version [AKK^{+}24, AAT^{+}20, ASPDL^{+}21, BC22, GLPG^{+}23, HQF^{+}20, KKM^{+}20, Pos22, SJP21a, ASS^{+}24, BMS^{+}23, CB23, CCG21, CZB^{+}23, DS22, DACA^{+}22, DP21, Dio23, FTG23, GST21, GOS^{+}22, GG24, Hor24, Jab22b, Koł22, MS24, MBJB24, MSN^{+}22, MPZB^{+}24, NAZ^{+}22, Org22, RZY^{+}24, SMO^{+}20a, SLL^{+}24, SVJ^{+}24, SAS20, YZW22, ZXT23]. versions [BBH^{+}23, IKM^{+}24]. vertex [WLCF21]. Very [Koz23, KAB^{+}21, OCK^{+}24]. VHEGEN [LHZ20]. via [AUO^{+}22, ABD^{+}23, DR21, EGKH24, GSV23, LG21a, LKK23, LM23, NAZ^{+}21, RV20, RV21, Roh22, VGGP^{+}21, YK21b, ZZC20]. vibrational [GTMB21, KSG22, MEC^{+}24]. vibronic [LHZ20]. viewSq [MKPW21, MKPW22]. Vincia [BP21]. violation [KKSV24]. Virtual [LLL24, BSG^{+}21, LHC20]. VIRTUS [MMP20]. viscous [RCW^{+}20]. VISROC [CTPS22]. Visual [MKPW22, Gro22, MKPW21]. visualisation [NRG24]. Visualization [CTPS22, EL24, NRG22, RZWW23, WGGC22]. visualizing [MKPW21, MKPW22, TQGE23, vTDGCR21]. Vlasov [LSW^{+}20, CE22, Ein20, KMD^{+}21, PMK^{+}23, TCSA21, TKC^{+}21, YKY^{+}22, ZLL^{+}24]. VMD [MKPW21, MKPW22, WKBW21]. Vofi [CCL^{+}22]. VOFTools [LHG^{+}20]. volcano [ZLV23]. volume [ADF^{+}22, CCL^{+}22, Dau23, DS22, FZD^{+}24, FA20a, FA20b, KFC^{+}20, LXY^{+}21, LY24a, LHG^{+}20, LH22, LH24, MSM24, PGYF21, TGBM22, VGGP^{+}21, VTB^{+}21, WH23]. volume-conservation-improved [LY24a]. Voro [LLR23]. Voronoi [LLR23, MRD23, Nis22, PCL24]. Vorotis [Nis22]. vortices [KSDH23]. VPA [PBK21]. VQE [XFGS24]. VTAnDeM [TQGE23].

walks [MW21]. wall [BKB^{+}21, DAC^{+}23, RZH22, YGSW21]. wall-bounded [DAC^{+}23]. wan2respack [KMY^{+}23]. Wang [AM21b, MDP22, SJ21]. Wannier [DSQ23, Kor23, ZXW^{+}22]. Wannier90 [KMY^{+}23]. WannSymm [ZXW^{+}22]. WanTiBEXOS [DSQ23]. warm [BCTS22, BRHT21, MHK24]. warp [RSC^{+}22]. water [DS22, HRG^{+}22, DARJ23, STA20]. watershed [HRG^{+}22]. wave [AJDS^{+}21, BGW^{+}22, CGV^{+}22, Dio23, FASD20, FWL^{+}24a, GBD^{+}22, GMZ^{+}20, JZW^{+}22, LL23, MSHP02, MSHP20, NS20, NBB^{+}21, OYC24, Pos22, RCB^{+}20, RRC^{+}24, SJP21a, TTM22, TL20, Yan23, YJLW21, YR21, ZZZ^{+}20, WWDM20, AYI^{+}24]. wave-packet [Dio23]. waveform [DRZ^{+}21]. waveforms [GF23]. wavefunctions [MONW21, RCP^{+}24]. wavelet [LDGN24, WXZH24]. wavelet-learning [LDGN24]. waves [AB21, BR20a, CFPS23, HM24, MDZ24, WXY20]. way [DEdM24]. ways [GSBN22]. weak [LGBJ20]. Weakly [CLVV22, WLF^{+}23, HMR22]. weakly-compressible [HMR22]. web [BMR^{+}23, BMR^{+}24b]. week [BR20a]. weight [NBC^{+}24]. Weighted [MGX24, MRD23]. weighter [AAA^{+}21]. well [MA24, ZZ24]. well-balanced [MA24]. well-posed [ZZ24]. WENO [EOR21]. wettability [NLS24]. Weyl [SLBR22]. Wheeler [HA23]. whistler [NS20]. whole [CL21]. whole-core [CL21]. widths [ABH^{+}23]. WIEN2k [GBR23]. wiggler [ASA^{+}22]. Wigner [DLD^{+}21, PDD24, XWYQ21]. Wilson [Dür23, FKK^{+}21]. WIMP [GJJN23, JKST22]. WIMP-nucleus [GJJN23]. WimPyDD [JKST22]. wind [YCC22]. window [MDZ24]. within [BN20, BN23, Dan23, GOCSS^{+}23, Mar22, NVC20, TWW22, WZZ^{+}23, XSL^{+}22, YHY^{+}21]. without [Ham20b, NA20, SLZG20, SKM^{+}21]. WloopPHI [SLBR22]. workflow [DSW^{+}23, GO23, LACL24, SS22, TPS^{+}24, vTDGCR21]. workflows [DIAA21, ZGZW23]. WPF [GOS^{+}22]. wrapper [FCSP20]. writing [KSIL22]. wurtzite [HLCD20].

X [MKPW22, CML^{+}24, HTN21, Hua24, Jab22a, MKPW21, MSU^{+}21, SPF21, WKJB23]. X-ray [MKPW22, CML^{+}24, Hua24, Jab22a, MKPW21, SPF21, WKJB23]. x86 [CZTF23]. XANES [BDP^{+}21, MGG^{+}20]. Xatu [UÁEPGBP24]. XCHEM [BFMA^{+}24]. XCHEM-2.0 [BFMA^{+}24]. XLB [AS24b]. Xsorb [PRR23].

XTOR [NDFL24]. XTOR-K [NDFL24].

YADE [KGT22, CCC20]. YAM2 [Par21]. Yan [AFJ^{+}23]. years [SHB^{+}20]. yield [HA23]. ython [VMRFC23].

Z’ [LSZ23]. Z’-explorer [LSZ23]. Z1 [KDHL23]. Z2 [GBR23]. Zapdos [DIK^{+}23]. ZEFR [RCW^{+}20]. zero [DH20, MOMO24]. zero-dimensional [MOMO24]. Zhang [PWC24]. zinc [HLCD20]. zinc-blende [HLCD20]. ZM-Cintegral [CZ21, WZPW20, ZW20]. ZMCintegral-v5 [ZW20]. ZMCintegralv5.1 [CZ21]. ZZPolyCalc [PWC24].

References

Aiello:2020:CUK

[AAA^{+}20] S. Aiello, F. Ameli, M. Andre, G. Androulakis, M. Anghinolfi, G. Anton, M. Ardid, J. Aublin, C. Bagatelas, G. Barbarino, B. Baret, S. Basegmez du Pree, M. Bendahman, E. Berbee, A. M. van den Berg, V. Bertin, V. van Beveren, S. Biagi, A. Biagioni, M. Bissinger, J. Boumaaza, S. Bourret, M. Bouta, G. Bouvet, M. Bouwhuis, C. Bozza, H. Brânzaş, M. Bruchner, R. Bruijn, J. Brunner, E. Buis, R. Buompane, J. Busto, D. Calvo, A. Capone, S. Celli, M. Chabab, N. Chau, S. Cherubini, V. Chiarella, T. Chiarusi, M. Circella, R. Cocimano, J. A. B. Coelho, A. Coleiro, M. Colomer Molla, S. Colonges, R. Coniglione, P. Coyle, A. Creusot, G. Cuttone, A. D’Onofrio, R. Dallier, M. De Palma, I. Di Palma, A. F. Díaz, D. Diego-Tortosa, C. Distefano, A. Domi, R. Donà, C. Donzaud, D. Dornic, M. Dörr, M. Durocher, T. Eberl, I. El Bojaddaini, H. Eljarrari, D. Elsaesser, A. Enzenhöfer, P. Fermani, G. Ferrara, M. D. Filipović, A. Franco, L. A. Fusco, T. Gal, A. Garcia Soto, F. Garufi, L. Gialanella, E. Giorgio, S. R. Gozzini, R. Gracia, K. Graf, D. Grasso, T. Grégoire, G. Grella, D. Guderian, C. Guidi, S. Hallmann, H. Hamdaoui, H. van Haren, A. Heijboer, A. Hekalo, J. J. Hernández-Rey, J. Hofestädt, F. Huang, G. Illuminati, C. W. James, P. Jansweijer, M. de Jong, P. de Jong, M. Kadler, P. Kalaczyński, O. Kalekin, U. F. Katz, N. R. Khan Chowdhury, F. van der Knaap, E. N. Koffeman, P. Kooijman, A. Kouchner, V. Kulikovskiy, R. Lahmann, G. Larosa, R. Le Breton, F. Leone, E. Leonora, G. Levi, M. Lincetto, M. Lindsey Clark, A. Lonardo, F. Longhitano, D. Lopez-Coto, G. Maggi, J. Mańczak, K. Mannheim, A. Margiotta, A. Marinelli, C. Markou, G. Martignac, L. Martin, J. A. Martínez-Mora, A. Martini, F. Marzaioli, S. Mazzou, R. Mele, K. W. Melis, P. Migliozzi, E. Migneco, P. Mijakowski, L. S. Miranda, C. M. Mollo, M. Morganti, M. Moser, A. Moussa, R. Muller, M. Musumeci, L. Nauta, S. Navas, C. A. Nicolau, C. Nielsen, B. Ó Fearraigh, M. Organokov, A. Orlando, V. Panagopoulos, G. Papalashvili, R. Papaleo, C. Pastore, G. E. Păvălaş, C. Pellegrino, M. Perrin-Terrin, P. Piattelli, C. Pieterse, K. Pikounis, O. Pisanti, C. Poirè, G. Polydefki, V. Popa, M. Post, T. Pradier, G. Pühlhofer, S. Pulvirenti, L. Quinn, F. Raffaelli, N. Randazzo, A. Rapicavoli, S. Razzaque, D. Real, S. Reck, J. Reubelt, G. Riccobene, M. Richer, L. Rigalleau, A. Rovelli, I. Salvadori, D. F. E. Samtleben, A. Sánchez Losa, M. Sanguineti, A. Santangelo, D. Santonocito, P. Sapienza, J. Schnabel, V. Sciacca, J. Seneca, I. Sgura, R. Shanidze, A. Sharma, F. Simeone, A. Sinopoulou, B. Spisso, M. Spurio, D. Stavropoulos, J. Steijger, S. M. Stellacci, B. Strandberg, D. Stransky, M. Taiuti, Y. Tayalati, E. Tenllado, T. Thakore, S. Tingay, E. Tzamariudaki, D. Tzanetatos, V. Van Elewyck, G. Vannoye, F. Versari, S. Viola, D. Vivolo, G. de Wasseige, J. Wilms, R. Wojaczyński, E. de Wolf, D. Zaborov, A. Zegarelli, J. D. Zornoza, and J. Zúñiga. The control unit of the KM3NeT data acquisition system. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301909.

Abbasi:2021:LLN

[AAA^{+}21] R. Abbasi, M. Ackermann, J. Adams, J. A. Aguilar, M. Ahlers, M. Ahrens, C. Alispach, A. A. Alves, N. M. Amin, R. An, K. Andeen, T. Anderson, I. Ansseau, G. Anton, C. Argüelles, S. Axani, X. Bai, A. Balagopal, A. Barbano, S. W. Barwick, B. Bastian, V. Basu, V. Baum, S. Baur, R. Bay, J. J. Beatty, K.-H. Becker, J. Becker Tjus, C. Bellenghi, S. BenZvi, D. Berley, E. Bernardini, D. Z. Besson, G. Binder, D. Bindig, E. Blaufuss, S. Blot, S. Böser, O. Botner, J. Böttcher, E. Bourbeau, J. Bourbeau, F. Bradascio, J. Braun, S. Bron, J. Brostean-Kaiser, A. Burgman, R. S. Busse, M. A. Campana, C. Chen, D. Chirkin, S. Choi, B. A. Clark, K. Clark, L. Classen, A. Coleman, G. H. Collin, J. M. Conrad, P. Coppin, P. Correa, D. F. Cowen, R. Cross, P. Dave, C. De Clercq, J. J. DeLaunay, H. Dembinski, K. Deoskar, S. De Ridder, A. Desai, P. Desiati, K. D. de Vries, G. de Wasseige, M. de With, T. DeYoung, S. Dharani, A. Diaz, J. C.

Díaz-Vélez, H. Dujmovic, M. Dunkman, M. A. DuVernois, E. Dvorak, T. Ehrhardt, P. Eller, R. Engel, J. Evans, P. A. Evenson, S. Fahey, A. R. Fazely, S. Fiedlschuster, A. T. Fienberg, K. Filimonov, C. Finley, L. Fischer, D. Fox, A. Franckowiak, E. Friedman, A. Fritz, P. Fürst, T. K. Gaisser, J. Gallagher, E. Ganster, S. Garrappa, L. Gerhardt, A. Ghadimi, C. Glaser, T. Glauch, T. Glüsenkamp, A. Goldschmidt, J. G. Gonzalez, S. Goswami, D. Grant, T. Grégoire, Z. Griffith, S. Griswold, M. Gündüz, C. Haack, A. Hallgren, R. Halliday, L. Halve, F. Halzen, M. Ha Minh, K. Hanson, J. Hardin, A. A. Harnisch, A. Haungs, S. Hauser, D. Hebecker, K. Helbing, F. Henningsen, E. C. Hettinger, S. Hickford, J. Hignight, C. Hill, G. C. Hill, K. D. Hoffman, R. Hoffmann, T. Hoinka, B. Hokanson-Fasig, K. Hoshina, F. Huang, M. Huber, T. Huber, K. Hultqvist, M. Hünnefeld, R. Hussain, S. In, N. Iovine, A. Ishihara, M. Jansson, G. S. Japaridze, M. Jeong, B. J. P. Jones, R. Joppe, D. Kang, W. Kang, X. Kang, A. Kappes, D. Kappesser, T. Karg, M. Karl, A. Karle, U. Katz, M. Kauer, M. Kellermann, J. L. Kelley, A. Kheirandish, J. Kim, K. Kin, T. Kintscher, J. Kiryluk, S. R. Klein, R. Koirala, H. Kolanoski, L. Köpke, C. Kopper, S. Kopper, D. J. Koskinen, P. Koundal, M. Kovacevich, M. Kowalski, K. Krings, G. Krückl, N. Kurahashi, A. Kyriacou, C. Lagunas Gualda, J. L. Lanfranchi, M. J. Larson, F. Lauber, J. P. Lazar, K. Leonard, A. Leszczyńska, Y. Li, Q. R. Liu, E. Lohfink, C. J. Lozano Mariscal, L. Lu, F. Lucarelli, A. Ludwig, W. Luszczak, Y. Lyu, W. Y. Ma, J. Madsen, K. B. M. Mahn, Y. Makino, P. Mallik, S. Mancina, I. C. Mariş, R. Maruyama, K. Mase, F. McNally, K. Meagher, A. Medina, M. Meier, S. Meighen-Berger, J. Merz, J. Micallef, D. Mockler, G. Momenté, T. Montaruli, R. W. Moore, R. Morse, M. Moulai, R. Naab, R. Nagai, U. Naumann, J. Necker, L. V. Nguyêñ, H. Niederhausen, M. U. Nisa, S. C. Nowicki, D. R. Nygren, A. Obertacke Pollmann, M. Oehler, A. Olivas, E. O’Sullivan, H. Pandya, D. V. Pankova, N. Park, G. K. Parker, E. N. Paudel, P. Peiffer, C. Pérez de los Heros, S. Philippen, D. Pieloth, S. Pieper, A. Pizzuto, M. Plum, Y. Popovych, A. Porcelli, M. Prado Rodriguez, P. B. Price, B. Pries, G. T. Przybylski, C. Raab, A. Raissi, M. Rameez, K. Rawlins, I. C. Rea, A. Rehman, R. Reimann, M. Renschler, G. Renzi, E. Resconi, S. Reusch, W. Rhode, M. Richman, B. Riedel, S. Robertson, G. Roellinghoff, M. Rongen, C. Rott, T. Ruhe, D. Ryckbosch, D. Rysewyk Cantu, I. Safa, S. E. Sanchez Herrera, A. Sandrock, J. Sandroos, M. Santander, S. Sarkar, S. Sarkar, K. Satalecka, M. Scharf, M. Schaufel, H. Schieler, P. Schlunder, T. Schmidt, A. Schneider, J. Schneider, F. G. Schröder, L. Schumacher, S. Sclafani, D. Seckel, S. Seunarine, A. Sharma, S. Shefali, M. Silva, B. Skrzypek, B. Smithers, R. Snihur, J. Soedingrekso, D. Soldin, G. M. Spiczak, C. Spiering, J. Stachurska, M. Stamatikos, T. Stanev, R. Stein, J. Stettner, A. Steuer, T. Stezelberger, R. G. Stokstad, T. Stuttard, G. W. Sullivan, I. Taboada, F. Tenholt, S. Ter-Antonyan, S. Tilav, F. Tischbein, K. Tollefson, L. Tomankova, C. Tönnis, S. Toscano, D. Tosi, A. Trettin, M. Tselengidou, C. F. Tung, A. Turcati, R. Turcotte, C. F. Turley, J. P. Twagirayezu, B. Ty, M. A. Unland Elorrieta, N. Valtonen-Mattila, J. Vandenbroucke, D. van Eijk, N. van Eijndhoven, D. Vannerom, J. van Santen, S. Verpoest, M. Vraeghe, C. Walck, A. Wallace, T. B. Watson, C. Weaver, A. Weindl, M. J. Weiss, J. Weldert, C. Wendt, J. Werthebach, M. Weyrauch, B. J. Whelan, N. Whitehorn, K. Wiebe, C. H. Wiebusch, D. R. Williams, M. Wolf, K. Woschnagg, G. Wrede, J. Wulff, X. W. Xu, Y. Xu, J. P. Yanez, S. Yoshida, T. Yuan,

and Z. Zhang. LeptonInjector and LeptonWeighter: a neu-

trino event generator and weighter for neutrino observatories. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521001302.

Aiola:2021:HSS

[AAB^{+}21] S. Aiola, Y. Amhis, P. Billoir, B. Kishor Jashal, L. Henry, A. Oyanguren Campos, C. Marin Benito, F. Polci, R. Quagliani, M. Schiller, and M. Wang. Hybrid seeding: a standalone track reconstruction algorithm for scintillating fibre tracker at LHCb. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303520.

Aniszewski:2021:PRI

[AACE^{+}21] W. Aniszewski, T. Arrufat, M. Crialesi-Esposito, S. Dabiri, D. Fuster, Y. Ling, J. Lu, L. Malan, S. Pal, R. Scardovelli, G. Tryggvason, P. Yecko, and S. Zaleski. PArallel, Robust, Interface Simulator (PARIS). Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000175.

Alameddine:2024:ICL

[AAD^{+}24] Jean-Marco Alameddine, Johannes Albrecht, Hans Dembinski, Pascal Gutjahr, Karl-Heinz Kampert, Wolfgang Rhode, Maximilian Sackel, Alexander Sandrock, and Jan Soedingrekso. Improvements in charged lepton and photon propagation for the software PROPOSAL. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524001668.

Aiello:2020:GKG

[AAG^{+}20] S. Aiello, A. Albert, S. Alves Garre, Z. Aly, F. Ameli, M. Andre, G. Androulakis, M. Anghinolfi, M. Anguita, G. Anton, M. Ardid, J. Aublin, C. Bagatelas, G. Barbarino, B. Baret, S. Basegmez du Pree, M. Bendahman, E. Berbee, A. M. van den Berg, V. Bertin, S. Biagi, A. Biagioni, M. Bissinger, M. Boettcher, J. Boumaaza, M. Bouta, M. Bouwhuis, C. Bozza, H. Brânzaş, M. Bruchner, R. Bruijn, J. Brunner, E. Buis, R. Buompane, J. Busto, D. Calvo, A. Capone, V. Carretero, P. Castaldi, S. Celli, M. Chabab, N. Chau, A. Chen, S. Cherubini, V. Chiarella, T. Chiarusi, M. Circella, R. Cocimano, J. A. B. Coelho, A. Coleiro, M. Colomer Molla, R. Coniglione, I. Corredoira, P. Coyle, A. Creusot, G. Cuttone, A. D’Onofrio, R. Dallier, M. De Palma, I. Di Palma, A. F. Díaz, D. Diego-Tortosa, C. Distefano, A. Domi, R. Donà, C. Donzaud, D. Dornic, M. Dörr, D. Drouhin, M. Durocher, T. Eberl, D. van Eijk, I. El Bojaddaini, D. Elsaesser, A. Enzenhöfer, V. Espinosa Roselló, P. Fermani, G. Ferrara, M. D. Filipović, F. Filippini, A. Franco, L. A. Fusco, O. Gabella, T. Gal, A. Garcia Soto, F. Garufi, Y. Gatelet, N. Geißelbrecht, L. Gialanella, E. Giorgio, S. R. Gozzini, R. Gracia, K. Graf, D. Grasso, G. Grella, D. Guderian, C. Guidi, S. Hallmann, H. Hamdaoui, H. van Haren, A. Heijboer, A. Hekalo, J. J. Hernández-Rey, J. Hofestädt, F. Huang, W. Idrissi Ibnsalih, G. Illuminati, C. W. James, M. de Jong, P. de Jong, B. J. Jung, M. Kadler, P. Kalaczyński, O. Kalekin, U. F. Katz, N. R. Khan Chowdhury, F. van der Knaap, E. N. Koffeman, P. Kooijman, A. Kouchner, M. Kreter, V. Kulikovskiy, R. Lahmann, G. Larosa, R. Le Breton, O. Leonardi, F. Leone, E. Leonora, G. Levi, M. Lincetto, M. Lindsey Clark, T. Lipreau, A. Lonardo, F. Longhitano, D. Lopez-Coto, L. Maderer, J. Mańczak, K. Mannheim, A. Margiotta, A. Marinelli, C. Markou, L. Martin, J. A. Martínez-Mora, A. Martini, F. Marzaioli, S. Mastroianni, S. Mazzou, K. W. Melis, G. Miele, P. Migliozzi, E. Migneco, P. Mijakowski, L. S. Miranda, Z. Modebadze, C. M. Mollo, M. Morganti, M. Moser, A. Moussa, R. Muller, M. Musumeci, L. Nauta, S. Navas, C. A. Nicolau, B. Ó Fearraigh, M. Organokov, A. Orlando, G. Papalashvili, R. Papaleo, C. Pastore, A. M. Paun, G. E. Păvălaş, C. Pellegrino, M. Perrin-Terrin, P. Piattelli, C. Pieterse, K. Pikounis, O. Pisanti, C. Poirè, V. Popa, M. Post, T. Pradier, G. Pühlhofer, S. Pulvirenti, L. Quinn, O. Rabyang, F. Raffaelli, N. Randazzo, A. Rapicavoli, S. Razzaque, D. Real, S. Reck, J. Reubelt, G. Riccobene, M. Richer, S. Rivoire, A. Rovelli, F. Salesa Greus, D. F. E. Samtleben, A. Sánchez Losa, M. Sanguineti, A. Santangelo, D. Santonocito, P. Sapienza, J. Schnabel, V. Sciacca, J. Seneca, I. Sgura, R. Shanidze, A. Sharma, F. Simeone, A. Sinopoulou, B. Spisso, M. Spurio, D. Stavropoulos, J. Steijger, S. M. Stellacci, M. Taiuti, Y. Tayalati, E. Tenllado, T. Thakore, S. Tingay, E. Tzamariudaki, D. Tzanetatos, V. Van Elewyck, G. Vannoye, G. Vasileiadis, F. Versari, S. Viola, D. Vivolo, G. de Wasseige, J. Wilms, R. Wojaczyński, E. de Wolf, D. Zaborov, S. Zavatarelli, A. Zegarelli, J. D. Zornoza, J. Zúñiga, and N. Zywucka. gSeaGen: the KM3NeT GENIE-based code for neutrino telescopes. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302241.

Aiello:2024:ESK

[AAG^{+}24] S. Aiello, A. Albert, S. Alves Garre, Z. Aly, A. Ambrosone, F. Ameli, M. Andre, E. Androutsou, M. Anghinolfi, M. Anguita, L. Aphecetche, M. Ardid, S. Ardid, H. Atmani, J. Aublin, C. Bagatelas, L. Bailly-Salins, Z. Bardačová, B. Baret, S. Basegmez du Pree, Y. Becherini, M. Bendahman, F. Benfenati, M. Benhassi, D. M. Benoit, E. Berbee, V. Bertin, V. van Beveren, S. Biagi, M. Boettcher, J. Boumaaza, M. Bouta, M. Bouwhuis, C. Bozza, R. M. Bozza, H. Brânzaş, F. Bretaudeau, R. Bruijn, J. Brunner, R. Bruno, E. Buis, R. Buompane, J. Busto, B. Caiffi, D. Calvo, S. Campion, A. Capone, F. Carenini, V. Carretero, T. Cartraud, P. Castaldi, V. Cecchini, S. Celli, L. Cerisy, M. Chabab, M. Chadolias, A. Chen, S. Cherubini, T. Chiarusi, M. Circella, R. Cocimano, J. A. B. Coelho, A. Coleiro, R. Coniglione, P. Coyle, A. Creusot, A. Cruz, G. Cuttone, R. Dallier, Y. Darras, A. De Benedittis, B. De Martino, V. Decoene, R. Del Burgo, L. S. Di Mauro, I. Di Palma, A. F. Díaz, D. Diego-Tortosa, C. Distefano, A. Domi, C. Donzaud, D. Dornic, M. Dörr, E. Drakopoulou, D. Drouhin, R. Dvornický, T. Eberl, E. Eckerová, A. Eddymaoui, T. van Eeden, M. Eff, D. van Eijk, I. El Bojaddaini, S. El Hedri, A. Enzenhöfer, G. Ferrara, M. D. Filipović, F. Filippini, L. A. Fusco, O. Gabella, J. Gabriel, S. Gagliardini, T. Gal, J. García Méndez, A. Garcia Soto, C. Gatius Oliver, N. Geißelbrecht, H. Ghaddari, L. Gialanella, B. K. Gibson, E. Giorgio, A. Girardi, I. Goos, D. Goupilliere, S. R. Gozzini, R. Gracia, K. Graf, C. Guidi, B. Guillon, M. Gutiérrez, H. van Haren, A. Heijboer, A. Hekalo, L. Hennig, J. J. Hernández-Rey, F. Huang, W. Idrissi Ibnsalih, G. Illuminati, C. W. James, P. Jansweijer, M. de Jong, P. de Jong, B. J. Jung, P. Kalaczyński, O. Kalekin, U. F. Katz, N. R. Khan Chowdhury, A. Khatun, G. Kistauri, C. Kopper, A. Kouchner, V. Kulikovskiy, R. Kvatadze, M. Labalme, R. Lahmann, G. Larosa, C. Lastoria, A. Lazo, S. Le Stum and G. Lehaut, E. Leonora, N. Lessing, G. Levi, M. Lindsey Clark, F. Longhitano, J. Majumdar, L. Malerba, F. Mamedov, J. Mańczak, A. Manfreda, M. Marconi, A. Margiotta, A. Marinelli, C. Markou, L. Martin, J. A. Martínez-Mora, F. Marzaioli, M. Mastrodicasa, S. Mastroianni, S. Miccichè, G. Miele, P. Migliozzi, E. Migneco, S. Minutoli, M. L. Mitsou, C. M. Mollo, L. Morales-Gallegos, C. Morley-Wong, A. Mosbrugger, A. Moussa, I. Mozun Mateo, R. Muller, M. R. Musone, M. Musumeci, L. Nauta, S. Navas, A. Nayerhoda, C. A. Nicolau, B. Nkosi, B. Ó Fearraigh, V. Oliviero, A. Orlando, E. Oukacha, J. Palacios González, G. Papalashvili, E. J. Pastor Gomez, A. M. Păun, G. E. Păvălaş, S. Peña Martínez, M. Perrin-Terrin, J. Perronnel, V. Pestel, R. Pestes, P. Piattelli, C. Poirè, V. Popa, T. Pradier, S. Pulvirenti, G. Quéméner, C. Quiroz, U. Rahaman, N. Randazzo, S. Razzaque, I. C. Rea, D. Real, S. Reck, G. Riccobene, J. Robinson, A. Romanov, A. Saina, F. Salesa Greus, D. F. E. Samtleben, A. Sánchez Losa, M. Sanguineti, C. Santonastaso, D. Santonocito, P. Sapienza, Y. Scarpetta, J. Schnabel, M. F. Schneider, J. Schumann, H. M. Schutte, J. Seneca, B. Setter, I. Sgura, R. Shanidze, Y. Shitov, F. Šimkovic, A. Simonelli, A. Sinopoulou, M. V. Smirnov, B. Spisso, M. Spurio, D. Stavropoulos, I. Štekl, M. Taiuti, Y. Tayalati, H. Tedjditi, H. Thiersen, I. Tosta e Melo, B. Trocme, S. Tsagkli, V. Tsourapis, E. Tzamariudaki, A. Vacheret, V. Valsecchi, V. Van Elewyck, G. Vannoye, G. Vasileiadis, F. Vazquez de Sola, C. Verilhac, A. Veutro, S. Viola, D. Vivolo, H. Warnhofer, J. Wilms, E. de Wolf, H. Yepes-Ramirez, G. Zarpapis, S. Zavatarelli, A. Zegarelli, D. Zito, J. D. Zornoza, J. Zúñiga, and N. Zywucka. Embedded software of the KM3NeT central logic board. Computer Physics Communications, 296(??): ??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003818.

Arbey:2020:AVP

[AAHJ20] A. Arbey, J. Auffinger, K. P. Hickerson, and E. S. Jenssen. AlterBBN v2: a public code for calculating Big-Bang nucleosynthesis constraints in alternative cosmologies. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303236.

Alexeev:2020:MHP

[AALK20] Dmitry Alexeev, Lucas Amoudruz, Sergey Litvinov, and Petros Koumoutsakos. Mirheo: High-performance mesoscale simulations for microfluidics. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552030093X.

Alperen:2023:HEN

[AAMY23] Abdullah Alperen, Hasan Metin Aktulga, Pieter Maris, and Chao Yang. Hybrid eigensolvers for nuclear configuration interaction calculations. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002333.

Allanach:2014:NMS

[AAT^{+}14] B. C. Allanach, P. Athron, Lewis C. Tunstall, A. Voigt, and A. G. Williams. Next-to-minimal SOFTSUSY. Computer Physics Communications, 185(9):2322–2339, September 2014. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465514001386. See corri-

gendum [AAT^{+}20].

Allanach:2020:CNM

[AAT^{+}20] B. C. Allanach, P. Athron, Lewis C. Tunstall, A. Voigt, and A. G. Williams. Corrigendum to “Next-to-Minimal SOFTSUSY” [Comput. Phys. Comm. 185 (2014) 2322–2339] (New Version Announcement). Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303753.

See [AAT^{+}14].

Arellano:2021:SSW

[AB21] H. F. Arellano and G. Blanchon. SWANLOP: Scattering waves off nonlocal optical potentials in the presence of Coulomb interactions. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302599.

Andersen:2022:HHE

[ABB^{+}22] Jeppe R. Andersen, James Black, Helen Brooks, Bertrand Ducloué, Marian Heil, Andreas Maier, and Jennifer M. Smillie. HEJ 2.1: High-energy resummation with vector bosons and next-to-leading logarithms. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522001230.

Alguero:2024:MCD

[ABB^{+}24] G. Alguero, G. Bélanger, F. Boudjema, S. Chakraborti, A. Goudelis, S. Kraml, A. Mjallal, and A. Pukhov. micrOMEGAs 6.0: N-component dark matter. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524000560.

Ananthanarayan:2023:FMP

[ABBD23] B. Ananthanarayan, Sumit Banik, Souvik Bera, and Sudeepan Datta. FeynGKZ: a Mathematica package for solving Feynman integrals using GKZ hypergeometric systems. Computer Physics Communications, 287(??):??, June 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000449.

Armadillo:2023:EFI

[ABD^{+}23] Tommaso Armadillo, Roberto Bonciani, Simone Devoto, Narayan Rana, and Alessandro Vicini. Evaluation of Feynman integrals with arbitrary complex masses via series expansions. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002648.

Ananthanarayan:2023:EAD

[ABF^{+}23a] B. Ananthanarayan, Souvik Bera, S. Friot, O. Marichev, and Tanay Pathak. On the evaluation of the Appell F_{2} double hypergeometric function. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003083.

Araujo:2023:LSG

[ABF^{+}23b] M. V. Araújo, M. Begalli, W. S. Freund, G. I. Gonçalves, M. Khandoga, B. Laforge, A. Leopold, J. L. Marin, B. S-M. Peralva, J. V. F. Pinto, M. S. Santos, J. M. Seixas, E. F. Simas Filho, and E. E. P. Souza. Lorenzetti Showers — a general-purpose framework for supporting signal reconstruction and triggering with calorimeters. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000164.

Ananthanarayan:2024:OWR

[ABFP24] B. Ananthanarayan, Souvik Bera, S. Friot, and Tanay Pathak. Olsson.wl and ROC2.wl: Mathematica packages for transformations of multivariable hypergeometric functions and regions of convergence for their series representations in the two variables case. Computer Physics Communications, 300(??): ??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000857.

AlAtoum:2020:ETG

[ABGD^{+}20] B. Al Atoum, S. F. Biagi, D. González-Díaz, B. J. P. Jones, and A. D. McDonald. Electron transport in gaseous detectors with a Python-based Monte Carlo simulation code. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520301533.

Athron:2023:FAC

[ABH^{+}23] Peter Athron, Adam Büchner, Dylan Harries, Wojciech Kotlarski, Dominik Stöckinger, and Alexander Voigt. FlexibleDecay: an automated calculator of scalar decay widths. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003034.

Appleby:2022:MFF

[ABK^{+}22] Robert B. Appleby, Roger J. Barlow, Dirk Krücker, James Molson, Scott Rowan, Sam Tygier, Haroon Rafique, Nicholas Walker, and Andrzej Wolski. Merlin++, a flexible and featurerich accelerator physics and particle tracking library. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003167.

Antropov:2023:TUD

[ABWZ23] S. Antropov, Sw. Banerjee, Z. Was, and J. Zaremba. TAUOLA update for decay channels with e^{+}e^{−} pairs in the final state. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522003113.

Altenmuller:2022:RPR

[ACD^{+}22] Konrad Altenmüller, Susana Cebrián, Theopisti Dafni, David Díez-Ibáñez, Javier Galán, Javier Galindo, Juan Antonio García, Igor G. Irastorza, Gloria Luzón, Cristina Margalejo, Hector Mirallas, Luis Obis, Oscar Pérez, Ke Han, Kaixiang Ni, Yann Bedfer, Barbara Biasuzzi, Esther Ferrer-Ribas, Damien Neyret, Thomas Papaevangelou, Cristian Cogollos, and Eduardo Picatoste. REST-for-physics, a ROOT-based framework for event oriented data analysis and combined Monte Carlo response. Computer Physics Communications, 273(??): ??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003933.

Alpat:2023:MCN

[ACKB23] Ali Behcet Alpat, Abdullah Coban, Hakan Kaya, and Giovanni Bartolini. MRADSIM-Converter: a new software for STEP to GDML conversion. Computer Physics Communications, 286 (??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000334.

Amaricci:2022:EPE

[ACS^{+}22] A. Amaricci, L. Crippa, A. Scazzola, F. Petocchi, G. Mazza, L. de Medici, and M. Capone. EDIpack: a parallel exact diagonalization package for quantum impurity problems. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003738.

Adam:2022:GGE

[Ada22] Jaroslav Adam. GETaLM: a generator for electron tagger and luminosity monitor for electron–proton and ion collisions. Computer Physics Communications, 272(??):??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003635.

Abreu:2021:CCF

[ADC^{+}21] S. Abreu, J. Dormans, F. Febres Cordero, H. Ita, M. Kraus, B. Page, E. Pascual, M. S. Ruf, and V. Sotnikov. Caravel: a C++ framework for the computation of multi-loop amplitudes with numerical unitarity. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001818.

Alves:2020:GGF

[ADdM20] P. R. L. Alves, L. G. S. Duarte, and L. A. C. P. da Mota. The goodness of global fitting in the reconstruction scheme. Computer Physics Communications, 248(??):??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303765.

Antoniadis:2022:UOS

[ADF^{+}22] Antonis F. Antoniadis, Dimitris Drikakis, Pericles S. Farmakis, Lin Fu, Ioannis Kokkinakis, Xesús Nogueira, Paulo A. S. F. Silva, Martin Skote, Vladimir Titarev, and Panagiotis Tsoutsanis. UCNS3D: an open-source highorder finite-volume unstructured CFD solver. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001722.

Ambrogi:2020:SVL

[ADH^{+}20] Federico Ambrogi, Juhi Dutta, Jan Heisig, Sabine Kraml, Suchita Kulkarni, Ursula Laa, Andre Lessa, Philipp Neuhuber, Humberto Reyes-González, Wolfgang Waltenberger, and Matthias Wolf. SModelS v1.2: Long-lived particles, combination of signal regions, and other novelties. Computer Physics Communications, 251(??):??, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302255.

Alexander:2023:HPS

[ADW^{+}23] M. H. Alexander, P. J. Dagdigian, H.-J. Werner, J. Kłos, B. Desrousseaux, G. Raffy, and F. Lique. Hibridon: a program suite for time-independent non-reactive quantum scattering calculations. Computer Physics Communications, 289(??): ??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001066.

Asadollahi:2023:TSE

[AE23] T. Asadollahi and N. Golshan Ebrahimi. A transient solution for electric field driven orientation and deformation of the angled droplet. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002114.

Alvarez:2021:EST

[AES21] Ezequiel Alvarez, Mariel Estévez, and Rosa María Sandá Seoane. Z′-explorer: a simple tool to probe Z′ models against LHC data. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002563.

Africa:2024:LCO

[AFB^{+}24a] Pasquale Claudio Africa, Ivan Fumagalli, Michele Bucelli, Alberto Zingaro, Marco Fedele, Luca Dede’, and Alfio Quarteroni. lifex-cfd: an open-source computational fluid dynamics solver for cardiovascular applications. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003843.

Amoroso:2024:PTT

[AFB^{+}24b] A. Amoroso, R. Baldini Ferroli, I. Balossino, M. Bertani, D. Bettoni, F. Bianchi, A. Bortone, A. Calcaterra, S. Cerioni, W. Cheng, G. Cibinetto, A. Cotta Ramusino, G. Cotto, F. Cossio, M. Da Rocha Rolo, F. De Mori, M. Destefanis, J. Dong, F. Evangelisti, R. Farinelli, L. Fava, G. Felici, I. Garzia, M. Gatta, G. Giraudo, S. Gramigna, M. Greco, L. Lavezzi, M. Maggiora, R. Malaguti, A. Mangoni, S. Marcello, M. Melchiorri, G. Mezzadri, E. Pace, S. Pacetti, P. Patteri, J. Pellegrino, A. Rivetti, M. Scodeggio, S. Sosio, and S. Spataro. PARSIFAL: a toolkit for triple-GEM parametrized simulation. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003454.

Allwicher:2023:HTH

[AFJ^{+}23] Lukas Allwicher, Darius A. Faroughy, Florentin Jaffredo, Olcyr Sumensari, and Felix Wilsch. HighPT: a tool for highp_{T} Drell–Yan tails beyond the standard model. Computer Physics Communications, 289(??):??, August 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000942.

Akbari:2021:EMT

[AG21] Amir Akbari and Dennis Giannacopoulos. An efficient multithreaded Newton–Raphson algorithm for strong coupling modeling of multi-physics problems. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302708.

Altinli:2023:ERA

[AG23] Murat Altinli and Halil Gamsizkan. ERNIE: a reactor antineutrino inverse beta decay event generator. Computer Physics Communications, 282(??):??, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002624.

Ayyad:2021:SVF

[AGH21] Mahmoud Ayyad, Amr Guaily, and Maha A. Hassanein. Stabilized variational formulation of an Oldroyd-B fluid flow equations on a Graphic Processing Unit (GPU) architecture. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302332.

Akhtar:2023:CEU

[AGJ^{+}23] Jahanzeb Akhtar, Imran Ghous, Muhammad Jawad, Zhaoxia Duan, Ikram Ullah Khosa, and Saim Ahmed. A computationally efficient unscented Kalman smoother for ameliorated tracking of subatomic particles in high energy physics experiments. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003046.

Arbuzov:2023:CPD

[AGK^{+}23] A. Arbuzov, J. Gluza, L. Kalinovskaya, S. Riemann, T. Riemann, and V. Yermolchyk. Computer package DIZET v. 6.45. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001911.

Aguado:2023:QPP

[AGMFGE23] Daniel Gómez Aguado, Vicent Gimeno, Julio José Moyano-Fernández, and Juan Carlos Garcia-Escartin. QOptCraft: a Python package for the design and study of linear optical quantum systems. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002302.

Amrozik:2023:PMC

[AHM^{+}23] Piotr Amrozik, Krzysztof Hałagan, Paulina Maczugowska, Grzegorz Jabłoński, Rafał Kiełbik, and Kamil Rudnicki. Parallel model of chemical reactions on a grained molecular level. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002291.

Adam:2021:CLB

[AHP21] Saad Adam, Farzaneh Hajabdollahi, and Kannan N. Premnath. Cascaded lattice Boltzmann modeling and simulations of three-dimensional non-Newtonian fluid flows. Computer Physics Communications, 262(??):??, May 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000230.

Albanese:2023:SGS

[AIZ23] R. Albanese, A. Iaiunese, and P. Zumbolo. Solution of Grad– Shafranov equation with linear triangular finite elements providing magnetic field continuity with a second order accuracy. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001492.

Alonso-Jorda:2021:EUD

[AJDS^{+}21] Pedro Alonso-Jordá, Davor Davidović, Marin Sapunar, José R. Herrero, and Enrique S. Quintana-Ortí. Efficient update of determinants for many-electron wave function overlaps. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302459.

Arbuzov:2021:MCP

[AJW^{+}21] A. Arbuzov, S. Jadach, Z. Was, B. F. L. Ward, and S. A. Yost. The Monte Carlo program KKMC, for the lepton or quark pair production at LEP/SLC energies-updates of electroweak calculations. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303635.

Aiko:2024:HCV

[AKK^{+}24] Masashi Aiko, Shinya Kanemura, Mariko Kikuchi, Kodai Sakurai, and Kei Yagyu. H-COUP Version 3: a program for one-loop corrected decays of any Higgs bosons in non-minimal Higgs models. Computer Physics Communications, 301(??): ??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001541.

Austin:2021:PAM

[AKL^{+}21] Anthony P. Austin, Mohan Krishnamoorthy, Sven Leyffer, Stephen Mrenna, Juliane Müller, and Holger Schulz. Practical algorithms for multivariate rational approximation. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303222.

Alguero:2021:SIP

[AKW21] Gaël Alguero, Sabine Kraml, and Wolfgang Waltenberger. A SModelS interface for pyhf likelihoods. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000485.

Albert:2022:FFE

[ALB22] Christopher G. Albert, Patrick Lainer, and Oszkár Bíró. 2D Fourier finite element formulation for magnetostatics in curvilinear coordinates with a symmetry direction. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001205.

Arrigoni:2021:SPP

[AM21a] Marco Arrigoni and Georg K. H. Madsen. Spinney: Postprocessing of first-principles calculations of point defects in semiconductors with Python. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000709.

Atisattapong:2021:WLS

[AM21b] Wanyok Atisattapong and Pasin Marupanthorn. Wang– Landau sampling for estimation of the reliability of physical networks. Computer Physics Communications, 262(??): ??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000059.

Amicarelli:2020:SVC

[AMA^{+}20] Andrea Amicarelli, Sauro Manenti, Raffaele Albano, Giordano Agate, Marco Paggi, Laura Longoni, Domenica Mirauda, Latifa Ziane, Giacomo Viccione, Sara Todeschini, Aurelia Sole, Lara Martina Baldini, Davide Brambilla, Monica Papini, Mohamed Cherif Khellaf, Bonaventura Tagliafierro, Luca Sarno, and Guido Pirovano. SPHERA v.9.0.0: a computational fluid dynamics research code, based on the smoothed particle hydrodynamics mesh-less method. Computer Physics Communications, 250(??):??, May 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300187.

Adler:2024:PQE

[AMK24] Ran Adler, Corey Melnick, and Gabriel Kotliar. Portobello — quantum embedding in correlated materials made accessible. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002527.

Aguilar:2021:MAA

[AMP^{+}21] C. Aguilar, P. Martin, E. Pio, C. Salvo, and G. O. Neves. Materials analysis applying thermodynamic (MAAT) software: a friendly and free tool to analyze the formation of solid solutions, amorphous phases and intermetallic compounds. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302757.

Anonymous:2020:Aa

[Ano20a] Anonymous. April 2020. Computer Physics Communications, 249(??):??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. August 2020. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:CBP

[Ano20c] Anonymous. Corrigendum to “Bratu’s problem: a novel approach using fixed-point iterations and Green’s function” [Comput. Phys. Comm. 198 (2016) 97–104]. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519304382. See [KK16].

Anonymous:2020:D

[Ano20d] Anonymous. December 2020. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:EBa

[Ano20e] Anonymous. Editorial Board. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303273.

Anonymous:2020:EBb

[Ano20f] Anonymous. Editorial Board. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303625.

Anonymous:2020:EBc

[Ano20g] Anonymous. Editorial Board. Computer Physics Communications, 248(??):??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519304254.

Anonymous:2020:EBd

[Ano20h] Anonymous. Editorial Board. Computer Physics Communications, 249(??):??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520300138.

Anonymous:2020:EBe

[Ano20i] Anonymous. Editorial Board. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552030059X.

Anonymous:2020:EBf

[Ano20j] Anonymous. Editorial Board. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301077.

Anonymous:2020:EBg

[Ano20k] Anonymous. Editorial Board. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301302.

Anonymous:2020:EBh

[Ano20l] Anonymous. Editorial Board. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301715.

Anonymous:2020:EBi

[Ano20m] Anonymous. Editorial Board. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301958.

Anonymous:2020:EBj

[Ano20n] Anonymous. Editorial Board. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302034.

Anonymous:2020:EBk

[Ano20o] Anonymous. Editorial Board. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302496.

Anonymous:2020:EBl

[Ano20p] Anonymous. Editorial Board. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552030309X.

Anonymous:2020:F

[Ano20q] Anonymous. February 2020. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:Ja

[Ano20r] Anonymous. January 2020. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:Jc

[Ano20s] Anonymous. July 2020. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:Jb

[Ano20t] Anonymous. June 2020. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:Ma

[Ano20u] Anonymous. March 2020. Computer Physics Communications, 248(??):??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:Mb

[Ano20v] Anonymous. May 2020. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:N

[Ano20w] Anonymous. November 2020. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:O

[Ano20x] Anonymous. October 2020. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2020:S

[Ano20y] Anonymous. September 2020. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:Ab

[Ano21b] Anonymous. August 2021. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:EBa

[Ano21d] Anonymous. Editorial Board. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303349.

Anonymous:2021:EBb

[Ano21e] Anonymous. Editorial Board. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303751.

Anonymous:2021:EBc

[Ano21f] Anonymous. Editorial Board. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303970.

Anonymous:2021:EBd

[Ano21g] Anonymous. Editorial Board. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000102.

Anonymous:2021:EBe

[Ano21h] Anonymous. Editorial Board. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000552.

Anonymous:2021:EBf

[Ano21i] Anonymous. Editorial Board. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000874.

Anonymous:2021:EBg

[Ano21j] Anonymous. Editorial Board. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001260.

Anonymous:2021:EBh

[Ano21k] Anonymous. Editorial Board. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001491.

Anonymous:2021:EBi

[Ano21l] Anonymous. Editorial Board. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001636.

Anonymous:2021:EBj

[Ano21m] Anonymous. Editorial Board. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002228.

Anonymous:2021:EBk

[Ano21n] Anonymous. Editorial Board. Computer Physics Communications, 268(??):??, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002307.

Anonymous:2021:EB

[Ano21o] Anonymous. Editorial Board. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002733.

Anonymous:2021:F

[Ano21p] Anonymous. February 2021. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:Ja

[Ano21q] Anonymous. January 2021. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:Jc

[Ano21r] Anonymous. July 2021. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:Jb

[Ano21s] Anonymous. June 2021. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:Ma

[Ano21t] Anonymous. March 2021. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:Mb

[Ano21u] Anonymous. May 2021. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:N

[Ano21v] Anonymous. November 2021. Computer Physics Communications, 268(??):??, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:O

[Ano21w] Anonymous. October 2021. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2021:S

[Ano21x] Anonymous. September 2021. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Computer Physics Communications, 281(??):??, December 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003234.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003532.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Computer Physics Communications, 272(??):??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003842.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000145.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Computer Physics Communications, 274(??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000406.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000704.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001114.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001503.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001655.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002028.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002430.

Anonymous:2022:EBl

[Ano22o] Anonymous. Editorial Board. Computer Physics Communications, 281(??):??, December 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552200251X.

Anonymous:2022:F

[Ano22p] Anonymous. February 2022. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:Ja

[Ano22q] Anonymous. January 2022. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:Jc

[Ano22r] Anonymous. July 2022. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:Jb

[Ano22s] Anonymous. June 2022. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:Ma

[Ano22t] Anonymous. March 2022. Computer Physics Communications, 272(??):??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:Mb

[Ano22u] Anonymous. May 2022. Computer Physics Communications, 274(??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:N

[Ano22v] Anonymous. November 2022. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:Oa

[Ano22w] Anonymous. October 2022. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2022:S

[Ano22x] Anonymous. September 2022. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Computer Physics Communications, 289(??):??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Computer Physics Communications, 293(??):??, December 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002946.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003393.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003708.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000231.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000528.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000784.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001133.

Anonymous:2023:EBh

[Ano23k] Anonymous. Editorial Board. Computer Physics Communications, 289(??):??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552300139X.

Anonymous:2023:EBi

[Ano23l] Anonymous. Editorial Board. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001613.

Anonymous:2023:EBj

[Ano23m] Anonymous. Editorial Board. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002230.

Anonymous:2023:EBk

[Ano23n] Anonymous. Editorial Board. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552300262X.

Anonymous:2023:EBl

[Ano23o] Anonymous. Editorial Board. Computer Physics Communications, 293(??):??, December 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002692.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:Jc

[Ano23r] Anonymous. July 2023. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:Jb

[Ano23s] Anonymous. June 2023. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:Ma

[Ano23t] Anonymous. March 2023. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:Mb

[Ano23u] Anonymous. May 2023. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:PN

[Ano23v] Anonymous. November 2023. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2023:S

[Ano23x] Anonymous. September 2023. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:EBa

[Ano24c] Anonymous. Editorial Board. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003235.

Anonymous:2024:EBb

[Ano24d] Anonymous. Editorial Board. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003880.

Anonymous:2024:EBc

[Ano24e] Anonymous. Editorial Board. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000043.

Anonymous:2024:EBd

[Ano24f] Anonymous. Editorial Board. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000286.

Anonymous:2024:EBe

[Ano24g] Anonymous. Editorial Board. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000675.

Anonymous:2024:EBf

[Ano24h] Anonymous. Editorial Board. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001000.

Anonymous:2024:EBg

[Ano24i] Anonymous. Editorial Board. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001474.

Anonymous:2024:EBh

[Ano24j] Anonymous. Editorial Board. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001784.

Anonymous:2024:EBi

[Ano24k] Anonymous. Editorial Board. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001954.

Anonymous:2024:F

[Ano24l] Anonymous. February 2024. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:Ja

[Ano24m] Anonymous. January 2024. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:Jc

[Ano24n] Anonymous. July 2024. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:Jb

[Ano24o] Anonymous. June 2024. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:Ma

[Ano24p] Anonymous. March 2024. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:Mb

[Ano24q] Anonymous. May 2024. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:2024:S

[Ano24r] Anonymous. September 2024. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Angelidakis:2021:SAP

[ANU21] Vasileios Angelidakis, Sadegh Nadimi, and Stefano Utili.

SHape Analyser for Particle Engineering (SHAPE): Seam-

less characterisation and simplification of particle morphology from imaging data. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000953.

Allmann-Rahn:2024:MIC

[ARLDG24] F. Allmann-Rahn, S. Lautenbach, M. Deisenhofer, and R. Grauer. The muphy II code: Multiphysics plasma simulation on large HPC systems. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523004095.

Arnoldus:2020:NES

[Arn20] Henk F. Arnoldus. Numerical evaluation of Sommerfeldtype integrals for reflection and transmission of dipole radiation. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302411.

Amsler:2020:FLA

[ART^{+}20] Maximilian Amsler, Samare Rostami, Hossein Tahmasbi, Ehsan Rahmatizad Khajehpasha, Somayeh Faraji, Robabe Rasoulkhani, and S. Alireza Ghasemi. FLAME: a library of atomistic modeling environments. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520301855.

Akhmatskaya:2022:NRP

[AS22]  Elena Akhmatskaya and Dmitri Sokolovski.  Numeri-

cal Regge pole analysis of resonance structures in stateto-state reactive differential cross sections. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000893.

AlAli:2024:EII

[AS24a] Musaddiq Al Ali and Masatoshi Shimoda. Exploring the influence of initial design domain dependencies in concurrent multiscale topology optimization for heat conductivity maximization. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003132.

Ataei:2024:XDM

[AS24b] Mohammadmehdi Ataei and Hesam Salehipour. XLB: a differentiable massively parallel lattice Boltzmann library in Python. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001103.

Alba:2022:BCE

[ASA^{+}22] Arnau Albà, Jimin Seok, Andreas Adelmann, Scott Doran, Gwanghui Ha, Soonhong Lee, Yinghu Piao, John Power, Maofei Qian, Eric Wisniewski, Joseph Xu, and Alexander Zholents. Benchmarking collective effects of electron interactions in a wiggler with OPAL-FEL. Computer Physics Communications, 280(??):??, November 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001941.

Ataei:2021:LOS

[ASC^{+}21] Mohammadmehdi Ataei, Vahid Shaayegan, Franco Costa, Sejin Han, Chul B. Park, and Markus Bussmann. LBfoam: an open-source software package for the simulation of foaming using the Lattice Boltzmann Method. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030343X.

Aunai:2024:PPH

[ASC^{+}24] Nicolas Aunai, Roch Smets, Andrea Ciardi, Philip Deegan, Alexis Jeandet, Thibault Payet, Nathan Guyot, and Loic Darrieumerlou. PHARE: Parallel hybrid particle-in-cell code with patch-based adaptive mesh refinement. Computer Physics Communications, 295(??):??, February 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003119.

Almansa:2021:PGP

[ASPDL^{+}21] Julio Almansa, Francesc Salvat-Pujol, Gloria Díaz-Londoño, Artur Carnicer, Antonio M. Lallena, and Francesc Salvat. pkgPENGEOM — a general-purpose geometry package for Monte Carlo simulation of radiation transport in complex material structures (new version announcement). Computer Physics Communications, 264(??):??, July 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000795.

Abhishek:2024:TCS

[ASS^{+}24] Abhishek, Paul Stevenson, Yue Shi, Esra Yüksel, and A. S. Umar. The TDHF code Sky3D version 1.2. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001620.

Alber:2021:NOS

[ASU^{+}21] Lukas Alber, Valentino Scalera, Vivek Unikandanunni, Daniel Schick, and Stefano Bonetti. NTMpy: an open source package for solving coupled parabolic differential equations in the framework of the three-temperature model. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001028.

Arguelles:2020:SQI

[ASW20] Carlos A. Argüelles, Jordi Salvado, and Christopher N. Weaver. A Simple Quantum Integro-Differential Solver (SQuIDS). Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301867.

Arguelles:2022:NTN

[ASW22] Carlos A. Argüelles, Jordi Salvado, and Christopher N. Weaver. nuSQuIDS: a toolbox for neutrino propagation. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000649.

Arun:2023:IAT

[ATC^{+}23] A. D. Arun, S. Thirunavukkarasu, Sharat Chandra, M. P. Rajiniganth, N. Malathi, and M. Sivaramakrishna. Investigating algebraic topological method for solving 3D Laplace equation in electrostatic boundary value problems. Computer Physics Communications, 289(??):??, August 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001042.

Arthuis:2021:AAG

[ATRD21] P. Arthuis, A. Tichai, J. Ripoche, and T. Duguet. ADG: Automated generation and evaluation of many-body diagrams II. Particle-number projected Bogoliubov many-body perturbation theory. Computer Physics Communications, 261(??): ??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303295.

Akamatsu:2024:GAC

[AUEO24] Takanori Akamatsu, Mitsuharu Uemoto, Yoshiyuki Egami, and Tomoya Ono. GPU acceleration of conjugate gradient method obtaining Green’s function for transport-property calculation. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552300334X.

Alldritt:2022:ATF

[AUO^{+}22] Benjamin Alldritt, Fedor Urtev, Niko Oinonen, Markus Aapro, Juho Kannala, Peter Liljeroth, and Adam S. Foster. Automated tip functionalization via machine learning in scanning probe microscopy. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003702.

Akkurt:2022:CBS

[AWV22] Semih Akkurt, Freddie Witherden, and Peter Vincent. Cache blocking strategies applied to flux reconstruction. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003052.

Akkoyun:2024:GFF

[AYB24] Serkan Akkoyun, Cafer Mert Yesilkanat, and Tuncay Bayram. Generation of fusion and fusion-evaporation reaction crosssections by two-step machine learning methods. Computer Physics Communications, 297(??):??, April 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523004009.

Aoyama:2024:HWP

[AYI^{+}24] Tatsumi Aoyama, Kazuyoshi Yoshimi, Kota Ido, Yuichi Motoyama, Taiki Kawamura, Takahiro Misawa, Takeo Kato, and Akito Kobayashi. H-wave — a Python package for the Hartree–Fock approximation and the random phase approximation. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524000109.

Au-Yeung:2024:QAS

[AYWKL24] R. Au-Yeung, A. J. Williams, V. M. Kendon, and S. J. Lind. Quantum algorithm for smoothed particle hydrodynamics. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002540.

Abdullahi:2024:DPB

[AZH^{+}24] Asli M. Abdullahi, Jaime Hoefken Zink, Matheus Hostert, Daniele Massaro, and Silvia Pascoli. DarkNews: a Pythonbased event generator for heavy neutral lepton production in neutrino-nucleus scattering. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523004204.

Bagci:2024:BDM

[BA24] A. Bagci and Gustavo A. Aucar. A bi-directional method for evaluating integrals involving higher transcendental functions. HyperRAF: a Julia package for new hyper-radial functions. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003351.

Brown:2020:RMT

[BAB^{+}20] Andrew C. Brown, Gregory S. J. Armstrong, Jakub Benda, Daniel D. A. Clarke, Jack Wragg, Kathryn R. Hamilton, Zdeněk Mašín, Jimena D. Gorfinkiel, and Hugo W. van der Hart. RMT: R-matrix with time-dependence. solving the semi-relativistic, time-dependent Schrödinger equation for general, multielectron atoms and molecules in intense, ultrashort, arbitrarily polarized laser pulses. Computer Physics Communications, 250(??):??, May 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303856.

Bagci:2022:JJP

[Bag22] Ali Bagci. JRAF: a Julia package for computation of relativistic molecular auxiliary functions. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552100388X.

Bakhvalov:2023:CCE

[Bak23] P. Bakhvalov. ColESo: Collection of exact solutions for verification of numerical algorithms for simulation of compressible flows. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002612.

Bardsley:2022:OBA

[Bar22a] Michael Bardsley. An optimisation based algorithm for finding the nucleation temperature of cosmological phase transitions. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003647.

Bartos:2022:NMI

[Bar22b] Erik Bartos. Numerical multidimensional integration with PyMikor. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002617.

Bartsch:2021:MMC

[BB21] Jan Bartsch and Alfio Borzì. MOCOKI: a Monte Carlo approach for optimal control in the force of a linear kinetic model. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521001429.

Bati:2024:RPP

[BB24a] Ajay Bati and Spencer H. Bryngelson. RoseNNa: a performant, portable library for neural network inference with application to computational fluid dynamics. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003971.

Bhat:2024:MSL

[BB24b] Nayana G. Bhat and S. Balaji. Modeling and simulation of Lac–Operon using reaction-diffusion master equation on heterogeneous parallel platforms. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003636.

Busa:2020:PPE

[BBA^{+}20] Ján Buša, Ján Buša, Edik Ayryan, Shura Hayryan, Chin-Kun Hu, Imrich Pokorný, and Jaroslav Skřivánek. PBCAVE: Program for exact classification of the mesh points of a protein with possible internal cavities and its application to Poisson–Boltzmann equation solution. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303406.

Bezgin:2023:JFF

[BBA23] Deniz A. Bezgin, Aaron B. Buhendwa, and Nikolaus A. Adams. JAX-Fluids: a fully-differentiable high-order computational fluid dynamics solver for compressible two-phase flows. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002466.

Bottin:2020:TTD

[BBB20] François Bottin, Jordan Bieder, and Johann Bouchet. a-TDEP: Temperature dependent effective potential for Abinit — lattice dynamic properties including anharmonicity. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301156.

Bertacchi:2021:TFB

[BBB^{+}21] Valerio Bertacchi, Tadeas Bilka, Nils Braun, Giulia Casarosa, Luigi Corona, Sam Cunliffe, Filippo Dattola, Gaetano De Marino, Michael De Nuccio, Giacomo De Pietro, Thanh Van Dong, Giulio Dujany, Patrick Ecker, Michael Eliachevitch, Tristan Fillinger, Oliver Frost, Rudolf Frühwirth, Uwe Gebauer, Sasha Glazov, Nicolas Gosling, Aiqiang Guo, Thomas Hauth, Martin Heck, Mateusz Kaleta, Jakub Kandra, Claus Kleinwort, Thomas Kuhr, Simon Kurz, Peter Kvasnicka, Jakob Lettenbichler, Thomas Lueck, Alberto Martini, Felix Metzner, Dmitrii Neverov, Carsten Niebuhr, Eugenio Paoloni, Sourav Patra, Leo Piilonen, Cyrille Praz, Markus Tobias Prim, Christian Pulvermacher, Sebastian Racs, Navid Rad, Petar Rados, Martin Ritter, Giuliana Rizzo, Armine Rostomyan, Bianca Scavino, Tobias Schlüter, Benjamin Schwenker, Stefano Spataro, Björn Spruck, Henrikas Svidras, Francesco Tenchini, Yuma Uematsu, James Webb, Christian Wessel, and Laura Zani. Track finding at Belle II. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302861.

Boguslawski:2024:PIF

[BBC^{+}24] Katharina Boguslawski, Filip Brzek, Rahul Chakraborty, Kacper Cieślak, Seyedehdelaram Jahani, Aleksandra Leszczyk, Artur Nowak, Emil Sujkowski, Julian Świerczyński, Somayeh Ahmadkhani, Dariusz Kedziera, Maximilian H. Kriebel, Piotr Szymon Żuchowski, and Paweł Tecmer. PyBEST: Improved functionality and enhanced performance. Computer Physics Communications, 297(??):??, April 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003946.

Bahl:2023:HBS

[BBH^{+}23] Henning Bahl, Thomas Biekötter, Sven Heinemeyer, Cheng Li, Steven Paasch, Georg Weiglein, and Jonas Wittbrodt. HiggsTools: BSM scalar phenomenology with new versions of HiggsBounds and HiggsSignals. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001480.

Bendle:2024:PPS

[BBH^{+}24] Dominik Bendle, Janko Boehm, Murray Heymann, Rourou Ma, Mirko Rahn, Lukas Ristau, Marcel Wittmann, Zihao Wu, Hefeng Xu, and Yang Zhang. pfd-parallel, a Singular/ GPI-Space package for massively parallel multivariate partial fractioning. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002874.

Ameur:2022:RAA

[BBV^{+}22] Firas Ben Ameur, Joachim Balis, Ray Vandenhoeck, Andrea Lani, and Stefaan Poedts. r-Adaptive algorithms for supersonic flows with high-order Flux Reconstruction methods. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000923.

Bhattacharjee:2023:AAC

[BBV23] Himaghna Bhattacharjee, Jackson Burns, and Dionisios G. Vlachos. AIMSim: an accessible cheminformatics platform for similarity operations on chemicals datasets. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002983.

Bhom:2020:HOS

[BC20] Jihyun Bhom and Marcin Chrzaszcz. HEPLike: an open source framework for experimental likelihood evaluation. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300692.

Barnes:2021:FSG

[BC21] D. C. Barnes and L. Chacón. Finite spatial-grid effects in energy-conserving particle-in-cell algorithms. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030268X.

Budiardja:2022:PGB

[BC22] Reuben D. Budiardja and Christian Y. Cardall. GenASiS Basics: Object-oriented utilitarian functionality for largescale physics simulations (version 4). Computer Physics Communications, 281(??):??, December 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002247.

Barontini:2024:PIH

[BCCM^{+}24] Andrea Barontini, Alessandro Candido, Juan M. Cruz-Martinez, Felix Hekhorn, and Christopher Schwan. Pineline: Industrialization of high-energy theory predictions. Computer Physics Communications, 297(??):??, April 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300406X.

Bondi:2021:FGC

[BCD^{+}21] M. Bondi, A. Celentano, R. R. Dusaev, D. V. Kirpichnikov, M. M. Kirsanov, N. V. Krasnikov, L. Marsicano, and D. Shchukin. Fully Geant4 compatible package for the simulation of Dark Matter in fixed target experiments. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002411.

Beever:2024:TMC

[BCF^{+}24] Z. Beever, D. Caratelli, A. Fava, F. Pietropaolo, F. Stocker, and J. Zettlemoyer. TRANSLATE — a Monte Carlo simulation of electron transport in liquid argon. Computer Physics Communications, 297(??):??, April 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523004010.

Blanes:2024:GSM

[BCGT24] S. Blanes, F. Casas, C. González, and M. Thalhammer. Generalisation of splitting methods based on modified potentials to nonlinear evolution equations of parabolic and Schrödinger type. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003521.

Bai:2021:IPC

[BCHE21] Jinwei Bai, Yong Cao, Xiaoming He, and Peng E. An implicit particle-in-cell model based on anisotropic immersed-finite-element method. Computer Physics Communications, 261(??): ??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303180.

Buchwald:2024:SCR

[BCSS24] Simon Buchwald, Gabriele Ciaramella, Julien Salomon, and Dominique Sugny. A SPIRED code for the reconstruction of spin distribution. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000493.

Broz:2020:GFN

[BCT20] M. Broz, J. G. Contreras, and J. D. Tapia Takaki. A generator of forward neutrons for ultra-peripheral collisions: nOOn. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300321.

Blanchet:2022:EFP

[BCTS22] A. Blanchet, J. Clérouin, M. Torrent, and F. Soubiran. Extended first-principles molecular dynamics model for high temperature simulations in the Abinit code: Application to warm dense aluminum. Computer Physics Communications, 271(??): ??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003271.

Blanchon:2020:SNN

[BDA^{+}20] G. Blanchon, M. Dupuis, H. F. Arellano, R. N. Bernard, and B. Morillon. SIDES: Nucleon–nucleus elastic scattering code for nonlocal potential. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520301478.

Braz:2020:GFM

[BDdM20] A. Braz, L. G. S. Duarte, and L. A. C. P. da Mota. A generalization of the S-function method applied to a Duffing–van der Pol forced oscillator. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552030117X.

Billoir:2021:PKF

[BDGS21] P. Billoir, M. De Cian, P. A. Günther, and S. Stemmle. A parametrized Kalman filter for fast track fitting at LHCb. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001387.

Bondarenko:2023:HHC

[BDK^{+}23] Serge Bondarenko, Yahor Dydyshka, Lidia Kalinovskaya, Renat Sadykov, and Vitaly Yermolchyk. Hadron–hadron collision mode in ReneSANCe-v1.3.0. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003654.

Benfatto:2021:MNP

[BDP^{+}21] Maurizio Benfatto, Stefano Della Longa, Elisabetta Pace, Giovanni Chillemi, Cristiano Padrin, Calogero R. Natoli, and Nico Sanna. MXAN: a new program for ab-initio structural quantitative analysis of XANES experiments. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001041.

Butler:2020:HPD

[BDR^{+}20] R. Butler, T. Dodwell, A. Reinarz, A. Sandhu, R. Scheichl, and L. Seelinger. High-performance dune modules for solving large-scale, strongly anisotropic elliptic problems with applications to aerospace composites. Computer Physics Communications, 249(??):??, April 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303364.

Belyaev:2024:CEA

[Bel24] Mikhail A. Belyaev. Causal explicit algorithm for heat conduction in a plasma. Computer Physics Communications, 294(??): ??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002795.

Bera:2024:MMP

[Ber24] Souvik Bera. MultiHypExp: a Mathematica package for expanding multivariate hypergeometric functions in terms of multiple polylogarithms. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523004058.

Beutel:2024:TMB

[BFCR24] Dominik Beutel, Ivan Fernandez-Corbaton, and Carsten Rockstuhl. treams — a T -matrix-based scattering code for nanophotonics. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523004216.

Bertini:2022:EGC

[BFD22] Marco Bertini, Francesco Ferrante, and Dario Duca. Empathes: a general code for nudged elastic band transition states search. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003362.

Bonfa:2021:UOS

[BFI^{+}21] Pietro Bonfà, Jonathan Frassineti, Muhammad Maikudi Isah, Ifeanyi John Onuorah, and Samuele Sanna. UNDI: an opensource library to simulate muon-nuclear interactions in solids. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303556.

Bierwage:2022:RMC

[BFL^{+}22] Andreas Bierwage, Michael Fitzgerald, Philipp Lauber, Mirko Salewski, Yevgen Kazakov, and Žiga Štancar. Representation and modeling of charged particle distributions in tokamaks. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000236.

Borras:2024:PCS

[BFMA^{+}24] Vicent J. Borràs, Pedro Fernández-Milán, Luca Argenti, Jesús González-Vázquez, and Fernando Martín. Photoionization cross sections and photoelectron angular distributions of molecules with XCHEM-2.0. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003788.

Bentine:2020:PLP

[BFT20] E. Bentine, C. J. Foot, and D. Trypogeorgos. (py)LIon: a package for simulating trapped ion trajectories. Computer Physics Communications, 253(??):??, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300369.

Barash:2022:CEM

[BGH22] Lev Barash, Stefan Güttel, and Itay Hen. Calculating elements of matrix functions using divided differences. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003313.

Blackmore:2023:DPP

[BGHC23] Jacob A. Blackmore, Philip D. Gregory, Jeremy M. Hutson, and Simon L. Cornish. Diatomic-py: a Python module for calculating the rotational and hyperfine structure of 1Σ molecules. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002314.

Byun:2023:ISS

[BGR23] Eun-Kyu Byun, Gibeom Gu, and Hoon Ryu. Improving strong scalability of electronic structure simulations with reduced overhead of communications. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000929.

Barnett:2022:RTC

[BGW^{+}22] R. L. Barnett, D. L. Green, C. L. Waters, J. D. Lore, D. N. Smithe, and J. R. Myra. RF-transpond: a 1D coupled cold plasma wave and plasma transport model for ponderomotive force driven density modification parallel to B_{0}. Computer Physics Communications, 274(??):??, May 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000042.

Bahl:2020:PCM

[BHH^{+}20] H. Bahl, T. Hahn, S. Heinemeyer, W. Hollik, S. Paßehr, H. Rzehak, and G. Weiglein. Precision calculations in the MSSM Higgs-boson sector with FeynHiggs 2.14. Computer Physics Communications, 249(??):??, April 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519304059.

Boddy:2021:MMA

[BHK^{+}21] Kimberly K. Boddy, Stephen Hill, Jason Kumar, Pearl Sandick, and Barmak Shams Es Haghi. MADHAT: Modelagnostic dark halo analysis tool. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520304100.

Biekotter:2021:MPP

[Bie21] Thomas Biekötter. munuSSM: a Python package for the μ-from-ν Supersymmetric Standard Model. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000655.

Buckley:2021:HER

[BIK^{+}21] Andy Buckley, Philip Ilten, Dmitri Konstantinov, Leif Lönnblad, James Monk, Witold Pokorski, Tomasz Przedzinski, and Andrii Verbytskyi. The HepMC3 event record library for Monte Carlo event generators. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301181.

Bae:2024:TCS

[BJL^{+}24] Cheonho Bae, Yifei Jin, Bo Lyu, Jia Fu, Fudi Wang, and Hongming Zhang. TransROTA: a code for solving collisionalityextended Braginskii’s closure formalism for toroidally-rotating plasmas with nonlinear successive overrelaxation. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003375.

Benjamin:2023:DTG

[BJS^{+}23] Stuart Benjamin, Henrik Järleblad, Mirko Salewski, Luke Stagner, Matthew Hole, and David Pfefferlé. Distribution transforms for guiding center orbit coordinates in axisymmetric tokamak equilibria. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523002382.

Barcelos:2021:CPP

[BKB^{+}21] Erika I. Barcelos, Shaghayegh Khani, Arman Boromand, Luiz F. Vieira, J. Alex Lee, Jeffrey Peet, Mônica F. Naccache, and Joao Maia. Controlling particle penetration and depletion at the wall using Dissipative Particle Dynamics. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302964.

Burmasov:2022:UMC

[BKBL22] Nazar Burmasov, Evgeny Kryshen, Paul Bühler, and Roman Lavicka. Upcgen: a Monte Carlo simulation program for dilepton pair production in ultra-peripheral collisions of heavy ions. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001072.

Byer:2023:SRE

[BKG^{+}23] Duane Byer, Vladimir Khachatryan, Haiyan Gao, Igor Akushevich, Alexander Ilyichev, Chao Peng, Alexei Prokudin, Stan Srednyak, and Zhiwen Zhao. SIDIS-RC EvGen: a Monte-Carlo event generator of semi-inclusive deep inelastic scattering with the lowest-order QED radiative corrections. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000474.

Blytt:2020:ETD

[BKO20] M. Blytt, M. Kachelrieß, and S. Ostapchenko. ELMAG 3.01: a three-dimensional Monte Carlo simulation of electromagnetic cascades on the extragalactic background light and in magnetic fields. Computer Physics Communications, 252(??): ??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300217.

Bak:2022:DPC

[BKP22] Soyoon Bak, Philsu Kim, and Sangbeom Park. Development of a parallel CUDA algorithm for solving 3D guiding center problems. Computer Physics Communications, 276(??): ??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000492.

Banger:2022:FFP

[BKRG22] Paramjeet Banger, Pardeep Kaur, Arko Roy, and Sandeep Gautam. FORTRESS: FORTRAN programs to solve coupled Gross–Pitaevskii equations for spin-orbit coupled spinf Bose–Einstein condensate with spin f = 1 or 2. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001618.

Bialas:2022:HAN

[BKS22] Piotr Białas, Piotr Korcyl, and Tomasz Stebel. Hierarchical autoregressive neural networks for statistical systems. Computer Physics Communications, 281(??):??, December 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002211.

Bialas:2024:TNF

[BKS24a] Piotr Białas, Piotr Korcyl, and Tomasz Stebel. Training normalizing flows with computationally intensive target probability distributions. Computer Physics Communications, 298(??): ??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000171.

Bondarenko:2024:MCT

[BKS24b] Sergey G. Bondarenko, Lidia V. Kalinovskaya, and Andrey A. Sapronov. Monte-Carlo tool SANCphot for polarized γ–γ collision simulation. Computer Physics Communications, 294(??): ??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002746.

BenAmeur:2021:AAH

[BL21] Firas Ben Ameur and Andrea Lani. r-Adaptive algorithms for high-speed flows and plasma simulations. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303441.

Brunekreef:2024:SCQ

[BLG24] Joren Brunekreef, Renate Loll, and Andrzej Görlich. Simulating CDT quantum gravity. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000936.

Bonaccorso:2022:LHP

[BLM^{+}22] Fabio Bonaccorso, Marco Lauricella, Andrea Montessori, Giorgio Amati, Massimo Bernaschi, Filippo Spiga, Adriano Tiribocchi, and Sauro Succi. LBcuda: a high-performance CUDA port of LBsoft for simulation of colloidal systems. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000996.

Boguslawski:2021:PBB

[BLN^{+}21] Katharina Boguslawski, Aleksandra Leszczyk, Artur Nowak, Filip Brzek, Piotr Szymon Żuchowski, Dariusz Kedziera, and Paweł Tecmer. Pythonic black-box electronic structure tool (PyBEST). An open-source Python platform for electronic structure calculations at the interface between chemistry and physics. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000643.

Badia:2020:TDI

[BM20] Santiago Badia and Alberto F. Martín. A tutorial-driven introduction to the parallel finite element library FEMPAR v1.0.0. Computer Physics Communications, 248(??):??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519303832.

Bishop-VanHorn:2022:PSO

[BM22] Logan Bishop-Van Horn and Kathryn A. Moler. SuperScreen: an open-source package for simulating the magnetic response of two-dimensional superconducting devices. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001837.

Bevan:2023:SSL

[BMI23] Kirk H. Bevan, Botong Miao, and Asif Iqbal. SLJCompact: a semiconductor-liquid junction solver for rapid band diagram insights into photoelectrochemical devices. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003575.

Basler:2021:BVT

[BMM21] Philipp Basler, Margarete Mühlleitner, and Jonas Müller. BSMPT v2 a tool for the electroweak phase transition and the baryon asymmetry of the universe in extended Higgs Sectors. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002368.

Benatto:2024:TSV

[BMP^{+}24] Leandro Benatto, Omar Mesquita, Kaike R. M. Pacheco, Lucimara S. Roman, Marlus Koehler, Rodrigo B. Capaz, and Graziâni Candiotto. TMM-Sim: a versatile tool for optical simulation of thin-film solar cells. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524001292.

Benatto:2023:FCF

[BMR^{+}23] Leandro Benatto, Omar Mesquita, João L. B. Rosa, Lucimara S. Roman, Marlus Koehler, Rodrigo B. Capaz, and Graziâni Candiotto. FRET-Calc: a free software and web server for Förster Resonance Energy Transfer Calculation. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000607.

Benatto:2024:PSC

[BMR^{+}24a] Leandro Benatto, Omar Mesquita, Lucimara S. Roman, Rodrigo B. Capaz, Graziâni Candiotto, and Marlus Koehler. PLQ-sim: a computational tool for simulating photoluminescence quenching dynamics in organic donor/acceptor blends. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003600.

Benatto:2024:RCU

[BMR^{+}24b] Leandro Benatto, Omar Mesquita, Lucimara S. Roman, Marlus Koehler, Rodrigo B. Capaz, and Graziâni Candiotto. RI-Calc: a user friendly software and web server for refractive index calculation. Computer Physics Communications, 298(??): ??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000237.

Barnes:2021:MDI

[BMREC21] Taylor A. Barnes, Eliseo Marin-Rimoldi, Samuel Ellis, and T. Daniel Crawford. The MolSSI Driver Interface Project: a framework for standardized, on-the-fly interoperability between computational molecular sciences codes. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303386.

Bernardini:2023:SST

[BMS^{+}23] Matteo Bernardini, Davide Modesti, Francesco Salvadore, Srikanth Sathyanarayana, Giacomo Della Posta, and Sergio Pirozzoli. STREAmS-2.0: Supersonic turbulent accelerated Navier–Stokes solver version 2.0. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522003630.

Bernardini:2021:SHF

[BMSP21] Matteo Bernardini, Davide Modesti, Francesco Salvadore, and Sergio Pirozzoli. STREAmS: a high-fidelity accelerated solver for direct numerical simulation of compressible turbulent flows. Computer Physics Communications, 263(??): ??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000473.

Bonaccorso:2020:LPO

[BMT^{+}20] Fabio Bonaccorso, Andrea Montessori, Adriano Tiribocchi, Giorgio Amati, Massimo Bernaschi, Marco Lauricella, and Sauro Succi. LBsoft: a parallel open-source software for simulation of colloidal systems. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520302137.

Borinsky:2023:TFI

[BMT23] Michael Borinsky, Henrik J. Munch, and Felix Tellander. Tropical Feynman integration in the Minkowski regime. Computer Physics Communications, 292(??):??, November 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002199.

Bjelcic:2020:IQF

[BN20] A. Bjelcić and T. Niksić. Implementation of the quasiparticle finite amplitude method within the relativistic selfconsistent mean-field framework: the program DIRQFAM. Computer Physics Communications, 253(??):??, August 2020.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300345.

Bjelcic:2023:IQF

[BN23] A. Bjelcić and T. Niksić. Implementation of the quasiparticle finite amplitude method within the relativistic selfconsistent mean-field framework (II): the program DIRQFAM v2.0.0. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000346.

Bjelcic:2022:CKP

[BND22] A. Bjelcić, T. Niksić, and Z. Drmac. Chebyshev kernel polynomial method for efficient calculation of the quasiparticle random phase approximation response function. Computer Physics Communications, 280(??):??, November 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001965.

Bala:2024:CPC

[BOSM24] K. A. Bala, M. R. Omar, John Y. H. Soo, and W. M. H. Wan Mokhtar. CNUCTRAN: a program for computing final nuclide concentrations using a direct simulation approach. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001814.

Brooks:2021:EMJ

[BP21] Helen Brooks and Christian T. Preuss. Efficient multijet merging with the Vincia sector shower. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000977.

Bernaschi:2024:QSH

[BPMMP24] Massimo Bernaschi, Isidoro González-Adalid Pemartín, Víctor Martín-Mayor, and Giorgio Parisi. The QISG suite: Highperformance codes for studying quantum Ising spin glasses.

Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000249.

Bircher:2020:WLT

[BR20a] Martin P. Bircher and Ursula Rothlisberger. From a week to less than a day: Speedup and scaling of coordinatescaled exact exchange calculations in plane waves. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302942.

Boffi:2020:PTD

[BR20b] Nicholas M. Boffi and Chris H. Rycroft. Parallel threedimensional simulations of quasi-static elastoplastic solids. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300795.

Ball:2023:ACC

[BRAC23] Akash Kumar Ball, Swati Rana, Gargi Agrahari, and Abhijit Chatterjee. Accelerated calculation of configurational free energy using a combination of reverse Monte Carlo and neural network models: Adsorption isotherm for 2D square and triangular lattices. Computer Physics Communications, 285(??): ??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003733.

Brouwer:2021:COP

[BRHT21] Nils Brouwer, Vanina Recoules, Natalie Holzwarth, and Marc Torrent. Calculation of optical properties with spinorbit coupling for warm dense matter. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521001417.

Bruno:2023:FCA

[BS23] Mattia Bruno and Rainer Sommer. On fits to correlated and auto-correlated data. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003629.

Bryngelson:2021:MOS

[BSC^{+}21] Spencer H. Bryngelson, Kevin Schmidmayer, Vedran Coralic, Jomela C. Meng, Kazuki Maeda, and Tim Colonius. MFC: an open-source high-order multi-component, multi-phase, and multi-scale compressible flow solver. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520301818.

Badal:2021:MBT

[BSG^{+}21] Andreu Badal, Diksha Sharma, Christian G. Graff, Rongping Zeng, and Aldo Badano. Mammography and breast tomosynthesis simulator for virtual clinical trials. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303891.

Boom:2022:PDE

[BSK^{+}22] Pieter D. Boom, Ashley Seepujak, Odysseas Kosmas, Lee Margetts, and Andrey Jivkov. Parallelized discrete exterior calculus for three-dimensional elliptic problems. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001758.

Bonaldo:2023:TIE

[BSS^{+}23] Luca Bonaldo, Terry Ethan Stearns, Ilaria Siloi, Nicholas A. Mecholsky, and Marco Fornari. Tuning and interpretation of electronic transport properties with. Computer Physics Communications, 292(??):??, November 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002023.

Brandao:2022:PQN

[BTG22] I. Brandão, D. Tandeitnik, and T. Guerreiro. QuGIT: a numerical toolbox for Gaussian quantum states. Computer Physics Communications, 280(??):??, November 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001904.

Berendeev:2024:ECC

[BTK24] E. A. Berendeev, I. V. Timofeev, and V. A. Kurshakov. Energy and charge conserving semi-implicit particle-in-cell model for simulations of high-pressure plasmas in magnetic traps. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552300365X.

Brock:2020:MBA

[BTW20] C. N. Brock, A. R. Tackett, and D. G. Walker. Metric based on the arctangents of the logderivatives for evaluating scattering properties of pseudopotentials. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303017.

Bulanchuk:2021:DFB

[Bul21] Pavlo Bulanchuk. On the delta function broadening in the Kubo–Greenwood equation. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303532.

Baeyens:2021:FAC

[BV21] Toon Baeyens and Marnix Van Daele. The fast and accurate computation of eigenvalues and eigenfunctions of timeindependent one-dimensional Schrödinger equations. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030271X.

Bay:2022:PSV

[BVV22] Mélanie M. Bay, Silvia Vignolini, and Kevin Vynck. PyLlama: a stable and versatile Python toolkit for the electromagnetic modelling of multilayered anisotropic media. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003684.

Bu:2023:ESG

[BW23] Ling-Ze Bu and Wei Wang. Efficient single-grid and multigrid solvers for real-space orbital-free density functional theory. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523001236.

Besier:2020:RSP

[BWW20] Marco Besier, Pascal Wasser, and Stefan Weinzierl. RationalizeRoots: Software package for the rationalization of square roots. Computer Physics Communications, 253(??):??, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300394.

Brandstetter:2021:KPP

[BYL^{+}21] Dominik Brandstetter, Xiaosheng Yang, Daniel Lüftner, F. Stefan Tautz, and Peter Puschnig. kMap.py: a Python program for simulation and data analysis in photoemission tomography. Computer Physics Communications, 263(??): ??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000461.

Bao:2023:PEN

[BYWW23] Nana Bao, Xingting Yan, Shiwen Wei, and Zihao Wang. Py-EFIT: a new Python package for plasma equilibrium reconstruction on EAST tokamak. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002685.

Bzowski:2021:TMP

[Bzo21] Adam Bzowski. TripleK: a Mathematica package for evaluating triple-K integrals and conformal correlation functions. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302563.

Campoli:2022:UUD

[CAC^{+}22] L. Campoli, A. Assonitis, M. Ciallella, R. Paciorri, A. Bonfiglioli, and M. Ricchiuto. UnDiFi-2D: an unstructured discontinuity fitting code for 2D grids. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521003143.

Castro:2024:QPQ

[Cas24] Alberto Castro. qocttools: a program for quantum optimal control calculations. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003284.

Cao:2022:OSC

[CAWK22] Z. Cao, M. B. Agir, C. White, and K. Kontis. An open source code for two-phase rarefied flows: rarefiedMultiphaseFoam. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000571.

Coles:2020:OSA

[CB20] Jonathan P. Coles and Rebekka Bieri. An optimizing symbolic algebra approach for generating fast multipole method operators. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303960.

Cardall:2023:GMO

[CB23] Christian Y. Cardall and Reuben D. Budiardja. GenASiS Mathematics: Object-oriented manifolds, operations, and solvers for large-scale physics simulations (version 2). Computer Physics Communications, 282(??):??, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002375.

Caulk:2020:AYP

[CCC20] Robert A. Caulk, Emanuele Catalano, and Bruno Chareyre. Accelerating Yade’s poromechanical coupling with matrix factorization reuse, parallel task management, and GPU computing. Computer Physics Communications, 248(??):??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519303340.

Chamon:2021:APP

[CCG21] L. C. Chamon, B. V. Carlson, and L. R. Gasques. São Paulo potential version 2 (SPP2) and Brazilian nuclear potential (BNP). Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001739.

Chen:2023:CSI

[CCK23] Ruyun Chen, Juan Chen, and Xuesong Kuang. Comment on ‘Symbolic integration of a product of two spherical Bessel functions with an additional exponential and polynomial factor’ by B. Gebremariam, T. Duguet and S. K. Bogner. Computer Physics Communications, 282(??):??, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200217X. See [GDB10].

Chierici:2022:OPV

[CCL^{+}22] A. Chierici, L. Chirco, V. Le Chenadec, R. Scardovelli, Ph. Yecko, and S. Zaleski. An optimized Vofi library to initialize the volume fraction field. Computer Physics Communications, 281(??):??, December 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002259.

Carrazza:2020:VAM

[CCM20] Stefano Carrazza and Juan M. Cruz-Martinez. VegasFlow: Accelerating Monte Carlo simulation across multiple hardware platforms. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301624.

Carrazza:2021:PPD

[CCMR21] Stefano Carrazza, Juan M. Cruz-Martinez, and Marco Rossi. PDFFlow: Parton distribution functions on GPU. Computer Physics Communications, 264(??):??, July 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001077.

Cui:2020:LIC

[CCW20] Jin Cui, Wenjun Cai, and Yushun Wang. A linearly-implicit and conservative Fourier pseudo-spectral method for the 3D Gross–Pitaevskii equation with angular momentum rotation. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300205.

Cardoso:2021:SAO

[CdBMdAS^{+}21] Halisson Alberdan Cavalcanti Cardoso, Silvio de Barros Melo, Ricardo Martins de Abreu Silva, Sidartha Azevedo Lobo de Carvalho, Silas Garrido Teixeira de Carvalho Santos, and Carlos Costa Dantas. Spectral analysis and optimization of the condition number problem. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302824.

Chabuda:2022:TTN

[CDD22] Krzysztof Chabuda and Rafał Demkowicz-Dobrzański. TNQMetro: Tensor-network based package for efficient quantum metrology computations. Computer Physics Communications, 274(??):

??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003945.

Coluci:2022:ACG

[CDT22] V. R. Coluci, S. O. Dantas, and V. K. Tewary. Accelerated causal Green’s function molecular dynamics. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000972.

Cassini:2022:EVS

[CE22] Fabio Cassini and Lukas Einkemmer. Efficient 6D Vlasov simulation using the dynamical low-rank framework Ensign. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002089.

Chernov:2024:SIR

[CEC^{+}24] V. Chernov, F. Escudero, J. J. Cruz, N. A. Eaves, and A. Fuentes. SootImage: an image recreation, post-processing validation procedure for sooting axisymmetric flames. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001279.

Crialesi-Esposito:2023:FGA

[CESD^{+}23] Marco Crialesi-Esposito, Nicolò Scapin, Andreas D. Demou, Marco Edoardo Rosti, Pedro Costa, Filippo Spiga, and Luca Brandt. FluTAS: a GPU-accelerated finite difference code for multiphase flows. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003216.

Costa:2021:SQM

[CF21] Guilherme S. Costa and Silvio C. Ferreira. Simple quasistationary method for simulations of epidemic processes with localized states. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001582.

Calvo-Fernandez:2024:PCN

[CFBRE24] Aitor Calvo-Fernández, María Blanco-Rey, and Asier Eiguren. The PointGroupNRG code for numerical renormalization group calculations with discrete point-group symmetries. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003776.

Chatterjee:2022:TBL

[CFL^{+}22] Suman Chatterjee, Nikolaus Frohner, Lukas Lechner, Robert Schöfbeck, and Dennis Schwarz. Tree boosting for learning EFT parameters. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001047.

Cen:2024:SPG

[CFLR24] Chunze Cen, Georgios Fourtakas, Steven Lind, and Benedict D. Rogers. A single-phase GPU-accelerated surface tension model using SPH. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003570.

Castillo-Felisola:2023:CPA

[CFPS23] Oscar Castillo-Felisola, Dominic T. Price, and Mattia Scomparin. Cadabra and Python algorithms in general relativity and cosmology II: Gravitational waves. Computer Physics Communications, 289(??):??, August 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000930.

Cheng:2023:DDM

[CFW^{+}23] Wenjie Cheng, Haiyang Fu, Liang Wang, Chuanfei Dong, Yaqiu Jin, Mingle Jiang, Jiayu Ma, Yilan Qin, and Kexin Liu. Data-driven, multi-moment fluid modeling of Landau damping. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002570.

Chekanov:2021:JDA

[CGG21] S. V. Chekanov, G. Gavalian, and N. A. Graf. Jas4pp — a data-analysis framework for physics and detector studies. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000199.

Chen:2021:CPT

[CGR21] Jiahui Chen, Weihua Geng, and Daniel R. Reynolds. Cyclically parallelized treecode for fast computations of electrostatic interactions on molecular surfaces. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303672.

Castagna:2020:TES

[CGSO20] Jony Castagna, Xiaohu Guo, Michael Seaton, and Alan O’Cais. Towards extreme scale dissipative particle dynamics simulations using multiple GPGPUs. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300199.

Chuluunbaatar:2022:KPC

[CGV^{+}22] O. Chuluunbaatar, A. A. Gusev, S. I. Vinitsky, A. G. Abrashkevich, P. W. Wen, and C. J. Lin. KANTBP 3.1: a program for computing energy levels, reflection and transmission matrices, and corresponding wave functions in the coupled-channel and adiabatic approaches. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522001163.

Chen:2020:TAA

[CGZ^{+}20] Xin Chen, Xing-Yu Gao, Ya-Fan Zhao, De-Ye Lin, Wei-Dong Chu, and Hai-Feng Song. TensorAlloy: an automatic atomistic neural network program for alloys. Computer Physics Communications, 250(??):??, May 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303820.

Champciaux:2021:BPD

[CHA21] Valentin Champciaux, Juan Carlos Garcia Hernandez, and Mathieu Agelou. A breakdown of the pseudo-deterministic transport variance reduction method: Formalization and usage considerations. Computer Physics Communications, 264(??): ??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000916.

Cheche:2023:AAS

[Che23] Tiberius O. Cheche. Atomistic approach to the strain field in finite-sized heterostructures. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523002126.

Chen:2024:MPS

[CHY^{+}24] Dandan Chen, Jingyuan Hu, Shaoxiong Yang, Xiao He, Yang Li, Shuai Ren, He Bai, and Jue Wang. A massively parallel spatially resolved stochastic cluster dynamics method for simulations of irradiated materials. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300382X.

Creech:2021:HLE

[CJ21] A. C. W. Creech and A. Jackson. Hybrid large eddy simulation for low-order discontinuous Galerkin methods using an explicit filter. Computer Physics Communications, 260(??): ??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303611.

Crandall:2020:MSC

[CJD^{+}20] P. Crandall, D. Jarema, H. Doerk, Q. Pan, G. Merlo, T. Görler, A. Bañón Navarro, D. Told, M. Maurer, and F. Jenko. Multi-species collisions for delta-f gyrokinetic simulations: Implementation and verification with GENE. Computer Physics Communications, 255(??):??, October 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301545.

Chen:2021:PID

[CJZ21] Wenqian Chen, Yaping Ju, and Chuhua Zhang. A parallel inverted dual time stepping method for unsteady incompressible fluid flow and heat transfer problems. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301375.

Cai:2020:NLT

[CK20a] X.-X. Cai and T. Kittelmann. NCrystal: a library for thermal neutron transport. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519302280.

Carrazza:2020:SRT

[CK20b] Stefano Carrazza and Daniel Krefl. Sampling the Riemanntheta Boltzmann machine. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302174.

Claes:2023:LIE

[CK23] Niels Claes and Rony Keppens. Legolas 2.0: Improvements and extensions to an MHD spectroscopic framework. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002011.

Choi:2021:DHF

[CKC^{+}21] Sooyoung Choi, Wonkyeong Kim, Jiwon Choe, Woonghee Lee, Hanjoo Kim, Bamidele Ebiwonjumi, Eun Jeong, Kyeongwon Kim, Dongmin Yun, Hyunsuk Lee, and Deokjung Lee. Development of high-fidelity neutron transport code STREAM. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000515.

Christopher:2022:HOS

[CKGW22] Peter J. Christopher, Andrew Kadis, George S. D. Gordon, and Timothy D. Wilkinson. HoloGen: an opensource toolbox for high-speed hologram generation. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002514.

Chikitkin:2021:NSB

[CKT21] A. V. Chikitkin, E. K. Kornev, and V. A. Titarev. Numerical solution of the Boltzmann equation with S-model collision integral using tensor decompositions. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000758.

Chang:2020:ACT

[CL20] Kuo-Chuan Chang and Chia-Jyi Liu. An algorithm of calculating transport parameters of thermoelectric materials using single band model with optimized integration methods. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302620.

Choi:2021:TDM

[CL21] Sooyoung Choi and Deokjung Lee. Three-dimensional method of characteristics/diamond-difference transport analysis method in STREAM for whole-core neutron transport calculation. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301430.

Chu:2022:CVG

[CL22a] Moody T. Chu and Matthew M. Lin. A complex-valued gradient flow for the entangled bipartite low rank approximation. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002976.

Cooke:2022:ISL

[CL22b] Joseph R. Cooke and Jennifer R. Lukes. An implicit spin lattice dynamics integrator in LAMMPS. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003155.

Cruz-Lopez:2024:ASF

[CLEP24] Carlos-Antonio Cruz-López and Gilberto Espinosa-Paredes. Analytical solution of the fractional neutron point kinetic equations using the Mittag-Leffler function. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003739.

Cruz-Lopez:2022:DGB

[CLEPF22] Carlos-Antonio Cruz-López, Gilberto Espinosa-Paredes, and Juan-Luis François. Development of the General Bateman Solution using fractional calculus: a theoretical and algorithmic approach. Computer Physics Communications, 273(??): ??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003805.

Cruz-Lopez:2023:NSA

[CLEPF23] Carlos-Antonio Cruz-López, Gilberto Espinosa-Paredes, and Juan-Luis François. A new simplified analytical solution to solve the neutron point kinetics equations using the Laplace transform method. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002831.

Chen:2022:TDS

[CLS^{+}22a] Mengzhi Chen, Tong Li, Bastian Schuetrumpf, Paul-Gerhard Reinhard, and Witold Nazarewicz. Three-dimensional Skyrme Hartree–Fock–Bogoliubov solver in coordinate-space representation. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000625.

Chen:2022:FND

[CLS22b] Zhao Chen, Yang Liu, and Hao Sun. Forecasting of nonlinear dynamics based on symbolic invariance. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001011.

Casati:2022:WNP

[CLVV22] M. Casati, P. Lorenzoni, D. Valeri, and R. Vitolo. Weakly nonlocal Poisson brackets: Tools, examples, computations. Computer Physics Communications, 274(??):??, May 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000029.

Cai:2022:FAB

[CLY22] Zhenning Cai, Jianfeng Lu, and Siyao Yang. Fast algorithms of bath calculations in simulations of quantum system-bath dynamics. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001369.

Chu:2021:MSH

[CLZ^{+}21] Genshen Chu, Yang Li, Runchu Zhao, Shuai Ren, Wen Yang, Xinfu He, Changjun Hu, and Jue Wang. MD simulation of hundred-billion-metal-atom cascade collision on Sunway Taihulight. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552100240X.

Castangna:2021:SLM

[CMJC21] Jony Castangna, Fausto Martelli, Kirk E. Jordan, and Jason Crain. Simulation of large molecular systems with electronically-derived forces. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000783.

Cliche:2024:IMM

[CML^{+}24] D. T. Cliche, M. E. Martin, R. A. London, H. A. Scott, and M. V. Patel. An improved methodology for modeling short pulse buried layer X-ray emission spectra. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000912.

Caprile:2022:PPP

[CMS22] Fernando Caprile, Luciano A. Masullo, and Fernando D. Stefani. PyFocus — a Python package for vectorial calculations of focused optical fields under realistic conditions. Application to toroidal foci. Computer Physics Communications, 275(??): ??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000339.

Candoli:2023:PSQ

[CNB^{+}23] Davide Candoli, Ilija K. Nikolov, Lucas Z. Brito, Stephen Carr, Samuele Sanna, and Vesna F. Mitrović. PULSEE: a software for the quantum simulation of an extensive set of magnetic resonance observables. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003174.

Curtis:2022:ARF

[CNS22] Nicholas J. Curtis, Kyle E. Niemeyer, and Chih-Jen Sung. Accelerating reactive-flow simulations using vectorized chemistry integration. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552200128X.

Catalan:2022:DRS

[COJ^{+}22] J. P. Catalan, F. Ogando, R. Juárez, P. Sauvan, G. Pedroche, J. Alguacil, and J. Sanz. Development of radiation sources for nuclear analysis beyond ITER bio-shield: SRC-UNED code. Computer Physics Communications, 275(??): ??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000273.

Costa:2022:FAM

[Cos22] Pedro Costa. A FFT-accelerated multi-block finite-difference solver for massively parallel simulations of incompressible flows. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003064.

Chen:2024:MOS

[COW24] Yuan Chen, Mahmut Sait Okyay, and Bryan M. Wong. MISTER-T: an open-source software package for quantum optimal control of multi-electron systems on arbitrary geometries. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001711.

Certik:2024:HOF

[ČPF^{+}24] Ondřej Čertík, John E. Pask, Isuru Fernando, Rohit Goswami, N. Sukumar, Lee. A. Collins, Gianmarco Manzini, and Jiří Vackář. High-order finite element method for atomic structure calculations. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552300396X.

Casagrande:2021:ADM

[CPL21] Heitor P. Casagrande, Dario Poletti, and Gabriel T. Landi. Analysis of a density matrix renormalization group approach for transport in open quantum systems. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001727.

Caliari:2020:GVM

[CR20] Marco Caliari and Stefan Rainer. GSGPEs-v1.1: a MAT-LAB code for computing the ground state of systems of Gross–Pitaevskii equations. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303170.

Chandar:2022:FCA

[CS22a] Dominic D. J. Chandar and Jayanarayanan Sitaraman. A flux correction approach for the pressure equation in incompressible flows on overset meshes in OpenFOAM. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100391X.

Crivellin:2022:CLS

[CS22b] Andreas Crivellin and Luc Schnell. Complete Lagrangian and set of Feynman rules for scalar leptoquarks. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003003.

Cosme:2023:TST

[CSBF23] Pedro Cosme, João S. Santos, João P. S. Bizarro, and Ivan Figueiredo. TETHYS: a simulation tool for graphene hydrodynamic models. Computer Physics Communications, 282(??): ??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002697.

Chen:2024:SOS

[CSW24] Yuan Chen, Simon N. Sandhofer, and Bryan M. Wong. SHORYUKEN: an open-source software package for calculating nonlocal exchange interactions in nanowires. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001206.

Clavier:2023:CTE

[CT23] Germain Clavier and Aidan P. Thompson. Computation of the thermal elastic constants for arbitrary manybody potentials in LAMMPS using the stress-fluctuation formalism. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300019X.

Christopoulos:2022:PVU

[CTPS22] Stavros-Richard G. Christopoulos, George I. Tsagiannis, Konstantina A. Papadopoulou, and Nicholas V. Sarlis. VISROC 2.0: Updated software for the visualization of the significance of Receiver Operating Characteristics based on confidence ellipses. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002119.

Chen:2023:FFP

[CTZW23] Yuxi Chen, Gábor Tóth, Hongyang Zhou, and Xiantong Wang. FLEKS: a flexible particle-in-cell code for multi-scale plasma simulations. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000590.

Cohen:2021:CKB

[CV21] Maximilian Cohen and Dionisios G. Vlachos. Chemical Kinetics Bayesian Inference Toolbox (CKBIT). Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001016.

Cherepanov:2024:PVD

[CV24] Vladimir Cherepanov and Christian Veelken. The polarimeter vector for τ → 3π_{τ} decays. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000766.

Chevalier:2022:HOS

[CW22] Cameron Chevalier and Bryan M. Wong. HADOKEN: an opensource software package for predicting electron confinement effects in various nanowire geometries and configurations. Computer Physics Communications, 274(??):??, May 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000170.

Chen:2021:MLE

[CWG^{+}21] Xin Chen, Li-Fang Wang, Xing-Yu Gao, Ya-Fan Zhao, De-Ye Lin, Wei-Dong Chu, and Hai-Feng Song. Machine learning enhanced empirical potentials for metals and alloys. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002447.

Cui:2021:AHO

[CWJ21] Jin Cui, Yushun Wang, and Chaolong Jiang. Arbitrarily high-order structure-preserving schemes for the Gross– Pitaevskii equation with angular momentum rotation. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303830.

Cao:2023:USB

[CXCZ23] Guohua Cao, Chuanqi Xu, Ping Cui, and Zhenyu Zhang. Universal scaling between on-site Coulomb repulsion and numbers of core and valence electrons in transition metal trichalcogenides. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000279.

Cao:2021:ZVS

[CZ21] Xiao-Yan Cao and Jun-Jie Zhang. ZMCintegral-v5.1: Support for multi-function integrations on GPUs. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001065.

Camino:2023:CPI

[CZA^{+}23] Bruno Camino, Huanyu Zhou, Eleonora Ascrizzi, Alberto Boccuni, Filippo Bodo, Alessandro Cossard, Davide Mitoli, Anna Maria Ferrari, Alessandro Erba, and Nicholas M. Harrison. CRYSTALpytools: a Python infrastructure for the Crystal code. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523001984.

Chen:2023:MNV

[CZB^{+}23] Wenqi Chen, Jingjing Zheng, Junwei Lucas Bao, Donald G. Truhlar, and Xuefei Xu. MSTor 2023: a new version of the computer code for multistructural torsional anharmonicity, now with automatic torsional identification using redundant internal coordinates. Computer Physics Communications, 288 (??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000851.

Carlson:2021:OSP

[CZS^{+}21] Max Carlson, Xiaoning Zheng, Hari Sundar, George Em Karniadakis, and Robert M. Kirby. An open-source parallel code for computing the spectral fractional Laplacian on 3D complex geometry domains. Computer Physics Communications, 261(??): ??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303416.

Chen:2023:HPI

[CZTF23] Hongwei Chen, Yujia Zhai, Joshua J. Turner, and Adrian Feiguin. A high-performance implementation of atomistic spin dynamics simulations on x86 CPUs. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001960.

Chen:2023:DKP

[CZWE23] Yixiao Chen, Linfeng Zhang, Han Wang, and Weinan E. DeePKS-kit: a package for developing machine learningbased chemically accurate energy and density functional models. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002399.

Chen:2021:MOO

[CZX^{+}21] Zhenping Chen, Zhenyu Zhang, Jinsen Xie, Qian Guo, Tao Yu, Pengcheng Zhao, Zijing Liu, and Chao Xie. Multi-objective optimization strategies for radiation shielding design with genetic algorithm. Computer Physics Communications, 260(??): ??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300850.

Chen:2020:ENS

[CZY20] Chuanjun Chen, Jun Zhang, and Xiaofeng Yang. Efficient numerical scheme for a new hydrodynamicallycoupled conserved Allen–Cahn type Ohta–Kawasaki phasefield model for diblock copolymer melt. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520301910.

Dosta:2024:COS

[DAA^{+}24] M. Dosta, D. Andre, V. Angelidakis, R. A. Caulk, M. A. Celigueta, B. Chareyre, J.-F. Dietiker, J. Girardot, N. Govender, C. Hubert, R. Kobyłka, A. F. Moura, V. Skorych, D. K. Weatherley, and T. Weinhart. Comparing open-source DEM frameworks for simulations of common bulk processes. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523004113.

DeVanna:2023:UGA

[DAC^{+}23] Francesco De Vanna, Filippo Avanzi, Michele Cogo, Simone Sandrin, Matt Bettencourt, Francesco Picano, and Ernesto Benini. URANOS: a GPU accelerated Navier–Stokes solver for compressible wall-bounded flows. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000620.

Denniston:2022:LLF

[DACA^{+}22] Colin Denniston, Navid Afrasiabian, M. G. Cole-André, F. E. Mackay, S. T. T. Ollila, and T. Whitehead. LAMMPS lb/fluid fix version 2: Improved hydrodynamic forces implemented into LAMMPS through a lattice-Boltzmann fluid. Computer Physics Communications, 275(??):??, June 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000364.

Daniluk:2023:CCS

[Dan23] Andrzej Daniluk. A C++ code for simulations of RHEED intensity oscillations within the kinematical approximation. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552200306X.

Daniluk:2024:FPC

[Dan24] Andrzej Daniluk. First-principles calculations of specular reflection of high-energy electrons during the two-dimensional crystal growth. Computer Physics Communications, 295(??): ??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003508.

Pietri:2023:DVW

[DARJ23] Marco De Pietri, Javier Alguacil, Eduardo Rodríguez, and Rafael Juárez. Development and validation in water of FLUNED, an open-source tool for fluid activation calculations. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001522.

Daude:2023:DDC

[Dau23] Frédéric Daude. A 1-D/3-D coupling approach for compressible non–equilibrium two-phase flows using the Baer-Nunziato model based on the finite-volume framework. Computer Physics Communications, 288(??):??, July 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000693.

Du:2023:OOP

[DBBP23] Dou Du, Taylor J. Baird, Sara Bonella, and Giovanni Pizzi. OSSCAR, an open platform for collaborative development of computational tools for education in science. Computer Physics Communications, 282(??):??, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200265X.

deBrugiere:2020:QCS

[dBBVA20] Timothée Goubault de Brugière, Marc Baboulin, Benoît Valiron, and Cyril Allouche. Quantum circuits synthesis using Householder transformations. Computer Physics Communications, 248(??):??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303388.

Dutra:2021:PEP

[DBdFdSR21] José Diogo L. Dutra, Thiago D. Bispo, Sabrina M. de Freitas, and Marcos V. dos S. Rezende. ParamGULP: an efficient Python code for obtaining interatomic potential parameters for General Utility Lattice Program. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001089.

Diederichs:2022:HPQ

[DBH^{+}22] S. Diederichs, C. Benedetti, A. Huebl, R. Lehe, A. Myers, A. Sinn, J.-L. Vay, W. Zhang, and M. Thévenet. HiPACE++: a portable, 3D quasi-static particle-in-cell code. Computer Physics Communications, 278(??):??, September 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001400.

Dey:2024:PGN

[DBM^{+}24] Tanay Dey, Purba Bhattacharya, Supratik Mukhopadhyay, Nayana Majumdar, Abhishek Seal, and Subhasis Chattopadhyay. Parallelization of Garfield++ and neBEM to simulate space-charge effects in RPCs. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002898.

Dung:2024:ITI

[DBR24] Nguyen Tuan Dung, Christophe Besse, and François Rogier. An implicit time integration approach for simulation of corona discharges. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002515.

Dhib:2024:DIH

[DBV^{+}24] Rayan Dhib, Firas Ben Ameur, Ray Vandenhoeck, Andrea Lani, and Stefaan Poedts. Development of an implicit highorder flux reconstruction solver for high-speed flows on simplex elements. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552300351X.

DiGiusto:2022:SAM

[DC22] Davide Di Giusto and Jony Castagna. A scalable algorithm for many-body dissipative particle dynamics using multiple general purpose graphic processing units. Computer Physics Communications, 280(??):??, November 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001916.

Dang:2023:LGP

[DCRF23] Khanh Dang, Jie Chen, Brian Rodgers, and Saryu Fensin. LAVA 1.0: a general-purpose Python toolkit for calculation of material properties with LAMMPS and VASP. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000127.

Dou:2023:FIM

[DCZ23] Wei Dou, Minhong Chen, and Shenggao Zhou. Fast iterative method for local steric Poisson–Boltzmann theories in biomolecular solvation. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001534.

Dias:2023:LKM

[DdCAG23] Tiago C. Dias, Antonio Tejero del Caz, Luís L. Alves, and Vasco Guerra. The LisbOn KInetics Monte Carlo solver. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002739.

Denner:2020:PMC

[DDM20] Ansgar Denner, Stefan Dittmaier, and Alexander Mück. Prophecy4f 403.0: a Monte Carlo program for Higgs-boson decays into four-fermion final states in and beyond the Standard Model. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301454.

Duarte:2024:EWD

[DEdM24] L. G. S. Duarte, J. C. Eiras, and L. A. C. P. da Mota. An efficient way to determine Liouvillian first integrals of rational second order ordinary differential equations. Computer Physics Communications, 298(??):??, May 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000110.

Dohet-Eraly:2020:COB

[DEV20] J. Dohet-Eraly and M. Viviani. Computing an orthonormal basis of symmetric or antisymmetric hyperspherical harmonics. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300333.

Deluzet:2023:EPS

[DFG^{+}23] Fabrice Deluzet, Gwenael Fubiani, Laurent Garrigues, Clément Guillet, and Jacek Narski. Efficient parallelization for 3D-3V sparse grid Particle-In-Cell: Single GPU architectures. Computer Physics Communications, 289(??):??, August 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001005.

Denk:2020:EEC

[DFP^{+}20] S. S. Denk, R. Fischer, E. Poli, O. Maj, S. K. Nielsen, J. Rasmussen, M. Stejner, and M. Willensdorfer. ECRad: an electron cyclotron radiation transport solver for advanced data analysis in thermal and non-thermal fusion plasmas. Computer Physics Communications, 253(??):??, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300291.

DiRenzo:2020:HSO

[DFU20] Mario Di Renzo, Lin Fu, and Javier Urzay. HTR solver: an open-source exascale-oriented task-based multi-GPU highorder code for hypersonic aerothermodynamics. Computer Physics Communications, 255(??):??, October 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300837.

Dominguez-Garcia:2020:BDL

[DG20] Pablo Domínguez-García. Brownian Disks Lab: Simulating time-lapse microscopy experiments for exploring microrheology techniques and colloidal interactions. Computer Physics Communications, 252(??):??, July 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519304333.

DAzevedo:2020:DGS

[DGM20] Eduardo D’Azevedo, David L. Green, and Lin Mu. Discontinuous Galerkin sparse grids methods for time domain Maxwell’s equations. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301880.

Du:2020:DDR

[DH20] Lipei Du and Ulrich Heinz. (3 + 1)-dimensional dissipative relativistic fluid dynamics at non-zero net baryon density. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519303996.

Deppisch:2022:RCL

[DH22] Thomas Deppisch and Florian Herren. RGE++: a C++ library to solve renormalisation group equations in quantum field theory. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002630.

Dinh:2024:RTT

[DHE^{+}24] P. M. Dinh, J. Heraud, A. Estaña, M. Vincendon, P.-G. Reinhard, and E. Suraud. The real-time TDDFT code “Quantum Dissipative Dynamics” on a GPU. Computer Physics Communications, 295(??):??, February 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002928.

Dingel:2021:AAP

[DHK^{+}21] Kristina Dingel, Rico Huhnstock, André Knie, Arno Ehresmann, and Bernhard Sick. AdaPT: Adaptable particle tracking for spherical microparticles in lab on chip systems. Computer Physics Communications, 262(??):??, May 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000205.

Renzo:2022:HSP

[Di 22] Mario Di Renzo. HTR-1.3 solver: Predicting electrified combustion using the hypersonic task-based research solver. Computer Physics Communications, 272(??):??, March 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003593.

Davidsson:2021:AAW

[DIAA21] Joel Davidsson, Viktor Ivády, Rickard Armiento, and Igor A. Abrikosov. ADAQ: Automatic workflows for magneto-optical properties of point defects in semiconductors. Computer Physics Communications, 269(??):??, December 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002034.

DeChant:2023:VVO

[DIK^{+}23] Corey DeChant, Casey Icenhour, Shane Keniley, Grayson Gall, Alexander Lindsay, Davide Curreli, and Steven Shannon. Verification and validation of the open-source plasma fluid code: Zapdos. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001820.

Diaz-Ibarra:2024:CPP

[DIKSN24] Oscar H. Díaz-Ibarra, Kyungjoo Kim, Cosmin Safta, and Habib N. Najm. CSPlib: a performance portable parallel software toolkit for analyzing complex kinetic mechanisms. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523004149.

Dion:2023:PQW

[Dio23] Claude M. Dion. Program for quantum wave-packet dynamics with time-dependent potentials (new version announcement). Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001558.

Djukanovic:2020:QCT

[Dju20] D. Djukanovic. Quark Contraction Tool — QCT. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303054.

Dudson:2024:HMC

[DKM^{+}24] Ben Dudson, Mike Kryjak, Hasan Muhammed, Peter Hill, and John Omotani. Hermes-3: Multi-component plasma simulations with BOUT++. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003363.

Duda:2022:TSP

[DKRSR22] R. Duda, J. Keski-Rahkonen, J. Solanpää, and E. Räsänen. tinie — a software package for electronic transport through two-dimensional cavities in a magnetic field. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002538.

Dytrych:2021:SCL

[DLD^{+}21] Tomás Dytrych, Daniel Langr, Jerry P. Draayer, Kristina D. Launey, and Daniel Gazda. SU3lib: a C++ library for accurate computation of Wigner and Racah coefficients of SU(3). Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002496.

Decyk:2023:ABP

[DML23] Viktor K. Decyk, Warren B. Mori, and Fei Li. An analytic Boris pusher for plasma simulation. Computer Physics Communications, 282(??):??, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002788.

deMiras:2020:MTC

[dMMLOS20] Juan Ruiz de Miras, Guillermo Martínez-Lledó, William Orwig, and Jorge Sepulcre. A MATLAB tool for computing the spherical harmonic fractal dimension of the cerebral cortex. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520301648.

Das:2022:PDF

[DMS^{+}22] Sambit Das, Phani Motamarri, Vishal Subramanian, David M. Rogers, and Vikram Gavini. DFT-FE 1.0: a massively parallel hybrid CPU–GPU density functional theory code using finite-element discretization. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001928.

Demesy:2020:NLE

[DNG^{+}20] Guillaume Demésy, André Nicolet, Boris Gralak, Christophe Geuzaine, Carmen Campos, and Jose E. Roman. Non-linear eigenvalue problems with GetDP and SLEPc: Eigenmode computations of frequency-dispersive photonic open structures. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302381.

DiRenzo:2021:HSH

[DP21] Mario Di Renzo and Sergio Pirozzoli. HTR-1.2 solver: Hypersonic Task-based Research solver version 1.2. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303623.

Dedes:2020:SFR

[DPR^{+}20] A. Dedes, M. Paraskevas, J. Rosiek, K. Suxho, and L. Trifyllis. SmeftFR 40-Feynman rules generator for the Standard Model Effective Field Theory. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303005.

DellAquila:2021:ACN

[DR21] D. Dell’Aquila and M. Russo. Automatic classification of nuclear physics data via a Constrained Evolutionary Clustering approach. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303234.

Desai:2022:INN

[DRB22] Saaketh Desai, Samuel Temple Reeve, and James F. Belak. Implementing a neural network interatomic model with performance portability for emerging exascale architectures. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100268X.

Dahmen:2020:MHA

[DRM20] Nour Dahmen, François Rogier, and Vincent Maget. On the modelling of highly anisotropic diffusion for electron radiation belt dynamic codes. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552030148X.

Dedes:2024:SVF

[DRR^{+}24] A. Dedes, J. Rosiek, M. Ryczkowski, K. Suxho, and L. Trifyllis. SmeftFR v3 — Feynman rules generator for the Standard Model Effective Field Theory. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002886.

Dandekar:2023:MCA

[DRSZ23] Rahul Dandekar, R. Rajesh, V. Subashri, and Oleg Zaboronski. A Monte Carlo algorithm to measure probabilities of rare events in cluster-cluster aggregation. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000723.

Dryzek:2021:APP

[Dry21] Jerzy Dryzek. Analysis of positron profiling data using e + DSc computer code. Computer Physics Communications, 264(??): ??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000667.

Deng:2021:TDE

[DRZ^{+}21] Jian Deng, Yves Rogez, Peimin Zhu, Alain Herique, Jinpeng Jiang, and Wlodek Kofman. 3D time-domain electromagnetic full waveform inversion in Debye dispersive medium accelerated by multi-GPU paralleling. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001144.

Dolgov:2020:PCI

[DS20] Sergey Dolgov and Dmitry Savostyanov. Parallel cross interpolation for high-precision calculation of high-dimensional integrals. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519302565.

Delmas:2022:MGI

[DS22] Vincent Delmas and Azzedine Soulaïmani. Multi-GPU implementation of a time-explicit finite volume solver using CUDA and a CUDA-aware version of OpenMPI with application to shallow water flows. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521003027.

deSouza:2022:MFD

[dSOZ22] R. F. de Souza, E. K. Omori, and R. S. Zola. Marlics: a finite difference liquid crystal simulation package. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000984.

Dias:2023:WWB

[DSQ23] Alexandre C. Dias, Julian F. R. V. Silveira, and Fanyao Qu. WanTiBEXOS: a Wannier based tight binding code for electronic band structure, excitonic and optoelectronic properties of solids. Computer Physics Communications, 285(??): ??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003551.

Delgado:2022:RSE

[DSSW22] Rafael L. Delgado, Sebastian Steinbeißer, Michael Strickland, and Johannes Heinrich Weber. The relativistic Schrödinger equation through FFTW 3: an extension of quantumfdtd. Computer Physics Communications, 272(??):??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003623.

Duff:2023:APD

[DSW^{+}23] Andrew Ian Duff, Ridwan Sakidja, Helen C. Walker, Russell A. Ewings, and David Voneshen. Automated potential development workflow: Application to BaZrO_{3}. Computer Physics Communications, 293(??):??, December 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002412.

Daniel:2020:FIS

[DTC20] Don Daniel, William T. Taitano, and Luis Chacón. A fully implicit, scalable, conservative nonlinear relativistic Fokker–Planck 0D-2P solver for runaway electrons. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301557.

Degiacomi:2021:DAC

[DTGE21] Matteo T. Degiacomi, Shansi Tian, H. Chris Greenwell, and Valentina Erastova. DynDen: Assessing convergence of molecular dynamics simulations of interfaces. Computer Physics Communications, 269(??):??, December 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002381.

Durr:2023:PCI

[Dür23] Stephan Dürr. Portable CPU implementation of Wilson, Brillouin and Susskind fermions in lattice QCD. Computer Physics Communications, 282(??):??, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002740.

delValle:2024:SEA

[dVAR^{+}24] Carlos Andrés del Valle, Vasileios Angelidakis, Sudeshna Roy, José Daniel Muñoz, and Thorsten Pöschel. SPIRAL: an efficient algorithm for the integration of the equation of rotational motion. Computer Physics Communications, 297(??): ??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523004228.

Dinh:2022:QDD

[DVC^{+}22] P. M. Dinh, M. Vincendon, F. Coppens, E. Suraud, and P.-G. Reinhard. Quantum Dissipative Dynamics (QDD): a real-time real-space approach to far-off-equilibrium dynamics in finite electron systems. Computer Physics Communications, 270(??): ??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002678.

Dautzenberg:2022:PGG

[DvHSdS22] Ellen Dautzenberg, Simon van Hurne, Maarten M. J. Smulders, and Louis C. P. M. de Smet. GraphIAST: a graphical user interface software for Ideal Adsorption Solution Theory (IAST) calculations. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002132.

Dong:2024:DPM

[DWD^{+}24] Qian Dong, Binglin Wang, Xiaojun Duan, Liang Yan, Ke Liu, Wen Luo, Fuqiu Shao, and Tongpu Yu. A dynamical particle merging and splitting algorithm for Particle-In-Cell simulations. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002588.

Daniluk:2021:EMC

[DZZ21] Andrzej Daniluk, Lucyna Zurawek, and Ryszard Zdyb. An effective method to calculate RHEED rocking curves from nanoheteroepitaxial systems. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303404.

Ercole:2022:PSC

[EBBB22] Loris Ercole, Riccardo Bertossa, Sebastiano Bisacchi, and Stefano Baroni. SporTran: a code to estimate transport coefficients from the cepstral analysis of (multivariate) current time series. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001898.

Eckel:2022:PPP

[EBNS22] Stephen Eckel, Daniel S. Barker, Eric B. Norrgard, and Julia Scherschligt. PyLCP: a Python package for computing laser cooling physics. Computer Physics Communications, 270(??): ??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002782.

Ethier:2023:PGA

[ECS23] Jeffrey G. Ethier, Andrés Córdoba, and Jay D. Schieber. pyDSM: GPU-accelerated rheology predictions for entangled polymers in Python. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001315.

Esmaeili:2024:NMR

[EES24] AmirFarzan Esmaeili, Arman Esmaili, and Pasquale Dario Serpico. Neutrinos from muon-rich ultra high energy electromagnetic cascades: the MUNHECA code. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524000778.

Efros:2020:PCC

[Efr20] Victor D. Efros. Program to calculate coefficients of transformations between three-particle hyperspherical harmonics. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300941.

Efros:2021:CCT

[Efr21a] Victor D. Efros. Calculation of coefficients of transformations between three-particle hyperspherical harmonics. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520304069.

Efros:2021:COT

[Efr21b] Victor D. Efros. Calculation of oscillator (Talmi–Moshinsky– Smirnov) brackets. Computer Physics Communications, 265 (??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552100117X.

Efros:2023:OBN

[Efr23] Victor D. Efros. Oscillator brackets at non equal particle masses. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001972.

Ejtehadi:2024:FCF

[EGKH24] Omid Ejtehadi, Aashish K. Gupta, Sorush Khajepor, and Sina Haeri. Force-chain finder: a software tool for the recursive detection of force-chains in granular materials via minor principal stress. Computer Physics Communications, 297(??): ??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523004150.

Eghtesad:2020:MGI

[EGLK20] Adnan Eghtesad, Kai Germaschewski, Ricardo A. Lebensohn, and Marko Knezevic. A multi-GPU implementation of a full-field crystal plasticity solver for efficient modeling of high-resolution microstructures. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300655.

Einkemmer:2020:SLV

[Ein20] Lukas Einkemmer. Semi-Lagrangian Vlasov simulation on GPUs. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301508.

Ekuma:2024:SSM

[Eku24] C. E. Ekuma. SMATool: Strength of materials analysis toolkit. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001127.

Ekuma:2024:EVE

[EL24] C. E. Ekuma and Z. L. Liu. ElasTool v3.0: Efficient computational and visualization toolkit for elastic and mechanical properties of materials. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524000845.

Elman:2024:MMC

[ELSV24] Howard C. Elman, Jiaxing Liang, and Tonatiuh Sánchez-Vizuet. Multilevel Monte Carlo methods for the Grad– Shafranov free boundary problem. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000225.

Elazab:2022:PLI

[EML22] Hossam Elazab, B. D. E. McNiven, and J. P. F. LeBlanc. LIBAMI: Implementation of algorithmic Matsubara integration. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001886.

Eichlersmith:2023:SDB

[EMM^{+}23] Tom Eichlersmith, Jeremiah Mans, Omar Moreno, Joseph Muse, Michael Revering, and Natalia Toro. Simulation of dark bremsstrahlung in Geant 4. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000358.

Esmaeilpour:2024:GAE

[ENK24] Morteza Esmaeilpour, Fabian Nitschke, and Thomas Kohl. GenEOS: an accurate equation of state for the fast calculation of two-phase geofluids properties based on gene expression programming. Computer Physics Communications, 297(??): ??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523004137.

Einkemmer:2021:EIW

[EOR21] Lukas Einkemmer, Alexander Ostermann, and Mirko Residori. An exponential integrator/WENO discretization for sonic-boom simulation on modern computer hardware. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002459.

Eichstadt:2023:EVK

[EPM23] Jan Eichstädt, Joaquim Peiró, and David Moxey. Efficient vectorised kernels for unstructured high-order finite element fluid solvers on GPU architectures in two dimensions. Computer Physics Communications, 284(??):??, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003435.

Vieira-e-Silva:2021:FSS

[eSdSBST21] André Luiz Buarque Vieira e Silva, Caio José dos Santos Brito, Francisco Paulo Magalhães Simões, and Veronica Teichrieb. A fluid simulation system based on the MPS method. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302745.

Ekstedt:2023:DPE

[EST23] Andreas Ekstedt, Philipp Schicho, and Tuomas V. I. Tenkanen. DRalgo: a package for effective field theory approach for thermal phase transitions. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552300070X.

Eroglu:2024:BCN

[ET24] Emre Eroglu and Oleg A. Tretyakov. The Bézier curve and neural network model of the time-domain transient signals. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001346.

Espinoza-Valverde:2023:CLI

[EVFRHR23] Jesus Espinoza-Valverde, Andreas Frommer, Gustavo Ramirez-Hidalgo, and Matthias Rottmann. Coarsest-level improvements in multigrid for lattice QCD on large-scale computers. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552300214X.

Eichstadt:2020:CSM

[EVMP20] Jan Eichstädt, Martin Vymazal, David Moxey, and Joaquim Peiró. A comparison of the shared-memory parallel programming models OpenMP, OpenACC and Kokkos in the context of implicit solvers for high-order FEM. Computer Physics Communications, 255(??):??, October 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300746.

Fan:2020:VMO

[FA20a] Wenyuan Fan and Henryk Anglart. varRhoTurbVOF 2: Modified OpenFOAM volume of fluid solvers with advanced turbulence modeling capability. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520302198.

Fan:2020:VNS

[FA20b] Wenyuan Fan and Henryk Anglart. varRhoTurbVOF: a new set of volume of fluid solvers for turbulent isothermal multiphase flows in OpenFOAM. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519302632.

Frey:2021:GSA

[FA21] Matthias Frey and Andreas Adelmann. Global sensitivity analysis on numerical solver parameters of Particle-In-Cell models in particle accelerator systems. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302770.

Frey:2020:APA

[FAL20] Matthias Frey, Andreas Adelmann, and Uldis Locans. On architecture and performance of adaptive mesh refinement in an electrostatics Particle-In-Cell code. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302905. See erratum [FAL21].

Frey:2021:EAP

[FAL21] Matthias Frey, Andreas Adelmann, and Uldis Locans. Erratum to: “On architecture and performance of adaptive mesh refinement in an electrostatics Particle-In-Cell code” [comput. phys. commun. 247 (2020) 106912]. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000928. See [FAL20].

Fathurrahman:2020:PJJ

[FASD20] Fadjar Fathurrahman, Mohammad Kemal Agusta, Adhitya Gandaryus Saputro, and Hermawan Kresno Dipojono. PWDFT.jl: a Julia package for electronic structure calculation using density functional theory and plane wave basis. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520301600.

Fitzgerald:2020:HFO

[FBA^{+}20] M. Fitzgerald, J. Buchanan, R. J. Akers, B. N. Breizman, and S. E. Sharapov. HALO: a full-orbit model of nonlinear interaction of fast particles with eigenmodes. Computer Physics Communications, 252(??):??, July 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519301183.

Furst:2021:OTO

[FBC^{+}21] Magnus Fürst, Andrea Bertolino, Alberto Cuoci, Tiziano Faravelli, Alessio Frassoldati, and Alessandro Parente. OptiSMOKE++: a toolbox for optimization of chemical kinetic mechanisms. Computer Physics Communications, 264(??): ??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000680.

Fontana:2020:FCM

[FBMD20] Mauro Fontana, Oscar P. Bruno, Pablo D. Mininni, and Pablo Dmitruk. Fourier continuation method for incompressible fluids with boundaries. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520302265.

Fadiga:2020:CCW

[FCSP20] Ettore Fadiga, Nicola Casari, Alessio Suman, and Michele Pinelli. CoolFOAM: the CoolProp wrapper for OpenFOAM. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303777.

Ferro-Costas:2020:PTR

[FCTFR20] David Ferro-Costas, Donald G. Truhlar, and Antonio Fernández-Ramos. Pilgrim: a thermal rate constant calculator and a chemical kinetics simulator. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520302149.

Fattebert:2023:TFP

[FDPT23] Jean-Luc Fattebert, Stephen DeWitt, Aurelien Perron, and John Turner. Thermo4PFM: Facilitating phase-field simulations of alloys with thermodynamic driving forces. Computer Physics Communications, 288(??):??, July 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300084X.

Francis:2020:MFS

[FFLR20] Anthony Francis, Patrick Fritzsch, Martin Lüscher, and Antonio Rago. Master-field simulations of O(a)-improved lattice QCD: Algorithms, stability and exactness. Computer Physics Communications, 255(??):??, October 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301521.

Figueroa:2023:MCL

[FFTV23] Daniel G. Figueroa, Adrien Florio, Francisco Torrenti, and Wessel Valkenburg. C osmo L attice: a modern code for lattice simulations of scalar and gauge field dynamics in an expanding universe. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003058.

Fingerhut:2021:MMS

[FGCN^{+}21] Robin Fingerhut, Gabriela Guevara-Carrion, Isabel Nitzke, Denis Saric, Joshua Marx, Kai Langenbach, Sergei Prokopev, David Celný, Martin Bernreuther, Simon Stephan, Maximilian Kohns, Hans Hasse, and Jadran Vrabec. ms2: a molecular simulation tool for thermodynamic properties, release 4.0. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000217.

Filipovich:2022:POS

[FH22] Matthew J. Filipovich and Stephen Hughes. PyCharge: an open-source Python package for self-consistent electrodynamics simulations of Lorentz oscillators and moving point charges. Computer Physics Communications, 274(??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000091.

Fischbacher:2024:BTH

[Fis24] Thomas Fischbacher. Bit-twiddling hacks for gamma matrices. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003041.

Fei:2022:HOU

[FJ22] Fei Fei and Patrick Jenny. A high-order unified stochastic particle method based on the Bhatnagar–Gross–Krook model for multi-scale gas flows. Computer Physics Communications, 274 (??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000212.

Frommer:2021:MAE

[FKK^{+}21] Andreas Frommer, Karsten Kahl, Francesco Knechtli, Matthias Rottmann, Artur Strebel, and Ian Zwaan. A multigrid accelerated eigensolver for the Hermitian Wilson–Dirac operator in lattice QCD. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302952.

Feger:2020:LMA

[FKS20] Robert Feger, Thomas W. Kephart, and Robert J. Saskowski. LieART 2.0 — a Mathematica application for Lie algebras and representation theory. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302290.

Friak:2020:MPE

[FLK^{+}20] M. Friák, D. Lago, N. Koutná, D. Holec, T. Rebok, and M. Šob. Multi-phase ELAStic Aggregates (MELASA) software tool for modeling anisotropic elastic properties of lamellar composites. Computer Physics Communications, 247(??): ??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519302504.

Floettmann:2024:ASF

[Flo24] K. Floettmann. Analytical solution of the field integrals of a cylindrical grid element. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002497.

Feng:2023:TLC

[FLW^{+}23] Xianghui Feng, Nan Li, Kai Wu, Yonghong Cheng, and Bing Xiao. Thermo-lp: a computational tool to evaluate reaction thermodynamics for synthesizing M_{n+1} AX_{n} (MAX) phases based on linear programming optimization method. Computer Physics Communications, 289(??):??, August 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300098X.

Fontana:2022:VPB

[FMBD22] Mauro Fontana, Pablo D. Mininni, Oscar P. Bruno, and Pablo Dmitruk. Vector potential-based MHD solver for nonperiodic flows using Fourier continuation expansions. Computer Physics Communications, 275(??):??, June 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000224.

Fernandez-Menchero:2020:PTI

[FMCB^{+}20] L. Fernández-Menchero, A. C. Conroy, C. P. Ballance, N. R. Badnell, D. M. Mitnik, T. W. Gorczyca, and M. J. Seaton. PSTGF: Time-independent R-matrix atomic electron-impact code. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302307.

Francisquez:2024:CDG

[FMHH24] Manaure Francisquez, Noah R. Mandell, Ammar Hakim, and Gregory W. Hammett. Conservative discontinuous Galerkin interpolation: Sheared boundary conditions. Computer Physics Communications, 298(??):??, May 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000328.

Fan:2021:ACC

[FN21] Deli Fan and Cheng Ning. Accurately charge-conserving scheme of current assignment based on the current continuity integral equation for particle-in-cell simulations. Computer Physics Communications, 263(??):??, June 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000400.

Forembski:2023:MCS

[FN23] Andrew Forembski and Lampros A. A. Nikolopoulos. MP-CITDSE: a set of ab-initio programs for the simulation of hydrogenic and helium-like atom-laser interactions. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001650.

Fan:2020:APC

[FO20] Tao Fan and Artem R. Oganov. AICON: a program for calculating thermal conductivity quickly and accurately. Computer Physics Communications, 251(??):??, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303923.

Fan:2021:APC

[FO21] Tao Fan and Artem R. Oganov. AICON2: a program for calculating transport properties quickly and accurately. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001399.

Fonseca:2021:GMP

[Fon21] Renato M. Fonseca. GroupMath: a Mathematica package for group theory calculations. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001971.

Forthomme:2022:CGC

[For22] Laurent Forthomme. CepGen — a generic central exclusive processes event generator for hadron–hadron collisions. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003374.

Fredericks:2021:PPL

[FPSZ21] Scott Fredericks, Kevin Parrish, Dean Sayre, and Qiang Zhu. PyXtal: a Python library for crystal structure generation and symmetry analysis. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520304057.

Figueroa:2022:UOL

[FQRR22] Diogenes Figueroa, Seth Quackenbush, Laura Reina, and Christian Reuschle. Updates to the one-loop provider NLOX. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002629.

Fontes:2020:FPF

[FR20] Duarte Fontes and Jorge C. Romão. FeynMaster: a plethora of Feynman tools. Computer Physics Communications, 256 (??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301223.

Fourtakas:2021:TPS

[FRN21] G. Fourtakas, B. D. Rogers, and A. M. A. Nasar. Towards pseudo-spectral incompressible smoothed particle hydrodynamics (ISPH). Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001405.

Ferrer:2023:PHH

[FRN^{+}23] E. Ferrer, G. Rubio, G. Ntoukas, W. Laskowski, O. A. Mariño, S. Colombo, A. Mateo-Gabín, H. Marbona, F. Manrique de Lara, D. Huergo, J. Manzanero, A. M. Rueda-Ramírez, D. A. Kopriva, and E. Valero. HORSE3D: a high-order discontinuous Galerkin solver for flow simulations and multi-physics applications. Computer Physics Communications, 287(??): ??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000450.

Farias:2024:DPF

[FSMM24] Tiago S. Farias, Vitor V. Schultz, José C. M. Mombach, and Jonas Maziero. A differentiable programming framework for spin models. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001577.

Feng:2023:HCL

[FTG23] Feng Feng, Shan-Rong Tang, and Ya-Di Gao. HepLib: a C++ library for high energy physics (version 1.1). Computer Physics Communications, 285(??):??, April 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003502.

Feng:2023:SOS

[FTZ^{+}23] Kaikai Feng, Peng Tian, Jun Zhang, Fei Fei, and Dongsheng Wen. SPARTACUS: an open-source unified stochastic particle solver for the simulation of multiscale nonequilibrium gas flows. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522003265.

Feng:2024:MPI

[FWL^{+}24a] Junwei Feng, Lingyun Wan, Jielan Li, Shizhe Jiao, Xinhui Cui, Wei Hu, and Jinlong Yang. Massively parallel implementation of iterative eigensolvers in large-scale plane-wave density functional theory. Computer Physics Communications, 299(??): ??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000584.

Wu:2024:FJJ

[fWL24b] Quan feng Wu and Zhao Li. FeAmGen.jl: a Julia program for Feynman amplitude generation. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552400153X.

Fan:2021:MMA

[FXQS21] Chao Fan, Xin-Lin Xia, Jian Qiu, and Chuang Sun. mcrtFOAM: a mesh-agglomeration Monte Carlo ray-tracing solver for radiative transfer in gray semitransparent solids. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302848.

Feng:2021:HCL

[FXZT21] Feng Feng, Yi-Fan Xie, Qiu-Chen Zhou, and Shan-Rong Tang. HepLib: a C++ library for high energy physics. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000941.

Fu:2022:ISC

[FYM^{+}22] Yuliang Fu, Juan Yang, Hao Mou, Renwei Tan, Xu Xia, and Zhenye Gao. Integrative simulation of a 2 cm electron cyclotron resonance ion source with full particle-in-cell method. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552200114X.

Feng:2023:CLB

[FYW23] Haoyang Feng, Xiaokui Yue, and Xuechuan Wang. A class of linearization-based collocation methods for initial value and boundary value engineering problems. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003204.

Falone:2024:LOF

[FZD^{+}24] Matteo Falone, Alon Zameret, Yann Delorme, Yanir Edri, Solal A. T. Amouyal, Mark Wasserman, Zhaohui Ding, and Valerio D’Alessandro. libFastMesh: an optimized finitevolume framework for computational aeroacoustics. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002849.

Gonze:2020:API

[GAA^{+}20] Xavier Gonze, Bernard Amadon, Gabriel Antonius, Frédéric Arnardi, Lucas Baguet, Jean-Michel Beuken, Jordan Bieder, François Bottin, Johann Bouchet, Eric Bousquet, Nils Brouwer, Fabien Bruneval, Guillaume Brunin, Théo Cavignac, Jean-Baptiste Charraud, Wei Chen, Michel Côté, Stefaan Cottenier, Jules Denier, Grégory Geneste, Philippe Ghosez, Matteo Giantomassi, Yannick Gillet, Olivier Gingras, Donald R. Hamann, Geoffroy Hautier, Xu He, Nicole Helbig, Natalie Holzwarth, Yongchao Jia, François Jollet, William Lafargue-Dit-Hauret, Kurt Lejaeghere, Miguel A. L. Marques, Alexandre Martin, Cyril Martins, Henrique P. C. Miranda, Francesco Naccarato, Kristin Persson, Guido Petretto, Valentin Planes, Yann Pouillon, Sergei Prokhorenko, Fabio Ricci, Gian-Marco Rignanese, Aldo H. Romero, Michael Marcus Schmitt, Marc Torrent, Michiel J. van Setten, Benoit Van Troeye, Matthieu J. Verstraete, Gilles Zérah, and Josef W. Zwanziger. The Abinit project: Impact, environment and recent developments. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303741.

Gimenez-Alventosa:2021:PEP

[GAGO21] V. Giménez-Alventosa, V. Giménez Gómez, and S. Oliver. PenRed: an extensible and parallel Monte-Carlo framework for radiation transport based on PENELOPE. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001776.

Gaigalas:2020:CPS

[Gai20] G. Gaigalas. Coupling: the program for searching optimal coupling scheme in atomic theory. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303157.

Ga:2023:PAO

[GAJK23] Seongbin Ga, Nahyeon An, Chonghyo Joo, and Junghwan Kim. pyAPEP: an all-in-one software package for the automated preparation of adsorption process simulations. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001753.

Galassi:2022:PPP

[Gal22] Riccardo Malpica Galassi. PyCSP: a Python package for the analysis and simplification of chemically reacting systems based on Computational Singular Perturbation. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000832.

Gardiner:2021:SLE

[Gar21] Steven Gardiner. Simulating low-energy neutrino interactions with MARLEY. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002356.

Groen:2023:FAT

[GAS^{+}23] Derek Groen, Hamid Arabnejad, Diana Suleimenova, Wouter Edeling, Erwan Raffin, Yani Xue, Kevin Bronik, Nicolas Monnier, and Peter V. Coveney. FabSim3: an automation toolkit for verified simulations using high performance computing. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522003150.

Gu:2020:FFN

[GB20] Chuan Gu and Lorenzo Botto. FIPI: a fast numerical method for the simulation of particle-laden fluid interfaces. Computer Physics Communications, 256(??):??, November 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302095.

Gorobets:2022:HCP

[GB22] Andrey Gorobets and Pavel Bakhvalov. Heterogeneous CPU+GPU parallelization for high-accuracy scale-resolving simulations of compressible turbulent flows on hybrid supercomputers. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552100343X.

Gorni:2022:PTC

[GBD^{+}22] Tommaso Gorni, Oscar Baseggio, Pietro Delugas, Stefano Baroni, and Iurii Timrov. turboMagnon — a code for the simulation of spin-wave spectra using the Liouville-Lanczos approach to time-dependent density-functional perturbation theory. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002193.

Gupta:2020:CDD

[GBH20] Lalit Gupta, Lev Barash, and Itay Hen. Calculating the divided differences of the exponential function by addition and removal of inputs. Computer Physics Communications, 254 (??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301673.

Grossu:2021:CMB

[GBJ^{+}21] I. V. Grossu, C. Besliu, Al. Jipa, D. Felea, and T. Esanu. Chaos many-body engine module for estimating pentaquark production in proton-proton collisions at CBM energies. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302678.

Gomez-Bastidas:2023:SIC

[GBR23] A. F. Gomez-Bastidas and O. Rubel. Software implementation for calculating Chern and Z2 topological invariants of crystalline solids with WIEN2k all-electron density functional package. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002096.

Giorgiani:2020:HON

[GBS^{+}20] G. Giorgiani, H. Bufferand, F. Schwander, E. Serre, and P. Tamain. A high-order non field-aligned approach for the discretization of strongly anisotropic diffusion operators in magnetic fusion. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301612.

Gontchar:2021:DCC

[GCK21] Igor I. Gontchar, Maria V. Chushnyakova, and Natalya A. Khmyrova. DFMSPH22: a C-code for the double folding interaction potential of two spherical nuclei. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303398.

Gu:2020:GQF

[GCWZ20] Jie Gu, Jia Chen, Yang Wang, and X.-G. Zhang. Generalized quadrature for finite temperature Green’s function methods. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552030031X.

Gebremariam:2010:SIP

[GDB10] B. Gebremariam, T. Duguet, and S. K. Bogner. Symbolic integration of a product of two spherical Bessel functions with an additional exponential and polynomial factor. Computer Physics Communications, 181(6):1136–1143, June 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465510000408. See comment

[CCK23].

Gibbons:2023:EOS

[GDJG23] Nicholas N. Gibbons, Kyle A. Damm, Peter A. Jacobs, and Rowan J. Gollan. Eilmer: an open-source multi-physics hypersonic flow solver. Computer Physics Communications, 282(??): ??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002703.

Gerlero:2021:EVO

[GDK21] Gabriel S. Gerlero, Santiago Márquez Damián, and Pablo A.

Kler. electroMicroTransport v2107: Open-source tool-

box for paper-based electromigrative separations. Computer Physics Communications, 269(??):??, December 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002551.

Gharibnejad:2021:MCQ

[GDS^{+}21] H. Gharibnejad, N. Douguet, B. I. Schneider, J. Olsen, and L. Argenti. A multi-center quadrature scheme for the molecular continuum. Computer Physics Communications, 263(??): ??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000382.

Gariazzo:2022:PR

[GdSPC22] S. Gariazzo, P. F. de Salas, O. Pisanti, and R. Consiglio. PArthENoPE revolutions. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003179.

Gaigalas:2021:ACS

[GF21] Gediminas Gaigalas and Stephan Fritzsche. Angular coefficients for symmetry-adapted configuration states in jjcoupling. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001983.

Giscard:2023:EST

[GF23] Pierre-Louis Giscard and Mohammadali Foroozandeh. Exact solutions for the time-evolution of quantum spin systems under arbitrary waveforms using algebraic graph theory. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002806.

Guo:2024:MOS

[GFD^{+}24] Yu-Chen Guo, Fan Feng, An Di, Shi-Qi Lu, and Ji-Chong Yang. MLAnalysis: an open-source program for high energy physics analyses. Computer Physics Communications, 294(??): ??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003028.

Gehringer:2023:MCC

[GFH23] Dominik Gehringer, Martin Friák, and David Holec. Models of configurationally-complex alloys made simple. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000097.

Grossu:2024:NVH

[GG24] I. V. Grossu and J. L. Garrett. New version of Hyper-Fractal Analysis application for estimating the fuzzy fractal dimension of hyperspectral satellite ocean color images. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001383.

Gu:2024:CAH

[GGCW24] Xuelong Gu, Yuezheng Gong, Wenjun Cai, and Yushun Wang. A class of arbitrarily high-order energy-preserving method for nonlinear Klein–Gordon–Schrödinger equations. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000821.

Gordon:2021:SUP

[GH21] D. F. Gordon and B. Hafizi. Special unitary particle pusher for extreme fields. Computer Physics Communications, 258(??): ??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303015.

Gunther:2020:CHM

[GHK20] Michael Günther, Roman Höllwieser, and Francesco Knechtli. Constrained hybrid Monte Carlo algorithms for gauge-Higgs models. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300382.

Gorjao:2022:MEM

[GHKW22] Leonardo Rydin Gorjão, Galib Hassan, Jürgen Kurths, and Dirk Witthaut. MFDFA: Efficient multifractal detrended fluctuation analysis in Python. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003660.

Green:2022:EHO

[GHL^{+}22] David Green, Xiaozhe Hu, Jeremy Lore, Lin Mu, and Mark L. Stowell. An efficient high-order numerical solver for diffusion equations with strong anisotropy. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000510.

Gerlach:2023:TTT

[GHL23] Marvin Gerlach, Florian Herren, and Martin Lang. tapir: a tool for topologies, amplitudes, partial fraction decomposition and input for reductions. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002636.

Goodsell:2024:BSF

[GJ24] Mark D. Goodsell and Ari Joury. BSMArt: Simple and fast parameter space scans. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523004022.

Gilbreth:2021:RCF

[GJA21] C. N. Gilbreth, S. Jensen, and Y. Alhassid. Reducing the complexity of finite-temperature auxiliary-field quantum Monte Carlo. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000746.

Gorton:2023:DFP

[GJJN23] Oliver C. Gorton, Calvin W. Johnson, Changfeng Jiao, and Jonathan Nikoleyczik. dmscatter: a fast program for WIMP-nucleus scattering. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003162.

Gaigalas:2021:CFM

[GK21] G. Gaigalas and D. Kato. Crystal field module for the general relativistic atomic structure package. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303854.

Georgescu:2021:BSS

[GKIB21] Alexandru B. Georgescu, Minjung Kim, and Sohrab Ismail-Beigi. Boson Subsidiary Solver (BoSS) v1.1. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100103X.

Green:2024:NOS

[GKT^{+}24] M. D. Green, K. S. Kirilov, M. Turner, J. Marcon, J. Eichstädt, E. Laughton, C. D. Cantwell, S. J. Sherwin, J. Peiró, and D. Moxey. NekMesh: an open-source high-order mesh generation framework. Computer Physics Communications, 298(??): ??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000122.

Glosser:2021:ATS

[GLB^{+}21] C. Glosser, E. Lu, T. J. Bertus, C. Piermarocchi, and B. Shanker. Acceleration techniques for semiclassical Maxwell– Bloch systems: an application to discrete quantum dot ensembles. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302356.

Graziosi:2023:ECF

[GLN23] Patrizio Graziosi, Zhen Li, and Neophytos Neophytou. ElecTra code: Full-band electronic transport properties of materials. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000152.

Granelli:2023:UUL

[GLPG^{+}23] A. Granelli, C. Leslie, Y. F. Perez-Gonzalez, H. Schulz, B. Shuve, J. Turner, and R. Walker. ULYSSES, universal LeptogeneSiS equation solver: Version 2. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001790.

Gan:2021:ESA

[GLSH21] Chee Kwan Gan, Yun Liu, Tze Chien Sum, and Kedar Hippalgaonkar. Efficacious symmetry-adapted atomic displacement method for lattice dynamical studies. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303052.

George:2023:CCO

[GLW^{+}23] Manu George, Chun-Yu Lin, Meng-Ru Wu, Tony G. Liu, and Zewei Xiong. COSEν: a collective oscillation simulation engine for neutrinos. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003071.

Golub:2020:CPO

[GM20] Pavlo Golub and Sergei Manzhos. CONUNDrum: a program for orbital-free density functional theory calculations. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301570.

Gorfinkiel:2024:ACS

[GM24] Jimena D. Gorfinkiel and Fernando Martín. Attosecond chemistry special issue. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000638.

Granelli:2021:UUL

[GMPG^{+}21] A. Granelli, K. Moffat, Y. F. Perez-Gonzalez, H. Schulz, and J. Turner. ULYSSES: Universal LeptogeneSiS Equation Solver. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520304082.

Guo:2020:LFD

[GMZ^{+}20] Shimin Guo, Liquan Mei, Zhengqiang Zhang, Can Li, Mingjun Li, and Ying Wang. A linearized finite difference/spectral-Galerkin scheme for three-dimensional distributed-order timespace fractional nonlinear reaction-diffusion-wave equation: Numerical simulations of Gordon-type solitons. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300059.

Guada:2020:FPM

[GNP20] Victor Guada, Miha Nemevsek, and Matevz Pintar. FindBounce: Package for multi-field bounce actions. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520302253.

Galasso:2023:AAW

[GO23] Michele Galasso and Artem R. Oganov. Automag: an automatic workflow software for calculating the ground magnetic state of a given structure and estimating its critical temperature. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002909.

Gomez-Ortiz:2023:CDN

[GOCSS^{+}23] Fernando Gómez-Ortiz, Nayara Carral-Sainz, James Sifuna, Virginia Monteseguro, Ramón Cuadrado, Pablo García-Fernández, and Javier Junquera. Compatibility of DFT+U with non-collinear magnetism and spin-orbit coupling within a framework of numerical atomic orbitals. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000292.

Gonoskov:2022:ACS

[Gon22] A. Gonoskov. Agnostic conservative down-sampling for optimizing statistical representations and PIC simulations. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100312X.

Grossu:2022:NVH

[GOS^{+}22] I. V. Grossu, M. Opritescu, O. Savencu, A. I. Miron, M. Verga, and N. Verga. A new version of Hyper-Fractal Analysis: .Net WPF module for RGB 3D reconstruction of medical threechannel images. Computer Physics Communications, 276(??): ??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000534.

Gil:2023:EGF

[GOST23] Amparo Gil, Andrzej Odrzywołek, Javier Segura, and Nico M. Temme. Evaluation of the generalized Fermi–Dirac integral and its derivatives for moderate/large values of the parameters. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552200282X.

Gusto:2022:HAM

[GP22] Brandon Gusto and Tomasz Plewa. A hybrid adaptive multiresolution approach for the efficient simulation of reactive flows. Computer Physics Communications, 274(??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000182.

Gao:2022:ACP

[GPD^{+}22] Leyun Gao, Jing Peng, Zilin Dai, Sitian Qian, Tao Li, Qiang Li, and Meng Lu. Animating collider processes with event-time-frame format. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001801.

Galvis:2021:BBE

[GPM^{+}21] Andres F. Galvis, Daniel M. Prada, Lucas S. Moura, Cecilia Zavaglia, Jamie M. Foster, Paulo Sollero, and Luiz C. Wrobel. BESLE: Boundary element software for 3D linear elasticity. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521001211.

Giraldo:2020:EES

[GPN20] Juan-Carlos Giraldo, Néstor M. Peña, and Michel M. Ney. Encoding the electrodynamics in spatiotemporal boundaries. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519302358.

Guan:2022:IAE

[GQ22] Xin Guan and Chong Qi. An iterative approach for the exact solution of the pairing Hamiltonian. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000285.

Garzon:2020:AME

[GRCT20] Alejandro Garzón, Wilmar Rodriguez, Fernando Cristancho, and Molei Tao. AhKin: a modular and efficient code for the Doppler shift attenuation method. Computer Physics Communications, 246(??):??, January 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302310.

Grossu:2022:MHF

[Gro22] I. V. Grossu. Migration of hyper-fractal analysis from Visual Basic 6 to C# .Net. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003015.

Grozin:2023:DFD

[Gro23] Andrey Grozin. Drawing Feynman diagrams with GLE. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003095.

Gratl:2022:WSS

[GSBN22] Fabio Alexander Gratl, Steffen Seckler, Hans-Joachim Bungartz, and Philipp Neumann. N ways to simulate shortrange particle systems: Automated algorithm selection with the node-level library AutoPas. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552100374X.

Giangaspero:2023:RTS

[GSL^{+}23] Vincent F. Giangaspero, Vatsalya Sharma, Johannes Laur, Jan Thoemel, Alessandro Munafò, Andrea Lani, and Stefaan Poedts. 3D ray tracing solver for communication blackout analysis in atmospheric entry missions. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000085.

Gasparino:2024:SGE

[GSL24] L. Gasparino, F. Spiga, and O. Lehmkuhl. SOD2D: a GPUenabled Spectral Finite Elements Method for compressible scale-resolving simulations. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523004125.

Gharibnejad:2020:CNA

[GSLS20] H. Gharibnejad, B. I. Schneider, M. Leadingham, and H. J. Schmale. A comparison of numerical approaches to the solution of the time-dependent Schrödinger equation in one dimension. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519301778.

Grossu:2022:MMH

[GSM^{+}22] I. V. Grossu, O. Savencu, A. I. Miron, C. Besliu, and N. Verga. Medical module for hyper-fractal analysis. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003672.

Gil:2021:GPI

[GST21] Amparo Gil, Javier Segura, and Nico M. Temme. GammaCHI: a package for the inversion and computation of the gamma and chi-square cumulative distribution functions (central and noncentral). new version announcement. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001958.

Gusev:2023:APC

[GSV23] Alexander A. Gusev, Evgeni A. Solov’ev, and Sergue I. Vinitsky. ARSENY: a program for computing inelastic transitions via hidden crossings in one-electron atomic ion-ion collisions with classical description of nuclear motion. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000073.

Gaidamour:2021:BHS

[GTA21] Jérémie Gaidamour, Qinglin Tang, and Xavier Antoine. BEC2HPC: a HPC spectral solver for nonlinear Schrödinger and rotating Gross–Pitaevskii equations. stationary states computation. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521001193.

Gajdics:2021:EMC

[GTE21] Bence Gajdics, János J. Tomán, and Zoltán Erdélyi. An effective method to calculate atomic movements in 3D objects with tuneable stochasticity (3DO-SKMF). Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302885.

Gayday:2021:SPP

[GTMB21] Igor Gayday, Alexander Teplukhin, Jonathan Moussa, and Dmitri Babikov. SpectrumSDT: a program for parallel calculation of coupled rotational-vibrational energies and lifetimes of bound states and scattering resonances in triatomic systems. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552100196X.

Golovatiuk:2022:DLD

[GUAD22] Artem Golovatiuk, Andrey Ustyuzhanin, Andrey Alexandrov, and Giovanni De Lellis. Deep learning for direct Dark Matter search with nuclear emulsions. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000303.

Gulevich:2020:MMT

[Gul20] D. R. Gulevich. MiTMoJCo: Microscopic tunneling model for Josephson contacts. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519304011.

Garcia:2022:IIG

[GUW^{+}22] B. M. Garcia, M. V. Umansky, J. Watkins, J. Guterl, and O. Izacard. INGRID: an interactive grid generator for 2D edge plasma modeling. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000340.

Gins:2024:SUP

[GvdBdGN24] W. Gins, B. van den Borne, R. P. de Groote, and G. Neyens. SATLAS2: an update to the package for analysis of counting data. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003983.

Gorbunov:2022:PBC

[GVV22] S. A. Gorbunov, A. E. Volkov, and R. A. Voronkov. Periodic boundary conditions effects on atomic dynamics analysis. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001734.

Gong:2023:CCE

[GWA^{+}23] Zheng Gong, Zi Wu, Chenge An, Bangwen Zhang, and Xudong Fu. CP3d: a comprehensive Euler–Lagrange solver for direct numerical simulation of particle-laden flows. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000115.

Gao:2021:IOI

[GWPW21] Jiacheng Gao, Quansheng Wu, Clas Persson, and Zhijun Wang. Irvsp: To obtain irreducible representations of electronic states in the VASP. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303805.

Gong:2020:AHO

[GZW20] Yuezheng Gong, Jia Zhao, and Qi Wang. Arbitrarily highorder unconditionally energy stable SAV schemes for gradient flow models. Computer Physics Communications, 249(??): ??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303716.

Hossain:2021:ASP

[HA21] M. Alamgir Hossain and Jahrul M. Alam. Assessment of a symmetry-preserving JFNK method for atmospheric convection. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002253.

He:2023:ACE

[HA23] Y. He and A. Arefiev. Algorithm for computing the electron-positron yield from the linear Breit–Wheeler process in high-intensity laser-plasma interactions. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000024.

Hughey:2020:FSE

[HAA^{+}20] S. Hughey, A. Alsnayyan, H. M. Aktulga, T. Gao, and B. Shanker. Fast and scalable evaluation of pairwise potentials. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300758.

Halder:2021:JLS

[Hal21] Prithish Halder. JaSTA-3: Light scattering simulations for heterogeneous aggregate. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520304112.

Hahner:2020:DSF

[HAM^{+}20a] Urs R. Hähner, Gonzalo Alvarez, Thomas A. Maier, Raffaele Solcà, Peter Staar, Michael S. Summers, and Thomas C. Schulthess. DCA++: a software framework to solve correlated electron problems with modern quantum cluster methods. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519300086.

Hammond:2020:PPD

[Ham20b] Karl D. Hammond. Parallel point defect identification in molecular dynamics simulations without post-processing: a compute and dump style for LAMMPS. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302498.

Hazel:2023:PSS

[Haz23] Andrew Hazel. Professor N. Stan Scott. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003290.

Houfek:2024:USP

[HBM^{+}24] Karel Houfek, Jakub Benda, Zdeněk Mašín, Alex Harvey, Thomas Meltzer, Vincent Graves, and Jimena D. Gorfinkiel. UKRmol-scripts: a Perl-based system for the automated operation of the photoionization and electron/positron scattering suite UKRmol+. Computer Physics Communications, 298(??): ??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000365.

Hetherington:2024:MAD

[HCAH^{+}24] Ashton Hetherington, Adrián Corrochano, Rodrigo Abadía-Heredia, Eneko Lazpita, Eva Muñoz, Paula Díaz, Egoitz Maiora, Manuel López-Martín, and Soledad Le Clainche. ModelFLOWs-app: Data-driven post-processing and reduced order modelling tools. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001401.

Huber:2020:DFE

[HCP20] Markus Q. Huber, Anton K. Cyrol, and Jan M. Pawlowski. DoFun 3.0: Functional equations in Mathematica. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303844.

Hoppe:2021:DFK

[HEF21] Mathias Hoppe, Ola Embreus, and Tünde Fülöp. DREAM: a fluid-kinetic framework for tokamak disruption runaway electron simulations. Computer Physics Communications, 268 (??):??, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002101.

Heylmun:2021:QBM

[HFP21] Jeffrey C. Heylmun, Rodney O. Fox, and Alberto Passalacqua. A quadrature-based moment method for the evolution of the joint size-velocity number density function of a particle population. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001843.

Humeniuk:2022:ECG

[HG22] Alexander Humeniuk and William J. Glover. Efficient CPU and GPU implementations of multicenter integrals over longrange operators using Cartesian Gaussian functions. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001862.

Huang:2020:BTF

[HGS20] Yunfei Huang, Gerhard Gompper, and Benedikt Sabass. A Bayesian traction force microscopy method with automated denoising in a user-friendly software package. Computer Physics Communications, 256(??):??, November 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030120X.

Huddart:2022:PMP

[HHMH^{+}22] B. M. Huddart, A. Hernández-Melián, T. J. Hicken, M. Gomilšek, Z. Hawkhead, S. J. Clark, F. L. Pratt, and T. Lancaster.

MuFinder: a program to determine and analyse muon stopping sites. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002077.

Horvath:2022:CFT

[HHT22] D. X. Horváth, K. Hódsági, and G. Takács. Chirally factorised truncated conformal space approach. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000959.

Hung:2024:QOS

[HHT^{+}24] Nguyen Tuan Hung, Jianqi Huang, Yuki Tatsumi, Teng Yang, and Riichiro Saito. QERaman: an open-source program for calculating resonance Raman spectra based on Quantum ESPRESSO. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003120.

He:2021:TPP

[HHVB21] Xu He, Nicole Helbig, Matthieu J. Verstraete, and Eric Bousquet. TB2J: a Python package for computing magnetic interaction parameters. Computer Physics Communications, 264(??): ??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000679.

Hidding:2021:DMP

[Hid21] Martijn Hidding. DiffExp, a Mathematica package for computing Feynman integrals in terms of one-dimensional series expansions. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552100237X.

Hu:2022:SPD

[HJGL22] Shuyao Hu, Chongwen Jiang, Zhenxun Gao, and Chun-Hian Lee. Spatial parallel disturbance region update method with OpenMP for steady compressible flows. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000789.

Heinrich:2022:ERP

[HJJ^{+}22] G. Heinrich, S. Jahn, S. P. Jones, M. Kerner, F. Langer, V. Magerya, A. Põldaru, J. Schlenk, and E. Villa. Expansion by regions with pySecDec. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003799.

Heinrich:2024:NSA

[HJK^{+}24] G. Heinrich, S. P. Jones, M. Kerner, V. Magerya, A. Olsson, and J. Schlenk. Numerical scattering amplitudes with pySecDec. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003016.

Himanen:2020:DLD

[HJM^{+}20] Lauri Himanen, Marc O. J. Jäger, Eiaki V. Morooka, Filippo Federici Canova, Yashasvi S. Ranawat, David Z. Gao, Patrick Rinke, and Adam S. Foster. DScribe: Library of descriptors for machine learning in materials science. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303042.

Hammond:2024:ISP

[HK24] K. C. Hammond and A. A. Kaptanoglu. Improved stellarator permanent magnet designs through combined discrete and continuous optimizations. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552400050X.

Harlander:2020:F

[HKL20] R. V. Harlander, S. Y. Klein, and M. Lipp. FeynGame. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302186.

Hoshi:2021:OSL

[HKY^{+}21] Takeo Hoshi, Mitsuaki Kawamura, Kazuyoshi Yoshimi, Yuichi Motoyama, Takahiro Misawa, Youhei Yamaji, Synge Todo, Naoki Kawashima, and Tomohiro Sogabe. Kω — open-source library for the shifted Krylov subspace method of the form (zI − H)x = b. Computer Physics Communications, 258(??): ??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302551.

Haldane:2021:MGM

[HL21] Allan Haldane and Ronald M. Levy. Mi3-GPU: MCMC-based inverse Ising inference on GPUs for protein covariation analysis. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520301193.

Hsiao:2020:EBO

[HLCD20] Fu-Chen Hsiao, Ching-Tarng Liang, Yia-Chung Chang, and John M. Dallesasse. Effective bond-orbital model of III-nitride wurtzite structures based on modified interaction parameters of zinc-blende structures. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300023.

Hoang:2022:RAR

[HLM22] André H. Hoang, Christopher Lepenik, and Vicent Mateu. REvolver: Automated running and matching of couplings and masses in QCD. Computer Physics Communications, 270(??): ??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002575.

Herraiz:2024:MPO

[HLMB24] Joaquin L. Herraiz, Alejandro Lopez-Montes, and Andreu Badal. MCGPU-PET: an open-source real-time Monte Carlo PET simulator. Computer Physics Communications, 296(??): ??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003533.

Huhn:2020:GAA

[HLzY^{+}20] William P. Huhn, Björn Lange, Victor Wen zhe Yu, Mina Yoon, and Volker Blum. GPU acceleration of allelectron electronic structure theory using localized numeric atom-centered basis functions. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520301168.

Haule:2024:AEG

[HM24] Kristjan Haule and Subhasish Mandal. All electron GW with linearized augmented plane waves for metals and semiconductors. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003314.

Hunley:2022:EIC

[HMM22] Christian Hunley, Md Mohsin, and Marcelo Marucho. Electrical impulse characterization along actin filaments in pathological conditions. Computer Physics Communications, 275(??): ??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000352.

Haftu:2022:PAW

[HMR22] Asmelash Haftu, Abhinav Muta, and Prabhu Ramachandran. Parallel adaptive weakly-compressible SPH for complex moving geometries. Computer Physics Communications, 277(??): ??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000960.

Heurtel:2023:SSL

[HMSV23] Nicolas Heurtel, Shane Mansfield, Jean Senellart, and Benoît Valiron. Strong simulation of linear optical processes. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001935.

Hanada:2022:STR

[HMYH22] Takashi Hanada, Yuichi Motoyama, Kazuyoshi Yoshimi, and Takeo Hoshi. sim-trhepd-rheed — open-source simulator of total-reflection high-energy positron diffraction (TRHEPD) and reflection high-energy electron diffraction (RHEED). Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200090X.

Hohenester:2024:NRS

[Hoh24] Ulrich Hohenester. Nanophotonic resonators in stratified media with the nanobem toolbox. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002941.

Horn:2023:PTD

[Hor23a] Logan Bishop-Van Horn. pyTDGL: Time-dependent Ginzburg– Landau in Python. Computer Physics Communications, 291 (??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001443.

Horvat:2023:FFP

[Hor23b] Vladimir Horvat. FITEVT: a FORTRAN program for arrival-time analysis of nuclear-decay events. Computer Physics Communications, 284(??):??, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003198.

Horvat:2024:EUV

[Hor24] Vladimir Horvat. ERCS24: an updated version of the ERCS08 program for calculations of the cross sections for atomic electron removal based on the ECPSSR theory and its variants. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524000705.

Han:2021:FBE

[HPAW21] K. S. Han, B. H. Park, A. Y. Aydemir, and M. H. Woo. A free-boundary equilibrium solver with a hybrid iteration method in a semi-bounded computational domain. Computer Physics Communications, 264(??):??, July 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000370.

Herschberg:2023:CPM

[HPP23] Tom Herschberg, Kyle Pifer, and Eleni Panagiotou. A computational package for measuring Topological Entanglement in Polymers, Proteins and Periodic systems (TEPPP). Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003587.

Hammad:2023:EPS

[HPRS23] A. Hammad, Myeonghun Park, Raymundo Ramos, and Pankaj Saha. Exploration of parameter spaces assisted by machine learning. Computer Physics Communications, 293(??):??, December 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002473.

Ha:2021:MGM

[HPY21] Sanghyun Ha, Junshin Park, and Donghyun You. A multi-GPU method for ADI-based fractional-step integration of incompressible Navier–Stokes equations. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001119.

Hollmer:2020:JVP

[HQF^{+}20] Philipp Höllmer, Liang Qin, Michael F. Faulkner, A. C. Maggs, and Werner Krauth. JeLLyFysh — version 1.0 — a Python application for all-atom event-chain Monte Carlo. Computer Physics Communications, 253(??):??, August 2020.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300254.

Honeywell:2020:NOL

[HQRR20] Steve Honeywell, Seth Quackenbush, Laura Reina, and Christian Reuschle. NLOX, a one-loop provider for Standard Model processes. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300965.

Horgue:2022:PVO

[HRG^{+}22] P. Horgue, F. Renard, G. S. Gerlero, R. Guibert, and

G. Debenest. porousMultiphaseFoam v2107: an open-source

tool for modeling saturated/unsaturated water flows and solute transfers at watershed scale. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003908.

Hohenester:2022:NRM

[HRU22] Ulrich Hohenester, Nikita Reichelt, and Gerhard Unger. Nanophotonic resonance modes with the nanobem toolbox. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000558.

Hagh:2022:RRA

[HS22] Varda F. Hagh and Mahdi Sadjadi. rigidPy: Rigidity analysis in Python. Computer Physics Communications, 275(??): ??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000248.

Harlander:2024:AAO

[HS24] Robert V. Harlander and Magnus C. Schaaf. AutoEFT: Automated operator construction for effective field theories. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001218.

Hwang:2022:HPL

[HSA22] Myungwon Hwang, Carlo Scalo, and Andres F. Arrieta. Highperformance large-scale simulation of multi-stable metastructures. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000844.

Hansen:2024:TOS

[HSB^{+}24] C. Hansen, I. G. Stewart, D. Burgess, M. Pharr, S. Guizzo, F. Logak, A. O. Nelson, and C. Paz-Soldan. TokaMaker: an open-source time-dependent Grad-Shafranov tool for the design and modeling of axisymmetric fusion devices. Computer Physics Communications, 298(??):??, May 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000341.

Hunt-Smith:2024:AMC

[HSMR^{+}24] N. T. Hunt-Smith, W. Melnitchouk, F. Ringer, N. Sato, A. W. Thomas, and M. J. White. Accelerating Markov chain Monte Carlo sampling with diffusion models. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523004046.

Hoshi:2022:DDS

[HSO^{+}22] Takeo Hoshi, Daishiro Sakata, Shotaro Oie, Izumi Mochizuki, Satoru Tanaka, Toshio Hyodo, and Koji Hukushima. Datadriven sensitivity analysis in surface structure determination using total-reflection high-energy positron diffraction (TRHEPD). Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002988.

Havasi-Toth:2020:NGP

[HT20] Balázs Havasi-Tóth. Nauticle: a general-purpose particlebased simulation tool. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519302322.

Hagita:2024:TDP

[HT24] Katsumi Hagita and Tetsuo Tominaga. Two-dimensional pattern reverse Monte Carlo analysis of nanoparticles in polymer matrices using a combination of OpenACC and cuFFT. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003168.

Herath:2020:PPL

[HTH^{+}20] Uthpala Herath, Pedram Tavadze, Xu He, Eric Bousquet, Sobhit Singh, Francisco Muñoz, and Aldo H. Romero. PyProcar: a Python library for electronic structure pre/post-processing. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303935.

Hao:2022:OGB

[HTL^{+}22] Changwei Hao, Yuan Tian, Ping Lin, Yunzhen Du, Lijuan Yang, Sheng Zhang, Lei Yang, Qingguo Zhou, and Wenshan Duan. Optimizing the GPU based method calculating energy deposition of beams coupling with discrete materials in dynamical and thermal simulations for higher computing efficiency. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552200145X.

Ho:2021:TTQ

[HTN21] Le Bin Ho, Kieu Quang Tuan, and Hung Q. Nguyen. tqix: a toolbox for Quantum in X: X: Quantum measurement, quantum tomography, quantum metrology, and others. Computer Physics Communications, 263(??):??, June 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100045X.

Hajinazar:2021:MCN

[HTS^{+}21] Samad Hajinazar, Aidan Thorn, Ernesto D. Sandoval, Saba Kharabadze, and Aleksey N. Kolmogorov. MAISE: Construction of neural network interatomic models and evolutionary structure optimization. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520303301.

Huang:2023:AOS

[Hua23] Li Huang. ACFlow: an open source toolkit for analytic continuation of quantum Monte Carlo data. Computer Physics Communications, 292(??):??, November 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002084.

Hua:2024:MCC

[Hua24] Weijie Hua. MCNOX: a code for computing and interpreting ultrafast nonlinear X-ray spectra of molecules at the multiconfigurational level. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003612.

Heller:2022:MGP

[HvM22] Matthias Heller and Andreas von Manteuffel. MultivariateApart: Generalized partial fractions. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521002861.

Herb:2022:PTI

[HW22] Konstantin Herb and Pol Welter. Parallel time integration using Batched BLAS (Basic Linear Algebra Subprograms) routines. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002939.

Hoppe:2022:ALS

[HWAA22] Nils Hoppe, Josef M. Winter, Stefan Adami, and Nikolaus A. Adams. ALPACA — a level-set based sharp-interface multiresolution solver for conservation laws. Computer Physics Communications, 272(??):??, March 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003581.

He:2023:HHO

[HWL^{+}23] Xiaofeng He, Kun Wang, Tiegang Liu, Yiwei Feng, Bin Zhang, Weixiong Yuan, and Xiaojun Wang. HODG: highorder discontinuous Galerkin methods for solving compressible Euler and Navier–Stokes equations — an opensource component-based development framework. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552300005X.

Hong:2024:CDD

[HWZ24] Xue Hong, Qianrui Wei, and Xiaofei Zhao. Comparison of different discontinuous Galerkin methods based on various reformulations for gKdV equation: Soliton dynamics and blowup. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001036.

Hasbestan:2020:POS

[HXS20] Jaber J. Hasbestan, Cheng-Nian Xiao, and Inanc Senocak. PittPack: an open-source Poisson’s equation solver for extreme-scale computing with accelerators. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300898.

Hatano:2023:AIL

[HY23] Yasuyo Hatano and Shigeyoshi Yamamoto. AIHFLTF: Integrals in Laguerre function bases for electronic structure calculations in atoms. Computer Physics Communications, 284(??): ??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003423.

Han:2022:FEM

[HYL^{+}22] Zherui Han, Xiaolong Yang, Wu Li, Tianli Feng, and Xiulin Ruan. FourPhonon: an extension module to ShengBTE for computing four-phonon scattering rates and thermal conductivity. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002915.

Harris:2023:GGE

[HyLF23] Bradley Harris, Gang yu Liu, and Roland Faller. GenEvaPa: a generic evaporation package for modeling evaporation in molecular dynamics simulations. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002582.

Iteney:2024:PTB

[IGL^{+}24] Hugo Iteney, Javier Antonio Gonzalez Joa, Christophe Le Bourlot, Thomas W. Cornelius, Olivier Thomas, and Jonathan Amodeo. Pyrough: a tool to build 3D samples with rough surfaces for atomistic and finite-element simulations. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300303X.

Ivanov:2021:DEM

[IJVJ21] Aleksei V. Ivanov, Elvar Ö. Jónsson, Tejs Vegge, and Hannes Jónsson. Direct energy minimization based on exponential transformation in density functional calculations of finite and extended systems. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001594.

Ishikawa:2023:PLQ

[IKM^{+}23] Ken-Ichi Ishikawa, Issaku Kanamori, Hideo Matsufuru, Ikuo Miyoshi, Yuta Mukai, Yoshifumi Nakamura, Keigo Nitadori, and Miwako Tsuji. 102 PFLOPS lattice QCD quark solver on Fugaku. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002296.

Ido:2024:UHN

[IKM^{+}24] Kota Ido, Mitsuaki Kawamura, Yuichi Motoyama, Kazuyoshi Yoshimi, Youhei Yamaji, Synge Todo, Naoki Kawashima, and Takahiro Misawa. Update of HΦ: Newly added functions and methods in versions 2 and 3. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552400016X.

Ilten:2021:CFM

[Ilt21] Philip Ilten. CIMBA: Fast Monte Carlo generation using cubic interpolation. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302988.

Iraola:2022:ISE

[IMB^{+}22] Mikel Iraola, Juan L. Mañes, Barry Bradlyn, Matthew K. Horton, Titus Neupert, Maia G. Vergniory, and Stepan S. Tsirkin. IrRep: Symmetry eigenvalues and irreducible representations of ab initio band structures. Computer Physics Communications, 272(??):??, March 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003386.

Ihssen:2024:LDG

[IPSW24] Friederike Ihssen, Jan M. Pawlowski, Franz R. Sattler, and Nicolas Wink. Local discontinuous Galerkin for the functional renormalisation group. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552400105X.

Ivanov:2021:FRA

[IUJ21] A. V. Ivanov, V. M. Uzdin, and H. Jónsson. Fast and robust algorithm for energy minimization of spin systems applied in an analysis of high temperature spin configurations in terms of skyrmion density. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303696.

Jablonski:2020:IAC

[Jab20] A. Jablonski. Improved algorithm for calculating high accuracy values of the Chandrasekhar function. Computer Physics Communications, 251(??):??, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300709.

Jablonski:2022:AFC

[Jab22a] A. Jablonski. Analytical formalism for calculations of parameters needed for quantitative analysis by X-ray photoelectron spectroscopy. Computer Physics Communications, 272(??):??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003453.

Jablonski:2022:MES

[Jab22b] A. Jablonski. Multiple elastic scattering of electrons in condensed matter (new version announcement). Computer Physics Communications, 278(??):??, September 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001217.

Jablonski:2024:SRD

[Jab24] A. Jablonski. Series representation of differential elastic scattering cross section for electrons: Applications in theoretical models of electron transport in condensed matter. Computer Physics Communications, 297(??):??, April 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000079.

Jalalvand:2020:SDP

[JCM20] Morteza Jalalvand, Mohammad A. Charsooghi, and Sarah Mohammadinejad. Smoothed dissipative particle dynamics package for LAMMPS. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520300825.

Juhasz:2021:EGI

[JĎD^{+}21] Zoltan Juhasz, Ján Ďurian, Aranka Derzsi, Štefan Matejčík, Zoltán Donkó, and Peter Hartmann. Efficient GPU implementation of the Particle-in-Cell/Monte-Carlo collisions method for 1D simulation of low-pressure capacitively coupled plasmas. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000503.

Jankovic:2020:GHP

[JDS20] Marija R. Janković, V. Dmitrasinović, and Milovan Suvakov. A guide to hunting periodic three-body orbits with non-vanishing angular momentum. Computer Physics Communications, 250 (??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303807.

Jensen:2020:EID

[JGJ20] Lucas Frese Grønbech Jensen and Niels Grønbech-Jensen. The effects of intrinsic dynamical ghost modes in discretetime Langevin simulations. Computer Physics Communications, 249(??):??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303455.

Jia:2020:CCA

[Jia20] L. Y. Jia. Computer code for applying the variational principle to a coherent-pair condensate: the BCS case. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303133.

Joswig:2023:PPF

[JKKN23] Fabian Joswig, Simon Kuberski, Justus T. Kuhlmann, and Jan Neuendorf. pyerrors: a Python framework for error analysis of Monte Carlo data. Computer Physics Communications, 288 (??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000954.

Jeong:2022:WOO

[JKST22] Injun Jeong, Sunghyun Kang, Stefano Scopel, and Gaurav Tomar. WimPyDD: an object-oriented Python code for the calculation of WIMP direct detection signals. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000601.

Jo:2022:DGH

[JKSY22] Gahyung Jo, Jae-Min Kwon, Janghoon Seo, and Eisung Yoon. Development of a gyrokinetic hyperbolic solver based on discontinuous Galerkin method in tokamak geometry. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003775.

Jiang:2024:ANM

[JLL^{+}24] Tao Jiang, Yu-Hang Liu, Qiang Li, Jin-Lian Ren, and Deng-Shan Wang. An accelerated novel meshless coupled algorithm for non-local nonlinear behavior in 2D/3D space-fractional GPEs. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003685.

Jiang:2024:UPF

[JLW24] Wenjing Jiang, Ziling Lu, and Jian Wang. Uniform patterns formation based on Gray–Scott model for 3D printing. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003193.

Janssens:2024:PTM

[JM24] N. Janssens and J. Meyers. Parallel-in-time multiple shooting for optimal control problems governed by the Navier–Stokes equations. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003648.

Jiang:2021:GAF

[JMOC21] Fei Jiang, Kazuki Matsumura, Junji Ohgi, and Xian Chen. A GPU-accelerated fluid-structure-interaction solver developed by coupling finite element and lattice Boltzmann methods. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303210.

Jin:2023:PEP

[JPJ^{+}23] Gan Jin, Hongsheng Pang, Yuyang Ji, Zujian Dai, and Lixin He. PYATB: an efficient Python package for electronic structure calculations using ab initio tight-binding model. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001893.

Jalowiecki:2021:BFS

[JRG21] Konrad Jałowiecki, Marek M. Rams, and Bartłomiej Gardas. Brute-forcing spin-glass problems with CUDA. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030360X.

Jin:2021:LPS

[JRS^{+}21] Shi Jin, Kenneth J. Roche, Ionel Stetcu, Ibrahim Abdurrahman, and Aurel Bulgac. The LISE package: Solvers for static and time-dependent superfluid local density approximation equations in three dimensions. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521002423.

Johnson:2022:PSL

[JS22] P. R. Johnson and C. E. Sosolik. SAFARI — a low and hyperthermal energy ion scattering simulation. Computer Physics Communications, 280(??):??, November 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001989.

Jha:2024:GAT

[JS24] Raghav G. Jha and Abhishek Samlodia. GPU-acceleration of tensor renormalization with PyTorch using CUDA. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002862.

Jarleblad:2024:FSD

[JSS^{+}24] H. Järleblad, L. Stagner, M. Salewski, J. Eriksson, M. Nocente, B. S. Schmidt, and M. Rud Larsen. A framework for synthetic diagnostics using energetic-particle orbits in tokamaks. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002758.

Jakubczyk:2021:MPC

[JWK^{+}21] Paweł Jakubczyk, Andrzej Wal, Michał Kaczor, Dorota Jakubczyk, Mirosław Łabuz, and Jan Milewski. A Maple package for combinatorial aspects of Bethe Ansatz. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303568.

Jadach:2023:MPM

[JWW^{+}23] S. Jadach, B. F. L. Ward, Z. Was, S. A. Yost, and A. Siodmok. Multi-photon Monte Carlo event generator KKMCee for lepton and quark pair production in lepton colliders. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002752.

Jambunathan:2023:TFP

[JYL^{+}23] Revathi Jambunathan, Zhi Yao, Richard Lombardini, Aaron Rodriguez, and Andrew Nonaka. Two-fluid physical modeling of superconducting resonators in the ARTEMIS framework. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001819.

Jiao:2022:KEM

[JZW^{+}22] Shizhe Jiao, Zhenlin Zhang, Kai Wu, Lingyun Wan, Huanhuan Ma, Jielan Li, Sheng Chen, Xinming Qin, Jie Liu, Zijing Ding, Jinlong Yang, Yingzhou Li, Wei Hu, Lin Lin, and Chao Yang. KSSOLV 2.0: an efficient MATLAB toolbox for solving the Kohn–Sham equations with plane-wave basis set. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001436.

Karthein:2021:AMC

[KAB^{+}21] J. Karthein, D. Atanasov, K. Blaum, D. Lunney, V. Manea, and M. Mougeot. Analysis methods and code for very highprecision mass measurements of unstable isotopes. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100182X.

Kroger:2024:GGP

[KAG24] Martin Kröger, Samarth Agrawal, and Sandra Galmarini. Generalized geometric pore size distribution code GPSD-3D for periodic systems composed of monodisperse spheres. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001358.

Kaltiaisenaho:2020:PTP

[Kal20] Toni Kaltiaisenaho. Photon transport physics in Serpent 2 Monte Carlo code. Computer Physics Communications, 252 (??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300035.

Kandemir:2023:OCT

[Kan23] Mustafa Kandemir. OPSimTool: a custom tool for optical photon simulation in Geant4. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523002187.

Kandemir:2024:CDC

[Kan24] Mustafa Kandemir. A custom detector construction pattern for Geant4 applications. Computer Physics Communications, 300 (??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001061.

Karamitros:2022:MMM

[Kar22] Dimitrios Karamitros. MiMeS: Misalignment mechanism solver. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000297.

Karakoc:2023:BPC

[Kar23a] Mesut Karakoç. BiFold: a Python code for the calculation of double-folded (bifold) potentials with density-in/dependent nucleon–nucleon interactions. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003320.

Karamitros:2023:NNS

[Kar23b] Dimitrios Karamitros. NSC++: Non-standard cosmologies in C++. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000887.

Kumar:2024:ITL

[KAS24] Gokul Raman Arumugam Kumar, James P. Andrews, and Ulf D. Schiller. Implementation of a ternary lattice Boltzmann model in LAMMPS. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002436.

Keita:2021:ESO

[KBB21] Sana Keita, Abdelaziz Beljadid, and Yves Bourgault. Efficient second-order semi-implicit finite element method for fourth-order nonlinear diffusion equations. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302812.

Kleiber:2024:EGG

[KBH^{+}24] R. Kleiber, M. Borchardt, R. Hatzky, A. Könies, H. Leyh, A. Mishchenko, J. Riemann, C. Slaby, J. M. García-Regaña, E. Sánchez, and M. Cole. EUTERPE: a global gyrokinetic code for stellarator geometry. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523003582.

Kulyk:2020:CFC

[KBSH20] Nadiia Kulyk, Daniel Berger, Ana-Suncana Smith, and Jens Harting. Catalytic flow with a coupled finite difference — lattice Boltzmann scheme. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552030206X.

Kim:2022:EDI

[KBSL22] Young Jun Kim, Benjamin Bouscasse, Sopheak Seng, and David Le Touzé. Efficiency of diagonally implicit Runge–Kutta time integration schemes in incompressible two-phase flow simulations. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001345.

Kittelmann:2021:ENS

[KC21] T. Kittelmann and X.-X. Cai. Elastic neutron scattering models for NCrystal. Computer Physics Communications, 267(??):

??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001946.

Khan:2024:LSS

[KCJX24] Niaz Ali Khan, Wen Chen, Munsif Jan, and Gao Xianlong. Linear-scale simulations of quench dynamics. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000559.

Kaye:2022:PLE

[KCS22] Jason Kaye, Kun Chen, and Hugo U. R. Strand. libdlr: Efficient imaginary time calculations using the discrete Lehmann representation. Computer Physics Communications, 280 (??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001771.

Kurnia:2023:NOS

[KD23] Ruddy Kurnia and Guillaume Ducrozet. NEMOH: Open-source boundary element solver for computation of first- and secondorder hydrodynamic loads in the frequency domain. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002308.

Kroger:2023:ZPS

[KDHL23] Martin Kröger, Joseph D. Dietz, Robert S. Hoy, and Clarisse Luap. The Z1+ package: Shortest multiple disconnected path for the analysis of entanglements in macromolecular systems. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002867.

Kim:2023:TPP

[KDIN^{+}23] Kyungjoo Kim, Oscar H. Díaz-Ibarra, Habib N. Najm, Judit Zádor, and Cosmin Safta. TChem: a performance portable parallel software toolkit for complex kinetic mechanisms. Computer Physics Communications, 285(??):??, April 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003472.

Knyazev:2023:IVM

[KDK23] A. R. Knyazev, M. Dorf, and S. I. Krasheninnikov. Implementation and verification of a model linearized multispecies collision operator in the COGENT code. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001741.

Kendl:2024:TGT

[Ken24] Alexander Kendl. TIFF: Gyrofluid turbulence in full-f and full-k. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002989.

Kawase:2023:FCS

[KF23] Yoshiaki Kawase and Keisuke Fujii. Fast classical simulation of Hamiltonian dynamics by simultaneous diagonalization using Clifford transformation with parallel computation. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000656.

Komen:2020:AND

[KFC^{+}20] E. M. J. Komen, E. M. A. Frederix, T. H. J. Coppen, V. D’Alessandro, and J. G. M. Kuerten. Analysis of the numerical dissipation rate of different Runge–Kutta and velocity interpolation methods in an unstructured collocated finite volume method in OpenFOAM(R). Computer Physics Communications, 253(??):??, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300047.

Kuhl:2024:ISV

[KFHR24] Niklas Kühl, Hendrik Fischer, Michael Hinze, and Thomas Rung. An incremental singular value decomposition approach for large-scale spatially parallel and distributed but temporally serial data — applied to technical flows. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003673.

Krek:2021:BKC

[KFPV21] Janez Krek, Yangyang Fu, Guy M. Parsey, and John P. Verboncoeur. Benchmark of the KGMf with a coupled Boltzmann equation solver. Computer Physics Communications, 260(??): ??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303702.

Kuhn:2021:POB

[KG21] Christoph Kühn and Rodion Groll. picFoam: an Open-FOAM based electrostatic Particle-in-Cell solver. Computer Physics Communications, 262(??):??, May 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000187.

Klapetek:2021:GOS

[KGN^{+}21] Petr Klapetek, Petr Grolich, David Nezval, Miroslav Valtr, Radek Šlesinger, and David Nečas. GSvit — an open source FDTD solver for realistic nanoscale optics simulations. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001375.

Kozicki:2022:IHP

[KGT22] Janek Kozicki, Anton Gladky, and Klaus Thoeni. Implementation of high-precision computation capabilities into the open-source dynamic simulation framework YADE. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002794.

Kaufmann:2023:APP

[KH23] Josef Kaufmann and Karsten Held. ana cont: Python package for analytic continuation. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002387.

Kaidisch:2024:PFC

[KHKL24] Siegfried Kaidisch, Thomas U. Hilger, Andreas Krassnigg, and Wolfgang Lucha. Pole-fitting for complex functions: Enhancing standard techniques by artificial-neural-network classifiers and regressors. Computer Physics Communications, 295(??): ??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003430.

Kielbik:2023:SDD

[KHR^{+}23] Rafał Kiełbik, Krzysztof Hałagan, Kamil Rudnicki, Grzegorz Jabłoński, Piotr Polanowski, and Jarosław Jung. Simulation of diffusion in dense molecular systems on ARUZ

— massively-parallel FPGA-based machine. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003101.

Kestner:2020:FAC

[KIK20] Dan Kestner, Glenn Ierley, and Alex Kostinski. A fast algorithm for computing a matrix transform used to detect trends in noisy data. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552030165X.

Kim:2023:SSA

[KJL^{+}23] Geon Kim, Yunsong Jung, Myeongkyu Lee, Eisung Yoon, and Sangjoon Ahn. SANTA: a safety analysis code for neutron absorbers in spent nuclear fuel pools. Computer Physics Communications, 282(??):??, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002326.

Kafri:2016:BPN

[KK16] H. Q. Kafri and S. A. Khuri. Bratu’s problem: a novel approach using fixed-point iterations and Green’s functions. Computer Physics Communications, 198(??):97–104, January 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465515003446. See corrigen-

dum [Ano20c].

Korcyl:2020:TLQ

[KK20]  Grzegorz Korcyl and Piotr Korcyl.  Towards Lattice

Quantum Chromodynamics on FPGA devices. Computer Physics Communications, 249(??):??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303686.

Kim:2024:EPB

[KKB^{+}24] Minhyo Kim, Pilsung Kim, Riccardo Bassiri, Kiran Prasai, Martin M. Fejer, and Kyung ha Lee. ePDFpy: a Python-based interactive GUI tool for electron pair distribution function analysis of amorphous materials. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000602.

Klappert:2021:IDS

[KKL21] Jonas Klappert, Sven Yannick Klein, and Fabian Lange. Interpolation of dense and sparse rational functions and other improvements in FireFly. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000825.

Kan:2023:GPI

[KKLZ23a] Yi-Kai Kan, Franz X. Kärtner, Sabine Le Borne, and Jens-Peter M. Zemke. A GPU-parallelized interpolation-based fast multipole method for the relativistic space-charge field calculation. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001704.

Kan:2023:RSC

[KKLZ23b] Yi-Kai Kan, Franz X. Kärtner, Sabine Le Borne, and Jens-Peter M. Zemke. Relativistic space-charge field calculation by interpolation-based treecode. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000139.

Kanemura:2020:HCV

[KKM^{+}20] Shinya Kanemura, Mariko Kikuchi, Kentarou Mawatari, Kodai Sakurai, and Kei Yagyu. H-COUP Version 2: a program for one-loop corrected Higgs boson decays in non-minimal Higgs sectors. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552030240X.

Kronheim:2022:TBI

[KKP22] B. S. Kronheim, M. P. Kuchera, and H. B. Prosper. TensorBNN: Bayesian inference for neural networks using Tensor-Flow. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002800.

Kim:2021:PLP

[KKPC21] Ki-Ha Kim, Ji-Hoon Kang, Xiaomin Pan, and Jung-Il Choi. PaScaL TDMA: a library of parallel and scalable solvers for massive tridiagonal systems. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030357X.

Kim:2023:PVS

[KKPC23] Ki-Ha Kim, Ji-Hoon Kang, Xiaomin Pan, and Jung-Il Choi. PaScaL TCS: a versatile solver for large-scale turbulent convective heat transfer problems with temperature-dependent fluid properties. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001248.

Khasianevich:2024:FEA

[KKSV24] Uladzimir Khasianevich, Wojciech Kotlarski, Dominik Stöckinger, and Alexander Voigt. FlexibleSUSY extended to automatically compute physical quantities in any beyond the standard model theory: Charged lepton flavor violation processes, Higgs decays, and user-defined observables. Computer Physics Communications, 302(??):??, September 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552400167X.

Kim:2024:NIG

[KKY24] Innyoung Kim, Sejin Kim, and Donghyun You. Non-iterative generation of an optimal mesh for a blade passage using deep reinforcement learning. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003077.

Klappert:2020:RRF

[KL20] Jonas Klappert and Fabian Lange. Reconstructing rational functions with FireFly. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303066.

Kerbizi:2022:SAS

[KL22]  Albi Kerbizi and Leif Lönnblad.  StringSpinner —

adding spin to the PYTHIA string fragmentation. Computer Physics Communications, 272(??):??, March 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003465.

Kerbizi:2023:ESH

[KL23a] Albi Kerbizi and Leif Lönnblad. Extending StringSpinner to handle vector-meson spin. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552300231X.

Kim:2023:TGA

[KL23b] Changhyun Kim and Byoungho Lee. TORCWA: GPU-accelerated Fourier modal method and gradient-based optimization for metasurface design. Computer Physics Communications, 282 (??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002715.

Klaes:2021:MPI

[KLD^{+}21] Benjamin Klaes, Rodrigue Lardé, Fabien Delaroche, Stefan Parviainen, Nicolas Rolland, Shyam Katnagallu, Baptiste Gault, and François Vurpillot. A model to predict image formation in the three-dimensional field ion microscope. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520301235.

Kumar:2022:DAS

[KLM^{+}22] Anmol Kumar, Rafael López, Frank Martínez, Guillermo Ramírez, Ignacio Ema, David Zorrilla, Sachin D. Yeole, and Shridhar R. Gadre. DAMQT 3: Advanced suite for the analysis of molecular density and related properties in large systems. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001795.

Klappert:2021:IRK

[KLMU21] Jonas Klappert, Fabian Lange, Philipp Maierhöfer, and Johann Usovitsch. Integral reduction with Kira 2.0 and finite field methods. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001363.

Krause:2020:EPC

[KM20] Marcel Krause and Margarete Mühlleitner. ewN2HDECAY — a program for the calculation of electroweak one-loop corrections to Higgs decays in the Next-to-Minimal Two-Higgs-Doublet Model including state-of-the-art QCD corrections. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046551930298X.

Krivenko:2022:PTS

[KM22] Igor Krivenko and Andrey S. Mishchenko. TRIQS/SOM 2.0: Implementation of the stochastic optimization with consistent constraints for analytic continuation. Computer Physics Communications, 280(??):??, November 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002107.

Kohno:2023:FEP

[KM23] H. Kohno and J. R. Myra. A finite element procedure for time-dependent radio-frequency sheaths based on a two-dimensional microscale fluid model. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001868.

Kenneweg:2024:QMT

[KMBP24] Tristan Kenneweg, Stefan Mueller, Tobias Brixner, and Walter Pfeiffer. QDT — a Matlab toolbox for the simulation of coupled quantum systems and coherent multidimensional spectroscopy. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003764.

Koshkarov:2021:MDH

[KMD^{+}21] O. Koshkarov, G. Manzini, G. L. Delzanno, C. Pagliantini, and V. Roytershteyn. The multi-dimensional Hermitediscontinuous Galerkin method for the Vlasov–Maxwell equations. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000266.

Krishnamoorthy:2020:EMO

[KMG^{+}20] Aravind Krishnamoorthy, Ankit Mishra, Nicholas Grabar, Nitish Baradwaj, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. Evolutionary multi-objective optimization and Pareto-frontal uncertainty quantification of interatomic forcefields for thermal conductivity simulations. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301466.

Kloffel:2021:ISA

[KMM21] Tobias Klöffel, Gerald Mathias, and Bernd Meyer. Integrating state of the art compute, communication, and autotuning strategies to multiply the performance of ab initio molecular dynamics on massively parallel multi-core supercomputers. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303684.

Khan:2021:DFB

[KMN21] Shuaib Ahmad Khan, Jubin Mitra, and Tapan K. Nayak. Development of FPGA based phase alignment logic for the high speed protocol in HEP experiments. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303155.

Kivioja:2022:GAT

[KMR22] Markus Kivioja, Sanna Mönkölä, and Tuomo Rossi. GPUaccelerated time integration of Gross–Pitaevskii equation with discrete exterior calculus. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001461.

Krachkov:2020:ATP

[KMS20a] P. A. Krachkov, A. I. Milstein, and A. G. Shamov. Amplitudes of two-photon processes and extension of BDK generator. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302447.

Krause:2020:PCE

[KMS20b] Marcel Krause, Margarete Mühlleitner, and Michael Spira. 2HDECAY — a program for the calculation of electroweak one-loop corrections to Higgs decays in the Two-Higgs-Doublet Model including state-of-the-art QCD corrections. Computer Physics Communications, 246(??):??, January 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302292.

Kneur:2023:SCC

[KMU^{+}23] Jean-Loïc Kneur, Gilbert Moultaka, Michaël Ughetto, Dirk Zerwas, and Abdelhak Djouadi. SuSpect3: a C++ code for the supersymmetric and Higgs particle spectrum of the MSSM. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001509.

Kurita:2023:ITW

[KMY^{+}23] Kensuke Kurita, Takahiro Misawa, Kazuyoshi Yoshimi, Kota Ido, and Takashi Koretsune. Interface tool from

Wannier90 to RESPACK: wan2respack. Computer Physics

Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523001996.

Kumar:2023:FGA

[KNJ^{+}23] Prabhat Kumar, Andrew Nonaka, Revathi Jambunathan, Girish Pahwa, Sayeef Salahuddin, and Zhi Yao. FerroX: a GPU-accelerated, 3D phase-field simulation framework for modeling ferroelectric devices. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523001029.

Kolmes:2021:MMI

[KOF21] E. J. Kolmes, I. E. Ochs, and N. J. Fisch. MITNS: Multiple-Ion Transport Numerical Solver for magnetized plasmas. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030237X.

Kolodziej:2022:CNV

[Koł22] Karol Kołodziej. carlomat 4.0, a new version of the general purpose Monte Carlo program. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000480.

Kolodziej:2023:PGP

[Koł23] Karol Kołodziej. PSGen, a generator of phase space parameterizations for the multichannel Monte Carlo integration. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002151.

Koretsune:2023:CML

[Kor23] Takashi Koretsune. Construction of maximally-localized Wannier functions using crystal symmetry. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522003642.

Kachelriess:2023:AIR

[KOT23] M. Kachelrieß, S. Ostapchenko, and J. Tjemsland. AAfrag 2.01: interpolation routines for Monte Carlo results on secondary production including light antinuclei in hadronic interactions. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000437.

Kozicki:2023:VAT

[Koz23] Janek Kozicki. Very accurate time propagation of coupled Schrödinger equations for femto- and attosecond physics and chemistry, with C++ source code. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001844.

Kincl:2023:GTR

[KP23] Ondrej Kincl and Michal Pavelka. Globally time-reversible fluid simulations with smoothed particle hydrodynamics. Computer Physics Communications, 284(??):??, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003125.

Kobayashi:2021:QTS

[KPL^{+}21] Michikazu Kobayashi, Philippe Parnaudeau, Francky Luddens, Corentin Lothodé, Luminita Danaila, Marc Brachet, and Ionut Danaila. Quantum turbulence simulations using the Gross–Pitaevskii equation: High-performance computing and new numerical benchmarks. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302769.

Karnes:2024:PBS

[KPR^{+}24] John J. Karnes, Andrew J. Pascall, Christoph Rehbock, Vaijayanthi Ramesh, Marcus A. Worsley, Stephan Barcikowski, Elaine Lee, and Brian Giera. Particle-based simulations of electrophoretic deposition with adaptive physics models. Computer Physics Communications, 297(??):??, April 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523004071.

Kang:2021:EFT

[KPST21] Zhong-Bo Kang, Alexei Prokudin, Nobuo Sato, and John Terry. Efficient Fourier transforms for transverse momentum dependent distributions. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302903.

Kim:2023:CCD

[KR23] Minkyung Kim and Junsuk Rho. CDPDS: Coupled dipole method-based photonic dispersion solver. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465522002120.

Kim:2020:MLM

[KRC^{+}20] Deok-Soo Kim, Joonghyun Ryu, Youngsong Cho, Mokwon Lee, Jehyun Cha, Chanyoung Song, Sang Wha Kim, Roman A. Laskowski, Kokichi Sugihara, Jong Bhak, and Seong Eon Ryu. MGOS: a library for molecular geometry and its operating system. Computer Physics Communications, 251(??): ??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519304047.

Kraposhin:2022:NAB

[KRE22] Matvey V. Kraposhin, Daniil A. Ryazanov, and Tatiana G. Elizarova. Numerical algorithm based on regularized equations for incompressible flow modeling and its implementation in OpenFOAM. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003283.

Kaur:2021:FFP

[KRG21] Pardeep Kaur, Arko Roy, and Sandeep Gautam. FORTRESS: FORTRAN programs for solving coupled Gross–Pitaevskii equations for spin-orbit coupled spin-1 Bose–Einstein condensate. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552030326X.

Karimi:2023:IDE

[KRJ23] Javad Karimi, Faezeh Rahmani, and S. Bijan Jia. Improving the detection efficiency of IRAND based on machine learning. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523001789.

Koide:2024:MPA

[KRL^{+}24] Akihiro Koide, Sara Rabouli, Pierre Le Meur, Sylvain Tricot, Philippe Schieffer, Didier Sébilleau, and Calogero R. Natoli. MASAP: a package for atomic scattering amplitude in solids. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003338.

Kopper:2023:FHF

[KSC^{+}23] Patrick Kopper, Anna Schwarz, Stephen M. Copplestone, Philip Ortwein, Stephan Staudacher, and Andrea Beck. A framework for high-fidelity particle tracking on massively parallel systems. Computer Physics Communications, 289(??):??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001078.

Kalt:2023:IVQ

[KSDH23] Victor Kalt, Georges Sadaka, Ionut Danaila, and Frédéric Hecht. Identification of vortices in quantum fluids: Finite element algorithms and programs. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003253.

Khanwale:2022:FCF

[KSF^{+}22] Makrand A. Khanwale, Kumar Saurabh, Milinda Fernando, Victor M. Calo, Hari Sundar, James A. Rossmanith, and Baskar Ganapathysubramanian. A fully-coupled framework for solving Cahn–Hilliard Navier–Stokes equations: Secondorder, energy-stable numerical methods on adaptive octree based meshes. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552200220X.

Khire:2022:MSC

[KSG22] Subodh S. Khire, Nityananda Sahu, and Shridhar R. Gadre. MTASpec software for calculating the vibrational IR and Raman spectra of large molecules at ab initio level. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002873.

Korkin:2022:PGW

[KSIL22] S. Korkin, A. M. Sayer, A. Ibrahim, and A. Lyapustin. A practical guide to writing a radiative transfer code. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003106.

Kim:2022:NFP

[KSJ^{+}22] Dongkyu Kim, Janghoon Seo, Gahyung Jo, Jae-Min Kwon, and Eisung Yoon. Nonlinear Fokker–Planck collision operator in Rosenbluth form for gyrokinetic simulations using discontinuous Galerkin method. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001783.

Kolotinskii:2023:OFG

[KT23] D. Kolotinskii and A. Timofeev. OpenDust: a fast GPUaccelerated code for the calculation of forces acting on microparticles in a plasma flow. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000917.

Kandemir:2022:NGB

[KTF22] Mustafa Kandemir, Emrah Tiras, and Vincent Fischer. NuSD: a Geant4 based simulation framework for segmented antineutrino detectors. Computer Physics Communications, 277 (??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001060.

Kuberski:2024:LMD

[Kub24] Simon Kuberski. Low-mode deflation for twisted-mass and RHMC reweighting in lattice QCD. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000961.

Kulikov:2020:NCN

[Kul20] Igor Kulikov. A new code for the numerical simulation of relativistic flows on supercomputers by means of a low-dissipation scheme. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302538.

Kursten:2023:AAP

[Kür23] Rüdiger Kürsten. Aligning active particles py package. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001194.

Kutepov:2020:SCG

[Kut20] A. L. Kutepov. Self-consistent GW method: O(N) algorithm for polarizability and self energy. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520302368.

Kutnii:2024:DCL

[Kut24] Sergii Kutnii. Dirac: a command-line γ-matrix calculator. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002461.

Kadkhodaei:2020:STT

[KvdW20] Sara Kadkhodaei and Axel van de Walle. Software tools for thermodynamic calculation of mechanically unstable phases from first-principles data. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519300141.

Konopka:2021:DSM

[KvH21] Piotr Konopka and Barthélémy von Haller. Data sampling methods in the ALICE O^{2} distributed processing system. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302794.

Klymenko:2021:NEP

[KVSC21] M. V. Klymenko, J. A. Vaitkus, J. S. Smith, and J. H. Cole. NanoNET: an extendable Python framework for semi-empirical tight-binding models. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520303283.

Kumar:2023:QEG

[KW23] Manoj Kumar and Martin Weigel. Quasi-exact groundstate extrapolation for the random-field Potts model. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000309.

K:2021:CSF

[KWK^{+}21] Akshay Krishna A. K., Eddie Wadbro, Christof Köhler, Pavlin Mitev, Peter Broqvist, and Jolla Kullgren. CCS: a software framework to generate two-body potentials using Curvature Constrained Splines. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302915.

Karzhaubayev:2024:DGE

[KWZ24] Kairzhan Karzhaubayev, Lian-Ping Wang, and Dauren Zhakebayev. DUGKS-GPU: an efficient parallel GPU code for 3D turbulent flow simulations using Discrete Unified Gas Kinetic Scheme. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001395.

Kudo:2024:FEC

[KYH24] Shuhei Kudo, Yusaku Yamamoto, and Takeo Hoshi. A fast and efficient computation method for reflective diffraction simulations. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003740.

Lazzarin:2021:MTS

[LAC21] Marco Lazzarin, Simone Alioli, and Stefano Carrazza. MCNNTUNES: Tuning shower Monte Carlo generators with machine learning. Computer Physics Communications, 263(??): ??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000448.

Li:2024:EPC

[LACL24] Zhenglu Li, Gabriel Antonius, Yang-Hao Chan, and Steven G. Louie. Electron–phonon coupling from GW perturbation theory: Practical workflow combining BerkeleyGW, ABINIT, and EPW. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552300348X.

Li:2021:QSP

[LAD^{+}21] Fei Li, Weiming An, Viktor K. Decyk, Xinlu Xu, Mark J. Hogan, and Warren B. Mori. A quasi-static particle-in-cell algorithm based on an azimuthal Fourier decomposition for highly efficient simulations of plasma-based acceleration: QPAD. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303921.

Latosh:2023:F

[Lat23] B. Latosh. FeynGrav 2.0. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002163.

Liu:2024:SER

[LB24] Yang Liu and Dietmar Block. Stokes–Einstein relation for binary mixtures. Computer Physics Communications, 300(??): ??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001073.

Lauriello:2023:DAR

[LBM^{+}23] N. Lauriello, G. Boccardo, D. Marchisio, M. Lísal, and A. Buffo. Development of an automated reliable method to compute transport properties from DPD equilibrium simulations: Application to simple fluids. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001881.

Lalescu:2022:EPT

[LBRW22] Cristian C. Lalescu, Bérenger Bramas, Markus Rampp, and Michael Wilczek. An efficient particle tracking algorithm for large-scale parallel pseudo-spectral simulations of turbulence. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001254.

Lahooti:2023:DFS

[LBS^{+}23] Mohsen Lahooti, Yan Bao, David Scott, Rafael Palacios, and Spencer J. Sherwin. LES/DNS fluid-structure interaction simulation of non-linear slender structures in Nektar++ framework. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002478.

Liu:2024:OPQ

[LC24] Hengjie Liu and Ziguang Chen. On the one-point quadrature discretization in peridynamics: a novel perspective from Monte Carlo integration. Computer Physics Communications, 298 (??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000389.

Li:2023:ACM

[LCL^{+}23] Lei Li, Ryan A. Ciufo, Jiyoung Lee, Chuan Zhou, Bo Lin, Jaeyoung Cho, Naman Katyal, and Graeme Henkelman. Atomcentered machine-learning force field package. Computer Physics Communications, 292(??):??, November 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300228X.

Liu:2021:SPI

[LCZ^{+}21] Gui-Bin Liu, Miao Chu, Zeying Zhang, Zhi-Ming Yu, and Yugui Yao. SpaceGroupIrep: a package for irreducible representations of space group. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001053.

Linghu:2024:SOW

[LDGN24] Jiale Linghu, Hao Dong, Weifeng Gao, and Yufeng Nie. Self-optimization wavelet-learning method for predicting nonlinear thermal conductivity of highly heterogeneous materials with randomly hierarchical configurations. Computer Physics Communications, 295(??):??, February 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003144.

Luo:2021:CPC

[LDW^{+}21] Chenxing Luo, Xin Deng, Wenzhong Wang, Gaurav Shukla, Zhongqing Wu, and Renata M. Wentzcovitch. cij: a Python code for quasiharmonic thermoelasticity. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100179X.

Lahdour:2021:OIM

[LEE^{+}21a] M. Lahdour, T. El Bardouni, O. El Hajjaji, E. Chakir, H. Ziani, Jamal Al Zain, E. Chham, and M. El Barbari. OpenNTP: Implementation of the S_{N} method in Cartesian 2D geometry and the CP method in cylindrical and spherical 1D geometry. Computer Physics Communications, 261(??): ??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520304070.

Lee:2021:LPT

[Lee21b] Roman N. Lee. Libra: a package for transformation of differential systems for multiloop integrals. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001703.

Lahdour:2024:EOP

[LEE^{+}24] M. Lahdour, T. El Bardouni, O. El Hajjaji, J. El Bakkali, J. Al-Zain, S. Oulad-Belayachi, H. Ziani, Abdelghani Idrissi, and S. El Maliki El Hlaibi. ERSN-OpenMC-Py: a Python-based open-source software for OpenMC Monte Carlo code. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000444.

Liu:2022:EAT

[LEL^{+}22] Zhong-Li Liu, C. E. Ekuma, Wei-Qi Li, Jian-Qun Yang, and Xing-Ji Li. ElasTool: an automated toolkit for elastic constants calculation. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002927.

Lopez:2020:TAA

[LF20] Carlos Antonio Cruz López and Juan Luis François. Two alternative approaches to the solution of cyclic chains in transmutation and decay problems. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300643.

Liang:2024:NHO

[LF24] Tian Liang and Lin Fu. A new high-order shock-capturing TENO scheme combined with skew-symmetric-splitting method for compressible gas dynamics and turbulence simulation. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001590.

Liu:2024:PMM

[LFL^{+}24] Weier Liu, Yangde Feng, Ruilin Li, Chenhan Bai, and Beifang Niu. Peridynamic modeling for multiscale heat transport of phonon Boltzmann transport equation. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000808.

Luo:2020:GAA

[LFZ20] Xiao-Lin Luo, Jie Feng, and Hong-Hao Zhang. A genetic algorithm for astroparticle physics studies. Computer Physics Communications, 250(??):??, May 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551930195X.

Lapolla:2021:BEC

[LG21a] Alessio Lapolla and Aljaz Godec. BetheSF: Efficient computation of the exact tagged-particle propagator in singlefile systems via the Bethe eigenspectrum. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302733.

Lapolla:2021:BVP

[LG21b] Alessio Lapolla and Aljaz Godec. BetheSF V2: 3-point propagator and additional external potentials. Computer Physics Communications, 269(??):??, December 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002435.

Lin:2023:TGA

[LG23] Chih-Chuen Lin and Vikram Gavini. TTDFT: a GPU accelerated Tucker tensor DFT code for large-scale Kohn–Sham DFT calculations. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002351.

Li:2020:HPC

[LGBJ20] Wenxian Li, Jon Grumer, Tomas Brage, and Per Jönsson. Hfszeeman 95 — a program for computing weak and intermediate magnetic-field- and hyperfine-induced transition rates. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300485.

Li:2020:DFR

[LGDF20] Jingwei Li, Zhiming Gao, Zihuan Dai, and Xinlong Feng. Divergence-free radial kernel for surface Stokes equations based on the surface Helmholtz decomposition. Computer Physics Communications, 256(??):??, November 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301879.

Lingis:2022:MSL

[LGK^{+}22] Danielius Lingis, Mindaugas Gaspariunas, Vitalij Kovalevskij, Arturas Plukis, and Vidmantas Remeikis. A model to simulate large angle Rutherford backscattering spectra in GEANT4. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100299X.

Lopez:2022:GOS

[LH22] Joaquín López and Julio Hernández. gVOF: an open-source package for unsplit geometric volume of fluid methods on arbitrary grids. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001199.

Lopez:2024:PIH

[LH24] Joaquín López and Julio Hernández. On polytope intersection by half-spaces and hyperplanes for unsplit geometric volume of fluid methods on arbitrary grids. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000900.

Liu:2020:FDV

[LHC20] Xin Liu, Zhengkang He, and Zhangxin Chen. A fully discrete virtual element scheme for the Cahn–Hilliard equation in mixed form. Computer Physics Communications, 246(??): ??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519302577.

Lopez:2020:VEN

[LHG^{+}20] Joaquín López, Julio Hernández, Pablo Gómez, Claudio Zanzi, and Rosendo Zamora. VOFTools 5: an extension to nonconvex geometries of calculation tools for volume of fluid methods. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300928.

Liu:2024:GID

[LHG24] Peiyao Liu, Changsheng Huang, and Zhaoli Guo. GPU implementation of the discrete unified gas kinetic scheme for low-speed isothermal flows. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002539.

Ly:2024:TDH

[LHTP^{+}24] Duy-Nhat Ly, Duong D. Hoang-Trong, Ngoc-Hung Phan, Duy-Anh P. Nguyen, and Van-Hoang Le. Two-dimensional heliumlike atom in a homogeneous magnetic field: Numerically exact solutions. Computer Physics Communications, 298(??): ??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552400033X.

Luo:2024:NDV

[LHWX24] Hao Luo, Shanfang Huang, Kan Wang, and Xiang Xiao. A newly developed and verified transport capacity of Monte Carlo photon particles in RMC. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002801.

Lang:2020:VVH

[LHZ20] Robert A. Lang, Riley J. Hickman, and Tao Zeng. VHEGEN: a vibronic Hamiltonian expansion generator for trigonal and tetragonal polyatomic systems. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303029.

Lindblad:2024:NDC

[LIG^{+}24] Daniel Lindblad, João Isler, Margarida Moragues Ginard, Spencer J. Sherwin, and Chris D. Cantwell. Nektar++: Development of the compressible flow solver for jet aeroacoustics. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001267.

Lin:2023:DRU

[LJH^{+}23] Hongyuan Lin, Chongwen Jiang, Shuyao Hu, Zhenxun Gao, and Chun-Hian Lee. Disturbance region update method with preconditioning for steady compressible and incompressible flows. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552200354X.

Liu:2022:HSE

[LJQ^{+}22] Chang Liu, Stephen C. Jardin, Hong Qin, Jianyuan Xiao, Nathaniel M. Ferraro, and Joshua Breslau. Hybrid simulation of energetic particles interacting with magnetohydrodynamics using a slow manifold algorithm and GPU acceleration. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000315.

Lusher:2021:OAC

[LJS21] David J. Lusher, Satya P. Jammy, and Neil D. Sandham. OpenSBLI: Automated code-generation for heterogeneous computing architectures applied to compressible fluid dynamics on structured grids. Computer Physics Communications, 267(??): ??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001752.

Lee:2023:PPP

[LKK23] Kyunghoon Lee, Jun Hyeong Kim, and Woo Youn Kim. pyMCD: Python package for searching transition states via the multicoordinate driven method. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001765.

Lee:2021:PAP

[LKP21] DongGun Lee, TaeHyung Kim, and Q-Han Park. Performance analysis of parallelized PSTD-FDTD method for large-scale electromagnetic simulation. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520303039.

Li:2024:MCB

[LKP24] Dongming Li, James Kestyn, and Eric Polizzi. A method of calculating bandstructure in real-space with application to all-electron and full potential. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523003594.

Lekdadri:2023:HMC

[LL23] Abdelmajid Lekdadri and Hassan Lassri. A hybrid model to calculate the spin wave excitations in ferromagnetic/nonmagnetic multilayers. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000280.

Li:2023:MPC

[LLC^{+}23] Xiangyue Li, Xiaojing Liu, Xiang Chai, Hui He, Bin Zhang, and Tengfei Zhang. Multi-physics coupling simulation of small mobile nuclear reactor with finite element-based models. Computer Physics Communications, 293(??):??, December 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552300245X.

Li:2021:ERA

[LLH21a] Haiyan Li, Guiming Liang, and Yunbao Huang. An efficient radiation analysis approach through compressive model for laser driven inertial confinement fusion. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030312X.

Lyu:2021:BLS

[LLH^{+}21b] Xing-Long Lyu, Tiexiang Li, Tsung-Ming Huang, Wen-Wei Lin, and Heng Tian. The bi-Lebedev scheme for the Maxwell eigenvalue problem with 3D bi-anisotropic complex media. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303842.

Liang:2024:VLM

[LLL24] Jingang Liang, Ruihan Li, and Zhaoyuan Liu. Virtual lattice method for efficient Monte Carlo transport simulation of dispersion nuclear fuels. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523003302.

Li:2023:ETS

[LLQ^{+}23] Baoxia Li, Yang Liu, Haoyu Qi, Hanyu Tang, and Feng Huang. Effects of temperature on the structures and dynamics of a binary complex plasma. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002727.

Lu:2023:EVM

[LLR23] Jiayin Lu, Emanuel A. Lazar, and Chris H. Rycroft. An extension to Voro++ for multithreaded computation of Voronoi cells. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001777.

Liao:2021:ETC

[LLS^{+}21] Mingqing Liao, Yong Liu, Shun-Li Shang, Fei Zhou, Nan Qu, Yichuan Chen, Zhonghong Lai, Zi-Kui Liu, and Jingchuan Zhu. Elastic3rd: a tool for calculating third-order elastic constants from first-principles calculations. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303908.

Li:2023:EDP

[LLT^{+}23] Baoxia Li, Yang Liu, Xiaojiang Tang, Guannan Shi, Haoyu Qi, Xin Liu, Eric Robert, and Feng Huang. Effects of dust particle number on the structure and dynamics in a binary complex plasma system by Langevin dynamics simulation. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522003319.

Liu:2022:HOE

[LLY^{+}22] Haibo Liu, Tianran Liu, Peixun Yang, Siqi Wang, Xingjian Liu, Te Li, and Yongqing Wang. Higher-order elastic constitutive relation: Micro mechanism and application to acoustoelasticity. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552200100X.

Liao:2022:HES

[LLZ^{+}22] Mingqing Liao, Yong Liu, Fei Zhou, Tianyi Han, Danni Yang, Nan Qu, Zhonghong Lai, Zi-Kui Liu, and Jingchuan Zhu. A high-efficient strain-stress method for calculating higher-order elastic constants from first-principles. Computer Physics Communications, 280(??):??, November 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001977.

Li:2023:EHP

[LLZ^{+}23] Runhua Li, Jie Liu, Guangchun Zhang, Chunye Gong, Bo Yang, and Yuechao Liang. An efficient heterogeneous parallel algorithm of the 3D MOC for multizone heterogeneous systems. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523001510.

Lira:2020:SFS

[LM20] Pedro Alberto Resendiz Lira and Richard Marchand. Selfforce subtraction in particle in cell simulations. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300473.

Li:2021:EDS

[LM21a] Qi Li and Liquan Mei. Efficient, decoupled, and second-order unconditionally energy stable numerical schemes for the coupled Cahn–Hilliard system in copolymer/homopolymer mixtures. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300990.

Lukowski:2021:BAA

[LM21b] Tomasz Lukowski and Robert Moerman. Boundaries of the amplituhedron with amplituhedron boundaries. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303179.

Liu:2022:IMN

[LM22] Guangdong Liu and Richard Marchand. Inference of m-NLP data using radial basis function regression with center-evolving algorithm. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002168.

Liu:2023:AMP

[LM23] Xiao Liu and Yan-Qing Ma. AMFlow: a Mathematica package for Feynman integrals computation via auxiliary mass flow. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002843.

Li:2020:TOS

[LMHL20] Dongyue Li, Daniele Marchisio, Christian Hasse, and Dirk Lucas. twoWayGPBEFoam: an open-source Eulerian QBMM solver for monokinetic bubbly flows. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303728.

Lopez-Menchon:2023:GPR

[LMHUR23] Hector Lopez-Menchon, Alexander Heldring, Eduard Ubeda, and Juan M. Rius. A GPU parallel randomized CUR compression method for the Method of Moments. Computer Physics Communications, 287(??):??, June 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000413.

Lopez-Miralles:2023:AJF

[LMMP23] Jose López-Miralles, Jose María Martí, and Manel Perucho. On the application of Jacobian-free Riemann solvers for relativistic radiation magnetohydrodynamics under M1 closure. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522003496.

Liu:2023:HAT

[LMQ^{+}23] Wen-Bin Liu, Dong-Jun Ma, Jian-Zhen Qian, Ming-Yu Zhang, An-Min He, Nan-Sheng Liu, and Pei Wang. Highaccuracy three-dimensional surface detection in smoothed particle hydrodynamics for free-surface flows. Computer Physics Communications, 290(??):??, September 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001340.

Lashomb:2024:HDP

[LMWW24a] Paul Lashomb, Ronald B. Morgan, Travis Whyte, and Walter Wilcox. High-degree polynomial noise subtraction for disconnected loops. Computer Physics Communications, 298(??): ??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000432.

Lashomb:2024:MMC

[LMWW24b] Paul Lashomb, Ronald B. Morgan, Travis Whyte, and Walter Wilcox. Multipolynomial Monte Carlo for trace estimation in lattice QCD. Computer Physics Communications, 300(??): ??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000869.

Li:2021:NFS

[LMX^{+}21] Fei Li, Kyle G. Miller, Xinlu Xu, Frank S. Tsung, Viktor K. Decyk, Weiming An, Ricardo A. Fonseca, and Warren B. Mori. A new field solver for modeling of relativistic particle-laser interactions using the particle-in-cell algorithm. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302782.

Li:2023:MPM

[LN23] Yunguo Li and Huaiwei Ni. MD2D: a Python module for accurate determination of diffusion coefficient from molecular dynamics. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003186.

Lutsyshyn:2023:SPM

[LNB23] Y. Lutsyshyn, F. Navarrete, and D. Bauer. The N-shaped partition method: a novel parallel implementation of the Crank Nicolson algorithm. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000589.

Lange:2024:TLI

[LNP^{+}24] Torben Lange, Saswati Nandan, Joosep Pata, Laurits Tani, and Christian Veelken. Tau lepton identification and reconstruction: a new frontier for jet-tagging ML algorithms. Computer Physics Communications, 298(??):??, May 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000183.

Lanti:2020:OGE

[LOT^{+}20] E. Lanti, N. Ohana, N. Tronko, T. Hayward-Schneider, A. Bottino, B. F. McMillan, A. Mishchenko, A. Scheinberg, A. Biancalani, P. Angelino, S. Brunner, J. Dominski, P. Donnel, C. Gheller, R. Hatzky, A. Jocksch, S. Jolliet, Z. X. Lu, J. P. Martin Collar, I. Novikau, E. Sonnendrücker, T. Vernay, and L. Villard. Orb 5: a global electromagnetic gyrokinetic code using the PIC approach in toroidal geometry. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519303911.

Lombardi:2024:QFE

[LP24] Guido Lombardi and Davide Papapicco. Quadrature of functions with endpoint singular and generalised polynomial behaviour in computational physics. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552400047X.

Lot:2020:PPA

[LPSK20] Ruggero Lot, Franco Pellegrini, Yusuf Shaidu, and Emine Küçükbenli. PANNA: Properties from artificial neural network architectures. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520301843.

Liu:2021:PED

[LQ21] Xiao-Yu Liu and Chong Qi. PairDiag: an exact diagonalization program for solving general pairing Hamiltonians. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520301442.

Liu:2021:PGE

[LQGL21] Xiao-Yu Liu, Chong Qi, Xin Guan, and Zhong Liu. PairDiagSph: Generalization of the exact pairing diagonalization program for spherical systems. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000424.

Lodeiro:2022:DPP

[LR22] Lucas Lodeiro and Tomás Rauch. DensityTool: a postprocessing tool for space- and spin-resolved density of states from VASP. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001035.

Li:2021:INM

[LS21a] Haolin Li and Anbang Sun. Issues in the numerical modelling of positive ion extraction. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520303027.

Litim:2021:AAR

[LS21b] Daniel F. Litim and Tom Steudtner. ARGES — Advanced Renormalisation Group Equation Simplifier. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001338.

Lindvall:2022:CCG

[LS22] Kristoffer Lindvall and Jan Scheffel. 2D continuous Chebyshev– Galerkin time-spectral method. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003295.

Lee:2023:CSS

[LS23] In-Ho Lee and Seoleun Shin. Crystal structure search with principal invariants. Computer Physics Communications, 292 (??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002345.

Liu:2023:NDN

[LSF23] Weishuo Liu, Ziming Song, and Jian Fang. NNPred: Deploying neural networks in computational fluid dynamics codes to facilitate data-driven modeling studies. Computer Physics Communications, 290(??):??, September 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001200.

Lenton:2020:OTS

[LSNRD20] Isaac C. D. Lenton, Alexander B. Stilgoe, Timo A. Nieminen, and Halina Rubinsztein-Dunlop. OTSLM toolbox for structured light methods. Computer Physics Communications, 253(??): ??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300412.

Lynch:2024:ESM

[LSS24] Joel E. Lynch, Travis R. Sippel, and Shankar Subramaniam. An efficient solution of the multi-term multi-harmonic electron Boltzmann equation for use in global models. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000791.

Liu:2020:DUG

[LSW^{+}20] Hongtao Liu, Feng Shi, Jie Wan, Xiaoming He, and Yong Cao. Discrete unified gas kinetic scheme for a reformulated BGK–Vlasov–Poisson system in all electrostatic plasma regimes. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552030182X.

Lozano:2023:ERD

[LSZ23] Víctor Martín Lozano, Rosa María Sandá Seoane, and Jose Zurita. Z'-explorer 2.0: Reconnoitering the dark matter landscape. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000747.

Li:2021:MEC

[LTMK21] Botao Li, Synge Todo, A. C. Maggs, and Werner Krauth. Multithreaded event-chain Monte Carlo with local times. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303453.

Lang:2024:EPN

[LTT^{+}24] Logan Lang, Pedram Tavadze, Andres Tellez, Eric Bousquet, He Xu, Francisco Muñoz, Nicolas Vasquez, Uthpala Herath, and Aldo H. Romero. Expanding PyProcar for new features, maintainability, and reliability. Computer Physics Communications, 297(??):??, April 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523004083.

Lobanov:2021:LMS

[LU21] I. S. Lobanov and V. M. Uzdin. The lifetime of micron scale topological chiral magnetic states with atomic resolution. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002484.

Li:2022:ETT

[LVB22] Yunguo Li, Lidunka Vocadlo, and John P. Brodholt. ElasT: a toolkit for thermoelastic calculations. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521003921.

Lopez-Villellas:2023:AEC

[LVMGF^{+}23] Lorién López-Villellas, Carl Christian Kjelgaard Mikkelsen, Juan José Galano-Frutos, Santiago Marco-Sola, Jesús Alastruey-Benedé, Pablo Ibáñez, Miquel Moretó, Javier Sancho, and Pablo García-Risueño. Accurate and efficient constrained molecular dynamics of polymers using Newton’s method and special purpose code. Computer Physics Communications, 288 (??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000875.

Liang:2024:GEP

[LW24] Dong Liang and Cong Wang. Global energy-preserving local mesh-refined S-FDTD schemes for two dimensional Maxwell’s equations in Drude metamaterials. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002813.

Lu:2021:PDP

[LWC^{+}21] Denghui Lu, Han Wang, Mohan Chen, Lin Lin, Roberto Car, Weinan E, Weile Jia, and Linfeng Zhang. 86 PFLOPS Deep Potential Molecular Dynamics simulation of 100 million atoms with ab initio accuracy. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552030299X.

Ladshaw:2020:AAS

[LWhK^{+}20] Austin Ladshaw, Alexander I. Wiechert, Yong ha Kim, Costas Tsouris, and Sotira Yiacoumi. Algorithms and algebraic solutions of decay chain differential equations for stable and unstable nuclide fractionation. Computer Physics Communications, 246(??):??, January 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302887.

Li:2023:RCL

[LWS^{+}23] Yan Ting Li, Kai Wang, Ran Si, Michel Godefroid, Gediminas Gaigalas, Chong Yang Chen, and Per Jönsson. Reducing the computational load — atomic multiconfiguration calculations based on configuration state function generators. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002818.

Lym:2020:PMT

[LWV20] Jonathan Lym, Gerhard R. Wittreich, and Dionisios G. Vlachos. A Python Multiscale Thermochemistry Toolbox (pMuTT) for thermochemical and kinetic parameter estimation. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519302516.

Li:2021:SEV

[LXY^{+}21] Yibao Li, Qing Xia, Sungha Yoon, Chaeyoung Lee, Bingheng Lu, and Junseok Kim. Simple and efficient volume merging method for triply periodic minimal structures. Computer Physics Communications, 264(??):??, July 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100076X.

Lorin:2022:TDD

[LY22] Emmanuel Lorin and Xu Yang. Time-dependent Dirac equation with physics-informed neural networks: Computation and properties. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552200193X.

Li:2024:LVC

[LY24a] Jianqing Li and Junxiang Yang. Local volume-conservation-improved diffuse interface model for simulation of Rayleigh– Plateau fluid instability. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003958.

Lorin:2024:QOD

[LY24b] Emmanuel Lorin and Xu Yang. Quasi-optimal domain decomposition method for neural network-based computation of the time-dependent Schrödinger equation. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000523.

Liou:2020:SIP

[LYC20] Kai-Hsin Liou, Chao Yang, and James R. Chelikowsky. Scalable implementation of polynomial filtering for density functional theory calculation in PARSEC. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301429.

Liang:2022:HEP

[LYX22] Jiuyang Liang, Jiaxing Yuan, and Zhenli Xu. HSMA: an O(N) electrostatics package implemented in LAMMPS. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000509.

Liu:2024:MMP

[LYZL24] Jiahao Liu, Xiufeng Yang, Zhilang Zhang, and Moubin Liu. A massive MPI parallel framework of smoothed particle hydrodynamics with optimized memory management for extreme mechanics problems. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523003156.

Li:2021:SGS

[LZ21] Haolong Li and Ping Zhu. Solving the Grad-Shafranov equation using spectral elements for tokamak equilibrium with toroidal rotation. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520300849.

Li:2023:TTB

[LZK^{+}23] Yunhai Li, Zhen Zhan, Xueheng Kuang, Yonggang Li, and Shengjun Yuan. TBPLaS: a tight-binding package for largescale simulation. Computer Physics Communications, 285(??): ??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003514.

Lin:2024:QPC

[LZP^{+}24] Hai Lin, Yan Zhang, Soroosh Pezeshki, Adam W. Duster, Bo Wang, Xin-Ping Wu, Shi-Wen Zheng, Laura Gagliardi, and Donald G. Truhlar. QMMM 2023: a program for combined quantum mechanical and molecular mechanical modeling and simulations. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003326.

Liu:2023:MPC

[LZYY23] Gui-Bin Liu, Zeying Zhang, Zhi-Ming Yu, and Yugui Yao. MSGCorep: a package for corepresentations of magnetic space groups. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552300067X.

Malusa:2024:WBA

[MA24] Sandro Malusà and Alessandro Alaia. A well-balanced all-Mach scheme for compressible two-phase flow. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000547.

Michel:2020:TSM

[MAJ20] N. Michel, H. M. Aktulga, and Y. Jaganathen. Toward scalable many-body calculations for nuclear open quantum systems using the Gamow Shell Model. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303200.

Mijin:2021:SKF

[MAMK21] S. Mijin, A. Antony, F. Militello, and R. J. Kingham. SOL– KiT — fully implicit code for kinetic simulation of parallel electron transport in the tokamak Scrape-Off Layer. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302873.

Mangan:2024:FES

[Man24]  James Mangan.  FiniteFieldSolve: Exactly solving

large linear systems in high-energy theory. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524000948.

Martinez:2022:CJE

[Mar22] Leandro Martínez. CellListMap.jl: Efficient and customizable cell list implementation for calculation of pairwise particle properties within a cutoff. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001710.

Maurin:2020:USA

[Mau20] David Maurin. usine: Semi-analytical models for galactic cosmic-ray propagation. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519302930.

May:2021:MLG

[May21] Simon May. minimal-lagrangians: Generating and studying dark matter model Lagrangians with just the particle content. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303878.

Muruganandam:2021:OSR

[MBA21] Paulsamy Muruganandam, Antun Balaz, and Sadhan K. Adhikari. OpenMP solver for rotating spin-1 spin-orbit-and Rabi-coupled Bose–Einstein condensates. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000618.

Mazur:2024:SSM

[MBC^{+}24] Lukas Mazur, Dennis Bollweg, David A. Clarke, Luis Altenkort, Olaf Kaczmarek, Rasmus Larsen, Hai-Tao Shu, Jishnu Goswami, Philipp Scior, Hauke Sandmeyer, Marius Neumann, Henrik Dick, Sajid Ali, Jangho Kim, Christian Schmidt, Peter Petreczky, and Swagato Mukherjee. SIMULATeQCD: a simple multi-GPU lattice code for QCD calculations. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000870.

Mioduszewski:2023:CBM

[MBCC23] Lukasz Mioduszewski, Jakub Bednarz, Mateusz Chwastyk, and Marek Cieplak. Contact-based molecular dynamics of structured and disordered proteins in a coarse-grained model: Fixed contacts, switchable contacts and those described by pseudo-improper-dihedral angles. Computer Physics Communications, 284(??):??, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003307.

Muraviev:2021:SPR

[MBE^{+}21] A. Muraviev, A. Bashinov, E. Efimenko, V. Volokitin, I. Meyerov, and A. Gonoskov. Strategies for particle resampling in PIC simulations. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000035.

Masin:2020:USM

[MBG^{+}20] Zdenek Masín, Jakub Benda, Jimena D. Gorfinkiel, Alex G. Harvey, and Jonathan Tennyson. UKRmol+: a suite for modelling electronic processes in molecules interacting with electrons, positrons and photons using the R-matrix method. Computer Physics Communications, 249(??):??, April 2020.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303972.

McNelis:2021:AFD

[MBH21] Mike McNelis, Dennis Bazow, and Ulrich Heinz. Anisotropic fluid dynamical simulations of heavy-ion collisions. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001892.

Mandal:2024:MPC

[MBJB24] Bikramaditya Mandal, Dulat Bostan, Carolin Joy, and Dmitri Babikov. MQCT 2024: a program for calculations of inelastic scattering of two molecules (new version announcement). Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002837.

Marcolongo:2021:QOS

[MBTB21] Aris Marcolongo, Riccardo Bertossa, Davide Tisi, and Stefano Baroni. QEHeat: an open-source energy flux calculator for the computation of heat-transport coefficients from first principles. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002022.

Moxey:2020:NEC

[MCB^{+}20] David Moxey, Chris D. Cantwell, Yan Bao, Andrea Cassinelli, Giacomo Castiglioni, Sehun Chun, Emilia Juda, Ehsan Kazemi, Kilian Lackhove, Julian Marcon, Gianmarco Mengaldo, Douglas Serson, Michael Turner, Hui Xu, Joaquim Peiró, Robert M. Kirby, and Spencer J. Sherwin. Nektar++: Enhancing the capability and application of high-fidelity spectral/hp element methods. Computer Physics Communications, 249(??):??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519304175.

Munoz:2024:FFE

[MCMS24] Sergio Sevillano Muñoz, Edmund J. Copeland, Peter Millington, and Michael Spannowsky. FeynMG: a FeynRules extension for scalar-tensor theories of gravity. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003806.

Movilla:2023:GMI

[MCP23] Jose L. Movilla, Juan I. Climente, and Josep Planelles. Generalized method of image dyons for quasi-two dimensional slabs with ordinary — topological insulator interfaces. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001716.

Ma:2020:CAE

[MD20a] Pui-Wai Ma and S. L. Dudarev. CALANIE: Anisotropic elastic correction to the total energy, to mitigate the effect of periodic boundary conditions. Computer Physics Communications, 252 (??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519304370.

Macias-Diaz:2020:EIP

[MD20b] J. E. Macías-Díaz. An easy-to-implement parallel algorithm to simulate complex instabilities in three-dimensional (fractional) hyperbolic systems. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301661.

Matyka:2021:MEL

[MD21] Maciej Matyka and Michał Dzikowski. Memory-efficient lattice Boltzmann method for low Reynolds number flows. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001569.

Melchert:2022:PFP

[MD22] O. Melchert and A. Demircan. py-fmas: a Python package for ultrashort optical pulse propagation in terms of forward models for the analytic signal. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521003696.

Municchi:2021:HMR

[MDDI21] Federico Municchi, Nicodemo Di Pasquale, Marco Dentz, and Matteo Icardi. Heterogeneous multi-rate mass transfer models in OpenFOAM(R). Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303817.

Moreno:2022:PFI

[MDP22] Felipe Moreno, Sergio Davis, and Joaquín Peralta. A portable and flexible implementation of the Wang–Landau algorithm in order to determine the density of states. Computer Physics Communications, 274(??):??, May 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000017.

Motamarri:2020:DFM

[MDR^{+}20] Phani Motamarri, Sambit Das, Shiva Rudraraju, Krishnendu Ghosh, Denis Davydov, and Vikram Gavini. DFT-FE — a massively parallel adaptive finite-element code for large-scale density functional theory calculations. Computer Physics Communications, 246(??):??, January 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302309.

Murzov:2024:AMW

[MDZ24] S. A. Murzov, S. A. Dyachkov, and V. V. Zhakhovsky. Adaptive moving window technique for SPH simulation of stationary shock waves. Computer Physics Communications, 298(??): ??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000390.

Mankodi:2024:NSN

[MEC^{+}24] Tapan K. Mankodi, Omid Ejtehadi, Tushar Chourushi, Amin Rahimi, and R. S. Myong. nccrFOAM suite: Nonlinear coupled constitutive relation solver in the OpenFOAM framework for rarefied and microscale gas flows with vibrational nonequilibrium. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003697.

Martinez-Estevez:2023:CSB

[MEDT^{+}23] I. Martínez-Estévez, J. M. Domínguez, B. Tagliafierro, R. B. Canelas, O. García-Feal, A. J. C. Crespo, and M. Gómez-Gesteira. Coupling of an SPH-based solver with a multiphysics library. Computer Physics Communications, 283(??): ??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003009.

McNelis:2021:PFD

[MEH21] Mike McNelis, Derek Everett, and Ulrich Heinz. Particlization in fluid dynamical simulations of heavy-ion collisions: the iS3D module. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302836.

Maffioli:2024:TTS

[MES^{+}24] Luca Maffioli, James P. Ewen, Edward R. Smith, Sleeba Varghese, Peter J. Daivis, Daniele Dini, and B. D. Todd. TTCF4LAMMPS: a toolkit for simulation of the non-equilibrium behaviour of molecular fluids at experimentally accessible shear rates. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001280.

Martinez:2023:BOS

[MFB23] Emanuel A. Martínez, Juan I. Beltrán Fínez, and Flavio Y. Bruno. BinPo: an open-source code to compute the band structure of two-dimensional electron systems. Computer Physics Communications, 284(??):??, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003149.

Mann:2022:PEE

[MFS^{+}22] Sean Mann, Eric Fadel, Samuel S. Schoenholz, Ekin D. Cubuk, Steven G. Johnson, and Giuseppe Romano. ∂PV: an end-to-end differentiable solar-cell simulator. Computer Physics Communications, 272(??):??, March 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003441.

Ma:2022:SSC

[MG22] J. Ma and W. Guo. SCELT (symbolic computation aided eigenvalue and linear code for tokamaks): a full MHD eigenvalue code in toroidal geometry developed with the use of a symbolic computation technique. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552200131X.

Medina:2023:VBR

[MGC^{+}23] Brandon M. Medina, Pierre Grua, Keith L. Cartwright, David Hébert, Nicolas Szalek, Clément Caizergues, Israel Owens, Elaine L. Rhoades, Jacques Gardelle, and Christopher H. Moore. Verification and benchmarking relativistic electron beam transport through a background gas. Computer Physics Communications, 288(??):??, July 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000668.

Martini:2020:PSQ

[MGG^{+}20] A. Martini, S. A. Guda, A. A. Guda, G. Smolentsev, A. Algasov, O. Usoltsev, M. A. Soldatov, A. Bugaev, Yu. Rusalev, C. Lamberti, and A. V. Soldatov. PyFitit: the software for quantitative analysis of XANES spectra using machinelearning algorithms. Computer Physics Communications, 250 (??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046551930387X.

Ma:2024:DTM

[MGX24] J. Ma, W. Guo, and Y. Xie. Development of a tokamak magnetohydrodynamic code with the discontinuous Galerkin and weighted essentially non-oscillatory methods. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000626.

Madshaven:2021:CSS

[MHÅ21] I. Madshaven, O. L. Hestad, and P.-O. Åstrand. Cerman: Software for simulating streamer propagation in dielectric liquids based on the Townsend–Meek criterion. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000990.

Mihaylov:2024:DMG

[MHK24] Deyan I. Mihaylov, S. X. Hu, and Valentin V. Karasiev. Dragon: a multi-GPU orbital-free density functional theory molecular dynamics simulation package for modeling of warm dense matter. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552300276X.

Mignone:2023:GCI

[MHP23] A. Mignone, H. Haudemand, and E. Puzzoni. A guiding center implementation for relativistic particle dynamics in the PLUTO code. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003447.

MacFadden:2022:EMP

[MK22] Nathaniel J. L. MacFadden and Ara N. Knaian. Efficient modeling of particle transport through aerosols in GEANT4. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001023.

Molnar:2020:ACT

[MKHT20] Ferenc Molnár, Shubha R. Kharel, Xiaobo Sharon Hu, and Zoltán Toroczkai. Accelerating a continuous-time analog SAT solver using GPUs. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302204.

Mackoy:2021:VVM

[MKPW21] Travis Mackoy, Bharat Kale, Michael E. Papka, and Ralph A. Wheeler. viewSq, a Visual Molecular Dynamics (VMD) module for calculating, analyzing, and visualizing X-ray and neutron structure factors from atomistic simulations. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0010465521000345. See corrigendum [MKPW22].

Mackoy:2022:CVV

[MKPW22] Travis Mackoy, Bharat Kale, Michael E. Papka, and Ralph A. Wheeler. Corrigendum to “viewSq, a Visual Molecular Dynamics (VMD) module for calculating, analyzing, and visualizing X-ray and neutron structure factors from atomistic simulations” [Comput. Phys. Commun. 264 (2021) 107881]. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000777. See [MKPW21].

Mercenne:2022:ESA

[MLD^{+}22] A. Mercenne, K. D. Launey, T. Dytrych, J. E. Escher, S. Quaglioni, G. H. Sargsyan, D. Langr, and J. P. Draayer. Efficacy of the symmetry-adapted basis for ab initio nucleon– nucleus interactions for light- and intermediate-mass nuclei. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001953.

Miller:2021:DLB

[MLT^{+}21] Kyle G. Miller, Roman P. Lee, Adam Tableman, Anton Helm, Ricardo A. Fonseca, Viktor K. Decyk, and Warren B. Mori. Dynamic load balancing with enhanced sharedmemory parallelism for particle-in-cell codes. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303040.

Mao:2023:DDL

[MLZ^{+}23] Runze Mao, Minqi Lin, Yan Zhang, Tianhan Zhang, Zhi-Qin John Xu, and Zhi X. Chen. DeepFlame: a deep learning empowered open-source platform for reacting flow simulations. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552300187X.

Mostajeran:2022:DDN

[MM22] F. Mostajeran and R. Mokhtari. DeepBHCP: Deep neural network algorithm for solving backward heat conduction problems. Computer Physics Communications, 272(??):??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003489.

Maloney:2023:PMF

[MM23] Samuel A. Maloney and Ben F. McMillan. A partially mesh-free scheme for representing anisotropic spatial variations along field lines: Conservation, quadrature, and the delta property. Computer Physics Communications, 284(??): ??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003484.

Mandia:2021:AIM

[MMC^{+}21] Anup Kumar Mandia, Bhaskaran Muralidharan, Jung-Hae Choi, Seung-Cheol Lee, and Satadeep Bhattacharjee. AMMCR: Ab initio model for mobility and conductivity calculation by using Rode Algorithm. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303428.

Manteigas:2022:EBE

[MMC^{+}22] V. Manteigas, L. Martins, J. Cruz, M. Fonseca, and A. P. Jesus. ERYA-Bulk and ERYA-Profiling: an application for quantitative PIGE analysis. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552200025X.

Martin:2022:HUF

[MMCC^{+}22] P. Martin, C. E. Madrid-Cortes, C. Cáceres, N. Araya, C. Aguilar, and J. M. Cabrera. HEAPS: a user-friendly tool for the design and exploration of high-entropy alloys based on semi-empirical parameters. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522001175.

Macias-Medri:2021:AVS

[MMFdL21] A. E. Macias-Medri, Carlos E. Fiore, and M. G. E. da Luz. Analyzing and validating simulated tempering implementations at phase transition regimes. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300801.

Marepalli:2020:ADA

[MMM20] Prabhakar Marepalli, Sanjay R. Mathur, and Jayathi Y. Murthy. Automatic differentiation approach for property computations in nanoscale thermal transport. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300011.

Mabillard:2023:UMT

[MMM23] Joël Mabillard, Isha Malhotra, and Bortolo Matteo Mognetti. Using Markov transition matrices to generate trial configurations in Markov chain Monte Carlo simulations. Computer Physics Communications, 285(??):??, April 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003605.

Magarotto:2020:VEC

[MMP20] M. Magarotto, D. Melazzi, and D. Pavarin. 3D-VIRTUS: Equilibrium condition solver of radio-frequency magnetized plasma discharges for space applications. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551930308X.

Marini:2024:EOS

[MMP^{+}24] Giovanni Marini, Guglielmo Marchese, Gianni Profeta, Jelena Sjakste, Francesco Macheda, Nathalie Vast, Francesco Mauri, and Matteo Calandra. EPIq: an open-source software for the calculation of electron-phonon interaction related properties. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002953.

Miller:2024:RGB

[MMV^{+}24] Benjamin N. Miller, David H. Meyer, Teemu Virtanen, Christopher M. O’Brien, and Kevin C. Cox. RydIQule: a graph-based paradigm for modeling Rydberg and atomic sensors. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002977.

Morishita:2022:ADA

[MMYU22] Y. Morishita, S. Murakami, M. Yokoyama, and G. Ueno. ASTI: Data assimilation system for particle and heat transport in toroidal plasmas. Computer Physics Communications, 274 (??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000054.

Mohankumar:2021:IOR

[MN21] N. Mohankumar and A. Natarajan. InvFD, an OCTAVE routine for the numerical inversion of the Fermi–Dirac integral. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001740.

Mykhailenko:2024:SMC

[MNS^{+}24] Viacheslav Mykhailenko, Martin Nguyen, Michal Solanik, Ján Genči, Yuriy Kolesnyk, and Pavol Bobik. SDE method for cosmic rays modulation in the heliosphere statistical error and solution uniqueness. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003715.

Mirramezani:2024:RAC

[MOA24] Mehran Mirramezani, Deniz Oktay, and Ryan P. Adams. A rapid and automated computational approach to the design of multistable soft actuators. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000134.

Morgan:2024:RZD

[MOMO24] Brandon E. Morgan, Maxwell Osawe, Marty Marinak, and Britton J. Olson. RANSBox: a zero-dimensional modular software package for Reynolds-averaged Navier–Stokes modeling. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524000201.

Matsugatani:2021:QOS

[MONW21] Akishi Matsugatani, Seishiro Ono, Yusuke Nomura, and Haruki Watanabe. qeirreps: an open-source program for quantum ESPRESSO to compute irreducible representations of Bloch wavefunctions. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000722.

Moreno:2021:EUG

[MOV21] José A. Moreno, Eduardo Oliva, and Pedro Velarde. EMcLAW: an unsplit Godunov method for Maxwell’s equations including polarization, metals, divergence control and AMR. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300862.

Motoyama:2022:TTN

[MOY^{+}22] Yuichi Motoyama, Tsuyoshi Okubo, Kazuyoshi Yoshimi, Satoshi Morita, Takeo Kato, and Naoki Kawashima. TeNeS: Tensor network solver for quantum lattice systems. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001564.

Mendez:2021:MPS

[MP21] Juan Pedro Mendez and Mauricio Ponga. MXE: a package for simulating long-term diffusive mass transport phenomena in nanoscale systems. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301211.

Mijin:2024:RFB

[MPH^{+}24] Stefan Mijin, Dominic Power, Ryan Holden, William Hornsby, David Moulton, and Fulvio Militello. ReMKiT1D — a framework for building reactive multi-fluid models of the tokamak scrape-off layer with coupled electron kinetics in 1D. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001188.

Mortazavi:2021:AFP

[MPN^{+}21] Bohayra Mortazavi, Evgeny V. Podryabinkin, Ivan S. Novikov, Timon Rabczuk, Xiaoying Zhuang, and Alexander V. Shapeev.

Accelerating first-principles estimation of thermal conductivity by machine-learning interatomic potentials: a MTP/ ShengBTE solution. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302800.

Muir:2022:NIP

[MPQ^{+}22] Dean Muir, David Pfefferlé, Zhisong Qu, Matthew Hole, and Markus Hegland. Numerical integration of particle orbits in discontinuous fields using VENUS-LEVIS and SPEC. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003039.

Massaro:2024:CFE

[MPS^{+}24] D. Massaro, A. Peplinski, R. Stanly, S. Mirzareza, V. Lupi, T. Mukha, and P. Schlatter. A comprehensive framework to enhance numerical simulations in the spectral-element code Nek5000. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001723.

Moustafa:2021:PVP

[MPSK21] Sabry G. Moustafa, Apoorva Purohit, Andrew J. Schultz, and David A. Kofke. pyHMA: a VASP post-processor for precise measurement of crystalline anharmonic properties using harmonically mapped averaging. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302654.

Meana-Paneda:2024:PCP

[MPZB^{+}24] Rubén Meana-Pañeda, Jingjing Zheng, Junwei Lucas Bao, Shuxia Zhang, Benjamin J. Lynch, José C. Corchado, Yao-Yuan Chuang, Patton L. Fast, Wei-Ping Hu, Yi-Ping Liu, Gillian C. Lynch, Kiet A. Nguyen, Charles F. Jackels, Antonio Fernández-Ramos, Benjamin A. Ellingson, Vasilios S. Melissas, Jordi Villà, Ivan Rossi, Elena L. Coitiño, Jingzhi Pu, Titus V. Albu, Rui Ming Zhang, Xuefei Xu, Artur Ratkiewicz, Rozeanne Steckler, Bruce C. Garrett, Alan D. Isaacson, and Donald G. Truhlar. Polyrate 2023: a computer program for the calculation of chemical reaction rates for polyatomics. New version announcement. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002783.

Magalhaes:2022:PPI

[MR22] Tiago E. C. Magalhães and José M. Rebordão. PyWolf: a Py-OpenCL implementation for simulating the propagation of partially coherent light. Computer Physics Communications, 276 (??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000546.

Muratov:2023:DDD

[MRD23] R. V. Muratov, P. N. Ryabov, and S. A. Dyachkov. Dynamic domain decomposition method based on weighted Voronoi diagrams. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001352.

Mukut:2022:MAF

[MRG22] Khaled Mosharraf Mukut, Somesh Roy, and Eirini Goudeli. Molecular arrangement and fringe identification and analysis from molecular dynamics (MAFIA-MD): a tool for analyzing the molecular structures formed during reactive molecular dynamics simulation of hydrocarbons. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000431.

Meredith:2023:HHA

[MRH^{+}23] L. T. Meredith, M. Rezazadeh, M. F. Huq, J. Drobny, V. V. Srinivasaragavan, O. Sahni, and D. Curreli. hPIC2: a hardware-accelerated, hybrid particle-in-cell code for dynamic plasma-material interactions. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002880.

Muta:2020:EOS

[MRN20] Abhinav Muta, Prabhu Ramachandran, and Pawan Negi. An efficient, open source, iterative ISPH scheme. Computer Physics Communications, 255(??):??, October 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300953.

Meng:2022:PPL

[MRT^{+}22] Fanwang Meng, Michael Richer, Alireza Tehrani, Jonathan La, Taewon David Kim, Paul W. Ayers, and Farnaz Heidar-Zadeh. Procrustes: a Python library to find transformations that maximize the similarity between matrices. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000522.

Magalhaes:2024:NVP

[MS24] Tiago E. C. Magalhães and Daniela O. Santos. A new version of PyWolf for the propagation of partially coherent light in media other than free space. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002448.

Mi:2021:ESD

[MSG^{+}21] Wenhui Mi, Xuecheng Shao, Alessandro Genova, Davide Ceresoli, and Michele Pavanello. eQE 2.0: Subsystem DFT beyond GGA functionals. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521002344.

Markel:2022:FMC

[MSH22] Vadim A. Markel, Markus Schöbinger, and Karl Hollaus. A fast method to compute dispersion diagrams of three-dimensional photonic crystals with rectangular geometry. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001606.

Mostofi:2002:TEC

[MSHP02] Arash A. Mostofi, Chris-Kriton Skylaris, Peter D. Haynes, and Mike C. Payne. Total-energy calculations on a real space grid with localized functions and a plane-wave basis. Computer Physics Communications, 147(3):788–802, September 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004617. See corri-

gendum [MSHP20].

Mostofi:2020:CTE

[MSHP20] Arash A. Mostofi, Chris-Kriton Skylaris, Peter D. Haynes, and Mike C. Payne. Corrigendum to “Total-energy calculations on a real space grid with localised functions and a plane-wave basis” [Comput. Phys. Comm. 147/3 (2002) 788– 802]. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300680. See [MSHP02].

Morab:2024:FFV

[MSM24] Sumant R. Morab, Atul Sharma, and Janani S. Murallidharan. Fully finite volume method on a curvilinear grid-based arbitrary Lagrangian Eulerian approach for computational fluid flexible-structure interaction. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003995.

Marevic:2022:ADS

[MSN^{+}22] P. Marević, N. Schunck, E. M. Ney, R. Navarro Pérez, M. Verriere, and J. O’Neal. Axially-deformed solution of the Skyrme– Hartree–Fock–Bogoliubov equations using the transformed harmonic oscillator basis (IV) hfbtho (v4.0): a new version of the program. Computer Physics Communications, 276(??): ??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000868.

Michels:2021:GXF

[MSU^{+}21] Dominik Michels, Andreas Stegmeir, Philipp Ulbl, Denis Jarema, and Frank Jenko. GENE-X: a full-f gyrokinetic turbulence code based on the flux-coordinate independent approach. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000989.

Melnick:2021:AIS

[MSY^{+}21] Corey Melnick, Patrick Sémon, Kwangmin Yu, Nicholas D’Imperio, André-Marie Tremblay, and Gabriel Kotliar. Accelerated impurity solver for DMFT and its diagrammatic extensions. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001879.

Montefiori:2023:SHP

[MT23] Samuele Montefiori and Matteo Tamburini. SFQEDtoolkit: a high-performance library for the accurate modeling of strongfield QED processes in PIC and Monte Carlo codes. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300200X.

Motoyama:2022:BOP

[MTY^{+}22] Yuichi Motoyama, Ryo Tamura, Kazuyoshi Yoshimi, Kei Terayama, Tsuyoshi Ueno, and Koji Tsuda. Bayesian optimization package: PHYSBO. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001242.

Moratilla-Vega:2022:OSC

[MVAXP22] M. A. Moratilla-Vega, M. Angelino, H. Xia, and G. J. Page. An open-source coupled method for aeroacoustics modelling. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001394.

Milasinovic:2020:DOS

[MVF20] Danko Z. Milasinovic, Arso M. Vukicevic, and Nenad D. Filipovic. dfemtoolz: an open-source C++ framework for efficient imposition of material and boundary conditions in finite element biomedical simulations. Computer Physics Communications, 249(??):??, April 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303376.

Martinez:2024:MGI

[MVK^{+}24] Pablo Antonio Martínez, Theresa Vock, Liliane Racha Kharchi, Jesus Nain Pedroza-Montero, Xiaojing Wu, Karim Hasnaoui, and Aurélien de la Lande. A multi-GPU implementation of real-time time-dependent auxiliary density functional theory for the investigation of nanosystems irradiations. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002916.

Matwiejew:2021:QFP

[MW21] Edric Matwiejew and Jingbo Wang. QSW MPI: a framework for parallel simulation of quantum stochastic walks. Computer Physics Communications, 260(??):??, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303581.

Ma:2023:SOS

[MWJL23] Kuang Ma, Yaning Wang, Maoqiang Jiang, and Zhaohui Liu. A simple one-step index algorithm for implementation of lattice Boltzmann method on GPU. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003228.

Min:2024:RMR

[MXH^{+}24] Qi Min, Ziyang Xu, Siqi He, Haidong Lu, Xingbang Liu, Ruizi Shen, Yanhong Wu, Qikun Pan, Chongxiao Zhao, Fei Chen, Maogen Su, and Chenzhong Dong. RHDLPP: a multigroup radiation hydrodynamics code for laser-produced plasmas. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001656.

Mohanty:2023:ETM

[MYKC23] Shaswat Mohanty, SangHyuk Yoo, Keonwook Kang, and Wei Cai. Evaluating the transferability of machine-learned force fields for material property modeling. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000681.

Motoyama:2022:DAS

[MYM^{+}22] Yuichi Motoyama, Kazuyoshi Yoshimi, Izumi Mochizuki, Harumichi Iwamoto, Hayato Ichinose, and Takeo Hoshi. Dataanalysis software framework 2DMAT and its application to experimental measurements for two-dimensional material structures. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001849.

Motoyama:2021:DDS

[MYMK^{+}21] Yuichi Motoyama, Kazuyoshi Yoshimi, Akiko Masaki-Kato, Takeo Kato, and Naoki Kawashima. DSQSS: Discrete space quantum systems solver. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000692.

Michel:2022:TPT

[MZ22] Marco Michel and Sebastian Zell. TimeEvolver: a program for time evolution with improved error bound. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000935.

Morawiec:2020:PSN

[MZD^{+}20] Krzysztof Morawiec, Wiktoria Zajkowska, Piotr Dłuzewski, Makoto Shiojiri, and Jan Kusiński. PyHoLo software, a new tool for electron hologram analysis and magnetic investigation. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302216.

Midtbo:2021:NSI

[MZL^{+}21] Jørgen E. Midtbø, Fabio Zeiser, Erlend Lima, Ann-Cecilie Larsen, Gry M. Tveten, Magne Guttormsen, Frank Leonel Bello Garrote, Anders Kvellestad, and Therese Renstrøm. A new software implementation of the Oslo method with rigorous statistical uncertainty propagation. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552030401X.

Mystilidis:2023:OEN

[MZV23] Christos Mystilidis, Xuezhi Zheng, and Guy A. E. Vandenbosch. OpenSANS: a Semi-Analytical solver for Nonlocal plasmonic S. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003289.

Nagornov:2020:ANE

[NA20] Yuri S. Nagornov and Ryosuke Akashi. AtomREM: Nonempirical seeker of the minimum energy escape paths on many-dimensional potential landscapes without coarse graining. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300813.

Nieves:2021:MME

[NAZ^{+}21] P. Nieves, S. Arapan, S. H. Zhang, A. P. Kadzielawa, R. F. Zhang, and D. Legut. MAELAS: MAgneto-ELAStic properties calculation via computational high-throughput approach. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000801.

Nieves:2022:MNV

[NAZ^{+}22] P. Nieves, S. Arapan, S. H. Zhang, A. P. Kadzielawa, R. F. Zhang, and D. Legut. MAELAS 2.0: a new version of a computer program for the calculation of magneto-elastic properties. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552100309X.

Novikau:2021:IET

[NBB^{+}21] I. Novikau, A. Biancalani, A. Bottino, A. Di Siena, Ph. Lauber, E. Poli, E. Lanti, L. Villard, N. Ohana, and S. Briguglio. Implementation of energy transfer technique in ORB5 to study collisionless wave–particle interactions in phase-space. Computer Physics Communications, 262(??):??, May 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303704.

Neto:2024:PFO

[NBC^{+}24] Atilio Minotto Neto, Otávio Bianchi, Leonardo Bresciani Canto, Janete Eunice Zorzi, and Cláudio Antônio Perottoni. PolyWeight: a free and open-source program for determination of molecular weight distribution of linear polymers. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524000833.

Niess:2020:TLO

[NBCMH20] Valentin Niess, Anne Barnoud, Cristina Cârloganu, and Olivier Martineau-Huynh. TURTLE: a C library for an optimistic stepping through a topography. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303078.

Nevay:2020:BAT

[NBS^{+}20] L. J. Nevay, S. T. Boogert, J. Snuverink, A. Abramov, L. C. Deacon, H. Garcia-Morales, H. Lefebvre, S. M. Gibson, R. Kwee-Hinzmann, W. Shields, and S. D. Walker. BDSIM: an accelerator tracking code with particle-matter interactions. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300400.

Neveu:2023:CMO

[NCF^{+}23] Nicole Neveu, Tyler H. Chang, Paris Franz, Stephen Hudson, and Jeffrey Larson. Comparison of multiobjective optimization methods for the LCLS-II photoinjector. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002855.

Nicolas:2024:BFI

[NDFL24] T. Nicolas, V. Dubois, Q. Fang, and H. Lütjens. A binary filter inspired from the PIC sparse grid technique — illustration on the XTOR-K code. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524000742.

Nagy:2021:IBS

[NG21] Gabor Nagy and Helmut Grubmuller. Implementation of a Bayesian secondary structure estimation method for the SESCA circular dichroism analysis package. Computer Physics Communications, 266(??):??, September 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100134X.

Nazarov:2022:PMF

[NI22] Alexey E. Nazarov and Anatoly I. Ivanov. Principles of modeling the fluorescence spectral dynamics of dye molecules in solutions. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002903.

Nielsen:2020:GAS

[Nie20] C. F. Nielsen. GPU accelerated simulation of channeling radiation of relativistic particles. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519304369.

Nielsen:2022:GAM

[Nie22a] C. F. Nielsen. GPU accelerated Monte Carlo simulation of high-intensity pulsed laser-electron interaction. Computer Physics Communications, 278(??):??, September 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001448.

Niess:2022:PL

[Nie22b] Valentin Niess. The PUMAS library. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465522001576.

Niess:2023:AYA

[Nie23] Valentin Niess. Alouette: Yet another encapsulated TAUOLA, but revertible. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002272.

Nishio:2022:VSV

[Nis22] Kengo Nishio. Vorotis: Software for Voronoi tessellation analysis using the polyhedron code. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522001370.

Nishio:2023:CCC

[Nis23] Kengo Nishio. Cluster classification by chemi-topology. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000048.

Nguyen:2022:RFT

[NJSY22] Danh Nam Nguyen, Ki Sung Jung, Jae Won Shim, and Chun Sang Yoo. Real-fluid thermophysicalModels: an OpenFOAM-based library for reacting flow simulations at high pressure. Computer Physics Communications, 273(??): ??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003763.

Nakhaee:2020:TBS

[NKP20] Mohammad Nakhaee, S. Ahmad Ketabi, and Francois M. Peeters. Tight-Binding Studio: a technical software package to find the parameters of tight-binding Hamiltonian. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520301636.

Namvar:2023:LLB

[NL23a] Morteza Namvar and Sébastien Leclaire. LaBCof: Lattice Boltzmann boundary condition framework. Computer Physics Communications, 285(??):??, April 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003666.

Nguyen:2023:GCG

[NL23b] Tung Dong Cao Nguyen and Deokjung Lee. Group constants generation by Monte Carlo code MCS for LWR analysis. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522003617.

Norouzisadeh:2024:LMS

[NLS24] Mojtaba Norouzisadeh, Philippe Leroy, and Cyprien Soulaine. A lubrication model with slope-dependent disjoining pressure for modeling wettability alteration. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000377.

Naik:2022:TCS

[NNMJ22] Saismit Naik, Mit H. Naik, Indrajit Maity, and Manish Jain. Twister: Construction and structural relaxation of commensurate Moiré superlattices. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002964.

Nogueira:2021:FGG

[Nog21a] P. Nogueira. Feynman graph generation and propagator mixing, I. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002150.

Nogueira:2021:FRC

[Nog21b] P. Nogueira. From Feynman rules to conserved quantum numbers, III. Computer Physics Communications, 260(??): ??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303660.

Norouzi:2023:PPM

[Nor23] H. R. Norouzi. PhasicFlow: a parallel, multi-architecture open-source code for DEM simulations. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465523001662.

Negi:2021:AUP

[NR21] Pawan Negi and Prabhu Ramachandran. Algorithms for uniform particle initialization in domains with complex boundaries. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552100120X.

Nowakowski:2022:DSV

[NRG22] Piotr Nowakowski, Przemysław Rokita, and Lukasz Graczykowski. Distributed simulation and visualization of the ALICE detector magnetic field. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003180.

Nowakowski:2024:GPV

[NRG24] Piotr Nowakowski, Przemysław Rokita, and Lukasz Graczykowski. GPU propagation and visualisation of particle collisions with accurate model of ALICE detector magnetic field. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002576.

Nadtochy:2021:PEM

[NRK^{+}21] P. N. Nadtochy, E. G. Ryabov, A. V. Karpov, D. V. Vanin, and G. D. Adeev. Potential energy models of excited compound nucleus. Computer Physics Communications, 258(??): ??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302939.

Nadtochy:2022:TCM

[NRKA22] P. N. Nadtochy, E. G. Ryabov, A. V. Karpov, and G. D. Adeev. Transport coefficients for modeling fission dynamics. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000261.

Nandal:2020:NSG

[NS20] P. Nandal and R. P. Sharma. Numerical simulation on GPUs with CUDA to study nonlinear dynamics of whistler wave and its turbulent spectrum in radiation belts. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300497.

Naterop:2020:HRN

[NSU20] L. Naterop, A. Signer, and Y. Ulrich. handyG — rapid numerical evaluation of generalised polylogarithms in Fortran. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300230.

Nath:2023:UMB

[NSY^{+}23] Dhyanjyoti D. Nath, Vignesh V. Srinivasaragavan, Timothy R. Younkin, Gerrett Diamond, Cameron W. Smith, Alyssa Hayes, Mark S. Shephard, and Onkar Sahni. A 3D unstructured mesh based particle tracking code for impurity transport simulation in fusion tokamaks. Computer Physics Communications, 292(??):??, November 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002060.

Nazarov:2024:HLG

[NT24] N. A. Nazarov and V. V. Terekhov. High level GPUaccelerated 2D PIV framework in Python. Computer Physics Communications, 295(??):??, February 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003545.

Na:2024:UBR

[NTO24] Dong-Yeop Na, Fernando L. Teixeira, and Yuri A. Omelchenko. An unstructured body-of-revolution electromagnetic particlein-cell algorithm with radial perfectly matched layers and dual polarizations. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552400170X.

Nakamura:2022:CSU

[NUK^{+}22] H. Nakamura, K. Uchibori, S. Kawata, T. Karino, R. Sato, and A. I. Ogoyski. Code O-SUKI-N 3D: Upgraded directdrive fuel target 3D implosion code in heavy ion inertial fusion. Computer Physics Communications, 272(??):??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003350.

Novak:2020:EHF

[NVC20] Matyás Novák, Jirí Vackár, and Robert Cimrman. Evaluating Hellmann–Feynman forces within non-local pseudopotentials. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519303698.

Novak:2023:AAM

[NVCS23] Matyás Novák, Jirí Vackár, Robert Cimrman, and Ondrej Sipr. Adaptive Anderson mixing for electronic structure calculations. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002102.

Nakamura:2021:RIT

[NYN^{+}21] Kazuma Nakamura, Yoshihide Yoshimoto, Yusuke Nomura, Terumasa Tadano, Mitsuaki Kawamura, Taichi Kosugi, Kazuyoshi Yoshimi, Takahiro Misawa, and Yuichi Motoyama. RESPACK: an ab initio tool for derivation of effective low-energy model of material. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552030391X.

Ozay:2024:NOR

[ÖAÖ24] S. Özay, S. Akdemir, and E. Öztekin. New orthogonality relationships of the Gaunt coefficients. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524000419.

Ostanin:2024:RCM

[OAP^{+}24] Igor Ostanin, Vasileios Angelidakis, Timo Plath, Sahar Pourandi, Anthony Thornton, and Thomas Weinhart. Rigid clumps in the MercuryDPM particle dynamics code. Computer Physics Communications, 296(??):??, March 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300379X.

Oberhauser:2024:DFG

[OBC^{+}24] B. Banto Oberhauser, P. Bisio, A. Celentano, E. Depero, R. R. Dusaev, D. V. Kirpichnikov, M. M. Kirsanov, N. V. Krasnikov, A. Marini, L. Marsicano, L. Molina-Bueno, M. Mongillo, D. Shchukin, H. Sieber, and I. V. Voronchikhin. Development of the fully Geant4 compatible package for the simulation of Dark Matter in fixed target experiments. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552400122X.

Oliver:2024:EPT

[OBGA24] S. Oliver, S. Rodriguez Bosca, and V. Giménez-Alventosa. Enabling particle transport on CAD-based geometries for radiation simulations with penRed. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000146.

Ovtchinnikov:2020:SSI

[OBK^{+}20] Evgueni Ovtchinnikov, Richard Brown, Christoph Kolbitsch, Edoardo Pasca, Casper da Costa-Luis, Ashley G. Gillman, Benjamin A. Thomas, Nikos Efthimiou, Johannes Mayer, Palak Wadhwa, Matthias J. Ehrhardt, Sam Ellis, Jakob S. Jørgensen, Julian Matthews, Claudia Prieto, Andrew J. Reader, Charalampos Tsoumpas, Martin Turner, David Atkinson, and Kris Thielemans. SIRF: Synergistic Image Reconstruction Framework. Computer Physics Communications, 249(??):??, April 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303984.

Ostmeyer:2021:IMH

[OBL^{+}21] Johann Ostmeyer, Evan Berkowitz, Thomas Luu, Marcus Petschlies, and Ferenc Pittler. The Ising model with hybrid Monte Carlo. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000904.

Oliani:2023:IOL

[OCC23] Stefano Oliani, Nicola Casari, and Mauro Carnevale. ICSFoam: an OpenFOAM library for implicit coupled simulations of highspeed flows. Computer Physics Communications, 286(??): ??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000188.

Owkes:2023:BJF

[OCE^{+}23] Mark Owkes, Kai Coblentz, Austen Eriksson, Takumi Kammerzell, and Philip S. Stewart. Biofilm.jl: a fast solver for one-dimensional biofilm chemistry and ecology. Computer Physics Communications, 293(??):??, December 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002357.

Ochi:2023:CTF

[Och23a] Masayuki Ochi. Corrigendum to “TC++: First-principles calculation code for solids using the transcorrelated method” [Comput. Phys. Commun. 287 (2023) 108687]. Computer Physics Communications, 290(??):??, September 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001261.

Ochi:2023:TFP

[Och23b] Masayuki Ochi. TC++: First-principles calculation code for solids using the transcorrelated method. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000322.

Ocetkiewicz:2024:MGU

[OCK^{+}24] Krzysztof M. Ocetkiewicz, Cezary Czaplewski, Henryk Krawczyk, Agnieszka G. Lipska, Adam Liwo, Jerzy Proficz, Adam K. Sieradzan, and Paweł Czarnul. Multi-GPU UN-RES for scalable coarse-grained simulations of very large protein systems. Computer Physics Communications, 298(??): ??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000353.

OConnor:2022:QHD

[OCR^{+}22] Scott O’Connor, Zane D. Crawford, O. H. Ramachandran, John Luginsland, and B. Shanker. Quasi-Helmholtz decomposition, Gauss’ laws and charge conservation for finite element particle-in-cell. Computer Physics Communications, 276(??): ??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000637.

OConnor:2022:EIS

[ODR^{+}22] Joseph O’Connor, José M. Domínguez, Benedict D. Rogers, Steven J. Lind, and Peter K. Stansby. Eulerian incompressible smoothed particle hydrodynamics on multiple GPUs. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003751.

Omar:2022:QLE

[OEI^{+}22] Hisham Ba Omar, Miguel Ángel Escobedo, Ajaharul Islam, Michael Strickland, Sabin Thapa, Peter Vander Griend, and Johannes Heinrich Weber. QTRAJ 1.0: a Lindblad equation solver for heavy-quarkonium dynamics. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003787.

Ohana:2021:GSM

[OGL^{+}21] Noé Ohana, Claudio Gheller, Emmanuel Lanti, Andreas Jocksch, Stephan Brunner, and Laurent Villard. Gyrokinetic simulations on many- and multi-core architectures with the global electromagnetic Particle-In-Cell Code ORB5. Computer Physics Communications, 262(??):??, May 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300461.

ODonncha:2020:ATP

[OIA^{+}20] Fearghal O’Donncha, Roman Iakymchuk, Albert Akhriev, Philipp Gschwandtner, Peter Thoman, Thomas Heller, Xavier Aguilar, Kiril Dichev, Charles Gillan, Stefano Markidis, Erwin Laure, Emanuele Ragnoli, Vassilis Vassiliadis, Michael Johnston, Herbert Jordan, and Thomas Fahringer. AllScale toolchain pilot applications: PDE based solvers using a parallel development environment. Computer Physics Communications, 251(??):??, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551930400X.

Ozgun:2023:PFL

[OKBM23] Ozlem Ozgun, Mustafa Kuzuoglu, Hadrien Beriot, and Raj Mittra. Parametrization-free locally-conformal perfectly matched layer method for finite element solution of Helmholtz equation. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000863.

Ouro:2021:PHS

[OLNG21] Pablo Ouro, Unai Lopez-Novoa, and Martyn F. Guest. On the performance of a highly-scalable Computational Fluid Dynamics code on AMD, ARM and Intel processor-based HPC systems. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002174.

Olsen:2023:MMS

[Ols23] Richard Olsen. MolTwister — a molecular systems construction, manipulation and statistical mechanical calculation tool. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523001674.

Owen:2020:VBD

[ONH^{+}20] Benjamin Owen, Abouzied M. A. Nasar, Adrian R. G. Harwood, Sam Hewitt, Nicholas Bojdo, Bernard Keavney, Benedict D. Rogers, and Alistair Revell. Vector-based discrete element method for solid elastic materials. Computer Physics Communications, 254(??):??, September 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030151X.

Orgogozo:2022:RNV

[Org22] Laurent Orgogozo. RichardsFoam3: a new version of RichardsFoam for continental surfaces hydrogeology modelling. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002940.

Ostanin:2024:ACR

[OS24] Igor A. Ostanin and Matthias Sperl. Arbitrary controlled reorientation of a spinning body by evolving its tensor of inertia. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001048.

Ozgun:2020:PVP

[OSE^{+}20] Ozlem Ozgun, Volkan Sahin, Muhsin Eren Erguden, Gokhan Apaydin, Asim Egemen Yilmaz, Mustafa Kuzuoglu, and Levent Sevgi. PETOOL v2.0: Parabolic equation toolbox with evaporation duct models and real environment data. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302125.

Osorno:2021:CPH

[OSK^{+}21] M. Osorno, M. Schirwon, N. Kijanski, R. Sivanesapillai, H. Steeb, and D. Göddeke. A cross-platform, high-performance SPH toolkit for image-based flow simulations on the pore scale of porous media. Computer Physics Communications, 267(??): ??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001715.

Ochoa:2021:PPA

[OSLC21] Rodrigo Ochoa, Miguel A. Soler, Alessandro Laio, and Pilar Cossio. PARCE: Protocol for amino acid refinement through computational evolution. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303544.

Osca:2023:IPP

[OV23] Javier Osca and Jiri Vala. Implementation of photon partial distinguishability in a quantum optical circuit simulation. Computer Physics Communications, 289(??):??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523001182.

Orgogozo:2023:PMO

[OXOG23] L. Orgogozo, T. Xavier, H. Oulbani, and C. Grenier. Permafrost modelling with OpenFOAM(R): New advancements of the permaFoam solver. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002600.

Ogata:2024:PCP

[OYC24] Kazuyuki Ogata, Kazuki Yoshida, and Yoshiki Chazono. pikoe: a computer program for distorted-wave impulse approximation calculation for proton induced nucleon knockout reactions. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523004034.

Page:2024:FEA

[Pag24] James Page. Fast exact algorithm for neutrino oscillation in constant matter density. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001231.

Pashov:2020:QPE

[PAL^{+}20] Dimitar Pashov, Swagata Acharya, Walter R. L. Lambrecht, Jerome Jackson, Kirill D. Belashchenko, Athanasios Chantis, Francois Jamet, and Mark van Schilfgaarde. Questaal: a package of electronic structure methods based on the linear muffin-tin orbital technique. Computer Physics Communications, 249(??):??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303868.

Palmiotto:2024:DMS

[PAM24] Marco Palmiotto, Alexandre Arbey, and Farvah Mahmoudi. DarkPACK: a modular software to compute BSM squared amplitudes for particle physics and dark matter observables. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002503.

Park:2021:YYA

[Par21] Chan Beom Park. YAM2: Yet another library for the M_{2} variables using sequential quadratic programming. Computer Physics Communications, 264(??):??, July 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000813.

Pan:2022:SSM

[PAZ^{+}22] Qingquan Pan, Nan An, Tengfei Zhang, Xiaojing Liu, Yun Cai, Lianjie Wang, and Kan Wang. Single-step Monte Carlo criticality algorithm. Computer Physics Communications, 279(??): ??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001588.

Pandit:2023:FPM

[PB23] Abhiyan Pandit and Angelo Bongiorno. A first-principles method to calculate fourth-order elastic constants of solid materials. Computer Physics Communications, 288(??):??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000966.

Penazzi:2024:PIF

[PBC^{+}24] L. Penazzi, S. Blanco, C. Caliot, C. Coustet, M. El Hafi, R. Fournier, J. Gautrais, A. Golijanek-Jedrzejczyk, and M. Sans. Path integrals formulations leading to propagator evaluation for coupled linear physics in large geometric models. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002564.

Palov:2021:VCP

[PBK21] A. P. Palov and G. G. Balint-Kurti. VPA: Computer program for the computation of the phase shift in atom-atom potential scattering using the Variable Phase Approach. Computer Physics Communications, 263(??):??, June 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000412.

Planelles:2021:SVQ

[PC21] Josep Planelles and Juan I. Climente. A simple variational quantum Monte Carlo-effective mass approach for excitons and trions in quantum dots. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303866.

Pekmezi:2024:UBC

[PCL24] Gerald Pekmezi, Bruno Chareyre, and David Littlefield. Uniform boundary conditions on models of spherical particles through alpha shape surface tracking and Laguerre–Voronoi diagrams. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001371.

Prat:2020:ABM

[PCS^{+}20] Raphaël Prat, Thierry Carrard, Laurent Soulard, Olivier Durand, Raymond Namyst, and Laurent Colombet. AMR-based molecular dynamics for non-uniform, highly dynamic particle simulations. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300308.

Poulain:2023:BHO

[PCS^{+}23] Arthur Poulain, Cédric Content, Denis Sipp, Georgios Rigas, and Eric Garnier. BROADCAST: a high-order compressible CFD toolbox for stability and sensitivity using Algorithmic Differentiation. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002764.

Peng:2023:TDS

[PD23] Han Peng and Ralf Deiterding. A three-dimensional solver for simulating detonation on curvilinear adaptive meshes. Computer Physics Communications, 288(??):??, July 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000978.

Pan:2024:ECC

[PDD24] Feng Pan, Lianrong Dai, and Jerry P. Draayer. Elementary coupling coefficients for the Wigner supermultiplet symmetry. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003703.

Prearo:2022:RIH

[PFG22] Ivan Prearo, Arnaldo L. Lixandrão Filho, and Sandro Guedes. RDM: an R interface for high-throughput simulation of ion-material interactions using TRIM. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001709.

Polak:2023:FOA

[PG23] S. Polak and X. Gao. Fourth-order accurate numerical modeling of the multi-fluid plasma equations with adaptive mesh refinement. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001224.

Poberznik:2024:PPI

[PGS^{+}24] M. Poberznik, M. Gunde, N. Salles, A. Jay, A. Hemeryck, N. Richard, N. Mousseau, and L. Martin-Samos. pARTn: a plugin implementation of the Activation Relaxation Technique nouveau that takes over the FIRE minimisation algorithm. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003065.

Peng:2021:CCP

[PGYF21] Gang Peng, Zhiming Gao, Wenjing Yan, and Xinlong Feng. The cell-centered positivity-preserving finite volume scheme for 3D anisotropic diffusion problems on distorted meshes. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002113.

Park:2024:ESC

[PK24] Sae-Hoon Park and Yu-Seok Kim. Extraction simulation of a carbon ion beam with particle distribution in a three-electrode system. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002485.

Passalacqua:2020:SOR

[PLF20] Alberto Passalacqua, Frédérique Laurent, and Rodney O. Fox. A second-order realizable scheme for moment advection on unstructured grids. Computer Physics Communications, 248(??): ??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303352.

Pan:2024:HOC

[PLL24] Kejia Pan, Jin Li, and Zhilin Li. High order compact augmented methods for Stokes equations with different boundary conditions. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001565.

Pelusi:2022:TTL

[PLSB22] Francesca Pelusi, Matteo Lulli, Mauro Sbragaglia, and Massimo Bernaschi. TLBfind: a Thermal Lattice Boltzmann code for concentrated emulsions with FINite-size droplets. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003714.

Pan:2023:PPD

[PLT^{+}23] Qingquan Pan, Huanwen Lv, Songqian Tang, Jinbiao Xiong, and Xiaojing Liu. Pointing probability driven semi-analytic Monte Carlo method (PDMC) — Part I: Global variance reduction for large-scale radiation transport analysis. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001959.

Piris:2021:DOS

[PM21] Mario Piris and Ion Mitxelena. DoNOF: an open-source implementation of natural-orbital-functional-based methods for quantum chemistry. Computer Physics Communications, 259 (??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303167.

Palenzuela:2021:SUP

[PMA^{+}21] C. Palenzuela, B. Miñano, A. Arbona, C. Bona-Casas, C. Bona, and J. Massó. Simflowny 3: an upgraded platform for scientific modeling and simulation. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303271.

Pagliantini:2023:ECE

[PMK^{+}23] C. Pagliantini, G. Manzini, O. Koshkarov, G. L. Delzanno, and V. Roytershteyn. Energy-conserving explicit and implicit time integration methods for the multi-dimensional Hermite– DG discretization of the Vlasov–Maxwell equations. Computer Physics Communications, 284(??):??, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200323X.

Panaino:2020:FPR

[PMS^{+}20a] Costanza M. V. Panaino, Ranald I. Mackay, Marios Sotiropoulos, Karen J. Kirkby, and Michael J. Taylor. Full 3D position reconstruction of a radioactive source based on a novel hyperbolic geometrical algorithm. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519304345.

Piropo:2020:SCS

[PMS20b] Marcelo Pirôpo, Fernando Moraes, and Fernando A. N. Santos. Surfing on curved surfaces — the Maple package Surf. Computer Physics Communications, 249(??):??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551930339X.

Pablos-Marin:2023:SMH

[PMSHG23] José Miguel Pablos-Marín, Javier Serrano, and Carlos Hernández-García. Simulating macroscopic high-order harmonic generation driven by structured laser beams using artificial intelligence. Computer Physics Communications, 291(??): ??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001686.

Penkov:2024:TDE

[PNL^{+}24] Boyan Penkov, David Niedzwiecki, Nicolae Lari, Marija Drndić, and Kenneth Shepard. Time-domain event detection using single-instruction, multiple-thread gpGPU architectures in single-molecule biophysical data. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524001140.

Poskus:2022:BPW

[Pos22] Andrius Poskus. BREMS: Partial-wave calculation of spectra and angular distributions of electron-atom bremsstrahlung at electron energies less than 30 MeV (new version announcement). Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001333.

Potz:2020:PML

[Pöt20] Walter Pötz. Perfectly matched layers for Schrödinger-type equations with nontrivial energy-momentum dispersion. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302344.

Potvliege:2024:MCP

[Pot24] R. M. Potvliege. mqdtfit: a collection of Python functions for empirical multichannel quantum defect calculations. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552400095X.

Pereira:2021:HFI

[PP21] André Pereira and Alberto Proença. HEP-Frame: Improving the efficiency of pipelined data transformation and filtering for scientific analyses. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000151.

Pandey:2023:PNI

[PP23] Vivek Pandey and Sudhir K. Pandey. PY-Nodes: an abinitio Python code for searching nodes in a material using Nelder–Mead’s simplex approach. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002892.

Prokopeva:2022:GDA

[PPK22] Ludmila J. Prokopeva, Samuel Peana, and Alexander V. Kildishev. Gaussian dispersion analysis in the time domain: Efficient conversion with Padé approximants. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001321.

Peng:2021:GSE

[PPKK21] Bo Peng, Ajay Panyala, Karol Kowalski, and Sriram Krishnamoorthy. GFCCLib: Scalable and efficient coupled-cluster Green’s function library for accurately tackling many-body electronic structure problems. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001120.

Pathak:2021:MVR

[PPR^{+}21] Abhishek Pathak, Avinash Pawnday, Aditya Prasad Roy, Amjad J. Aref, Gary F. Dargush, and Dipanshu Bansal. MCBTE: a variance-reduced Monte Carlo solution of the linearized Boltzmann transport equation for phonons. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001156.

Pedretti:2023:XSI

[PRR23] Enrico Pedretti, Paolo Restuccia, and M. Clelia Righi. Xsorb: a software for identifying the most stable adsorption configuration and energy of a molecule on a crystal surface. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001728.

Pfeiler:2020:CMC

[PRS^{+}20] Carl-Martin Pfeiler, Michele Ruggeri, Bernhard Stiftner, Lukas Exl, Matthias Hochsteger, Gino Hrkac, Joachim Schöberl, Norbert J. Mauser, and Dirk Praetorius. Computational micromagnetics with Commics. Computer Physics Communications, 248(??):??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303145.

Postnicov:2024:ECC

[PSK^{+}24] Vasily Postnicov, Aleksei Samarin, Marina V. Karsanina, Mathieu Gravey, Aleksey Khlyupin, and Kirill M. Gerke. Evaluation of classical correlation functions from 2/3D images on CPU and GPU architectures: Introducing CorrelationFunctions.jl. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000572.

Pardal:2023:RES

[PSMRS^{+}23] M. Pardal, A. Sainte-Marie, A. Reboul-Salze, R. A. Fonseca, and J. Vieira. RaDiO: an efficient spatiotemporal radiation diagnostic for particle-in-cell codes. Computer Physics Communications, 285(??):??, April 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003538.

Pawlowski:2023:QDM

[PSW23] Jan M. Pawlowski, Coralie S. Schneider, and Nicolas

Wink. QMeS-Derivation: Mathematica package for the

symbolic derivation of functional equations. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000565.

Plekhanov:2020:SSP

[PTD20] E. Plekhanov, A. Tchougréeff, and R. Dronskowski. ΘΦ: Solid state package allowing Bardeen–Cooper–Schrieffer and magnetic superstructure electronic states. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303947.

Papp:2024:PRD

[PTH24] Paszkál Papp, Ágota Tóth, and Dezső Horváth. PaRe-DiSo: a reaction–diffusion solver coupled with OpenMPI and CVODE. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001164.

Podeszwa:2024:ZOS

[PWC24] Rafał Podeszwa, Henryk A. Witek, and Chien-Pin Chou. ZZPolyCalc: an open-source code with fragment caching for determination of Zhang–Zhang polynomials of carbon nanostructures. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001334.

Peng:2022:REI

[PWD22] Shiyu Peng, Hongming Weng, and Xi Dai. RTGW2020: an efficient implementation of the multi-orbital Gutzwiller method with general local interactions. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000662.

Pan:2024:PPC

[PYT^{+}24] Zi-Yi Pan, Ni Yang, Ming Tang, Peixun Shen, and Xiao-Xiao Cai. Prompt: Probability-conserved cross section biasing Monte Carlo particle transport system. Computer Physics Communications, 295(??):??, February 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003491.

Qian:2023:SBD

[QCZ23] Ye-Cheng Qian, Rong-Rong Cai, and Li-Zhi Zhang. A spheropolyhedral-based discrete element lattice Boltzmann method for simulation of non-spherical adhesive particulate flow. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001546.

Qiao:2021:DME

[QJ21] Jie Qiao and Quanlin Jie. Density matrix embedding theory of excited states for spin systems. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520303519.

Qi:2021:EHO

[QWZ^{+}21] Hongxin Qi, Yuheng Wang, Jie Zhang, Xianghui Wang, and Jianguo Wang. Explicit high-order exponential time integrator for discontinuous Galerkin solution of Maxwell’s equations. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001922.

Ruiz:2024:OTB

[RAJ^{+}24] Christian Guzman Ruiz, Mario Acosta, Oriol Jorba, Eduardo Cesar Galobardes, Matthew Dawson, Guillermo Oyarzun, Carlos Pérez García-Pando, and Kim Serradell. Optimized thread-block arrangement in a GPU implementation of a linear solver for atmospheric chemistry mechanisms. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001632.

Romero:2020:HPI

[RBFB20] Joshua Romero, Mauro Bisson, Massimiliano Fatica, and Massimo Bernaschi. High performance implementations of the 2D Ising model on GPUs. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520302228.

Ren:2022:RSH

[RBV^{+}22] Owen Ren, Mohamed Ali Boussaidi, Dmitry Voytsekhovsky, Manabu Ihara, and Sergei Manzhos. Random Sampling High Dimensional Model Representation Gaussian Process Regression (RS-HDMR-GPR) for representing multidimensional functions with machine-learned lower-dimensional terms allowing insight with a general method. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003325.

Rodriguez-Borbon:2024:TGA

[RBWD^{+}24] José M. Rodríguez-Borbón, Xian Wang, Adrián P. Diéguez, Khaled Z. Ibrahim, and Bryan M. Wong. TRAVOLTA: GPU acceleration and algorithmic improvements for constructing quantum optimal control fields in photo-excited systems. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003624.

Ricketson:2023:PIP

[RC23] L. F. Ricketson and G. Chen. A pseudospectral implicit particle-in-cell method with exact energy and charge conservation. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552300156X.

Reinarz:2020:EEP

[RCB^{+}20] Anne Reinarz, Dominic E. Charrier, Michael Bader, Luke Bovard, Michael Dumbser, Kenneth Duru, Francesco Fambri, Alice-Agnes Gabriel, Jean-Matthieu Gallard, Sven Köppel, Lukas Krenz, Leonhard Rannabauer, Luciano Rezzolla, Philipp Samfass, Maurizio Tavelli, and Tobias Weinzierl. ExaHyPE: an engine for parallel dynamically adaptive simulations of wave problems. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552030076X.

Renda:2021:BMC

[RCB21] M. Renda, D. A. Ciubotaru, and C. I. Banu. Betaboltz: a Monte-Carlo simulation tool for gas scattering processes. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001697.

Reascos:2024:BCD

[RCP^{+}24] Leander Reascos, Fábio Carneiro, André Pereira, Nuno Filipe Castro, and Ricardo Mendes Ribeiro. Berry: a code for the differentiation of Bloch wavefunctions from DFT calculations. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552300317X.

Russkov:2021:ARR

[RCS21] Alexander Russkov, Roman Chulkevich, and Lev N. Shchur. Algorithm for replica redistribution in an implementation of the population annealing method on a hybrid supercomputer architecture. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303933.

Romero:2020:ZGA

[RCW^{+}20] J. Romero, J. Crabill, J. E. Watkins, F. D. Witherden, and A. Jameson. ZEFR: a GPU-accelerated high-order solver for compressible viscous flows using the flux reconstruction method. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300229.

Rao:2024:HCG

[RCX24] Sihang Rao, Bing Chen, and Xu Xu. Heterogeneous CPU–GPU parallelization for modeling supersonic reacting flows with detailed chemical kinetics. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524001115.

Reed:2020:DPR

[RDC^{+}20] Darren S. Reed, Tim Dykes, Rubén Cabezón, Claudio Gheller, and Lucio Mayer. DIAPHANE: a portable radiation transport library for astrophysical applications. Computer Physics Communications, 252(??):??, July 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300679.

Ramasubramani:2020:FSS

[RDH^{+}20] Vyas Ramasubramani, Bradley D. Dice, Eric S. Harper, Matthew P. Spellings, Joshua A. Anderson, and Sharon C. Glotzer. freud: a software suite for high throughput analysis of particle simulation data. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300916.

Rodriguez:2024:GES

[RdPS24] Alejandro Quirós Rodríguez, Miguel Fosas de Pando, and Taraneh Sayadi. Gradient-enhanced stochastic optimization of high-fidelity simulations. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000456.

Ragossnig:2020:IDC

[RDR^{+}20] Florian Ragossnig, Ernst A. Dorfi, Bernhard Ratschiner, Lukas Gehrig, Daniel Steiner, Alexander Stökl, and Colin P. Johnstone. 1 + 1D implicit disk computations. Computer Physics Communications, 256(??):??, November 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302071.

Ricardo:2024:SDP

[RdSH^{+}24] E. Ricardo, F. da Silva, S. Heuraux, A. Silva, and J. Santos. Simulation and data processing techniques to design optimized PPR systems on plasma fusion devices. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002904.

Rangel:2020:RGC

[RDV^{+}20] Tonatiuh Rangel, Mauro Del Ben, Daniele Varsano, Gabriel Antonius, Fabien Bruneval, Felipe H. da Jornada, Michiel J. van Setten, Okan K. Orhan, David D. O’Regan, Andrew Canning, Andrea Ferretti, Andrea Marini, Gian-Marco Rignanese, Jack Deslippe, Steven G. Louie, and Jeffrey B. Neaton. Reproducibility in G0W0 calculations for solids. Computer Physics Communications, 255(??):??, October 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300734.

Rusconi:2020:OSC

[RDZ^{+}20] Simone Rusconi, Denys Dutykh, Arghir Zarnescu, Dmitri Sokolovski, and Elena Akhmatskaya. An optimal scaling to computationally tractable dimensionless models: Study of latex particles morphology formation. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519302954.

Reinhardt:2021:GIF

[RG21] Martin Reinhardt and Helmut Grubmüller. GROMACS implementation of free energy calculations with non-pairwise variationally derived intermediates. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000631.

Richardson:2021:TLP

[RGS^{+}21] A. S. Richardson, D. F. Gordon, S. B. Swanekamp, I. M. Rittersdorf, P. E. Adamson, O. S. Grannis, G. T. Morgan, A. Ostenfeld, K. L. Phlips, C. G. Sun, G. Tang, and D. J. Watkins. TurboPy: a lightweight Python framework for computational physics. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302897.

Ramirez:2022:PPP

[RHLTG^{+}22] Erick Ramírez, Sergio Hernández-López, Enelio Torres-Garcia, Karla Reyes-Morales, and Jorge Balmaseda. pICNIK: a Python package with isoconversional computations for non-isothermal kinetics. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001357.

Raza:2021:NCO

[RHW^{+}21] Akber Raza, Chengkuan Hong, Xian Wang, Anshuman Kumar, Christian R. Shelton, and Bryan M. Wong. NIC-CAGE: an open-source software package for predicting optimal control fields in photo-excited chemical systems. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302587.

Riesch:2021:MOS

[RJ21] Michael Riesch and Christian Jirauschek. mbsolve: an opensource solver tool for the Maxwell–Bloch equations. Computer Physics Communications, 268(??):??, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002095.

Ryu:2021:CES

[RL21] Hoon Ryu and Seungmin Lee. Cost-efficient simulations of large-scale electronic structures in the standalone manycore architecture. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001909.

Radhakrishnan:2024:MSP

[RLW^{+}24] Anand Radhakrishnan, Henry Le Berre, Benjamin Wilfong, Jean-Sebastien Spratt, Mauro Rodriguez, Tim Colonius, and Spencer H. Bryngelson. Method for scalable and performant GPU-accelerated simulation of multiphase compressible flow. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001619.

Raya-Moreno:2022:PBB

[RMCC22] Martí Raya-Moreno, Xavier Cartoixà, and Jesús Carrete. BTE-Barna: an extension of almaBTE for thermal simulation of devices based on 2D materials. Computer Physics Communications, 281(??):??, December 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002235.

Rubel:2023:BUG

[RMFB23] Oleg Rubel, Jean-Baptiste Moussy, Paul Foulquier, and Véronique Brouet. Band unfolding with a general transformation matrix: From code implementation to interpretation of photoemission spectra. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001455.

Reyes:2021:SQM

[RMM21] Justin A. Reyes, Dan C. Marinescu, and Eduardo R. Mucciolo. Simulation of quantum many-body systems on Amazon cloud. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520303714.

Rannala:2022:MAR

[RMR^{+}22] S. E. Rannala, A. Meo, S. Ruta, W. Pantasri, R. W. Chantrell, P. Chureemart, and J. Chureemart. Models of advanced recording systems: a multi-timescale micromagnetic code for granular thin film magnetic recording systems. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001813.

Rodini:2022:ADN

[Rod22] Simone Rodini. Analytical derivatives of neural networks. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002812.

Rohrmoser:2022:TMC

[Roh22] Martin Rohrmoser. The TMDICE Monte Carlo shower program and algorithm for jet-fragmentation via coherent medium induced radiations and scattering. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000613.

Roy:2020:FFT

[RPG^{+}20] Arko Roy, Sukla Pal, S. Gautam, D. Angom, and P. Muruganandam. FACt: FORTRAN toolbox for calculating fluctuations in atomic condensates. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300989.

Ren:2024:SPI

[RRC^{+}24] Pu Ren, Chengping Rao, Su Chen, Jian-Xun Wang, Hao Sun, and Yang Liu. SeismicNet: Physics-informed neural networks for seismic wave modeling in semi-infinite domain. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003557.

Rossignolo:2023:QQO

[RRM^{+}23] Marco Rossignolo, Thomas Reisser, Alastair Marshall, Phila Rembold, Alice Pagano, Philipp J. Vetter, Ressa S. Said, Matthias M. Müller, Felix Motzoi, Tommaso Calarco, Fedor Jelezko, and Simone Montangero. QuOCS: the quantum optimal control suite. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001273.

Rittich:2020:TPS

[RS20] Hannah Rittich and Robert Speck. Time-parallel simulation of the Schrödinger equation. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301569.

Ravipati:2022:CTW

[RSC^{+}22] Srikanth Ravipati, Giannis D. Savva, Ilektra-Athanasia Christidi, Roland Guichard, Jens Nielsen, Romain Réocreux, and Michail Stamatakis. Coupling the time-warp algorithm with the graph-theoretical kinetic Monte Carlo framework for distributed simulations of heterogeneous catalysts. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002605.

Rakotondrandisa:2020:FET

[RSD20] Aina Rakotondrandisa, Georges Sadaka, and Ionut Danaila. A finite-element toolbox for the simulation of solid-liquid phase-change systems with natural convection. Computer Physics Communications, 253(??):??, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300357.

Riaz:2024:MMT

[RSHS24] Usman Riaz, E. Seegyoung Seol, Robert Hager, and Mark S. Shephard. Modeling and meshing for tokamak edge plasma simulations. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003272.

Reinhard:2021:AHF

[RSM21] P.-G. Reinhard, B. Schuetrumpf, and J. A. Maruhn. The Axial Hartree–Fock + BCS Code SkyAx. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302927.

Robertson:2021:AEP

[RSPJ21] E. J. Robertson, N. Sibalić, R. M. Potvliege, and M. P. A. Jones. ARC 3.0: an expanded Python toolbox for atomic physics calculations. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520304136.

Rubel:2021:PAI

[RTRB21] Oleg Rubel, Fabien Tran, Xavier Rocquefelte, and Peter Blaha. Perturbation approach to ab initio effective mass calculations. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303143.

Ruijl:2020:FFP

[RUV20] B. Ruijl, T. Ueda, and J. A. M. Vermaseren. Forcer, a Form program for the parametric reduction of four-loop massless propagator diagrams. Computer Physics Communications, 253(??):??, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520300424.

Rocha:2020:GCF

[RV20] C. M. R. Rocha and A. J. C. Varandas. A general code for fitting global potential energy surfaces via CHIPR method: Triatomic molecules. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519302917.

Rocha:2021:GCF

[RV21] C. M. R. Rocha and A. J. C. Varandas. A general code for fitting global potential energy surfaces via CHIPR method: Direct-fit diatomic and tetratomic molecules. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302666.

Ravisankar:2021:SSO

[RVM^{+}21] Rajamanickam Ravisankar, Dusan Vudragović, Paulsamy Muruganandam, Antun Balaz, and Sadhan K. Adhikari. Spin-1 spin-orbit- and Rabi-coupled Bose–Einstein condensate solver. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303192.

Ruano:2022:NGM

[RVRT22] Jesus Ruano, Aleix Baez Vidal, Joaquim Rigola, and Francesc Xavier Trias. A new general method to compute dispersion errors on Cartesian stretched meshes for both linear and non-linear operators. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521003040.

Rydzewski:2020:MHL

[Ryd20] Jakub Rydzewski. maze: Heterogeneous ligand unbinding along transient protein tunnels. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519302528.

Rogowski:2024:UMP

[RYS^{+}24] Marcin Rogowski, Brandon C. Y. Yeung, Oliver T. Schmidt, Romit Maulik, Lisandro Dalcin, Matteo Parsani, and Gianmarco Mengaldo. Unlocking massively parallel spectral proper orthogonal decompositions in the PySPOD package. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001693.

Rozanski:2023:EUC

[RZ23] Piotr T. Rózański and Michał Zieliński. Exploiting underlying crystal lattice for efficient computation of Coulomb matrix elements in multi-million atoms nanostructures. Computer Physics Communications, 287(??):??, June 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000383.

Rezavand:2022:GEW

[RZH22] Massoud Rezavand, Chi Zhang, and Xiangyu Hu. Generalized and efficient wall boundary condition treatment in GPU-accelerated smoothed particle hydrodynamics. Computer Physics Communications, 281(??):??, December 2022.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002260.

Ran:2023:VOS

[RZWW23] Zheng Ran, Chunming Zou, Zunjie Wei, and Hongwei Wang. VELAS: an open-source toolbox for visualization and analysis of elastic anisotropy. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002594.

Ren:2024:SNV

[RZY^{+}24] Zhongxue Ren, Bin Zhang, Yan Yang, Yalei Zhu, Guangru Bai, Jinlei Liu, Jing Zhao, and Zengxiu Zhao. SLIMP 2.0: a new version of strong laser interaction model package for atoms and molecules, now with molecular orbital tomography based on high-order harmonic spectra. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552400136X.

Salvat:2023:SSP

[SA23] Francesc Salvat and Pedro Andreo. SBETHE: Stopping powers of materials for swift charged particles from the corrected Bethe formula. Computer Physics Communications, 287(??): ??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000425.

Smets:2021:NMD

[SAC^{+}21] Roch Smets, Nicolas Aunai, Andrea Ciardi, Matthieu Drouin, Martin Campos-Pinto, and Philip Deegan. A new method to dispatch split particles in Particle-In-Cell codes. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303246.

Sitnik:2024:DFM

[SAN24] I. M. Sitnik, I. I. Alexeev, and D. V. Nevsky. Debugging the FUMILIM minimization package. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002138.

Sitnik:2020:FVF

[SAS20] I. M. Sitnik, I. I. Alexeev, and O. V. Selugin. The final version of the FUMILIM minimization package. Computer Physics Communications, 251(??):??, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300436.

Sato:2021:SSP

[Sat21] Ryosuke Sato. SimpleBounce: a simple package for the false vacuum decay. Computer Physics Communications, 258(??): ??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302721.

Stier:2024:VFE

[SBB^{+}24] A. Stier, A. Bottino, M. Boesl, M. Campos Pinto, T. Hayward-Schneider, D. Coster, A. Bergmann, M. Murugappan, S. Brunner, L. Villard, and F. Jenko. Verification of the Fourierenhanced 3D finite element Poisson solver of the gyrokinetic full-f code PICLS. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552400078X.

Solanik:2023:GPC

[SBG23] Michal Solanik, Pavol Bobík, and Ján Genci. Geliosphere — parallel CPU and GPU based models of cosmic ray modulation in the heliosphere. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001923.

Strobl:2020:RED

[SBP20] Severin Strobl, Marcus N. Bannerman, and Thorsten Pöschel. Robust event-driven particle tracking in complex geometries. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300667.

Stegmeir:2023:ALA

[SBZ23] A. Stegmeir, T. Body, and W. Zholobenko. Analysis of locally-aligned and non-aligned discretisation schemes for reactor-scale tokamak edge turbulence simulations. Computer Physics Communications, 290(??):??, September 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001467.

Smirnov:2020:FFI

[SC20] A. V. Smirnov and F. S. Chukharev. FIRE6: Feynman Integral REduction with modular arithmetic. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302644.

Shivpuje:2022:TTT

[SC22] Saurabh Shivpuje and Abhijit Chatterjee. TPMD toolkit: a toolkit for studying rate processes using molecular dynamics trajectories and performing temperature programmed molecular dynamics calculations. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002897.

Singhal:2020:IHC

[SCF20] V. Singhal, S. Chattopadhyay, and V. Friese. Investigation of heterogeneous computing platforms for realtime data analysis in the CBM experiment. Computer Physics Communications, 253(??):??, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300370.

Schick:2021:UPT

[Sch21] Daniel Schick. udkm1Dsim — a Python toolbox for simulating 1D ultrafast dynamics in condensed matter. Computer Physics Communications, 266(??):??, September 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001430.

Smith:2022:GAS

[SCL22] Benjamin D. Smith, Logan W. Cooke, and Lindsay J. LeBlanc. GPU-accelerated solutions of the nonlinear Schrödinger equation for simulating 2D spinor BECs. Computer Physics Communications, 275(??):??, June 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000327.

Sinpetru:2022:AMN

[SCMP^{+}22] Luciana A. Sinpetru, Nicholas H. Crisp, David Mostaza-Prieto, Sabrina Livadiotti, and Peter C. E. Roberts. ADB-Sat: Methodology of a novel panel method tool for aerodynamic analysis of satellites. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000443.

Scott:2021:E

[Sco21] N. S. Scott. Editorial. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520304148.

Scrinzi:2022:TES

[Scr22a] Armin Scrinzi. tRecX — an environment for solving timedependent Schrödinger-like problems. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465521002587.

Sinpetru:2022:AVV

[SCR^{+}22b] Luciana A. Sinpetru, Nicholas H. Crisp, Peter C. E. Roberts, Valeria Sulliotti-Linner, Virginia Hanessian, Georg H. Herdrich, Francesco Romano, Daniel Garcia-Almiñana, Sílvia Rodríguez-Donaire, and Simon Seminari. ADBSat: Verification and validation of a novel panel method for quick aerodynamic analysis of satellites. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000455.

Szriftgiser:2021:CAP

[SCT21] Pascal Szriftgiser and Edgardo S. Cheb-Terrab. Computer algebra in physics: the hidden SO(4) symmetry of the hydrogen atom. Computer Physics Communications, 268(??):??, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001880.

Saha:2024:MSN

[SD24] Avik Saha and Arup Kumar Das. Multiscale simulation of nanodrop over surfaces with varying hydrophilicity. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003089.

Szenes:2024:DSI

[SDBS24] Kalman Szenes, Niccolò Discacciati, Luca Bonaventura, and William Sawyer. Domain-specific implementation of highorder discontinuous Galerkin methods in spherical geometry. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003387.

Shang:2021:MCA

[SDL^{+}21] Honghui Shang, Xiaohui Duan, Fang Li, Libo Zhang, Zhiqian Xu, Kan Liu, Haiwen Luo, Yingrui Ji, Wenxuan Zhao, Wei Xue, Li Chen, and Yunquan Zhang. Many-core acceleration of the first-principles all-electron quantum perturbation calculations. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001570.

Sazzad:2024:PBT

[SDP24] Shah Wasif Sazzad, Sanjay Dharmavaram, and Luigi E. Perotti. A physics-based tessellation algorithm for particle assemblies on arbitrary surfaces. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000894.

Seto:2023:BEF

[SDXY23] Haruki Seto, Benjamin D. Dudson, Xue-Qiao Xu, and Masatoshi Yagi. A BOUT++ extension for full annular tokamak edge MHD and turbulence simulations. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002879.

Schaffer:2024:CFP

[SE24] Sebastian Schaffer and Lukas Exl. Constraint free physicsinformed machine learning for micromagnetic energy minimization. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001255.

Schiffmann:2020:PPC

[SFBG20] Sacha Schiffmann, Livio Filippin, Daniel Baye, and Michel Godefroid. POLALMM: a program to compute polarizabilities for nominal one-electron systems using the Lagrangemesh method. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302101.

Sierra:2020:ESA

[SFC20] Javier Sierra, David Fabre, and Vincenzo Citro. Efficient stability analysis of fluid flows using complex mapping techniques. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519304023.

Santi:2022:RNM

[SFKC22] Lucio Santi, Joaquín Fernández, Ernesto Kofman, and Rodrigo Castro. retQSS: a novel methodology for efficient modeling and simulation of particle systems in reticulated geometries. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002691.

Solovsky:2022:BMD

[SFS22] Jakub Solovský, Radek Fucík, and Jakub Sístek. BDDC for MHFEM discretization of unsteady two-phase flow in porous media. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003118.

Schuler:2020:NES

[SGM^{+}20] Michael Schüler, Denis Golež, Yuta Murakami, Nikolaj Bittner, Andreas Herrmann, Hugo U. R. Strand, Philipp Werner, and Martin Eckstein. NESSi: the N-on-Equilibrium Systems Simulation package. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302277.

Schneider:2023:IIS

[SGS23] Ryan Schneider, Heman Gharibnejad, and Barry I. Schneider. ITVOLT: an iterative solver for the time-dependent Schrödinger equation. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552300125X.

Shi:2024:GTP

[SGY24] Y. H. Shi, N. Guo, and Z. X. Yang. GeoTaichi: a Taichipowered high-performance numerical simulator for multiscale geophysical problems. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001425.

Shakoor:2021:FMO

[Sha21] Modesar Shakoor. FEMS — a mechanics-oriented finite element modeling software. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303593.

Scott:2020:CAC

[SHB^{+}20] N. S. Scott, A. Hibbert, J. Ballantyne, S. Fritzsche, A. L. Hazel, D. P. Landau, D. W. Walker, and Z. Was. CPC’s 50th anniversary: Celebrating 50 years of open-source software in computational physics. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520300886.

Strickstrock:2022:OLJ

[SHRK22] Robin Strickstrock, Marco Hülsmann, Dirk Reith, and Karl N. Kirschner. Optimizing Lennard-Jones parameters by coupling single molecule and ensemble target data. Computer Physics Communications, 274(??):??, May 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000030.

Shao:2022:PTM

[SHS22] Cheng Shao, Takuma Hori, and Junichiro Shiomi. P-TRANS: a Monte Carlo ray-tracing software to simulate phonon transport in arbitrary nanostructures. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000807.

Singh:2021:DOS

[SHW^{+}21] Vijay Singh, Uthpala Herath, Benny Wah, Xingyu Liao, Aldo H. Romero, and Hyowon Park. DMFTwDFT: an opensource code combining Dynamical Mean Field Theory with various density functional theory packages. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552030388X.

Sijoy:2023:CEC

[Sij23] C. D. Sijoy. Compatible and energy conserving multimaterial arbitrary Lagrangian Eulerian scheme for multigroup radiation hydrodynamics simulations. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523000401.

Sitnik:2023:HPN

[Sit23] I. M. Sitnik. 5D histogram package NORA. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200279X.

Schnabel:2020:APS

[SJ20] Stefan Schnabel and Wolfhard Janke. Accelerating polymer simulation by means of tree data-structures and a parsimonious Metropolis algorithm. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520301892.

Schnabel:2021:WLS

[SJ21] Stefan Schnabel and Wolfhard Janke. Wang–Landau simulations with non-flat distributions. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001831.

Sjostrand:2020:PEG

[Sjö20] Torbjörn Sjöstrand. The Pythia event generator: Past, present and future. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519302899.

Salvat:2021:EDP

[SJP21a] Francesc Salvat, Aleksander Jablonski, and Cedric J. Powell. elsepa — Dirac partial-wave calculation of elastic scattering of electrons and positrons by atoms, positive ions and molecules (new version announcement). Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303465.

Schuster:2021:MPM

[SJP^{+}21b] C. U. Schuster, T. Johnson, G. Papp, R. Bilato, S. Sipilä, J. Varje, and M. Hasenöhrl. Moment-preserving and meshadaptive reweighting method for rare-event sampling in Monte-Carlo algorithms. Computer Physics Communications, 267 (??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001533.

Stephens:2022:RRP

[SJWL22] Victoria B. Stephens, Sally Jensen, Isaac Wheeler, and David O. Lignell. RadLib: a radiative property model library for CFD. Computer Physics Communications, 272(??): ??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003398.

Szidarovszky:2018:LCP

[SJY18] Tamás Szidarovszky, Maho Jono, and Kaoru Yamanouchi. LIMAO: Cross-platform software for simulating laser-induced alignment and orientation dynamics of linear-, symmetricand asymmetric tops. Computer Physics Communications, 228(??):219–228, July 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465518300626.

See erratum [SJY20].

Szidarovszky:2020:ELC

[SJY20] Tamás Szidarovszky, Maho Jono, and Kaoru Yamanouchi. Erratum to “LIMAO: Cross-platform software for simulating laser-induced alignment and orientation dynamics of linear-, symmetric- and asymmetric tops” [Comput. Phys. Comm. 228 (2018) 219–228]. Computer Physics Communications, 252(??): ??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519304151. See [SJY18].

Shivpuje:2021:AEK

[SKC21] Saurabh Shivpuje, Manish Kumawat, and Abhijit Chatterjee. An algorithm for estimating kinetic parameters of atomistic rare events using finite-time temperature programmed molecular dynamics trajectories. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000047.

Sadaka:2024:FET

[SKDH24] Georges Sadaka, Victor Kalt, Ionut Danaila, and Frédéric Hecht. A finite element toolbox for the Bogoliubov–de Gennes stability analysis of Bose–Einstein condensates. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300293X.

Sanchez:2024:CBF

[SKEZ24a] A. D. Sanchez, S. Chaitanya Kumar, and M. Ebrahim-Zadeh. CUDA-based focused Gaussian beams secondharmonic generation efficiency calculator. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524001553.

Sanchez:2024:CBO

[SKEZ24b] A. D. Sanchez, S. Chaitanya Kumar, and M. Ebrahim-Zadeh. CUDA-based optical parametric oscillator simulator. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002552.

Stolpp:2021:CSS

[SKM^{+}21] Jan Stolpp, Thomas Köhler, Salvatore R. Manmana, Eric Jeckelmann, Fabian Heidrich-Meisner, and Sebastian Paeckel. Comparative study of state-of-the-art matrix-product-state methods for lattice models with large local Hilbert spaces without U(1) symmetry. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002186.

Shimamura:2024:CML

[SKS24] Kohei Shimamura, Akihide Koura, and Fuyuki Shimojo. Construction of machine-learning interatomic potential under heat flux regularization and its application to power spectrum analysis for silver chalcogenides. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002655.

Seo:2021:NHS

[SKYQ21] J. H. Seo, Jae-Min Kwon, Sumin Yi, and Lei Qi. A new hybrid simulation model for tokamak plasma turbulence. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303003.

Scheffel:2022:TSS

[SL22] Jan Scheffel and Kristoffer Lindvall. Temporal smoothing — a step forward for time-spectral methods. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100285X.

Shah:2023:QPB

[SLB^{+}23] S. A. Shah, Hao Li, Eric R. Bittner, Carlos Silva, and Andrei Piryatinski. QuDPy: a Python-based tool for computing ultrafast non-linear optical responses. Computer Physics Communications, 292(??):??, November 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002369.

Saini:2022:WTI

[SLBR22] Himanshu Saini, Magdalena Laurien, Peter Blaha, and Oleg Rubel. WloopPHI: a tool for ab initio characterization of Weyl semimetals. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002599.

Silva:2022:FFE

[SLC^{+}22] Alexandre Silva, Veniero Lenzi, Albano Cavaleiro, Sandra Carvalho, and Luís Marques. FELINE: Finite element solver for hydrodynamic lubrication problems using the inexact Newton method. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552200159X.

Singh:2021:MPL

[SLDF^{+}21] Sobhit Singh, Logan Lang, Viviana Dovale-Farelo, Uthpala Herath, Pedram Tavadze, François-Xavier Coudert, and Aldo H. Romero. MechElastic: a Python library for analysis of mechanical and elastic properties of bulk and 2D materials. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001806.

Schiffmann:2022:RRE

[SLE^{+}22] S. Schiffmann, J. G. Li, J. Ekman, G. Gaigalas, M. Godefroid, P. Jönsson, and J. Bieroń. Relativistic radial electron density functions and natural orbitals from GRASP2018. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001229.

Silva:2024:MCS

[SLIC24] B. H. Silva, D. Lacroix, M. Isaiev, and L. Chaput. Monte Carlo simulation of phonon transport from ab-initio data with nano-κ. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002990.

Shemyakin:2023:NFT

[SLK23] O. P. Shemyakin, P. R. Levashov, and P. A. Krasnova. New features of the TFmix code: Thermodynamic properties of electrons in mixtures. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523001212.

Spannaus:2021:MFC

[SLL^{+}21] Adam Spannaus, Kody J. H. Law, Piotr Luszczek, Farzana Nasrin, Cassie Putman Micucci, Peter K. Liaw, Louis J. Santodonato, David J. Keffer, and Vasileios Maroulas. Materials fingerprinting classification. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521001314.

Schmalzried:2022:EHE

[SLL22] Anthony Schmalzried, Alejandro Luque, and Nikolai Lehtinen. Enhancing higher-energy spectral resolution for electron particle simulations in air. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000856.

Shalev:2024:MUR

[SLL^{+}24] Artem Shalev, Konstantin Ladutenko, Igor Lobanov, Vassilios Yannopapas, and Alexander Moroz. Multem 3: an updated and revised version of the program for transmission and band calculations of photonic crystals. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001413.

Salvat:2022:ECC

[SLLA22] Francesc Salvat, Josep Llosa, Antonio M. Lallena, and Julio Almansa. ECCPA: Calculation of classical and quantum cross sections for elastic collisions of charged particles with atoms. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200087X.

Safa:2022:TPP

[SLP^{+}22] Ibrahim Safa, Jeffrey Lazar, Alex Pizzuto, Oswaldo Vasquez, Carlos A. Argüelles, and Justin Vandenbroucke. TauRunner: a public Python program to propagate neutral and charged leptons. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001412.

Schwersenz:2024:CTD

[SLZ^{+}24] Nico Schwersenz, Victor Loaiza, Tim Zimmermann, Javier Madroñero, and Sandro Wimberger. Comparison of two different integration methods for the (1 + 1)-dimensional Schrödinger–Poisson equation. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001152.

Schlipf:2020:SPC

[SLZG20] Martin Schlipf, Henry Lambert, Nourdine Zibouche, and Feliciano Giustino. SternheimerGW: a program for calculating GW quasiparticle band structures and spectral functions without unoccupied states. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519302334.

Shang:2021:EPL

[SLZY21] Honghui Shang, WanZhen Liang, Yunquan Zhang, and Jinlong Yang. Efficient parallel linear scaling method to get the response density matrix in all-electron realspace density-functional perturbation theory. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302940.

Shiroto:2024:MNM

[SMA24] Takashi Shiroto, Akinobu Matsuyama, and Nobuyuki Aiba. MUSES: a nonlinear magnetohydrodynamics discontinuous Galerkin code for fusion plasmas. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523004162.

Semenov:2020:MUR

[SMB20] Alexander Semenov, Bikramaditya Mandal, and Dmitri Babikov. MQCT: User-ready program for calculations of inelastic scattering of two molecules. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520300175.

Seleson:2022:DDS

[SMC^{+}22] Pablo Seleson, Mohammad Mustafa, Davide Curreli, Cory D. Hauck, Miroslav Stoyanov, and David E. Bernholdt. Datadriven surrogate modeling of hPIC ion energy-angle distributions for high-dimensional sensitivity analysis of plasma parameters’ uncertainty. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001552.

Stepsys:2021:CFP

[SMGK21] Augustinas Stepsys, Saulius Mickevicius, Darius Germanas, and Ramutis Kazys Kalinauskas. Calculation of fiveparticle harmonic-oscillator transformation brackets. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001351.

Sanna:2020:SNP

[SMO^{+}20a] N. Sanna, G. Morelli, S. Orlandini, M. Tacconi, and I. Baccarelli. SCELib4.0: the new program version for computing molecular properties in the Single Center Approach. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551930311X.

Shtabovenko:2020:FNF

[SMO20b] Vladyslav Shtabovenko, Rolf Mertig, and Frederik Orellana. FeynCalc 9.3: New features and improvements. Computer Physics Communications, 256(??):??, November 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030223X.

Salehi:2023:SIS

[SN23] Saeed Salehi and Håkan Nilsson. A semi-implicit slip algorithm for mesh deformation in complex geometries, implemented in OpenFOAM. Computer Physics Communications, 287(??): ??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000486.

Shinaoka:2020:EIC

[SNG20] Hiroshi Shinaoka, Yusuke Nomura, and Emanuel Gull. Efficient implementation of the continuous-time interactionexpansion quantum Monte Carlo method. Computer Physics Communications, 252(??):??, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465519302036.

Satake:2020:BNM

[SNP^{+}20] Shinsuke Satake, Motoki Nataka, Theerasarn Pianpanit, Hideo Sugama, Masanori Nunami, Seikichi Matsuoka, Seiji Ishiguro, and Ryutaro Kanno. Benchmark of a new multi-ion-species collision operator for δf Monte Carlo neoclassical simulation. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300771.

Soler:2023:MEP

[SOH^{+}23] Miguel Angel Soler, Ozge Ozkilinc, Yamanappa Hunashal, Paolo Giannozzi, Gennaro Esposito, and Federico Fogolari. Molecular electrostatics and pKa shifts calculations with the Generalized Born model. A tutorial through examples with Bluues2. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000619.

Sozykin:2021:GSO

[Soz21] S. A. Sozykin. GUI4dft — a SIESTA oriented GUI. Computer Physics Communications, 262(??):??, May 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100014X.

Schwendt:2020:TBC

[SP20] Mathias Schwendt and Walter Pötz. Transparent boundary conditions for higher-order finite-difference schemes of the Schrödinger equation in (1 + 1)D. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303789.

Sihi:2023:TPC

[SP23] Antik Sihi and Sudhir K. Pandey. TRACK: a Python code for calculating the transport properties of correlated electron systems using Kubo formalism. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522003599.

Smid:2021:XRS

[SPF21] Michal Smíd, Xiayun Pan, and Katerina Falk. X-ray spectrometer simulation code with a detailed support of mosaic crystals. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520304045.

Schmidmayer:2020:EOS

[SPLD20] Kevin Schmidmayer, Fabien Petitpas, Sébastien Le Martelot, and Éric Daniel. ECOGEN: an open-source tool for multiphase, compressible, multiphysics flows. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303959.

Shojaei:2023:STP

[SPMS23] Mostafa Faghih Shojaei, John E. Pask, Andrew J. Medford, and Phanish Suryanarayana. Soft and transferable pseudopotentials from multi-objective optimization. Computer Physics Communications, 283(??):??, February 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003137.

Sanchez-Puga:2021:BUD

[SPTPR21] Pablo Sánchez-Puga, Javier Tajuelo, Juan Manuel Pastor, and Miguel A. Rubio. BiconeDrag updated — a data processing application for the oscillating conical bob interfacial shear rheometer. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001867.

Santi:2021:EDE

[SRC21] Lucio Santi, Lucas Rossi, and Rodrigo Castro. Efficient discrete-event based particle tracking simulation for high energy physics. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302976.

Schild:2024:PPI

[SRE^{+}24] Nils Schild, Mario Räth, Sebastian Eibl, Klaus Hallatschek, and Katharina Kormann. A performance portable implementation of the semi-Lagrangian algorithm in six dimensions. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003181.

Safina:2024:IPG

[SRK^{+}24] Liliya R. Safina, Elizaveta A. Rozhnova, Karina A. Krylova, Ramil T. Murzaev, and Julia A. Baimova. Interatomic potentials for graphene reinforced metal composites: Optimal choice. Computer Physics Communications, 301(??):??, August 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001589.

Shevchenko:2020:MEA

[SRML20] Ivan I. Shevchenko, Guillaume Rollin, Alexander V. Melnikov, and José Lages. Massive evaluation and analysis of Poincaré recurrences on grids of initial data: a tool to map chaotic diffusion. Computer Physics Communications, 246(??):??, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519302553.

Sadaka:2020:PFE

[SRT^{+}20] Georges Sadaka, Aina Rakotondrandisa, Pierre-Henri Tournier, Francky Luddens, Corentin Lothodé, and Ionut Danaila. Parallel finite-element codes for the simulation of two-dimensional and three-dimensional solid-liquid phase-change systems with natural convection. Computer Physics Communications, 257 (??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302319.

Sartore:2021:P

[SS21a] Lohan Sartore and Ingo Schienbein. PyR@TE 3. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520304124.

Song:2021:ERA

[SS21b] Yang Song and Bhuvana Srinivasan. An efficient reconstruction algorithm for diffusion on triangular grids using the nodal discontinuous Galerkin method. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000308.

Siboni:2022:AWS

[SS22] Morteza H. Siboni and Mark S. Shephard. Adaptive workflow for simulation of RF heaters. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001539.

Singor:2023:CST

[SSB^{+}23] Adam Singor, Jeremy S. Savage, Igor Bray, Barry I. Schneider, and Dmitry V. Fursa. Continuum solutions to the two-center Coulomb problem in prolate spheroidal coordinates. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002338.

Silvestri:2022:SFP

[SSD^{+}22] Luciano G. Silvestri, Lucas J. Stanek, Gautham Dharuman, Yongjun Choi, and Michael S. Murillo. Sarkas: a fast pure-Python molecular dynamics suite for plasma physics. Computer Physics Communications, 272(??):??, March 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100357X.

Schilling:2023:BSR

[SSH^{+}23] Jonathan Schilling, Jakob Svensson, Udo Höfel, Joachim Geiger, and Henning Thomsen. Biot–Savart routines with minimal floating point error. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000371.

Sieber:2024:BPF

[SSPF24] Armand Baptiste Sieber, Henri Hugo Sieber, Davide Bernardo Preso, and Mohamed Farhat. BIMBAMBUM: a potential flow solver for single cavitation bubble dynamics. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000730.

Smirnov:2022:FNH

[SSV22] A. V. Smirnov, N. D. Shapurov, and L. I. Vysotsky. FIESTA5: Numerical high-performance Feynman integral evaluation. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001059.

Shimizu:2023:QAM

[ST23a] Noritaka Shimizu and Yusuke Tsunoda. SO(3) quadratures in angular-momentum projection. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522003022.

Singh:2023:PED

[ST23b] Jaswant Singh and Tobias Toll. Predicting the exclusive diffractive electron-ion cross section at small x with machine learning in Sartre. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002175.

Salinas:2020:FEI

[STA20] Álvaro Salinas, Claudio Torres, and Orlando Ayala. A fast and efficient integration of boundary conditions into a unified CUDA kernel for a shallow water solver lattice Boltzmann method. Computer Physics Communications, 249(??): ??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303443.

Smith:2020:CLM

[STRF^{+}20] E. R. Smith, D. J. Trevelyan, E. Ramos-Fernandez, A. Sufian, C. O’Sullivan, and D. Dini. CPL library — a minimal framework for coupled particle and continuum simulation. Computer Physics Communications, 250(??):??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519303819.

Svaneborg:2024:IPP

[Sva24] Carsten Svaneborg. Inverse primitive path analysis. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001322.

Shu:2024:PNV

[SVJ^{+}24] Yinan Shu, Zoltan Varga, Ahren Jasper, Joaquin Espinosa-Garcia, Jose C. Corchado, and Donald G. Truhlar. PotLib 2023: New version of a potential energy surface library for chemical systems. Computer Physics Communications, 294 (??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002825.

Sanchez-Vizuet:2020:AHD

[SVSC20] Tonatiuh Sánchez-Vizuet, Manuel E. Solano, and Antoine J. Cerfon. Adaptive hybridizable discontinuous Galerkin discretization of the Grad–Shafranov equation by extension from polygonal subdomains. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520300710.

Sun:2023:CCG

[SW23] Y. Z. Sun and S. T. Wang. CNOK: a C++ Glauber model code for single-nucleon knockout reactions. Computer Physics Communications, 288(??):??, July 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000711.

Steiniger:2023:EEC

[SWB^{+}23] Klaus Steiniger, Rene Widera, Sergei Bastrakov, Michael Bussmann, Sunita Chandrasekaran, Benjamin Hernandez, Kristina Holsapple, Axel Huebl, Guido Juckeland, Jeffrey Kelling, Matt Leinhauser, Richard Pausch, David Rogers, Ulrich Schramm, Jeff Young, and Alexander Debus. EZ: an efficient, charge conserving current deposition algorithm for electromagnetic particle-in-cell simulations. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001947.

Sun:2023:DOS

[SWTC23] Jie Sun, Yiqing Wang, Baolin Tian, and Zheng Chen. detonationFoam: an open-source solver for simulation of gaseous detonation based on OpenFOAM. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523002047.

Shi:2023:RIP

[SWZ23] Shuzhe Shi, Lingxiao Wang, and Kai Zhou. Rethinking the ill-posedness of the spectral function reconstruction — why is it fundamentally hard and how artificial neural networks can help. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002661.

Xie:2022:BRT

[sXBkB^{+}22] Hua sheng Xie, Debabrata Banerjee, Yu kun Bai, Han yue Zhao, and Jing chun Li. BORAY: a ray tracing code for various magnetized plasma configurations. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000820.

Shang:2020:DPD

[SXW^{+}20] Honghui Shang, Lei Xu, Baodong Wu, Xinming Qin, Yunquan Zhang, and Jinlong Yang. The dynamic parallel distribution algorithm for hybrid density-functional calculations in HON-PAS package. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520300448.

Shi:2024:GAT

[SXYD24] Fulong Shi, Jianjian Xin, Pandeng Yin, and Yehong Dong. A GPU accelerated three-dimensional ghost cell method with an improved implicit surface representation for complex rigid or flexible boundary flows. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524000213.

Sadykov:2020:PNE

[SY20] Renat Sadykov and Vitaly Yermolchyk. Polarized NLO EW e^{+}e^{−} cross section calculations with ReneSANCe-v1.0.0. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302083.

Sabau:2023:OGO

[SYFT23] Adrian S. Sabau, Lang Yuan, Jean-Luc Fattebert, and John A. Turner. An OpenMP GPU-offload implementation of a nonequilibrium solidification cellular automata model for additive manufacturing. Computer Physics Communications, 284(??): ??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003241.

Shen:2023:NVS

[SZ23] Yulong Shen and Nengji Zhou. Numerical variational studies of quantum phase transitions in the sub-Ohmic spinboson model with multiple polaron ansatz. Computer Physics Communications, 293(??):??, December 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002400.

Shang:2024:ETC

[SZ24a] Liangliang Shang and Yang Zhang. EasyScan HEP: a tool for connecting programs to scan the parameter space of physics models. Computer Physics Communications, 296(??): ??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003727.

Smirnov:2024:FFI

[SZ24b] Alexander V. Smirnov and Mao Zeng. FIRE 6.5: Feynman integral reduction with new simplification library. Computer Physics Communications, 302(??):??, September 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552400184X.

Song:2023:PPC

[SZNW23] Shanshan Song, Mingyu Zhu, Hongcheng Ni, and Jian Wu. PyStructureFactor: a Python code for the molecular structure factor in tunneling ionization rates. Computer Physics Communications, 292(??):??, November 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002278.

Shu:2024:APA

[SZT24] Yinan Shu, Linyao Zhang, and Donald G. Truhlar. ANT 2023: a program for adiabatic and nonadiabatic trajectories. Computer Physics Communications, 296(??):??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523003661.

She:2022:PUI

[SZY^{+}22] Zhi-Lei She, Dai-Mei Zhou, Yu-Liang Yan, Liang Zheng, Hong ge Xu, Gang Chen, and Ben-Hao Sa. PACIAE 2.2.1: an updated issue of the parton and hadron cascade model PACIAE 2.2. Computer Physics Communications, 274(??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000078.

Thompson:2022:LFS

[TAB^{+}22] Aidan P. Thompson, H. Metin Aktulga, Richard Berger, Dan S. Bolintineanu, W. Michael Brown, Paul S. Crozier, Pieter J. in ’t Veld, Axel Kohlmeyer, Stan G. Moore, Trung Dac Nguyen, Ray Shan, Mark J. Stevens, Julien Tranchida, Christian Trott, and Steven J. Plimpton. LAMMPS — a flexible simulation tool for particle-based materials modeling at the atomic, meso, and continuum scales. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002836.

Turiansky:2021:NCN

[TAE^{+}21] Mark E. Turiansky, Audrius Alkauskas, Manuel Engel, Georg Kresse, Darshana Wickramaratne, Jimmy-Xuan Shen, Cyrus E. Dreyer, and Chris G. Van de Walle. Nonrad: Computing nonradiative capture coefficients from first principles. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001685.

Thomadakis:2022:NDC

[TAGC22a] Polykarpos Thomadakis, Angelos Angelopoulos, Gagik Gavalian, and Nikos Chrisochoides. De-noising drift chambers in CLAS12 using convolutional auto encoders. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521003131.

Thomadakis:2022:UML

[TAGC22b] Polykarpos Thomadakis, Angelos Angelopoulos, Gagik Gavalian, and Nikos Chrisochoides. Using machine learning for particle track identification in the CLAS12 detector. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000790.

Tang:2023:FPM

[Tan23] Jun Tang. A full-potential and multiscale computational scheme for interactions between ultrafast intense laser pulses and condensed medium. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003526.

Tang:2024:HEA

[Tan24] Shujiang Tang. A high-efficiency adaptive TENO scheme with optimal accuracy order for compressible flow simulation. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552400081X.

Tramm:2024:EAM

[TAY^{+}24] John Tramm, Bryce Allen, Kazutomo Yoshii, Andrew Siegel, and Leighton Wilson. Efficient algorithms for Monte Carlo particle transport on AI accelerator hardware. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523004174.

Tulsky:2020:QFC

[TB20] Vasily Tulsky and Dieter Bauer. Qprop with faster calculation of photoelectron spectra. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519304035.

Tretiak:2021:PBI

[TBAR21] Krasymyr Tretiak, James Buchanan, Rob Akers, and Daniel Ruprecht. Performance of the BGSDC integrator for computing fast ion trajectories in nuclear fusion reactors. Computer Physics Communications, 264(??):??, July 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100031X.

Tripura:2024:DIL

[TC24] Tapas Tripura and Souvik Chakraborty. Discovering interpretable Lagrangian of dynamical systems from data. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003053.

Traczykowski:2023:SEB

[TCcN23] P. Traczykowski, L. T. Campbell, and B. W. J. M. c Neil. Upsampling of electron beam simulation particles with addition of shot-noise. Computer Physics Communications, 286(??): ??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000061.

Takahashi:2020:PIF

[TCD20] Toru Takahashi, Chao Chen, and Eric Darve. Parallelization of the inverse fast multipole method with an application to boundary element method. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303194.

Taitano:2021:CPS

[TCSA21] W. T. Taitano, L. Chacón, A. N. Simakov, and S. E. Anderson. A conservative phase-space moving-grid strategy for a 1D-2V Vlasov–Fokker–Planck solver. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302629.

Toffoli:2024:TCM

[TCSD24] Daniele Toffoli, Sonia Coriani, Mauro Stener, and Piero Decleva. Tiresia: a code for molecular electronic continuum states and photoionization. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003831.

Tan:2021:NDM

[TCY^{+}21] Nongdie Tan, Lei Chen, Xianglin Ye, Hao Zhou, and Hailing Xiong. A novel dichotomy model based on the traditional CCA. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000965.

Tan:2023:GAC

[TCY23] Zhijun Tan, Le Chen, and Junxiang Yang. Generalized Allen–Cahn-type phase-field crystal model with FCC ordering structure and its conservative high-order accurate algorithm. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000012.

Taylor:2020:AIR

[TDR^{+}20] Ned Thaddeus Taylor, Francis Huw Davies, Isiah Edward Mikel Rudkin, Conor Jason Price, Tsz Hin Chan, and Steven Paul Hepplestone. ARTEMIS: Ab initio restructuring tool enabling the modelling of interface structures. Computer Physics Communications, 257(??):??, December 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302423.

Tolle:2022:TCV

[TGBM22] Tobias Tolle, Dirk Gründing, Dieter Bothe, and Tomislav Marić. triSurfaceImmersion: Computing volume fractions and signed distances from triangulated surfaces immersed in unstructured meshes. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003611.

Thomadakis:2023:CPR

[TGGC23] Polykarpos Thomadakis, Kevin Garner, Gagik Gavalian, and Nikos Chrisochoides. Charged particle reconstruction in CLAS12 using machine learning. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000395.

Tyson:2023:DLL

[TGIM23] R. Tyson, G. Gavalian, D. G. Ireland, and B. McKinnon. Deep learning level-3 electron trigger for CLAS12. Computer Physics Communications, 290(??):??, September 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001285.

Tsiolakis:2020:NPG

[TGS^{+}20] Vasileios Tsiolakis, Matteo Giacomini, Ruben Sevilla, Carsten Othmer, and Antonio Huerta. Nonintrusive proper generalised decomposition for parametrised incompressible flow problems in OpenFOAM. Computer Physics Communications, 249(??): ??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303479.

Tatsumi:2021:SNE

[THH21] Ryoko Tatsumi, Kazuo Hoshino, and Akiyoshi Hatayama. Study of numerical error of a Eulerian–Lagrangian scheme in the presence of particle source. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000771.

Tantardini:2024:GGO

[TIG^{+}24] Christian Tantardini, Miroslav Iliaš, Matteo Giantomassi, Alexander G. Kvashnin, Valeria Pershina, and Xavier Gonze. Generating and grading 34 optimised norm-conserving Vanderbilt pseudopotentials for actinides and super-heavy elements in the PseudoDojo. Computer Physics Communications, 295(??): ??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003478.

Taitano:2021:EVF

[TKC^{+}21] W. T. Taitano, B. D. Keenan, L. Chacón, S. E. Anderson, H. R. Hammer, and A. N. Simakov. An Eulerian Vlasov–Fokker–Planck algorithm for spherical implosion simulations of inertial confinement fusion capsules. Computer Physics Communications, 263(??):??, June 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000242.

Taheri:2022:TDP

[TKS22] Sina Taheri, Jacob R. King, and Uri Shumlak. Timediscretization of a plasma-neutral MHD model with a semiimplicit leapfrog algorithm. Computer Physics Communications, 274(??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000066.

Tavakkol:2020:CBI

[TL20] Sasan Tavakkol and Patrick Lynett. Celeris Base: an interactive and immersive Boussinesq-type nearshore wave simulation software. Computer Physics Communications, 248(??): ??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303169.

Tian:2021:FAG

[TLC^{+}21] Yuan Tian, Ping Lin, Hanjie Cai, Yaling Zhang, Qiong Yang, Meiling Qi, Guanghui Yang, Xiaofei Gao, Xiaolong Chen, Lei Yang, and Qingguo Zhou. A fast and accurate GPU based method on simulating energy deposition for beam-target coupling with granular materials. Computer Physics Communications, 269(??):??, December 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002162.

Tanygin:2024:CES

[TM24] Bogdan Tanygin and Simone Melchionna. Comparison of effective and stable Langevin dynamics integrators. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000754.

Timrov:2022:HCC

[TMC22] Iurii Timrov, Nicola Marzari, and Matteo Cococcioni. HP — a code for the calculation of Hubbard parameters using density-functional perturbation theory. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001746.

Tritt:2023:SGO

[TMH^{+}23] Alex Tritt, Joshua Morris, Joel Hochstetter, R. P. Anderson, James Saunderson, and L. D. Turner. Spinsim: a GPU optimized Python package for simulating spin-half and spin-one quantum systems. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000462.

Truong:2022:SPC

[TNL^{+}22] Thu D. H. Truong, Hanh H. Nguyen, Hieu B. Le, Do Hung Dung, H.-M. Tran, Nguyen Duy Vy, Tran Duong Anh-Tai, and Vinh N. T. Pham. Soft parameters in Coulomb potential of noble atoms for nonsequential double ionization: Classical ensemble model and simulations. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000911.

Turnquist:2021:MSP

[TO21] Brian Turnquist and Mark Owkes. multiUQ: a software package for uncertainty quantification of multiphase flows. Computer Physics Communications, 268(??):??, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002009.

Torsello:2020:BMP

[Tor20] Francesco Torsello. bimEX: a Mathematica package for exact computations in 3 + 1 bimetric relativity. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303030.

Tretiak:2022:EGM

[TPCT22] Krasymyr Tretiak, Meredith Plumley, Michael Calkins, and Steven Tobias. Efficiency gains of a multi-scale integration method applied to a scale-separated model for rapidly rotating dynamos. Computer Physics Communications, 273(??): ??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003659.

Tekgul:2021:DOS

[TPK^{+}21] Bulut Tekgül, Petteri Peltonen, Heikki Kahila, Ossi Kaario, and Ville Vuorinen. DLBFoam: an open-source dynamic load balancing model for fast reacting flow simulations in Open-FOAM. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001855.

Tedeschi:2024:PRB

[TPS^{+}24] Tommaso Tedeschi, Vincenzo Eduardo Padulano, Daniele Spiga, Diego Ciangottini, Mirco Tracolli, Enric Tejedor Saavedra, Enrico Guiraud, and Massimo Biasotto. Prototyping a ROOT-based distributed analysis workflow for HL-LHC: the CMS use case. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003107.

Toriyama:2023:VPT

[TQGE23] Michael Y. Toriyama, Jiaxing Qu, Lídia C. Gomes, and Elif Ertekin. VTAnDeM: a Python toolkit for simultaneously visualizing phase stability, defect energetics, and carrier concentrations of materials. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046552300036X.

Tawfik:2022:PPP

[TR22] Sherif Abdulkader Tawfik and Salvy P. Russo. PyPhotonics: a Python package for the evaluation of luminescence properties of defects. Computer Physics Communications, 273(??): ??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003349.

Tom:2020:GRS

[TRB^{+}20a] Rithwik Tom, Timothy Rose, Imanuel Bier, Harriet O’Brien, Álvaro Vázquez-Mayagoitia, and Noa Marom. Genarris 2.0: a random structure generator for molecular crystals. Computer Physics Communications, 250(??):??, May 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300242.

Tyson:2020:NMI

[TRB20b] Jack J. Tyson, Tasmiat Rahman, and Stuart A. Boden. A new mathematical interpretation of disordered nanoscale material systems for computational modelling. Computer Physics Communications, 255(??):??, October 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301831.

Tian:2024:ACU

[TRN^{+}24] Kaili Tian, Ali Ramadhan, Marcel Nooijen, Stefan V. Pantazi, Reza Karimi, and Joseph H. Sanderson. Approaches for constraining uncertainty and degeneracy in geometry reconstruction of molecules from simulated Coulomb explosion data. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523004198.

Teunissen:2023:GMM

[TS23] Jannis Teunissen and Francesca Schiavello. Geometric multigrid method for solving Poisson’s equation on octree grids with irregular boundaries. Computer Physics Communications, 286 (??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000103.

Tilaver:2021:DLA

[TSAK21] H. Tilaver, M. Salti, O. Aydogdu, and E. E. Kangal. Deep learning approach to Hubble parameter. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520304033.

Talou:2021:FFD

[TSJ^{+}21] P. Talou, I. Stetcu, P. Jaffke, M. E. Rising, A. E. Lovell, and T. Kawano. Fission fragment decay simulations with the CGMF code. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001995.

Tavakkol:2021:ATO

[TSL21] Sasan Tavakkol, Sangyoung Son, and Patrick Lynett. Adaptive third order Adams–Bashforth time integration for extended Boussinesq equations. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001181.

Takahashi:2022:EFT

[TTM22] Toru Takahashi, Masaki Tanigawa, and Naoya Miyazawa. An enhancement of the fast time-domain boundary element method for the three-dimensional wave equation. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003416.

Tani:2024:CBP

[TV24] Laurits Tani and Christian Veelken. Comparison of Bayesian and particle swarm algorithms for hyperparameter optimisation in machine learning applications in high energy physics. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003004.

Tang:2021:FSE

[TW21a] Jian Tang and TseChun Wang. Flavour symmetry embedded — GLoBES (FaSE-GLoBES). Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000436.

Trojak:2021:IVC

[TW21b] W. Trojak and F. D. Witherden. Inline vector compression for computational physics. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302691.

Thoma:2021:IHC

[TWR21] C. Thoma, D. R. Welch, and D. V. Rose. Implicit highlycoupled single-ion Hall-MHD formulation for hybrid particlein-cell codes. Computer Physics Communications, 261(??): ??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000011.

Trojak:2022:HDF

[TWW22] W. Trojak, R. Watson, and F. D. Witherden. Hyperbolic diffusion in flux reconstruction: Optimisation through kernel fusion within tensor-product elements. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003477.

Uria-Alvarez:2024:ECO

[UÁEPGBP24] Alejandro José Uría-Álvarez, Juan José Esteve-Paredes, M. A. García-Blázquez, and Juan José Palacios. Efficient computation of optical excitations in two-dimensional materials with the Xatu code. Computer Physics Communications, 295(??): ??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003466.

Uskov:2022:MDM

[UAS22] Artem Uskov, Abdaljalel Alizzi, and Zurab Silagadze. MuMuPy: a dimuonium-matter interaction calculator. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000479.

Ullah:2024:MPM

[UD24] Arif Ullah and Pavlo O. Dral. MLQD: a package for machine learning-based quantum dissipative dynamics. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002850.

Uroic:2021:PSA

[UJ21] Tessa Uroić and Hrvoje Jasak. Parallelisation of selective algebraic multigrid for block-pressure-velocity system in Open-FOAM. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302526.

Uhlrich:2021:OAT

[UMA21] Grégoire Uhlrich, Farvah Mahmoudi, and Alexandre Arbey. C++ MARTY — Modern ARtificial Theoretical phYsicist: a C++ framework automating theoretical calculations beyond the Standard Model. Computer Physics Communications, 264 (??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552100062X.

Umeda:2022:MRC

[Ume22] Takayuki Umeda. Multicolor reordering for computing moments in particle-in-cell plasma simulations. Computer Physics Communications, 281(??):??, December 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002181.

Uranagase:2020:FCT

[UO20] Masayuki Uranagase and Shuji Ogata. FE-CLIP: a tool for the calculation of the solid–liquid interfacial free energy. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300783.

Ueda:2022:QSS

[UYS22] Hiroshi Ueda, Seiji Yunoki, and Tokuro Shimokawa. Quantum spin solver near saturation: QS^{3}. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000881.

Uhlirova:2022:CAI

[UZB22] T. Uhlírová, J. Zamastil, and J. Benda. Calculation of atomic integrals between relativistic functions by means of algebraic methods. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522002090.

Valle:2022:IFL

[VÁFG^{+}22] Nicolás Valle, Xavier Álvarez-Farré, Andrey Gorobets, Jesús Castro, Assensi Oliva, and F. Xavier Trias. On the implementation of flux limiters in algebraic frameworks. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003428.

Verdugo:2022:SDG

[VB22a] Francesc Verdugo and Santiago Badia. The software design of Gridap: a Finite Element package based on the Julia JIT compiler. Computer Physics Communications, 276(??): ??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522000595.

Vitrano:2022:DPT

[VB22b] Andrea Vitrano and Bertrand Baudouy. Double phase transition numerical modeling of superfluid helium for fixed nonuniform grids. Computer Physics Communications, 273(??): ??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003878.

Valiev:2022:CPP

[VCF22] Marat Valiev, Gennady N. Chuev, and Marina V. Fedotova. CDFTPY: a Python package for performing classical density functional theory calculations for molecular liquids. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552200056X.

Viet:2023:TPT

[VCHH23] Nguyen Tan Viet, Nguyen Thi Chuong, Vu Thi Ngoc Huyen, and Le Bin Ho. tqix.pis: a toolbox for quantum dynamics simulation of spin ensembles in Dicke basis. Computer Physics Communications, 286(??):??, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000310.

vanderHeide:2023:FSP

[vdHKB^{+}23] T. van der Heide, J. Kullgren, P. Broqvist, V. Bačić, T. Frauenheim, and B. Aradi. Fortnet, a software package for training Behler–Parrinello neural networks. Computer Physics Communications, 284(??):??, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522002995.

Vazquez-Escobar:2021:EML

[VEHCM21] Julia Vázquez-Escobar, J. M. Hernández, and Miguel Cárdenas-Montes. Estimation of machine learning model uncertainty in particle physics event classifiers. Computer Physics Communications, 268(??):??, November 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002125.

Velhinho:2023:GSD

[VFS23] José Velhinho, Elsa Fonseca, and Rogério Serôdio. General solutions to decay chain equations. Computer Physics Communications, 283(??):??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522003010.

Vigano:2021:MTE

[VGGP^{+}21] Daniele Viganò, Alberto Garcia-Garcia, José A. Pons, Clara Dehman, and Vanessa Graber. Magneto-thermal evolution of neutron stars with coupled Ohmic, Hall and ambipolar effects via accurate finite-volume simulations. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001132.

Valiante:2021:COL

[VHBK21] Elisabetta Valiante, Maritza Hernandez, Amin Barzegar, and Helmut G. Katzgraber. Computational overhead of locality reduction in binary optimization problems. Computer Physics Communications, 269(??):??, December 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002149.

Vandenhoeck:2022:DIH

[VLS22] Ray Vandenhoeck, Andrea Lani, and Johan Steelant. Development of an implicit high-order flux reconstruction solver for the Langtry–Menter Laminar-Turbulent Transition RANS model. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001278.

Vergara:2023:SSP

[VMRFC23] J. M. Vergara, M. E. Mora-Ramos, E. Flórez, and J. D. Correa. SPIN: P[S]imple [P]ython [I]pywidgets [N]otebook interface to obtain the optoelectronic properties of materials employing DFT. Computer Physics Communications, 284(??): ??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003332.

Valles-Perez:2021:UCV

[VPPQ21] David Vallés-Pérez, Susana Planelles, and Vicent Quilis. Unravelling cosmic velocity flows: a Helmholtz–Hodge decomposition algorithm for cosmological simulations. Computer Physics Communications, 263(??):??, June 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000394.

Vacalebre:2023:LRC

[VPS23] Antonino Vacalebre, Silvano Pitassi, and Ruben Specogna. Low-rank compression techniques in integral methods for eddy currents problems. Computer Physics Communications, 289(??):??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001017.

vanRoekeghem:2021:QSC

[vRCM21] Ambroise van Roekeghem, Jesús Carrete, and Natalio Mingo. Quantum self-consistent ab-initio lattice dynamics. Computer Physics Communications, 263(??):??, June 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521000710.

Vetter:2024:PSP

[VRI24] Roman Vetter, Steve V. M. Runser, and Dagmar Iber. PolyHoop: Soft particle and tissue dynamics with topological transitions. Computer Physics Communications, 299(??): ??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000511.

Viola:2022:FGG

[VSM^{+}22] Francesco Viola, Vamsi Spandan, Valentina Meschini, Joshua Romero, Massimiliano Fatica, Marco D. de Tullio, and Roberto Verzicco. FSEI-GPU: GPU accelerated simulations of the fluidstructure-electrophysiology interaction in the left heart. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100360X.

Voltolina:2021:OCB

[VTB^{+}21] Dimitri Voltolina, Riccardo Torchio, Paolo Bettini, Ruben Specogna, and Piergiorgio Alotto. Optimized cycle basis in volume integral formulations for large scale eddy-current problems. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521001168.

vonToussaint:2021:FSW

[vTDGCR21] Udo von Toussaint, F. J. Domínguez-Gutiérrez, Michele Compostella, and Markus Rampp. FaVAD: a software workflow for characterization and visualizing of defects in crystalline structures. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520304094.

Verma:2021:SOB

[VV21] Abhishek Kumar Verma and Ayyaswamy Venkattraman. SOMAFOAM: an OpenFOAM based solver for continuum simulations of low-temperature plasmas. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000229.

Vogiatzis:2020:FEC

[VvBTH20] Georgios G. Vogiatzis, Lambèrt C. A. van Breemen, Doros N. Theodorou, and Markus Hütter. Free energy calculations by molecular simulations of deformed polymer glasses. Computer Physics Communications, 249(??):??, April 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303431.

Vance:2023:CMS

[VXT^{+}23] James Vance, Zhen-Hao Xu, Nikita Tretyakov, Torsten Stuehn, Markus Rampp, Sebastian Eibl, Christoph Junghans, and André Brinkmann. Code modernization strategies for shortrange non-bonded molecular dynamics simulations. Computer Physics Communications, 290(??):??, September 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001054.

VandeVijver:2020:KAS

[VZ20] Ruben Van de Vijver and Judit Zádor. KinBot: Automated stationary point search on potential energy surfaces. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302978.

Watanabe:2021:LSF

[WA21] Seiya Watanabe and Takayuki Aoki. Large-scale flow simulations using lattice Boltzmann method with AMR following free-surface on multiple GPUs. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000291.

Wen:2022:KFD

[WAET22] Mingjian Wen, Yaser Afshar, Ryan S. Elliott, and Ellad B. Tadmor. KLIFF: a framework to develop physicsbased and machine learning interatomic potentials. Computer Physics Communications, 272(??):??, March 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003301.

Welker:2022:CPF

[WAK22] Simon Welker, Muhamed Amin, and Jochen Küpper. CMInject: Python framework for the numerical simulation of nanoparticle injection pipelines. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002502.

Walker:2022:POP

[WAN^{+}22] S. D. Walker, A. Abramov, L. J. Nevay, W. Shields, and S. T. Boogert. Pyg4ometry: a Python library for the creation of Monte Carlo radiation transport physical geometries. Computer Physics Communications, 272(??):??, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003404.

Wu:2023:MAE

[WAYL23] Tong Wu, Denis Arrivault, Wei Yan, and Philippe Lalanne. Modal analysis of electromagnetic resonators: User guide for the MAN program. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522003460.

Walton:2024:MPI

[WBF^{+}24] Noah Walton, Jesse Brown, William Fritsch, Dave Brown, Gustavo Nobre, and Vladimir Sobes. Methodology for physicsinformed generation of synthetic neutron time-of-flight measurement data. Computer Physics Communications, 294(??): ??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002722.

Wanzenbock:2024:CFS

[WBK^{+}24] Ralf Wanzenböck, Florian Buchner, Péter Kovács, Georg K. H. Madsen, and Jesús Carrete. Clinamen2: Functionalstyle evolutionary optimization in Python for atomistic structure searches. Computer Physics Communications, 297(??): ??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523004101.

Wu:2024:NPG

[WBM^{+}24] Zihao Wu, Janko Boehm, Rourou Ma, Hefeng Xu, and Yang Zhang. NeatIBP 1.0 , a package generating small-size integration-by-parts relations for Feynman integrals. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003442.

Wragg:2020:BPM

[WBvdH20] Jack Wragg, Connor Ballance, and Hugo van der Hart. Breit–Pauli R-matrix approach for the time-dependent investigation of ultrafast processes. Computer Physics Communications, 254(??):??, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465520300904.

Wang:2024:AMM

[WDMZ24] J. Wang, J. M. Duan, Z. W. Ma, and W. Zhang. An adaptive moving mesh finite difference scheme for tokamak magneto-hydrodynamic simulations. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002965.

Waide:2020:BPS

[WGG20] D. T. Waide, D. G. Green, and G. F. Gribakin. BSHF: a program to solve the Hartree–Fock equations for arbitrary central potentials using a B-spline basis. Computer Physics Communications, 250(??):??, May 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519304187.

Wang:2022:PMA

[WGGC22] Hao Wang, Y. C. Gan, Hua Y. Geng, and Xiang-Rong Chen. MyElas: an automatized tool-kit for high-throughput calculation, post-processing and visualization of elasticity and related properties of solids. Computer Physics Communications, 281(??):??, December 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002144.

Weismantel:2022:EGS

[WGS^{+}22] Oliver Weismantel, Aikaterini A. Galata, Morteza Sadeghi, Achim Kröger, and Martin Kröger. Efficient generation of self-avoiding, semiflexible rotational isomeric chain ensembles in bulk, in confined geometries, and on surfaces. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002885.

Wiesenberger:2023:FVF

[WH23] Matthias Wiesenberger and Markus Held. A finite volume flux coordinate independent approach. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001832.

Wais:2021:NST

[WHB21] M. Wais, K. Held, and M. Battiato. Numerical solver for the time-dependent far-from-equilibrium Boltzmann equation. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000321.

Wu:2024:VQE

[WHSG24] Guojian Wu, Dejian Huang, Feng Shuang, and Fang Gao. Variational quantum eigenvalue solver algorithm utilizing bridgeinspired quantum circuits and a gradient filter module. Computer Physics Communications, 298(??):??, May 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000407.

Westerbeek:2024:DDH

[WHSK24] S. Westerbeek, S. Hulshoff, H. Schuttelaars, and M. Kotsonis. DeHNSSo: the Delft Harmonic Navier–Stokes Solver for nonlinear stability problems with complex geometric features. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001735.

Wang:2022:TSF

[WICA22] Qing Wang, Matthias Ihme, Yi-Fan Chen, and John Anderson. A TensorFlow simulation framework for scientific computing of fluid flows on tensor processing units. Computer Physics Communications, 274(??):??, May 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000108.

Wadgaonkar:2021:NSF

[WJB21] I. Wadgaonkar, R. Jain, and M. Battiato. Numerical scheme for the far-out-of-equilibrium time-dependent Boltzmann collision operator: 1D second-degree momentum discretisation and adaptive time stepping. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000254.

Wang:2021:CDV

[WKBW21] Yuxiang Wang, Alper Kiziltas, Patrick Blanchard, and Tiffany R. Walsh. Calculation of 1D and 2D densities in VMD: a flexible and easy-to-use code. Computer Physics Communications, 266(??):??, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521001442.

Windisch:2023:RTD

[WKJB23] Dominic Windisch, Jeffrey Kelling, Guido Juckeland, and André Bieberle. Real-time data processing for ultrafast X-ray computed tomography using modular CUDA based pipelines. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000644.

Wilkie:2023:MBC

[WKR23] George J. Wilkie, Torsten Keßler, and Sergej Rjasanow. Multidisciplinary benchmarks of a conservative spectral solver for the nonlinear Boltzmann equation. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001571.

Wei:2021:NIM

[WLCF21] Yi-Wen Wei, Chao-Kai Li, Yuchuang Cao, and Ji Feng. A non-iterative method for vertex corrections of the Kubo formula for electric conductivity. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302642.

Wu:2023:WCS

[WLF^{+}23] Wentao Wu, Jian Liu, Nathaniel J. Fisch, Jianyuan Xiao, Huishan Cai, Zhaoyuan Liu, Ruili Zhang, and Yang He. Weakly convergent stochastic simulation of electron collisions in plasmas. Computer Physics Communications, 289(??):??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523001030.

Wang:2024:ISI

[WLF^{+}24] Jun Wang, Tao Li, Ziwen Fu, Baoqin Fu, and Chengjun Gou. Identification of self-interstitial atoms and vacancies in crystalline materials in atomistic simulation. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000614.

Wei:2022:NMC

[WLL22] Linfang Wei, Deokjung Lee, and Hyun Chul Lee. A novel S_{N} /MOC collaborative method for few-group crosssection generation with versatile neutron spectral adaptability in advanced reactor physics analysis. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000583.

Wang:2024:PMR

[WLZ^{+}24] Yunlan Wang, Yufeng Liu, Rui Zhang, Tianhai Zhao, Sha Liu, Congshan Zhuo, and Chengwen Zhong. Performance modeling and running strategy of parallel cdugksFOAM program. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524001097.

Wiecha:2022:PNF

[WMA^{+}22] Peter R. Wiecha, Clément Majorel, Arnaud Arbouet, Adelin Patoux, Yoann Brûlé, Gérard Colas des Francs, and Christian Girard. “pyGDM” — new functionalities and major improvements to the Python toolkit for nano-optics full-field simulations. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552100254X.

Weichman:2024:APT

[WMM^{+}24] K. Weichman, K. G. Miller, B. Malaca, W. B. Mori, J. R. Pierce, D. Ramsey, J. Vieira, M. Vranic, and J. P. Palastro. Analytic pulse technique for computational electromagnetics. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524000195.

Woo:2021:OSM

[WNS^{+}21] Mino Woo, Robert T. Nishida, Mario A. Schriefl, Marc E. J. Stettler, and Adam M. Boies. Open-source modelling of aerosol dynamics and computational fluid dynamics: Nodal method for nucleation, coagulation, and surface growth. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303829.

Weinhart:2020:FFP

[WOP^{+}20] Thomas Weinhart, Luca Orefice, Mitchel Post, Marnix P. van Schrojenstein Lantman, Irana F. C. Denissen, Deepak R. Tunuguntla, J. M. F. Tsang, Hongyang Cheng, Mohamad Yousef Shaheen, Hao Shi, Paolo Rapino, Elena Grannonio, Nunzio Losacco, Joao Barbosa, Lu Jing, Juan E. Alvarez Naranjo, Sudeshna Roy, Wouter K. den Otter, and Anthony R. Thornton. Fast, flexible particle simulations — an introduction to MercuryDPM. Computer Physics Communications, 249(??): ??, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519304357.

Worth:2020:QGP

[Wor20] G. A. Worth. Quantics: a general purpose package for quantum molecular dynamics simulations. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551930373X.

Wieckowski:2021:IME

[WPMK21] Andrzej Wieckowski, Andrzej Ptok, Marcin Mierzejewski, and Michał Kupczyński. Identification of the Majorana edge modes in tight-binding systems based on the Krylov method. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002472.

Wright:2024:DPP

[WRM^{+}24] Steven A. Wright, Christopher P. Ridgers, Gihan R. Mudalige, Zaman Lantra, Josh Williams, Andrew Sunderland, H. Sue Thorne, and Wayne Arter. Developing performance portable plasma edge simulations: a survey. Computer Physics Communications, 298(??):??, May 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000468.

Willen:2020:RPS

[WS20] Daniel P. Willen and Adam J. Sierakowski. Resolved particle simulations using the Physalis method on many GPUs. Computer Physics Communications, 250(??):??, May 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551930390X.

Wu:2023:MFE

[WS23] Yong-Chao Wu and Jian-Li Shao. mdapy: a flexible and efficient analysis software for molecular dynamics simulations. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523001091.

Woo:2022:OSM

[WSK^{+}22] Mino Woo, Mario A. Schriefl, Markus Knoll, Adam M. Boies, Marc E. J. Stettler, Simone Hochgreb, and Robert T. Nishida. Open-source modelling of aerosol dynamics and computational fluid dynamics: Bipolar and unipolar diffusion charging and photoelectric charging. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522001187.

Whyte:2024:OSS

[WSRO24] Travis Whyte, Andreas Stathopoulos, Eloy Romero, and Kostas Orginos. Optimizing shift selection in multilevel Monte Carlo for disconnected diagrams in lattice QCD. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002734.

Wang:2022:SAI

[WT22] Weiyi Wang and Zhijun Tan. A simple augmented IIM for 3D incompressible two-phase Stokes flows with interfaces and singular forces. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002666.

Wittreich:2022:PGA

[WV22] Gerhard R. Wittreich and Dionisios G. Vlachos. Python Group Additivity (pGrAdd) software for estimating species thermochemical properties. Computer Physics Communications, 273(??):??, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521003891.

Wilson:2021:GAF

[WVK21] Leighton Wilson, Nathan Vaughn, and Robert Krasny. A GPU-accelerated fast multipole method based on barycentric Lagrange interpolation and dual tree traversal. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001296.

Walden:2021:NEI

[WW21] Moritz Walden and Stefan Weinzierl. Numerical evaluation of iterated integrals related to elliptic Feynman integrals. Computer Physics Communications, 265(??):??, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521001326.

Wadgaonkar:2022:NSE

[WWB22] Indrajit Wadgaonkar, Michael Wais, and Marco Battiato. Numerical solver for the out-of-equilibrium time dependent Boltzmann collision operator: Application to 2D materials. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521003192.

Willsch:2020:SVM

[WWDM20] D. Willsch, M. Willsch, H. De Raedt, and K. Michielsen. Support vector machines on the D-Wave quantum annealer. Computer Physics Communications, 248(??):??, March 2020.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551930342X.

Willsch:2022:GAS

[WWJ^{+}22] Dennis Willsch, Madita Willsch, Fengping Jin, Kristel Michielsen, and Hans De Raedt. GPU-accelerated simulations of quantum annealing and the quantum approximate optimization algorithm. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522001308.

Wang:2022:GAP

[WWM^{+}22] Bin Wang, Ingo Wald, Nate Morrical, Will Usher, Lin Mu, Karsten Thompson, and Richard Hughes. An GPUaccelerated particle tracking method for Eulerian–Lagrangian simulations using hardware ray tracing cores. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003337.

Wang:2023:FGH

[WWZ^{+}23] Yujie Wang, Shengquan Wang, Xuerui Zhang, Guangyao Li, and Yong Cai. Fine-grained heterogeneous parallel direct solver for finite element problems. Computer Physics Communications, 284(??):??, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522003563.

Wang:2021:VUF

[WXL^{+}21] Vei Wang, Nan Xu, Jin-Cheng Liu, Gang Tang, and Wen-Tong Geng. VASPKIT: a user-friendly interface facilitating high-throughput computing and analysis using VASP code. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001454.

Wang:2021:SLT

[WXW^{+}21] Y. M. Wang, X. Q. Xu, Z. Wang, N. M. Li, X. D. Yang, Z. Sun, T. Y. Xia, L. Zhang, Z. H. Wang, and X. Gao. Simulation of lithium transport using the BOUT++ framework. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521002393.

Wu:2024:HFH

[WXX24] Mingyang Wu, Andong Xu, and Chijie Xiao. HYPIC: a fast hybrid EM PIC-MCC code for ion cyclotron resonance energization in cylindrical coordinate system. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001309.

Wang:2020:SHO

[WXY20] Pengde Wang, Zhiguo Xu, and Jia Yin. Simple high-order boundary conditions for computing rogue waves in the nonlinear Schrödinger equation. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519304163.

Wei:2024:BSC

[WXZH24] Qi Wei, Jiawei Xiang, Weiping Zhu, and Hongjiu Hu. Band structure calculations of three-dimensional solid-fluid coupling phononic crystals using dual reciprocity boundary element method and wavelet compression method. Computer Physics Communications, 299(??):??, June 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000882.

Wang:2024:HSO

[WYHW24] Peilin Wang, Kuangkuang Ye, Xuerui Hao, and Jike Wang. Handling shape optimization of superconducting cavities with DNMOGA. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524000596.

Wu:2023:HEV

[WYT23] Yanyao Wu, Junxiang Yang, and Zhijun Tan. A highly efficient variant of scalar auxiliary variable (SAV) approach for the phase-field fluid-surfactant model. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523002059.

Wu:2023:SSU

[WYZZ23] Zhetong Wu, Hongfei Ye, Hongwu Zhang, and Yonggang Zheng. Seq-SVF: an unsupervised data-driven method for automatically identifying hidden governing equations. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002321.

Wang:2021:FMM

[WZC21] Bo Wang, Wenzhong Zhang, and Wei Cai. Fast multipole method for 3-D Laplace equation in layered media. Computer Physics Communications, 259(??):??, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303131.

Wu:2024:STN

[WZC^{+}24] Zheyu Wu, Remmy Zen, Heitor P. Casagrande, Dario Poletti, and Stéphane Bressan. Supervised training of neuralnetwork quantum states for the next-nearest neighbor Ising model. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524000924.

Wojenski:2024:HAC

[WZK^{+}24] A. Wojenski, H. Zbroszczyk, M. Kruszewski, P. Szymanski, E. Wawrzyn, D. Wielanek, W. Zabolotny, D. Pawlowska, and T. Gniazdowski. Hardware acceleration of complex HEP algorithms with HLS and FPGAs: Methodology and preliminary implementation. Computer Physics Communications, 295(??): ??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003429.

Wu:2020:ZPM

[WZPW20] Hong-Zhong Wu, Jun-Jie Zhang, Long-Gang Pang, and Qun Wang. ZMCintegral: a package for multi-dimensional Monte Carlo integration on multi-GPUs. Computer Physics Communications, 248(??):??, March 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519303121.

Wang:2023:PPP

[WZZ^{+}23] Hongjin Wang, Jingyi Zhuang, Zhen Zhang, Qi Zhang, and Renata M. Wentzcovitch. pgm: a Python package for free energy calculations within the phonon gas model. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300190X.

Wang:2021:FEM

[WZZS21] Liqun Wang, Meiling Zhao, Yifan Zhang, and Liwei Shi. A finite element method for the band structure computation of photonic crystals with complex scatterer geometry. Computer Physics Communications, 263(??):??, June 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100028X.

Xavier:2022:MCM

[Xav22] George D. Xavier. M-CHIPR: a Mathematica program for constructing multi-state coupled adiabatic potential energy functions in triatomic molecule using many body partitioning approach. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001382.

Xia:2020:GAP

[XBL^{+}20] Yidong Xia, Ansel Blumers, Zhen Li, Lixiang Luo, Yu-Hang Tang, Joshua Kane, Jan Goral, Hai Huang, Milind Deo, and Matthew Andrew. A GPU-accelerated package for simulation of flow in nanoporous source rocks with many-body dissipative particle dynamics. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302619.

Xiao:2020:PPC

[XDF20] Xufeng Xiao, Zihuan Dai, and Xinlong Feng. A positivity preserving characteristic finite element method for solving the transport and convection–diffusion–reaction equations on general surfaces. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519302929.

Xu:2024:MVV

[XFGS24] Zhiqian Xu, Yi Fan, Chu Guo, and Honghui Shang. MPS-VQE: a variational quantum computational chemistry simulator with matrix product states. Computer Physics Communications, 294(??):??, January 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523002424.

Xie:2022:UDS

[XFH^{+}22] Guoda Xie, Ming Fang, Zhixiang Huang, Xingang Ren, and Xianliang Wu. A unified 3-D simulating framework for Debye-type dispersive media and PML technique based on recursive integral method. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522001825.

Xiong:2024:GSH

[XHY^{+}24] Qiyang Xiong, Shiyong Huang, Zhigang Yuan, Bharatkumar Sharma, Lvlin Kuang, Kui Jiang, and Lin Yu. GPIC: a set of high-efficiency CUDA Fortran code using GPU for particle-in-cell simulation in space physics. Computer Physics Communications, 295(??):??, February 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003399.

Xie:2023:SFA

[Xie23] Huasheng Xie. A simple and fast approach for computing the fusion reactivities with arbitrary ion velocity distributions. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523002072.

Xue:2023:JFD

[XLG^{+}23] Tianju Xue, Shuheng Liao, Zhengtao Gan, Chanwook Park, Xiaoyu Xie, Wing Kam Liu, and Jian Cao. JAX-FEM: a differentiable GPU-accelerated 3D finite element solver for automatic inverse design and mechanistic data science. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001479.

Xu:2021:ESP

[XLKX21] Chudong Xu, Shengdong Lu, Yongfeng Kong, and Wanjie Xiong. The enhanced sampling in parallel finitetime dynamics method with replica exchange. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000497.

Xiao:2022:HFM

[XLL^{+}22] Wei Xiao, Xiangyue Li, Peijie Li, Tengfei Zhang, and Xiaojing Liu. High-fidelity multi-physics coupling study on advanced heat pipe reactor. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002642.

Xu:2022:OIC

[XOTI22] RuQing G. Xu, Tsuyoshi Okubo, Synge Todo, and Masatoshi Imada. Optimized implementation for calculation and fastupdate of Pfaffians installed to the open-source fermionic variational solver mVMC. Computer Physics Communications, 277(??):??, August 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000947.

Xie:2024:ASF

[XPF^{+}24] Guoda Xie, Pan Pan, Naixing Feng, Ke Xu, Ming Fang, Xianliang Wu, Yingsong Li, and Zhixiang Huang. Application of spatial filtering high order FDTD method in the multi-physics field coupled equations. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001243.

Xiao:2021:SMC

[XQ21] Jianyuan Xiao and Hong Qin. Slow manifolds of classical Pauli particle enable structure-preserving geometric algorithms for guiding center dynamics. Computer Physics Communications, 265(??):??, August 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552100093X.

Xiao:2022:ASC

[XSL^{+}22] Wei Xiao, Qizheng Sun, Xiaojing Liu, Hui He, Donghao He, Qingquan Pan, and Tengfei Zhang. Application of stiffness confinement method within variational nodal method for solving time-dependent neutron transport equation. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001692.

Xie:2022:BFC

[XSM22] Yaoshu Xie, Kiyou Shibata, and Teruyasu Mizoguchi. A brute-force code searching for cell of non-identical displacement for CSL grain boundaries and interfaces. Computer Physics Communications, 273(??):??, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003726.

Xiang:2021:PSS

[XWYQ21] Shaohui Xiang, Liming Wang, Zong-Chao Yan, and Haoxue Qiao. A program for simplifying summation of Wigner 3jsymbols. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521000333.

Xiong:2020:GES

[XZLX20] Mo Xiong, Xiujian Zhao, Neng Li, and Hong Xu. General energy-strain scheme for accurate evaluation of the Born elasticity term for solid and liquid systems under finite temperature and pressure conditions. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303108.

Yang:2021:BCM

[Yan21a] W. M. Yang. On the boundary conditions of magnetic field in OpenFOAM and a magnetic field solver for multi-region applications. Computer Physics Communications, 263(??): ??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000357.

Yang:2021:NFD

[Yan21b] Xiaofeng Yang. On a novel fully-decoupled, linear and secondorder accurate numerical scheme for the Cahn–Hilliard– Darcy system of two-phase Hele–Shaw flow. Computer Physics Communications, 263(??):??, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000278.

Yang:2023:LFW

[Yan23] Pengliang Yang. libEMM: a fictious wave domain 3D CSEM modelling library bridging sequential and parallel GPU implementation. Computer Physics Communications, 288(??): ??, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000905.

Yang:2024:UES

[Yan24a] Junxiang Yang. Unconditionally energy-stable linear convex splitting algorithm for the L^{2} quasicrystals. Computer Physics Communications, 295(??):??, February 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003296.

Yang:2024:SIT

[Yan24b] Pengliang Yang. SMIwiz: an integrated toolbox for multidimensional seismic modelling and imaging. Computer Physics Communications, 295(??):??, February 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523003569.

Ye:2020:RSM

[YC20] Lei Ye and Yang Chen. Re-splitting δf method for electromagnetic gyrokinetic particle-in-cell (PIC) simulation of tokamak plasmas. Computer Physics Communications, 250(??): ??, May 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303790.

Yan:2022:LES

[YCC22] Zhengzheng Yan, Rongliang Chen, and Xiao-Chuan Cai. Large eddy simulation of the wind flow in a realistic full-scale urban community with a scalable parallel algorithm. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002824.

Yang:2023:IDA

[YCCW23] Guang Yang, Yu Chen, Simeng Chen, and Moran Wang. Implementation of a direct-addressing based lattice Boltzmann GPU solver for multiphase flow in porous media. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552300173X.

Yu:2020:IDC

[YD20] Xiang Yu and Meenakshi Dutt. Implementation of dynamic coupling in hybrid Molecular Dynamics–Lattice Boltzmann approach: Modeling aggregation of amphiphiles. Computer Physics Communications, 257(??):??, December 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300977.

Yevick:2021:CLS

[Yev21] David Yevick. Conservation laws and spin system modeling through principal component analysis. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521000060.

Yang:2022:COB

[YFL22] Wenming Yang, Boshi Fang, and Beiying Liu. coronaChargingFoam: an OpenFOAM based solver for multi-physical simulations of direct unipolar diffusion charging of aerosol particles. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001540.

Younkin:2021:GAG

[YGSW21] T. R. Younkin, D. L. Green, A. B. Simpson, and B. D. Wirth. GITR: an accelerated global scale particle tracking code for wall material erosion and redistribution in fusion relevant plasmamaterial interactions. Computer Physics Communications, 264 (??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000369.

Yamoto:2020:IFO

[YHH^{+}20] Shohei Yamoto, Yuki Homma, Kazuo Hoshino, Mitsunori Toma, and Akiyoshi Hatayama. IMPGYRO: the full-orbit impurity transport code for SOL/divertor and its successful application to tungsten impurities. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303212.

Yeu:2021:III

[YHY^{+}21] In Won Yeu, Gyuseung Han, Kun Hee Ye, Cheol Seong Hwang, and Jung-Hae Choi. InterPhon: Ab initio interface phonon calculations within a 3D electronic structure framework. Computer Physics Communications, 268(??):??, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521002010.

Yamaguchi:2022:FPL

[YI22] Naoya Yamaguchi and Fumiyuki Ishii. First-principles LCPAO approach for insulators under finite electric fields with forces. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002065.

Yu:2024:SSC

[YJ24] Bo Yu and Ruijiang Jing. SCTBEM: a scaled coordinate transformation boundary element method with 99line MATLAB code for solving Poisson’s equation. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001085.

Ye:2021:LSF

[YJLW21] Meng Ye, Xiangwei Jiang, Shu-Shen Li, and Lin-Wang Wang. Large-scale first-principles quantum transport simulations using plane wave basis set on high performance computing platforms. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520303647.

Yang:2021:VSS

[YK21a] Junxiang Yang and Junseok Kim. A variant of stabilizedscalar auxiliary variable (S-SAV) approach for a modified phase-field surfactant model. Computer Physics Communications, 261(??):??, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521000023.

You:2021:ABT

[YK21b] Hojun You and Chongam Kim. Architecture-based and target-oriented algorithm optimization of high-order methods via complete-search tensor contraction. Computer Physics Communications, 264(??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521001004.

Yang:2023:PMG

[YKK^{+}23a] Mingyu Yang, Ji-Hoon Kang, Ki-Ha Kim, Oh-Kyoung Kwon, and Jung-Il Choi. PaScaL TDMA 2.0: a multi-GPU-based library for solving massive tridiagonal systems. Computer Physics Communications, 290(??):??, September 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001303.

You:2023:DOS

[YKK23b] Hojun You, Juhyun Kim, and Chongam Kim. Deneb: an open-source high-performance multi-physical flow solver based on high-order DRM-DG method. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000176.

Yoon:2020:JOS

[YKSH20] Hongkee Yoon, Taek Jung Kim, Jae-Hoon Sim, and Myung Joon Han. Jx: an open-source software for calculating magnetic interactions based on magnetic force theory. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302991.

Yu:2022:SPS

[YKY^{+}22] Kwangmin Yu, Prabhat Kumar, Shaohua Yuan, Aiqi Cheng, and Roman Samulyak. SPACE: 3D parallel solvers for Vlasov–Maxwell and Vlasov–Poisson equations for relativistic plasmas with atomic transformations. Computer Physics Communications, 277(??):??, August 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001151.

You:2023:IAN

[YKYK23] Hojun You, Juhyun Kim, Kyeol Yune, and Chongam Kim. 3IDEA: Artificial neural network models for 11-species air properties at thermochemical equilibrium. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523001339.

Youn:2020:AFA

[YLH^{+}20] Yong Youn, Miso Lee, Changho Hong, Doyeon Kim, Sangtae Kim, Jisu Jung, Kanghoon Yim, and Seungwu Han. AMP^{2}: a fully automated program for ab initio calculations of crystalline materials. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302113.

Yamauchi:2022:IRP

[YLIO22] Masataka Yamauchi, Giovanni La Penna, Satoru G. Itoh, and Hisashi Okumura. Implementations of replica-permutation and replica sub-permutation methods into LAMMPS. Computer Physics Communications, 276(??):??, July 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522000819.

Yu:2020:CFF

[YLK^{+}20] Jiankai Yu, Hyunsuk Lee, Hanjoo Kim, Peng Zhang, and Deokjung Lee. Coupling of FRAPCON for fuel performance analysis in the Monte Carlo code MCS. Computer Physics Communications, 251(??):??, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519300712.

Yeo:2021:GAE

[YLK21] Beomki Yeo, MyeongJae Lee, and Yoshitaka Kuno. GPUaccelerated event reconstruction for the COMET Phase-I experiment. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552030285X.

Yoon:2022:IEB

[YLL^{+}22] H. G. Yoon, C. Lee, D. B. Lee, S. M. Park, J. W. Choi, H. Y. Kwon, and C. Won. Interpolation and extrapolation between the magnetic chiral states using autoencoder. Computer Physics Communications, 272(??):??, March 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003568.

Yang:2024:FQG

[YLLW24] Guang Yang, Tong Liu, Xukang Lu, and Moran Wang. Fast-QSGS: a GPU accelerated program for structure generation of granular disordered media. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524001644.

Yevick:2021:ARB

[YM21] David Yevick and Roger Melko. The accuracy of restricted Boltzmann machine models of Ising systems. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302435.

Yoon:2023:PFO

[YMCF23] Tae Jun Yoon, Katie A. Maerzke, Robert P. Currier, and Alp T. Findikoglu. PyOECP: a flexible open-source software library for estimating and modeling the complex permittivity based on the open-ended coaxial probe (OECP) technique. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002363.

Yenusah:2024:PPP

[YML^{+}24] Caleb O. Yenusah, Nathaniel R. Morgan, Ricardo A. Lebensohn, Miroslav Zecevic, and Marko Knezevic. A parallel and performance portable implementation of a full-field crystal plasticity model. Computer Physics Communications, 300(??): ??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001139.

Yousaf:2024:CSR

[YNMR24] Noman Yousaf, Rubina Nasir, Nasir M. Mirza, and Saima Rafique. A computational solution of relativistic Thomas– Fermi equation of an atom with exponential collocation genetic algorithm optimization. Computer Physics Communications, 297(??):??, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523003934.

Yalameha:2022:TTA

[YNV22] Shahram Yalameha, Zahra Nourbakhsh, and Daryoosh Vashaee. ElATools: a tool for analyzing anisotropic elastic properties of the 2D and 3D materials. Computer Physics Communications, 271(??):??, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521003076. See erratum [YNV23].

Yalameha:2023:ETT

[YNV23] Shahram Yalameha, Zahra Nourbakhsh, and Daryoosh Vashaee. Erratum: “ElATools: a tool for analyzing anisotropic elastic properties of the 2D and 3D materials” [comput. phys. commun. 271 (2022) 108195]. Computer Physics Communications, 288(??):??, July 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000735. See [YNV22].

Yoon:2021:MMO

[YR21] Gwanho Yoon and Junsuk Rho. MAXIM: Metasurfaces-oriented electromagnetic wave simulation software with intuitive graphical user interfaces. Computer Physics Communications, 264 (??):??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000163.

Yamashita:2022:PCM

[YS22] Takumi Yamashita and Tetsuya Sakurai. A parallel computing method for the higher order tensor renormalization group. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001424.

Yang:2024:OPB

[YSI^{+}24] Xin Yang, Elyssa Sliheet, Reece Iriye, Daniel Reynolds, and Weihua Geng. Optimized parallelization of boundary integral Poisson–Boltzmann solvers. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524000481.

Young-S:2023:OFP

[YSMBA23] Luis E. Young-S., Paulsamy Muruganandam, Antun Balaz, and Sadhan K. Adhikari. OpenMP Fortran programs for solving the time-dependent dipolar Gross–Pitaevskii equation. Computer Physics Communications, 286(??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465523000140.

Ye:2020:OHE

[YSX^{+}20] Huilin Ye, Zhiqiang Shen, Weikang Xian, Teng Zhang, Shan Tang, and Ying Li. OpenFSI: a highly efficient and portable fluid-structure simulation package based on immersed-boundary method. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302162.

Yurdusen:2022:SPH

[YT22] Ismet Yurdusen and O. Ogulcan Tuncer. Symmetrization of the product of Hermitian operators. Computer Physics Communications, 274(??):??, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522000194.

Yalcin:2023:HFE

[YT23] M. A. Yalçin and I. Temizer. Hybrid finite element/multipole expansion method for atomic Kohn–Sham density functional theory calculations. Computer Physics Communications, 286 (??):??, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000036.

Yi:2020:QFT

[YTC^{+}20] Wencai Yi, Gang Tang, Xin Chen, Bingchao Yang, and Xiaobing Liu. qvasp: a flexible toolkit for VASP users in materials simulations. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046552030254X.

Yang:2021:HMC

[YW21] Sheng-Chun Yang and Yong-Lei Wang. A hybrid MPI-CUDA approach for nonequispaced discrete Fourier transformation. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302393.

Yang:2023:AEF

[YWM23] Wei Yang, Tiancheng Wang, and Jiangqiong Mao. Adaptive edge finite element method and numerical design for metasurface cloak. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523002035.

Yang:2023:EIC

[YWTK23] Junxiang Yang, Jian Wang, Zhijun Tan, and Junseok Kim. Efficient IMEX and consistently energy-stable methods of diffuse-interface models for incompressible three-component flows. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002776.

Yu:2023:GEC

[YWX^{+}23] Shimin Yu, Hao Wu, Jingwen Xu, Yu Wang, Jiamao Gao, Zhijiang Wang, Wei Jiang, and Ya Zhang. A generalized external circuit model for electrostatic particle-in-cell simulations. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522001874.

Yang:2024:CCD

[YX24] Debo Yang and Hailing Xiong. Code for the cluster dichotomy model. Computer Physics Communications, 296(??): ??, March 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523003752.

Yin:2021:CBT

[YXX^{+}21] Junhui Yin, Li Xu, Peng Xie, Lan Zhu, Shucheng Huang, Hangxin Liu, Zhonghai Yang, and Bin Li. A curved boundary treatment for discontinuous Galerkin method applied to Euler equations on triangular and tetrahedral grids. Computer Physics Communications, 258(??):??, January 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520302630.

Yang:2023:GPC

[YYC^{+}23] Yuheng Yang, Lei Ye, Yang Chen, Nong Xiang, and Youwen Sun. Gyrokinetic particle-in-cell simulation of the effect of compressional magnetic perturbations on the microscopic instabilities in tokamak. Computer Physics Communications, 292(??):??, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523002370.

Yang:2021:VIT

[YYH21] Haijian Yang, Chao Yang, and Jizu Huang. Variational inequality transport model on the sphere by the active-set reduced-space algorithm. Computer Physics Communications, 260(??):??, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520303659.

Yang:2022:MFS

[YYY22] Nian Yang, Haijian Yang, and Chao Yang. Multilevel field-split preconditioners with domain decomposition for steady and unsteady flow problems. Computer Physics Communications, 280(??):??, November 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002156.

Yan:2022:DLL

[YYZ^{+}22] Keding Yan, Shuwei Yang, Yingran Zhao, Chao Ma, Ying Jin, and Shouyu Wang. Deep learning for light scattering computation: Reconstructing light scattering fields from 1-d randomly rough surfaces as an example. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002952.

Yao:2020:AAA

[YZ20a] B. N. Yao and R. F. Zhang. AADIS: an atomistic analyzer for dislocation character and distribution. Computer Physics Communications, 247(??):??, February 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302346.

Yuan:2020:CIS

[YZ20b] Ruifeng Yuan and Chengwen Zhong. A conservative implicit scheme for steady state solutions of diatomic gas flow in all flow regimes. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519303182.

Yin:2022:QTI

[YZHL22] Han Yin, Tengfei Zhang, Donghao He, and Xiaojing Liu. A quasi-transport integral variational nodal method for homogeneous nodes based on the 2D/1D method. Computer Physics Communications, 274(??):??, May 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200008X.

Yan:2023:RCD

[YZL^{+}23] Yu-Liang Yan, Dai-Mei Zhou, An-Ke Lei, Xiao-Mei Li, Xiao-Ming Zhang, Liang Zheng, Gang Chen, Xu Cai, and Ben-Hao Sa. Revisiting the centrality definition and observable centrality dependence of relativistic heavy-ion collisions in PA-CIAE model. Computer Physics Communications, 284(??): ??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003344.

Yang:2021:NCE

[YZW21] Hongli Yang, Xianyang Zeng, and Xinyuan Wu. A novel class of explicit divergence-free time-domain methods for efficiently solving Maxwell’s equations. Computer Physics Communications, 268(??):??, November 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465521002137.

Yang:2022:PBU

[YZW22] Zhi Yang, Xu-Chang Zheng, and Xing-Gang Wu. BEEC2.0: an upgraded version for the production of heavy quarkonium at electron–positron collider. Computer Physics Communications, 281(??):??, December 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465522002223.

Yu:2023:LSB

[YZZ^{+}23] Yongchuan Yu, Yujie Zhu, Chi Zhang, Oskar J. Haidn, and Xiangyu Hu. Level-set based pre-processing techniques for particle methods. Computer Physics Communications, 289(??):??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000899.

Zhang:2021:SBS

[ZAW^{+}21] Xiao Zhang, Sebastian Achilles, Jan Winkelmann, Roland Haas, André Schleife, and Edoardo Di Napoli. Solving the Bethe–Salpeter equation on massively parallel architectures. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001934.

Zhang:2023:PSP

[ZBS^{+}23] Chi Zhang, Guangle Bu, Md Symon Jahan Sajib, Lida Meng, Shiying Xu, Size Zheng, Lin Zhang, and Tao Wei. PXLink: a simulation program of polymer crosslinking to study of polyamide membrane. Computer Physics Communications, 291(??):??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001856.

Zhang:2022:JJE

[ZCP^{+}22] Jun-Jie Zhang, Jian-Nan Chen, Guo-Liang Peng, Tai-Jiao Du, and Hai-Yan Xie. JefiGPU: Jefimenko’s equations on GPU. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522000467.

Zhu:2024:JOS

[ZCS^{+}24] Junyan Zhu, Jiang Cao, Chen Song, Bo Li, and Zhengsheng Han. Jiezi: an open-source Python software for simulating quantum transport based on non-equilibrium Green’s function formalism. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465524001747.

Zhang:2020:ESA

[ZCWY20] Jun Zhang, Chuanjun Chen, Jiangxing Wang, and Xiaofeng Yang. Efficient, second order accurate, and unconditionally energy stable numerical scheme for a new hydrodynamics coupled binary phase-field surfactant system. Computer Physics Communications, 251(??):??, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519304321.

Zhang:2021:ESI

[ZDG^{+}21] Ang Zhang, Jinglian Du, Zhipeng Guo, Qigui Wang, and Shoumei Xiong. Evolution of specific interface area during solidification: a three-dimensional thermosolutal phase-field study. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001545.

Zhou:2021:TGT

[ZDLS21] Xingyu Zhou, Shuxian Du, Gang Li, and Chengping Shen. TopoAna: a generic tool for the event type analysis of inclusive Monte-Carlo samples in high energy physics experiments. Computer Physics Communications, 258(??):??, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520302575.

Zhang:2023:DUM

[ZDSS23] Chonglin Zhang, Gerrett Diamond, Cameron W. Smith, and Mark S. Shephard. Development of an unstructured mesh gyrokinetic particle-in-cell code for exascale fusion plasma simulations on GPUs. Computer Physics Communications, 291(??): ??, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523001698.

Zeb:2022:ELS

[Zeb22] M. Ahsan Zeb. Efficient linear scaling mapping for permutation symmetric Fock spaces. Computer Physics Communications, 276(??):??, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000650.

Zhong:2020:DLT

[ZGW20] Linlin Zhong, Qi Gu, and Bingyu Wu. Deep learning for thermal plasma simulation: Solving 1-D arc model as an example. Computer Physics Communications, 257(??):??, December 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302320.

Zhang:2023:EEH

[ZGZW23] Qi Zhang, Chaoxuan Gu, Jingyi Zhuang, and Renata M. Wentzcovitch. express: Extensible, high-level workflows for swifter ab initio materials modeling. Computer Physics Communications, 282(??):??, January 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200234X.

Zhang:2020:SSD

[Zha20] Chenguang Zhang. sdfibm: a signed distance field based discrete forcing immersed boundary method in OpenFOAM. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301594.

Zhang:2023:GAA

[ZHI23] Meng Zhang, Koki Hibi, and Junya Inoue. GPU-accelerated artificial neural network potential for molecular dynamics simulation. Computer Physics Communications, 285(??):??, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522003745.

Zhuravlev:2021:EEM

[ZHM21] V. Zhuravlev, M. Hermanowicz, and M. T. Michalewicz. EQMO: Equation of motion method for efficient electronic structure calculations. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521001764.

Zhang:2024:ONC

[ZHM^{+}24] Shidong Zhang, Steffen Hess, Holger Marschall, Uwe Reimer, Steven Beale, and Werner Lehnert. openFuelCell2: a new computational tool for fuel cells, electrolyzers, and other electrochemical devices and processes. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524000158.

Zhou:2023:JFN

[Zho23] Xiafeng Zhou. Jacobian-free Newton Krylov coarse mesh finite difference algorithm based on high-order nodal expansion method for three-dimensional nuclear reactor pin-by-pin multiphysics coupled models. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002284.

Zhong:2020:PRG

[ZHS^{+}20] Yang Zhong, Zhenpeng Hu, Tongqing Sun, Weiwei Wang, Yongfa Kong, and Jingjun Xu. Pauling’s rules guided Monte Carlo search (PAMCARS): a shortcut of predicting inorganic crystal structures. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520302289.

Zhu:2023:AAM

[ZHZ^{+}23] Yicong Zhu, Changnian Han, Peng Zhang, Guojing Cong, James R. Kozloski, Chih-Chieh Yang, Leili Zhang, and Yuefan Deng. AI-aided multiscale modeling of physiologicallysignificant blood clots. Computer Physics Communications, 287(??):??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523000632.

Zhu:2023:NMB

[ZHZG23] Guixun Zhu, Jason Hughes, Siming Zheng, and Deborah Greaves. A novel MPI-based parallel smoothed particle hydrodynamics framework with dynamic load balancing for free surface flow. Computer Physics Communications, 284(??):??, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522003277.

Zbikowski:2023:BBL

[ZJ23] Ryan M. Zbikowski and Calvin W. Johnson. Bootstrapped block Lanczos for large-dimension eigenvalue problems. Computer Physics Communications, 291(??):??, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523001807.

Zhang:2021:LNB

[ZJM^{+}21] W. Zhang, S. C. Jardin, Z. W. Ma, A. Kleiner, and H. W. Zhang. Linear and nonlinear benchmarks between the CLT code and the M3D-C1 code for the 2/1 resistive tearing mode and the 1/1 resistive kink mode. Computer Physics Communications, 269(??):??, December 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002460.

Zheng:2020:SOS

[ZJS^{+}20] Size Zheng, Leili Javidpour, Muhammad Sahimi, Katherine S. Shing, and Aiichiro Nakano. sDMD: an open source program for discontinuous molecular dynamics simulation of protein folding and aggregation. Computer Physics Communications, 247(??): ??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519302607.

Zenitani:2020:MBI

[ZK20] Seiji Zenitani and Tsunehiko N. Kato. Multiple Boris integrators for particle-in-cell simulation. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465519303091.

Zhang:2024:LCG

[ZKZ^{+}24] Ziji Zhang, Georgios Kementzidis, Peng Zhang, Leili Zhang, James Kozloski, Adam Hansen, Miriam Rafailovich, Marcia Simon, and Yuefan Deng. Learning coarse-grained force fields for fibrogenesis modeling. Computer Physics Communications, 295(??):??, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523003090.

Zia:2020:PPH

[ZL20] Haseeb Zia and Brice Lecampion. PyFrac: a planar 3D hydraulic fracture simulator. Computer Physics Communications, 255(??):??, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520301582.

Zhang:2024:VFP

[ZLL^{+}24] S. T. Zhang, X. M. Li, D. J. Liu, X. X. Li, R. J. Cheng, S. Y. Lv, Z. M. Huang, B. Qiao, Z. J. Liu, L. H. Cao, C. Y. Zheng, and X. T. He. Vlasov–Fokker–Planck–Maxwell simulations for plasmas in inertial confinement fusion. Computer Physics Communications, 294(??):??, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523002771.

Zhai:2023:PPT

[ZLLM23] Yu Zhai, You Li, Hui Li, and Frederick R. W. McCourt. Peng: a program for transport properties of low-density binary gas mixtures. Computer Physics Communications, 287(??): ??, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465523000577.

Zhang:2023:QTP

[ZLMH23] Zhao-Han Zhang, Yang Li, Yi-Jia Mao, and Feng He. QPC-TDSE: a parallel TDSE solver for atoms and small molecules in strong lasers. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001327.

Zoni:2022:HNS

[ZLS^{+}22] Edoardo Zoni, Remi Lehe, Olga Shapoval, Daniel Belkin, Neil Zaïm, Luca Fedeli, Henri Vincenti, and Jean-Luc Vay. A hybrid nodal-staggered pseudo-spectral electromagnetic particle-in-cell method with finite-order centering. Computer Physics Communications, 279(??):??, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552200176X.

Zong:2023:ADS

[ZLV23] Xue Zong, Jonathan Lym, and Dionisios G. Vlachos. Automated descriptor selection, volcano curve generation, and active site determination using the DescMAP software. Computer Physics Communications, 290(??):??, September 2023.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465523000991.

Zhang:2024:TNT

[ZLY24] Dinglei Zhang, Ying Li, and Shihui Ying. Trans-Net: a transferable pretrained neural networks based on temporal domain decomposition for solving partial differential equations. Computer Physics Communications, 299(??):??, June 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524000535.

Zacharoudiou:2023:DPH

[ZMC23] I. Zacharoudiou, J. W. S. McCullough, and P. V. Coveney. Development and performance of a HemeLB GPU code for humanscale blood flow simulation. Computer Physics Communications, 282(??):??, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522002673.

Zhou:2021:PSP

[ZPL^{+}21] Jin-Jian Zhou, Jinsoo Park, I-Te Lu, Ivan Maliyov, Xiao Tong, and Marco Bernardi. Perturbo: a software package for ab initio electron-phonon interactions, charge transport and ultrafast dynamics. Computer Physics Communications, 264(??): ??, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521000837.

Zhang:2024:HEL

[ZPL^{+}24] Zehua Zhang, Cheng Peng, Chengxiang Li, Hua Zhang, Tao Xian, and Lian-Ping Wang. A highly-efficient locally encoded boundary scheme for lattice Boltzmann method on GPU. Computer Physics Communications, 298(??):??, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465524000420.

Zhang:2020:RMD

[ZPZH20] Ya Zhang, Guang Pan, Yonghao Zhang, and Sina Haeri. A relaxed multi-direct-forcing immersed boundary-cascaded lattice Boltzmann method accelerated on GPU. Computer Physics Communications, 248(??):??, March 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519303224.

Zheng:2021:ASC

[ZRH21] Daye Zheng, Xinguo Ren, and Lixin He. Accurate stress calculations based on numerical atomic orbital bases: Implementation and benchmarks. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521001557.

Zhang:2021:SOS

[ZRZ^{+}21] Chi Zhang, Massoud Rezavand, Yujie Zhu, Yongchuan Yu, Dong Wu, Wenbin Zhang, Jianhang Wang, and Xiangyu Hu. SPHinXsys: an open-source multi-physics and multiresolution library based on smoothed particle hydrodynamics. Computer Physics Communications, 267(??):??, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465521001788.

Zhu:2021:DRL

[ZSqXY21] Ben Zhu, Haruki Seto, Xue qiao Xu, and Masatoshi Yagi. Drift reduced Landau fluid model for magnetized plasma turbulence simulations in BOUT++ framework. Computer Physics Communications, 267(??):??, October 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521001910.

Zuk:2022:OST

[ŻTR^{+}22] Paweł Jan Żuk, Bartosz Tużnik, Tadeusz Rymarz, Kamil Kwiatkowski, Marek Dudyński, Flavio C. C. Galeazzo, and Guenther C. Krieger Filho. OpenFOAM solver for thermal and chemical conversion in porous media. Computer Physics Communications, 278(??):??, September 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465522001266.

Zhang:2024:CED

[ZTV^{+}24] Ruochun Zhang, Bonaventura Tagliafierro, Colin Vanden Heuvel, Shlok Sabarwal, Luning Bakke, Yulong Yue, Xin Wei, Radu Serban, and Dan Negruţ. Chrono DEM-Engine: a discrete element method dual-GPU simulator with customizable contact forces and element shape. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046552400119X.

Zhang:2020:ZVS

[ZW20] Jun-Jie Zhang and Hong-Zhong Wu. ZMCintegral-v5: Support for integrations with the scanning of large parameter space on multi-GPUs. Computer Physics Communications, 251(??):??, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465520300722.

Zhang:2020:DGC

[ZWC^{+}20] Yuzhi Zhang, Haidi Wang, Weijie Chen, Jinzhe Zeng, Linfeng Zhang, Han Wang, and Weinan E. DP-GEN: a concurrent learning platform for the generation of reliable deep learning based potential energy models. Computer Physics Communications, 253(??):??, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046552030045X.

Zorn:2022:QPT

[ZWC22] Jacob A. Zorn, Bo Wang, and Long-Qing Chen. Q-POP-Thermo: a general-purpose thermodynamics solver for ferroelectric materials. Computer Physics Communications, 275(??):??, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465522000200.

Zhao:2024:PFA

[ZWG^{+}24] Guang Zhao, Linghui Wu, Francesco Grancagnolo, Nicola De Filippis, Mingyi Dong, and Shengsen Sun. Peak finding algorithm for cluster counting with domain adaptation. Computer Physics Communications, 300(??):??, July 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001310.

Zhang:2022:USP

[ZWP^{+}22] Qi Zhang, Yunlan Wang, Dongxin Pan, Jianfeng Chen, Sha Liu, Congshan Zhuo, and Chengwen Zhong. Unified X-space parallelization algorithm for conserved discrete unified gas kinetic scheme. Computer Physics Communications, 278(??):??, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001291.

Zhang:2022:LDP

[ZWZ^{+}22] Chenchen Zhang, MinHua Wen, Bin Zhang, James Lin, and Hong Liu. A load-decoupling parallel strategy based on shared memory architecture for DSMC to simulate near-continuum gases. Computer Physics Communications, 279(??):??, October 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465522001850.

Zhou:2023:ETS

[ZX23] Xiaolan Zhou and Chuanju Xu. Efficient time-stepping schemes for the Navier–Stokes–Nernst–Planck–Poisson equations. Computer Physics Communications, 289(??):??, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552300108X.

Zhang:2022:TTU

[ZXT22] Rui Ming Zhang, Xuefei Xu, and Donald G. Truhlar. TUMME: Tsinghua University Minnesota Master Equation program. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002526.

Zhang:2023:TTU

[ZXT23] Rui Ming Zhang, Xuefei Xu, and Donald G. Truhlar. TUMME 2023: Tsinghua University Minnesota Master Equation program. New version announcement. Computer Physics Communications, 293(??):??, December 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465523002394.

Zhi:2022:WSA

[ZXW^{+}22] Guo-Xiang Zhi, Chenchao Xu, Si-Qi Wu, Fanlong Ning, and Chao Cao. WannSymm: a symmetry analysis code for Wannier orbitals. Computer Physics Communications, 271(??):??, February 2022. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465521003088.

Yu:2020:EOI

[zYCD^{+}20] Victor Wen zhe Yu, Carmen Campos, William Dawson, Alberto García, Ville Havu, Ben Hourahine, William P. Huhn, Mathias Jacquelin, Weile Jia, Murat Keçeli, Raul Laasner, Yingzhou Li, Lin Lin, Jianfeng Lu, Jonathan Moussa, Jose E. Roman, Álvaro Vázquez-Mayagoitia, Chao Yang, and Volker Blum. ELSI — an open infrastructure for electronic structure solvers. Computer Physics Communications, 256(??):??, November 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465520302150.

Zhao:2021:DVB

[ZYG21] Zelong Zhao, Yongwei Yang, and Qingyu Gao. Development and validation of burn-up calculation code IMPC-Burnup2.0 for accelerator-driven sub-critical system. Computer Physics Communications, 261(??):??, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520301491.

Zhang:2023:MPQ

[ZYL^{+}23] Zeying Zhang, Zhi-Ming Yu, Gui-Bin Liu, Zhenye Li, Shengyuan A. Yang, and Yugui Yao. MagneticKP: a package for quickly constructing k · p models of magnetic and non-magnetic crystals. Computer Physics Communications, 290(??):??, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465523001297.

Zhang:2022:MPT

[ZYLY22] Zeying Zhang, Zhi-Ming Yu, Gui-Bin Liu, and Yugui Yao. MagneticTB: a package for tight-binding model of magnetic and non-magnetic materials. Computer Physics Communications, 270(??):??, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002654.

Yu:2021:GAE

[zYMK^{+}21] Victor Wen zhe Yu, Jonathan Moussa, Pavel Kůs, Andreas Marek, Peter Messmer, Mina Yoon, Hermann Lederer, and Volker Blum. GPU-acceleration of the ELPA2 distributed eigensolver for dense symmetric and Hermitian eigenproblems. Computer Physics Communications, 262(??):??, May 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520304021.

Zhang:2024:HMC

[ZYML24] Ningbo Zhang, Shiqiang Yan, Qingwei Ma, and Qian Li. A hybrid method combining ISPH with graph neural network for simulating free-surface flows. Computer Physics Communications, 301(??):??, August 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524001437.

Zhao:2021:TDM

[ZYX21] PengFei Zhao, Lei Ye, and Nong Xiang. Time diffusion method for gyrokinetic simulation of electrostatic turbulence with kinetic electrons. Computer Physics Communications, 269(??):??, December 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465521002265.

See corrigendum [ZYX22].

Zhao:2022:CTD

[ZYX22]  PengFei Zhao, Lei Ye, and Nong Xiang.  Corrigen-

dum to “Time diffusion method for gyrokinetic simulation of electrostatic turbulence with kinetic electrons” [Comput. Phys. Commun. 269 (2021) 108114]. Computer Physics Communications, 270(??):??, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465521002848. See [ZYX21].

Zhang:2024:DSI

[ZYXC24] Qifeng Zhang, Tong Yan, Dinghua Xu, and Yong Chen. Direct/split invariant-preserving Fourier pseudo-spectral methods for the rotation-two-component Camassa–Holm system with peakon solitons. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465524001607.

Zhao:2021:SUP

[ZZ21a] Shiwei Zhao and Jidong Zhao. SudoDEM: Unleashing the predictive power of the discrete element method on simulation for non-spherical granular particles. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303258.

Zheng:2021:MGE

[ZZ21b] Fawei Zheng and Ping Zhang. MagGene: a genetic evolution program for magnetic structure prediction. Computer Physics Communications, 259(??):??, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520303209.

Zhou:2024:USW

[ZZ24] Xiang-Ren Zhou and Li Zhang. An unconditionally-stable well-posed relativistic particle pusher. Computer Physics Communications, 302(??):??, September 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465524001863.

Zhou:2020:CSP

[ZZC20] Yingcheng Zhou, Zheng Zhao, and Daojian Cheng. Cluster structure prediction via revised particle-swarm optimization algorithm. Computer Physics Communications, 247(??):??, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465519302966.

Zhuang:2020:SIP

[ZZZ^{+}20] Mingwei Zhuang, Qiwei Zhan, Jianyang Zhou, Zichao Guo, Na Liu, and Qing Huo Liu. A simple implementation of PML for second-order elastic wave equations. Computer Physics Communications, 246(??):??, January 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302541.